A Complete Bibliography of Publications in the Journal of Systems and Software (2020–2029)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

06 March 2025 Version 3.31

Title word cross-reference

2 [ZWF^{+}25]. ^{+} [FvB24]. ′ [CCZ^{+}24]. k [LMVRA^{+}20].

-nearest [LMVRA^{+}20].

161 [NBGC20a]. 19 [GMCA21, MBKL24, STM^{+}22, Tok22].

2018 [EHB21]. 2020 [Ano20a, Ano20b, Ano20c, Ano20p, Ano20q, Ano20s, Ano20r, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x]. 2021 [Ano21a, Ano21b, Ano21c, Ano21q, Ano21r, Ano21t, Ano21s, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z]. 2022 [Ano22a, Ano22b, Ano22c, Ano22p, Ano22q, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x]. 2023 [Ano23a, Ano23b, Ano23c, Ano23p, Ano23q, Ano23s, Ano23r, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x]. 2024 [Ano24a, Ano24b, Ano24c, Ano24o, Ano24q, Ano24p, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v]. 2025 [Ano25a, Ano25i, Ano25j, Ano25l, Ano25k, Ano25n, Ano25m]. 25012 [GRV^{+}21].

3 [RSL^{+}21].

5 [GKAHMO22].

61131 [RSL^{+}21]. 61131-3 [RSL^{+}21].

A/B [QWGS24]. abstract [BN23, JJS24, KKE21, PSZ21]. abstracted [YML^{+}22]. abstractions [AGL21]. academia [DdM^{+}24, WSL^{+}20]. academic [FdSN^{+}20, WCH^{+}22]. acceptance [Bat20, BbASP23, FFV^{+}23, LNST21]. access [LSB^{+}22, SLL20, SHWR22, SV20, WCH^{+}22, WFM^{+}24]. Accessibility [LSFE21, PFdMF21, ZFC^{+}22]. ACCORDANT [CVC21]. Accountability [KR23]. accuracy [HELW20, SHLM24]. Accurate [HLW^{+}23, KDB^{+}21, KGL^{+}22, WDXX23]. ACEOB [PSL25]. achieve [HPF23, ISW25]. Achieving [BBND^{+}20]. acknowledged [SB23]. ACM [GVBC^{+}24]. across [DMAH24, GSF24, LCT22, MWY^{+}22]. action [APB20]. Actionable [MDRdP^{+}24]. actions [IZAD21, TTL20, DMD23]. active [CBW^{+}23, HGHK24]. activities [DYZ^{+}23, HELW20, PPMC22]. activity [CDM25, DCMR20, KCMD21, NJF20, WFR21]. ad [FR20]. Adaptable [SSS21a, SSS21b]. adaptation [AZB25, AGM^{+}25, APAF21, CLLJ24, QWG22]. adapted [HO22, LYG^{+}24]. Adapting [IBP21]. Adaptive [AHT^{+}21, ACD^{+}21, BAH24, LWSZ23, MBP20, PtBD^{+}25, SG20, TLXW23, YLLY25, AAW20, ACPM22, AMRS20, BH20, BFHC20, CCZ^{+}24, GMS22, MDML25, MN23, PGW^{+}23, QWG22, RATS21, VWSCH23, WCGS23]. additional [VM24a]. address [BCLN21, HTC^{+}23]. Addressing [MPP24, KGM24, WLdCM22, ACC^{+}20]. adequacy [LGKT22]. admitted [IZAD21, YZL^{+}22, ZYZ^{+}23]. Adopting [BCJI22, AGPR20]. Adoption [ACM25, SvdBHV24, AF22, BGL^{+}22, CPD20, GBT^{+}20, LSFE21, MRA22b]. advanced [HTC^{+}23]. Advancing [GF24, HM24]. adversarial [FWZZ25, LXLZ20, QZS^{+}24, QHC^{+}24, TQT^{+}25, WDL^{+}25]. advertised [RFB20]. advertisements [KGM24]. advice [AWA^{+}22]. Aerial [ACA^{+}23, VGS^{+}23]. Affect [VBMB20, LNST21, NdOdO^{+}22, SMdS^{+}23]. affecting [AF22, MGSC22, Tok22]. Affective [BCF^{+}22, OBCR23]. Affirmed [AMO21]. after [FKGN23, ZFGH^{+}22]. against [TMV25, WCZL25]. agenda [ACM25, ECMC20, MMB22]. Agent [LLL^{+}25, PZVN^{+}25, DLBE22]. agents [LLL^{+}25, YRC^{+}25]. aggregated [RV22]. Aggregating [VM24b]. aggressiveness [BSS^{+}25]. Agile [BMHR21, LOBD25, OBCR23, ZH22, APB20, Bat20, BCLN21, BRO^{+}22, BCJI22, GAS25, Jør24, KKH^{+}21, KMAB20, LK23, LBMF^{+}22, Mam23, Mec24, NdSR^{+}21, RJ23, RJ25, SHA21, UPP^{+}22, VM24a, VM24b, WSL^{+}20, ZH21, AF22, EMCN^{+}22, OWGS23, VML21]. agility [BBND^{+}20]. Aging [ZXH^{+}24, JYZ^{+}24]. Aging-related [ZXH^{+}24]. agnostic [RAGCSS^{+}20]. agreement [BSS^{+}25, LPS^{+}23, DPGGP23, GPPDLF23]. agriculture [SSO^{+}23]. AI [CFI^{+}24, DMN^{+}23, DL21, DBO24, FEBO22, HKP23, KMS^{+}25, LDT^{+}25, MRM^{+}21, RPC^{+}24, RLK^{+}24, RBvB^{+}24, TGX^{+}22, VKJ^{+}21, ZLZ^{+}25]. AI-based [RLK^{+}24]. AI-enabled [RPC^{+}24]. AI-generated [CFI^{+}24]. AI/ML [KMS^{+}25]. aide [CA20]. Aiding [PCC^{+}20]. Airflow [YWTA25]. Alcohol [BIT24]. algebraic [LL23]. Algorithm [MLPC20, CHLT23, CCZ^{+}24, FR20, Hu25, PPL^{+}25, SVVD21, XBS21, YLW^{+}21, YZS22, YLLY25, YDP22, ZHLR23, dSIdÁ23]. algorithms [AA22, EM20, LL25, RF24b]. aligned [VKJ^{+}21]. alignment [CCR24, WKP20]. Allocating [KMS^{+}25]. allocation [NJF20, Pie20, ZHLR23]. along [vRMG23]. alternative [HM25, MS25]. AMASS [dlVRB21]. ambiguity [AWP^{+}25]. amBiguous [FGS23]. AmbiTRUS [AWP^{+}25]. AML [SHLM24]. aMLLibrary [SGG^{+}25]. AMon [VWSCH23]. among [AWA^{+}22]. analogy [GDK24]. analyses [AWMW20, HGH^{+}23, MSDC25, SVVD21].

Analysing [PtBD^{+}25]. Analysis [ASK^{+}23, EHB21, LHF22, MDAHS25, TPGH20, AvdBG^{+}24, ACC^{+}20, ACH25, AHFP24, ACG^{+}25, ANC^{+}23, APR24, AKMS23, BS23, BQR24, BSH^{+}20, CFP^{+}21, CDLN21, CSLN23, DdM^{+}24, DD20, DCÁ^{+}24, ETY^{+}22, FMF^{+}24, GCF22, GCSHB20, GFS21, GdOB25, GGMH23, HLL21, HOCK22, HS21, HSJB23, KDB^{+}21, KAA^{+}21, LHY^{+}25, LATV22, LHG^{+}25, LPS^{+}23, LSF^{+}25, LZP^{+}20, LTZ^{+}21, LL21, MVGHPT22, MFBP20, MCM^{+}24, MKB25, MCSAGB20, NdOdO^{+}22, PKGA22, PBC^{+}23, PCJNP23, SCdPL24, SH20, SCN^{+}21, SGG22, ST24, WXL^{+}20, WR22, YJZZ23, ZJXG20, ZAS25, ZGW^{+}20a, ZXW20, ZLG25, dCMM^{+}22]. analysis-based [ZJXG20]. analytical [HTC^{+}23]. analytics [Ala21, AR24, CVC21, FEBO22, WCZL25]. Analyzing [MCY23, NLS^{+}20, PKT24, SZ23, BPJ^{+}22, ECS23, HCB^{+}20, MAK^{+}21, MSB23]. anatomy [LMZ^{+}23]. Android [AIH20, ACC^{+}20, AJFD25, BF22, CATA21, CFAT25, CDN^{+}22, HMR21, LXL^{+}23, MCY23, MWY^{+}22, MREVEA^{+}22, MCH22, OAH^{+}23, PLP^{+}20, TPGH20, WXL^{+}20, WTS23, WCL23, XCH^{+}20, dCMM^{+}22, VKB^{+}25]. Android-specific [WCL23]. Andromeda [RPS25]. animation [LM21]. annealing [PCJNP23]. annotation [LH24a, RF23]. annotation-based [RF23]. anomalies [RB23]. Anomaly [RFB20, ASKS20, DFC^{+}23, PHLHM24, PMM^{+}25, YLG^{+}25]. Anomaly-aware [RFB20]. Ansible [OZVRD21]. answer [AA24, STW23, TXW^{+}20]. Ant [KKRT24, BH20]. Ant-colony [KKRT24]. anti [ASCR23, BKG20, CAA^{+}23, LMGK22, SJH^{+}25]. anti-patterns [ASCR23, BKG20, CAA^{+}23, LMGK22]. anti-phishing [SJH^{+}25]. anticipatory [PTW22]. antipattern [vDDPZ24].

Antipatterns [SC22, ABT^{+}22, TAT^{+}23]. anywhere [ŠMKGH23]. Apache [BGL^{+}20, CYW21, ISKB20, LQY^{+}22, LCY^{+}24, WCZW22, YWTA25]. API [HBSV^{+}22, KFJA23, LGMP24, MWD24, MCY23, NMT^{+}23, QWHH23, SWZ^{+}20, SZ22, WZFD25]. APIs [POWGH22, THN20, WZZ21, WZFD25]. app [AF22, GGB^{+}22, LXL^{+}23, NdOdO^{+}22, ZZH^{+}25, EKK^{+}24]. applicability [Win24]. Application [WGMT24, WXZL23, ZH21, BM25, DFT24, EM24, HSVMB20, HKP23, LLL^{+}22, MS25, SBF20, SSS22, VGS^{+}23, XCH^{+}20, ZSCD22, BGM^{+}21]. Applications [GRdAL23, KRSW22, PC23, ACC^{+}20, AHP21, AXUO24, AWMW20, BAH24, BF22, CAF22, CGM24, DDPP23, DJS^{+}22, ECL^{+}22, FPB^{+}24, FA24, GKB^{+}21, GGB20, GGMB^{+}22, GM20, HGHK24, IIK21, ISKB20, KRD^{+}23, KMS^{+}25, LSB^{+}22, LDH22, LDT^{+}25, LBCG23, NBGC20a, NBGC20b, OAH^{+}23, PKB24, POWGH22, RAGCSS^{+}20, RLK^{+}24, SGG^{+}25, SS23, SSO^{+}23, SMKI23, SCPB25, TAF^{+}20, VSTK21, VKG^{+}23, WCZW22, WGMT24, WGMT25, WLH^{+}25, WLdCM22, WTS23, XCZ23, XDL^{+}22, YZS22, ZFC^{+}22, ZNPR^{+}23, dSFdC^{+}25, SSS21a, SSS21b]. applied [MAP^{+}20, SKG^{+}24, YOH^{+}23]. Applying [DPGGP23, SGG24, YLC25]. appraisal [CFAT25]. appreciate [Fei23]. Approach [AFJ^{+}20, SVV24, AHP21, AA22, AS22, ALZ^{+}20, AHT^{+}21, BMP22, BRSR21, BEM^{+}23, BL21a, BL21b, BFZC21, CDRV20, CVC21, CBDK23, CFP^{+}21, CDET22, DBB20, EBAR21, FMG^{+}25, FFV^{+}23, GCLB22, GBSO20, HGHK24, HZ24, HKP23, Hu25, IIK21, KZTS22, KL20, KGS^{+}25, KKL^{+}21, KFJA23, KKE21, LWL20, LC20, LCS^{+}24, Liu21, LLH^{+}25, MOP24, MHR^{+}25, NMT^{+}23, NEJ24, NBGC20a, NBGC20b, PT21, PHLHM24, PZVN^{+}25, RPL^{+}21, RATS21, RF23, RHJ25, SAZN22, TTB^{+}23, WNZ25, WC20, WMS23, WC23, YJZZ23, ZCLP21, ZYZ^{+}23, dCMM^{+}22]. Approaches [LBCG23, ASM^{+}21, CFNA25, DL21, FAFJ24, HH22, HH25, PMM^{+}25, SCPB25]. approximate [LBF^{+}21a, LBF^{+}21b]. approximation [LBG^{+}20]. Apps [AJFD25, AAZB23, FBV24, GCF22, LCT22, MCY23, MWY^{+}22, MREVEA^{+}22, NdOdO^{+}22, NRFS25, PLP^{+}20, RB23, WXL^{+}20, WGL^{+}22]. April [Ano20a, Ano21a, Ano22a, Ano23a, Ano24a, Ano25a]. Arabia [AF22]. Arabian [AAZB23]. ARC [RFB20]. Architecting [BNMW23, EKB^{+}23, NBP24, MLS^{+}21]. Architectural [ETD^{+}24, Hei20, ISW25, LLZ^{+}22, MAK23, SBC20, dSS^{+}22, BCOC24, BB22, CWGS23, HOAM23, JKD25, LLL^{+}25, MTA24, MSRR22, MHR^{+}25, MSB21, OEW22, PFW21, RPT23, RHJ25, VKLM21, dTMS21, SZ22]. Architecture [ÅKH^{+}22, AFJ^{+}20, BDLT21, BLTX21, BDIL25, GGB20, KZTS22, MMO22, WLS20, AKL25, AGM^{+}25, BBG^{+}25, BMB20, ESO^{+}24, FJvdW20, GW23, GSK24, HKP23, KAW^{+}23, KGS^{+}25, KRS^{+}21a, KRS^{+}21b, LZP^{+}20, MAK23, MdSKD22, PLL^{+}23, RPC^{+}24, RCA^{+}22, SCD^{+}24, TAF^{+}20, WHK^{+}23, WL24, ZLC^{+}23, dDLSK23]. architecture-based [AGM^{+}25]. Architectures [SVV24, ATWS25, AÇCT21, BSH^{+}20, CVC21, GMFO^{+}21a, GMFO^{+}21b, GGPR24, MVSV25, MCMA21, MW25, RA25, dORGCG23, SMB^{+}20, WNZ25, ZTK^{+}23]. Archive [MCSAGB20]. Arduino [RPM^{+}22]. areas [FGRF24]. argument [NMT^{+}23, SSP21]. ARist [NMT^{+}23]. ARM [MCMA21]. ARRAY [TLXW23]. arrays [JT20]. art [BLHS23, Gir21, LHS^{+}24, PTW22, UPP^{+}22, AHT^{+}21, CCZ^{+}24]. Artcode [XTF^{+}21]. ARTe [RPM^{+}22]. artifacts [FPB^{+}24, FMF^{+}24, GVBC^{+}24, LHH^{+}24]. Artificial [MMO22, GPF22, SFR23].

ARTINALI [AAG21]. Aspects

[LK23, WNZ25, ZLS23]. ASPLe [AAW20]. assess [EBSB23]. Assessing [MDWS^{+}23, OKG25, POWGH22, CDRV20, WNZ25]. Assessment [AFJ^{+}20, CND22, CGM24, CDO24, CFI^{+}24, FC20, KRS^{+}21a, KRS^{+}21b, LHF22, PDDD20, RBS^{+}23, RPS25, SHG24, SvdBHV24, SZ22, SYB^{+}23, WJX^{+}24, WMAL21]. Asset [DMN^{+}23]. Assets [ZFGH^{+}22, ZGHG^{+}23]. assign [EBAR21]. assignment [MC20, WJX^{+}24]. assimilating [CFAT25]. assist [WLCC25]. assistant [LCY23]. assisted [WZC^{+}20]. association [CPC^{+}23]. associations [YOH^{+}23, ZJY^{+}24]. Assurance [dlVRB21, NNG21, OBW^{+}25, RPS25, WFM^{+}24]. AST [ZLD^{+}25]. AST-based [ZLD^{+}25]. Atlas [NMV25]. attacks [HCB^{+}20, QHC^{+}24, TMV25, WCZL25]. Attention [FZT^{+}22, OHM25, ZYL^{+}22, CFF^{+}23, HPZ^{+}20, HLZ^{+}23, YXLH24, ZFS^{+}22, ZLX^{+}22, ZSZ^{+}22]. Attention-based [OHM25, YXLH24]. attentive [NM25]. Attribute [LYZ^{+}22, ASSH22, MRVS24, WCGS23]. Attribute-Based [LYZ^{+}22]. Attributes [BFM^{+}25, BLTX21, CAC20, ISW25, MVŞ23, OAH^{+}23]. auditing [LKP^{+}21]. Augmentation [YWW22, Hu25, WZL^{+}25]. Augmented [YZM^{+}25, LZJ20, YZC^{+}23]. Augmenting [BH20]. August [Ano20b, Ano21b, Ano22b, Ano23b, Ano24b]. Authentic [HM25]. Authentication [AJFD25]. autism [CGM24]. Auto [ZLW^{+}24, ATWS25]. Auto-Completion [ZLW^{+}24]. auto-scaling [ATWS25]. automate [HG22]. Automated [BF22, CJR22, FZT^{+}22, GRdAL23, GGPR24, GS21a, GS21b, LSB^{+}22, NSL^{+}21, NZA^{+}25, PZDG21, SSS22, ZJXG20, ZHM^{+}23, ABMV24, BB22, CLZ^{+}24, GSF24, HBN^{+}25, IIK21, KBPMJ20, LCC^{+}23b, LLK^{+}21, LXL^{+}23, LZL^{+}25, MCZX20, MDVB^{+}23, MKB23, SAZN22, VGS^{+}23, WCGS23, WMS23, ZHS^{+}24]. Automatic [AMS23, FFV^{+}23, HSW^{+}24, JWZM20, OBW^{+}25, RPR22, RMM24, SRTGA^{+}25, SS23, SH20, SYB^{+}23, ZTK^{+}23, ZSZ^{+}22, AMO21, BDMP21, CKM^{+}20, CLNQ24, GJW^{+}22, HELW20, NLS^{+}20, WTS23, XXC^{+}24, YMDM21]. Automatically [CZLN22, KBBD23, FGG^{+}20, KZK22, LC20, OEW22, BGE^{+}21, MFBP20]. Automating [CFI^{+}24, SBM23, KKRT24]. automation [ABT^{+}22, BDLR23, BVHHO23, BGE^{+}21, ESO^{+}24, KBPMJ20, WMLM22]. automotive [BB22, GCLB22, KRS^{+}21a, KRS^{+}21b, MW25, Vog20, WKP20]. autonomic [WFR21]. Autonomous [KRD^{+}23, ALZ^{+}20, CDP22, YRC^{+}25]. autonomy [ŠMF^{+}23]. autoregressive [XZZ^{+}25]. Autosar [ZTK^{+}23]. availability [SH20, SCB22, VSTK21]. avant [SM21]. avant-garde [SM21]. avionics [GBT^{+}20]. aware [AZR^{+}21, AMB^{+}25, BAH24, CNdLL20, GTT^{+}20, HTB21, Hu25, JSTW22, LSG^{+}24, LBHW25, MdSKD22, QWHH23, RSLC24, RFB20, SCN^{+}21, TSLHS21, TAF^{+}20, WZZ21, WMH^{+}24, WTS23, YWCX24, YAaO^{+}21, ZZY^{+}24, ZGW^{+}20a, ZLX^{+}22, ZLSY23]. Awareness [MBG^{+}25, FPB^{+}24, SJH^{+}24, Vog20, ZGLPA24, LSFE21]. awry [FAA22].

B [QWGS24]. Babel [PKT24]. backdoor [QHC^{+}24]. Backsourcing [MLJ23]. bad [CAA^{+}23]. badges [GVBC^{+}24]. badges-driven [GVBC^{+}24]. bag [YZS22]. bag-of-tasks [YZS22]. balance [WKP20]. Balancing [XZZ^{+}25, DTZ^{+}22, PDDD20]. bandits [SG20]. Bangladeshi [SSO^{+}23]. banking [Mam23]. Barriers [GW24, LSFE21, ZFC^{+}22]. BASBA [SSS21a, SSS21b]. Based [BJB^{+}21, FZT^{+}22, LYZ^{+}22, SVV24, SSS21a, SSS21b, ASCR23, AC25, Ala21, AGM^{+}25, AXUO24, AS22, AJFD25, ALZ^{+}20, AON^{+}24, AWMW20, AFJ^{+}20, AR24, APAF21, BKG20, BRSR21, BEM^{+}23, BFHC20, BSH^{+}20, BHP^{+}21, CJR22, CCS20, CAF22, CBW^{+}23, CAC20, CA20, CJZ^{+}20, CLC^{+}23, CFF^{+}23, CZW20, CQZ^{+}20, CDM25, CDP22, CATA21, CFNA25, CDET22, CPD20, DLX^{+}23, DWG^{+}24, DD20, DFT24, DYZ^{+}23, DL21, DSO20, DZY^{+}23, DBB20, DHZ^{+}24, DLBE22, ESO^{+}24, EN23, FR20, FPB^{+}24, FC20, FBV24, FFSB23, GGB^{+}22, GKB^{+}21, Ger25, GTT^{+}20, GS21a, GS21b, GDWK24, GM20, HDX^{+}23, HZ24, Hei20, HLS24, HXJ^{+}20, HSJB23, ILUN21, IIK21, KSG^{+}22, KGS^{+}25, KKL^{+}21, KMS^{+}25, KBW^{+}24, LHY^{+}25, LATV22, LSB^{+}22, LBF^{+}21a, LBF^{+}21b, LWL20, LC20, LDH22, LWL^{+}23, LLW^{+}24, LBH^{+}24, LL25, LCS^{+}24, LXLZ20, LCC^{+}23b, LDT22, LCT22, LFFW23, LLL^{+}25, LZL^{+}25, LLH^{+}25, LCW^{+}24]. based [MRVS24, MAP^{+}20, MMDL23, MSC20, MAC^{+}22, MMSM^{+}23, MKB25, MBG^{+}25, MRA22b, MRR^{+}20, MdOT23, NMG^{+}20, NAEAK25, NAZ22, NNG21, NNV^{+}24, NDD^{+}24, NRFS25, NWZ25, OHM25, OWGS23, PBV24, PKB24, PDDD20, POZ20, RSU^{+}24, RPR22, RF24a, RFB20, RF23, RSVW23, RLK^{+}24, dORGCG23, RF24b, SLL20, SGG^{+}25, SGV22, SB23, SV25, SMM23, THB22, TTL20, TAT^{+}23, Tok22, TMV25, TGX^{+}22, VSTK21, WGL^{+}22, WHK^{+}23, WGMT24, WYFZ24, WCZL25, WNZ25, WYG^{+}24, WC20, WGY^{+}20, WC23, XCH^{+}20, XBS21, XCZ23, XLY^{+}21, YJZZ23, YLHZ20, YZC^{+}23, YZZ^{+}24, YWCX24, YXLH24, YAaO^{+}21, YB24, YYW^{+}24, YLLY25, ZJXG20, ZX23, ZZK^{+}24, ZXH^{+}24, ZHS^{+}24, ZWF^{+}25, ZHLR23, ZGW^{+}20b, ZLD^{+}25, ZYZZ21, ZB22, dSIdÁ23, vdG25]. BaSFuzz [LHY^{+}25]. Basilisk [BCOC24]. Bayes [OEW22, TLXW23]. BDD [IBP21]. Be [RBvB^{+}24, DMAH24, MCZX20]. beauty [HBSV^{+}22]. becoming [BHS24]. Behavior [ALSA21, ACA^{+}23, BbASP23, JCNS^{+}22, MWY^{+}22, PSZ21, IBP21].

Behavior-Driven [ALSA21]. Behaviour [BM23, ZLS22]. behavioural [LB24]. being [Fei23, HO22, PPMC22, Tok22]. belief [HBK25]. Benchmark [FA24, BG24, DSB23, KH24, SM24, WUK^{+}21, YMDM21, ZMLZ23]. Benchmarking [HH24, ZYK^{+}25]. benchmarks [HH22]. Benefiting [FAG^{+}20]. Benefits [CRV23, BDLR23, KOPN22, TH24]. BERT [AXUO24]. Best [HH25, NBP24]. Between [THN20, AA23, BRSK24, BFHC20, CXH^{+}24, CPC^{+}23, CJNDZ21, DCÁ^{+}24, IZAD21, KdJPK^{+}23, MLPC20, PFW21, Spa24, SMM23, VM24b, WWC23, YOH^{+}23, ZZL^{+}23, ZXZS24, dSIdÁ23]. Beyond [BRSK24, WLdCM22]. BFT [NMV25]. bi [JSG25, ZPL21, dSIdÁ23]. bi-modal [JSG25]. bi-objective [ZPL21, dSIdÁ23]. bias [KGM24, OKG25, RFS^{+}21]. bibliometric [WMAL21]. bidirectional [BBW22, HZ24, ZZY^{+}24]. big [CBZZ24, CVC21, ISKB20, AX21]. big-data [CBZZ24]. big.LITTLE [MCMA21]. bilingual [YOH^{+}23]. binary [SK22a, ZLW^{+}24, ZXW20]. Bioconductor [Gu23]. Biometric [VBMB20]. bit [SGW^{+}22, HZ24]. BiTCN DRSN [CLC^{+}23]. Black [LXLZ20, GBSO20, LZ24]. Black-box [LXLZ20, LZ24]. bleeping [YOH^{+}23]. Blended [AC21]. BLEU [EBSB23]. block [AC25, FYH^{+}24]. block-based [AC25]. block-level [FYH^{+}24]. blockchain [BLHS23, BMP22, LLZ^{+}23, LKP^{+}21, VDVC21]. Board [Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano25b, Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h]. body [SCPB25]. boilerplate [LCS^{+}24]. Boolean [WYFZ24]. Boolean-specification [WYFZ24]. boost [VVBGL^{+}23]. Boosting [CXP^{+}23, HHZW23]. Boot [MCH22]. bot [CDM25]. both [GFS21]. bots [GDLM21, VCT20]. boundary [LCW^{+}24, RA25, ZB22]. boundary-condition [LCW^{+}24]. bounds [ACG^{+}21]. box [CBC^{+}24, GBSO20, LXLZ20, LZ24]. BPEL [DOG^{+}23]. BPMN [CMP^{+}20, CFP^{+}21, DPRD21, VTS22]. Brain [WFR21]. Brazil [LSFE21]. Brazilian [GBLP25, PFC^{+}23]. breadth [AES22]. Breaking [FEBO22, KVP23]. Bridging [GDK24]. Browsers [GRdAL23]. Bug [LCY^{+}24, TXW^{+}20, AXUO24, AWHS22, DLX^{+}23, EBAR21, FGG^{+}20, JCNS^{+}22, JST^{+}23, KRP24, KCMD21, KGL^{+}22, LYS^{+}23, LSSZ21, MCM^{+}24, MCRM25, NLS^{+}20, PPB20, QZS^{+}24, SK22a, TSLHS21, WLC^{+}20, WCH^{+}23, WGMT24, WJX^{+}24, WGMT25, WZC^{+}20, YWCX24, ZXH^{+}24, ZLS20, ZLSY23]. bug-specific [ZLS20]. BugOss [KH24]. bugs [ABMV24, AA22, JZW^{+}21, KH24, KZK22, MLYL25, NZA^{+}25, RF23, ZKDP22, ZMLZ23, ZRGJ21]. Bugs4Q [ZMLZ23]. build [JS22, MOP24, MAC^{+}22]. Building [VKLM21, AZ21, RAGCSS^{+}20, TAF^{+}20, WZZ21, SSS21a, SSS21b]. buildings [HCB^{+}20]. builds [BEAK21, JS22]. Burr [LDH22]. Burr-type [LDH22]. business [FEBO22, JVZ25, NQT25, RK20, TAF^{+}20, VTS22]. business-aware [TAF^{+}20]. businesses [BGL^{+}22]. BXtendDSL [BBW22]. bytecode [ACG^{+}25, HLW^{+}23, PBC^{+}23]. bytes [LHY^{+}25].

C [Ger25, PSGD23, SSP21]. C-based [Ger25]. CafeOBJ [RO22]. Calculus [ACD^{+}24, ADS^{+}22]. Call [SMH^{+}23, AGL21, VHSB22]. calls [CBMM20]. Can [BTSC^{+}23, HG22, MCZX20, WUK^{+}21, ABMV24, DMAH24, GAL20]. Candidate [AHT^{+}21, SK22a]. candidates [NMRS22]. capabilities [PSL25, PMM^{+}25, VM24a]. capability [VML21]. capture [EMCN^{+}22, JCNS^{+}22]. Capturing [DBB20, FJvdW20]. care [SK22a]. Carlo [HGH^{+}23]. cascading [HCB^{+}20]. Case [WFM^{+}24, AIH20, ASSH22, AAZB23, BRO^{+}22, BFZC21, BMHR21, BDMP21, BGL^{+}20, CACHA23, CHLT23, CQZ^{+}20, CDO24, CID25, ECL^{+}22, EKK^{+}24, FEBO22, FFV^{+}23, FKF^{+}23, FA24, GGP21, HG22, HLS24, HZT^{+}20, IMTS23, KCMD21, LH24b, LWYW20, LCS^{+}24, MAP^{+}20, Mam23, MLPC20, MAV24, MCH22, MAS23, PN21, SRTGA^{+}25, SH20, SMB^{+}20, SB23, aSLF^{+}22, TST^{+}21a, TST^{+}21b, WR22, YOH^{+}23, YWTA25, YU22, ZFS^{+}22, dTMS21, dSNJ25, vRMG23]. case-split [CHLT23]. case-study [YOH^{+}23]. cases [KBBD23, MAP^{+}20, NNG21, OBW^{+}25, SKP20, SSS22]. CASL [CBW^{+}23]. casts [PSGD23]. Catalog [CAA^{+}23, DDPT20]. Cataloging [LMGK22]. catalogue [ASCR23, LLL^{+}25]. categorization [MSRR22, NRRS20]. categorize [MICV23]. categorizing [AHL22]. Causal [XCZ23]. causality [Bro24, DWBA24]. CausalRCA [XCZ23]. Cause [ZWL^{+}24, BSH^{+}20, LCC^{+}23a, NLS^{+}20, XCZ23]. caused [LGC^{+}25]. causes [BVHHO23, FGRF24, JZW^{+}21, RMT^{+}22, TRD^{+}23, ZLZ^{+}25, TSP20]. CCStokener [WDXX23]. cell [HXJ^{+}20]. cells [MDWS^{+}23]. center [PKS25]. Centered [RBvB^{+}24]. centers [MH20]. Centralization [MW25]. Centric [WFM^{+}24, LBCG23, MDVB^{+}23, NNV^{+}24, RAGCSS^{+}20]. certificate [MBKL24, WXL^{+}20]. certification [GGP21, GRV^{+}21, dlVRB21]. CfExplainer [YZZ^{+}24]. CG [Hu25]. CG-FL [Hu25]. challenge [GAB20]. Challenges [BM25, GCSHB20, Hyr25, LGMP24, MR24, MBO^{+}22, NBP24, BLHS23, BGC20, BGL^{+}22, CDET21, Gir21, HWL^{+}25, KKH^{+}21, LPPG20, dPPdJFF21, SJH^{+}24, SJH^{+}25, SWH^{+}20, VDVC21, WCZW22, YWTA25, ZNPR^{+}23]. Change [KAA^{+}21, LHG^{+}25, MBP20, DFT24, JWZM20, JSG25, LWZ21, LCY^{+}24, MSRR22, NMRS22, OZVRD21, WCZL25, ZWP^{+}24]. changed [JZM21]. Changes [STM^{+}22, AKMS23, KVP23, KGS23, LST20, MCY23, RVAC^{+}25]. changing [TH24]. channels [ZWP^{+}24]. Characterising [HTN^{+}25]. Characteristics [BVHHO23, Jør24, MAS23, MSB23]. characterization [ABT^{+}22, OSJB21, WDF^{+}22, ZKDP22, BJB^{+}21]. characterize [ALB20]. Characterizing [Hor21, LGC^{+}25, dDLSK23, MSS21, dORGCG23]. CharM [dORGCG23]. charts [CWTL21]. chat [SHB21]. ChatGPT [FG24, NDD^{+}24]. checking [CBC^{+}24, CMP^{+}20, DSO20, DLBE22, GAL20, LZ24, MDVB^{+}23, PBV24, PPL^{+}25, ZWC^{+}24, dSS^{+}22]. Checklists [MBP23]. checkpoint [GSF24]. checks [CI24]. children [CGM24]. choices [ACH25]. Choosing [DBO24]. CIPAC [LZL^{+}25]. circle [FEBO22]. circuit [BDS^{+}24, OPF24]. CIT [MAC^{+}22]. CIT-daily [MAC^{+}22]. city [MLGBRC24, MCDP24, SCD^{+}24, dSFdC^{+}25]. class [ASSH22, BHHQ^{+}22, BN23, CZLN22, DZY^{+}23, LWYW20, PGR24, RV22, RPL^{+}21, SKHLS24, ZX23]. class-level [BN23, LWYW20, RV22]. classes [ZYZ^{+}23]. Classical [HPP24, DCÁ^{+}24, QWG22]. classification [AMO21, CKM^{+}20, CLNQ24, DFT24, GDLM21, HELW20, KBPMJ20, MHOM22, MCRM25, NDP^{+}21, NLS^{+}20, OZVRD21, RPL^{+}21, SC22, SKHLS24, XTF^{+}21, YLW^{+}21, YSB^{+}21, ZAS25, ZLX^{+}22, ALB20]. classifier [NDD^{+}24]. classifiers [ZB22]. Classifying [LFH^{+}22, LC20, RSU^{+}24]. click [XDL^{+}22]. client [HBSV^{+}22]. clients [HBSV^{+}22]. climate [VM24a, VM24b]. clinical [AAZB23]. Clique [FC20]. Clone [AS22, LLL^{+}22, MRS20, MRR^{+}20, NMRS22, NAEAK25, RSL^{+}21, WDXX23, ZNPR^{+}23, ZFG^{+}23, ZYK^{+}25, CTD^{+}23]. cloned [NMRS22]. clones [ES24, BSGN21, vBD21]. closure [Mas22]. Cloud [AMB^{+}25, RFB20, TTL20, BBF^{+}21, CNdLL20, CRX24, CDLN21, CSLN23, EM20, FPB^{+}24, GGB20, GTT^{+}20, HOAM23, HH24, ISKB20, KMS^{+}25, KBW^{+}24, MdSKD22, MYH^{+}24, NAV^{+}24, TGX^{+}22, WZZ21, WCZL25, ZPL21, MH20]. cloud-based [FPB^{+}24, KBW^{+}24, WCZL25]. Cloud-integrated [RFB20]. clouds [CYW21, HTB21, LLWL22, YZS22]. cluster [HXJ^{+}20]. Clustering [TCA22, FR20, MPGB22, SXL^{+}22, XLY^{+}21, YDP22, ZWF^{+}25]. clustering-based [XLY^{+}21]. clusters [SCB22]. CM [CBW^{+}23]. CM-CASL [CBW^{+}23]. CNN [LWP^{+}21]. Co [HTG^{+}24, AA22, GdOB25, IZAD21, JZM21, NMRS22, OAH^{+}23, RA25, THB22]. co-analysis [GdOB25]. co-change [NMRS22]. co-changed [JZM21]. co-engineering [RA25]. co-evolutionary [AA22]. Co-Location-Resistant [THB22]. co-occurrence [IZAD21]. co-occurrent [OAH^{+}23]. Co-simulation [HTG^{+}24]. Code [CND22, DXL^{+}24, EHB21, KA24, LPPG20, NNV^{+}24, SELS21, VBMB20, ZZK^{+}24, ZYL^{+}22, AA23, ASKS20, ACC^{+}20, ACH25, ACM25, AES22, ANC^{+}23, BSGN21, BRSK24, BN23, CLZ^{+}24, CKY25, CJZ^{+}20, CXH^{+}24, CTD^{+}23, CFI^{+}24, DN21, FPB^{+}24, FRC24, FFSB23, GJW^{+}22, GSM^{+}23, GDWK24, GIP^{+}24, HMR21, HS21, HBN^{+}25, HELW20, Hor21, HPZ^{+}20, HLZ^{+}23, HHJ^{+}20, HZT^{+}20, HM24, HYF^{+}24, HHZW23, JSTW22, KBPMJ20, LALM23, LLL^{+}22, LNST21, LWP^{+}21, LYG^{+}24, LDT22, LAL^{+}23, LZ25, MAP^{+}20, MVSV25, MFBP20, MCM^{+}24, MICV23, MCH22, MHOM22, MLBD21, MRR^{+}20, NAEAK25, NM25, NAD^{+}20, NDD^{+}24, OSM^{+}23, OEW22, DDPT20, PZDG21, PSL25, PDDD20, PCC^{+}20, PdS23, PÁngeles MoragaGC25, RBS^{+}23, RPC^{+}24, RB23, SK22b, SAZN22, SKG^{+}24, SRD^{+}21, SMVC24, TTB^{+}23, TOO^{+}23, TST24, WDXX23, WTG23, XZZ^{+}25, YLC25, YZC^{+}23, YWS23, YXLH24, YU22, YWLZ23, YLGS24, ZNPR^{+}23, ZCLP21, ZFG^{+}23, ZZY^{+}24]. code [ZPSW24, ZSZ^{+}22, ZWP^{+}24, ZZP21, ZYK^{+}25, ZLS25, dSNJ25, vBD21, FAG^{+}20, EBSB23, SWG^{+}20, WUK^{+}21, YWTA25]. Code-centric [NNV^{+}24]. code-driven [CTD^{+}23]. Code-quality [CND22]. codebases [AvdBG^{+}24]. CodeBERT [NDD^{+}24, YZC^{+}23]. CodeBERT-based [NDD^{+}24]. Coder [GPPDLF23]. CoDEvo [SOMSCT23]. coding [ACSJ23]. cognitions [LWC20]. Cognitive [LALM23, PdS23, AA23, EN23, KL20]. Cohen [PDGMT20]. coincidental [SKP20]. Colla [BEM^{+}23]. Colla-Config [BEM^{+}23]. collaborating [ACA^{+}23]. collaboration [CFP^{+}21, HSJB23, ROL21]. collaborations [CMP^{+}20, WSL^{+}20]. Collaborative [DAML23, APR24, BEM^{+}23, CBW^{+}23, DLX^{+}23, DPGGP23, FCW^{+}23, LXYL20, MBG^{+}25, RCA^{+}23, SGW^{+}22, SZS25]. collect [MN23]. collection [LLL^{+}25]. collective [ACD^{+}24, LZL^{+}25]. collectivism [ZLS22]. Colombia [MVGHPT22]. colony [BH20, KKRT24]. colored [APR24]. Column [SOMSCT23]. combination [LXYL20]. combinations [HZT^{+}20]. Combinatorial [MDML25, BMLH25, BG24, DYZ^{+}23, JT20, MPP24, MAC^{+}22]. Combine [CLZ^{+}23]. Combining [XXC^{+}24, BBW22, CYW21, TTB^{+}23]. COMET [MRS25]. comment [AC25, LWP^{+}21, RPL^{+}21, RBS^{+}23]. commenting [HHJ^{+}20]. comments [AC25, BGE^{+}21, BSGN21, BRSK24, GDLM21, RMOGA20, STW23]. Commit [SKHLS24, DCMR20, ETY^{+}22, HELW20, MRS25, MCRM25, PMDN20, WDL^{+}25]. Commit-time [SKHLS24]. commits [AES22, LQY^{+}22, ZXZS24]. common [ES24, BHS24, RMT^{+}22]. CommtPst [HHJ^{+}20]. Communication [LOBD25, AÇCT21, AHL22, BDLT21, GS20, LCT22, SGG22, SGG24]. Communitie [LCP^{+}24]. community [DdM^{+}24, ES23, FVDF21a, FVDF21b, GW23, KR23, PT21]. community-wide [FVDF21a, FVDF21b]. companies [DBO24, GFCLB^{+}24, VHB21, ZH21, ZH22]. companion [RPT23]. company [APB20, VM24b, dSNJ25]. comparative [XLY^{+}21, ZGW^{+}20b, vBD21]. compared [WGL^{+}22]. Comparing [CTD^{+}23, KGS23]. Comparison [CBW^{+}23, EFPC21, JLL23, KRLS^{+}20, LPS^{+}23, SGV22, TGG21, vdG25]. Comparison-based [CBW^{+}23]. comparisons [Tai24]. CoMPers [SZS25]. competence [AWA^{+}22]. competencies [AGP22, HNM^{+}24]. competing [BFHC20]. compilers [vdG25]. complement [JWZM20]. complete [ZGW^{+}20a]. Completion [ZLW^{+}24]. Complex [AAG21, EKB^{+}23, LWSZ23, MCSAGB20]. Complexity [LALM23, AM23, Tai20]. compliance [ARPDP24]. Component [CCS20, BRS23, DSO20, Hei20]. Component-based [CCS20, DSO20, Hei20]. components [BMB20, BGL^{+}22, LMZT22, MHR^{+}25]. composite [WZZ21]. composition [CZW20, RFB20, RFB^{+}22]. compositional [HKP23]. Comprehending [MCDP24]. comprehension [AGL21, GIP^{+}24]. comprehensive [BGC20, DdM^{+}24, NEJ24, RPS^{+}23, SXL^{+}22, XLY^{+}21, YWS23, YMDM21, MCSAGB20]. compression [GLCH25]. computation [Mas22]. Computational [BFL23, AM23, DCÁ^{+}24, SGW^{+}22]. computer [HM25, HTC^{+}23, OKG25, YOH^{+}23]. computing [AMM25, CDLN21, CSLN23, DCÁ^{+}24, DHM^{+}21, DTZ^{+}22, GTT^{+}20, GM20, KAW^{+}23, KI23, MPGB22, PBC^{+}23, PGR24, SGG^{+}25, TGX^{+}22]. Concept [WZZ21, AC21, ROL21]. Concepts [LSBG21]. conceptual [BEAK21, HGH^{+}23]. concern [ZZH^{+}25]. concerns [BFHC20, LFLS24, LTJ^{+}20]. concolic [CI24]. concrete [KKE21]. concurrency [KZK22, QZS^{+}24, ZLD^{+}25]. concurrent [DLCH25, ZWC^{+}24]. condition [LCW^{+}24]. conditional [ZXW20]. conditionals [FFV^{+}23]. conditions [AHP21]. confidence [BHP^{+}21, TGG21]. confidence-based [BHP^{+}21]. confidentiality [Ala21]. confidentiality-based [Ala21]. Config [BEM^{+}23]. configurable [CBDK23, FVDF21a, FVDF21b, FKF^{+}23, LABJ23, LSG^{+}24, LBH^{+}24, LBHW25, MAV24, SZS25]. Configuration [LBHW25, BEM^{+}23, CBZZ24, CYW21, PAM^{+}21, SWH^{+}20]. Configurations [AFJ^{+}20, MOP24]. Configuring [ACA^{+}23]. Confix [XXC^{+}24]. conflict [LWC20, LLW^{+}23, LCW^{+}24, SZS25]. ConflictBench [SM24]. conflicts [AMS23, DBM^{+}24]. Conformance [SZ22, dSS^{+}22]. confusing [TOO^{+}23]. connected [RPS25]. connectedness [AWA^{+}22]. connections [Jør24]. connectors [AKL25]. consensus [BLHS23, MN21b]. consequences [BVHHO23]. CONSERVE [JRM^{+}22]. Considerations [BGL^{+}22]. considering [DZY^{+}23, TPV25]. consistency [AR24, CI24, KKL^{+}21, ZWC^{+}24]. consistent [WLH^{+}25]. constants [FC20]. Constrained [JT20]. constraints [CHLT23, KMS^{+}25, MDVB^{+}23, Spa24]. construct [BFL23]. constructed [IMTS23]. construction [CLZ^{+}24, LZL^{+}25, MC20, RAGCSS^{+}20, RF24b, WZC^{+}20]. constructs [PSZ21]. consultation [EKHJ^{+}20]. consumption [DLV^{+}22, PÁngeles MoragaGC25, TBD^{+}23]. contact [GCF22]. contact-tracing [GCF22]. container [BQR24, KBP24, RAGCSS^{+}20, TBD^{+}23]. container-centric [RAGCSS^{+}20]. containers [JRM^{+}22, LWZ21, RRFD24]. content [CA20]. Context [ZZY^{+}24, AdlBGZ^{+}23, ECS23, FdSN^{+}20, HLL21, Hu25, MRS25, MMDL23, RF24a, YWCX24, YLGS24, ZWC^{+}24]. Context-aware [ZZY^{+}24, Hu25, YWCX24]. context-free [RF24a]. context-oriented [ECS23, MMDL23]. context-sensitive [AdlBGZ^{+}23, HLL21]. contexts [CFI^{+}24, OWGS23]. Contextual [ZYL^{+}22, DFB20, MHR^{+}25, RSU^{+}24, SG20]. Contextualizing [PMDN20]. continental [ZLG25]. continua [SGG^{+}25]. continue [Mam23]. Continuous [GAB20, SB23, TC22, ÅKH^{+}22, CDET22, ETD^{+}24, JS22, LWYW20, SFR23, WMLM22, YLHZ20]. continuum [KMS^{+}25]. contract [CLZ^{+}23, CLZ^{+}24, CFF^{+}23, LFFW23, NNG21, VDVC21, VCT20, WCH^{+}23, YAaO^{+}21, YWLZ23]. contract-based [NNG21]. Contracts [OHM25, ACG^{+}21, DLV^{+}22, HSW^{+}24, LKP^{+}21, PBC^{+}23, VIL24, YML^{+}22]. contribution [YCWD23, ZLS25]. contributions [CND22, KHEC^{+}23, SWG^{+}20]. contributors [ZLG25]. control [ALZ^{+}20, CWTL21, CLLJ24, CCR24, FMG^{+}25, HIDT21, KBPMJ20, LSB^{+}22, LSBG21, MLPC20, Mas22, SHWR22, WM24, PBC^{+}23]. Control-Flow [PBC^{+}23]. controlled [DFB20, KZK22, MDWS^{+}23, PSZ21, WMS23, ZMLZ23]. controller [KBP24, VSTK21]. controllers [NMG^{+}20]. controls [ARPDP24]. convergence [LLW^{+}23]. Conversation [WDT^{+}24]. conversion [Spa24, XZZ^{+}25]. convolution [JYZ^{+}24]. Convolutional [LFFW23, NDP^{+}21, AAB^{+}22, CLC^{+}23, CFF^{+}23]. cooperation [CA20]. cooperative [KRS^{+}21a, KRS^{+}21b]. coordinated [ZZTC23]. coordination [RCA^{+}22, SM20b]. Copenhagen [RBvB^{+}24]. Copilot [DMN^{+}23]. Coping [SRS^{+}25, TSP20]. COPS [YWCX24]. Corona [EKK^{+}24]. Corona-Warn-App [EKK^{+}24]. corporatization [NF24]. correct [SKP20, SAZN22]. correction [LTZ^{+}21]. Correctness [CMP^{+}20, CFI^{+}24]. correlates [KMAB20]. correlation [WWC23]. correlations [PFW21]. Corrigendum [NBGC20a]. cosine [ZHLR23]. Cost [AMB^{+}25, GGMB^{+}22, ISKB20, AHT^{+}21, BS23, EM20, FPB^{+}24, RJ23, TQT^{+}25, ZWF^{+}25]. Cost-effective [GGMB^{+}22, AHT^{+}21, ZWF^{+}25]. Cost-efficient [ISKB20, TQT^{+}25]. costs [BDLR23, WL24]. count [LDH22]. counterfactuals [YZZ^{+}24]. coupling [ZZL^{+}23]. course [EKHJ^{+}20, SMM23, YU22]. courses [HM25]. coursework [EKHJ^{+}20]. covariate [NJF20]. cover [BH20]. Coverage [WLL^{+}24, BMLH25, GF24, HZT^{+}20, KBG^{+}24, MAP^{+}20, RF24b, ZZTC23]. coverage-directed [RF24b]. Coverage-guided [WLL^{+}24]. covering [DWG^{+}24]. COVID [GCF22, GMCA21, MBKL24, STM^{+}22, Tok22]. COVID-19 [GMCA21, MBKL24, STM^{+}22, Tok22]. CPS [BHP^{+}21, LZB^{+}23]. CRAN [MCSAGB20, Gu23]. CRAN/Bioconductor [Gu23]. crashing [XZY^{+}20]. created [FGG^{+}20]. creating [HKP23]. creation [FFV^{+}23]. creative [DBB20]. Creativity [GBMF22]. criteria [DWH23, ÖS20, ST24]. critical [APR24, CDO24, HIDT21, JLL23, KBBD23, LPS^{+}23, LLK^{+}21, MSKS24, SJH^{+}24, WFM^{+}24, WR22]. criticality [CDO24]. Cross [CCC^{+}24, DJR^{+}22, QZS^{+}24, ES24, BS23, BL21a, BL21b, DXD25, LMR^{+}23, NRRS20, TQT^{+}25, TLXW23, ZLG25]. cross-continental [ZLG25]. Cross-Domain [DJR^{+}22]. cross-language [ES24, LMR^{+}23]. cross-modal [TQT^{+}25]. cross-platform [BL21a, BL21b]. Cross-project [CCC^{+}24, QZS^{+}24, BS23, DXD25, TLXW23]. crossover [LLW^{+}23, MPP24]. CrossRec [NDDD20]. Crowd [THN20, MRA22b, NIJ22]. crowd-based [MRA22b]. Crowdsourced [ALSA21, LFH^{+}22]. crowdsourcing [BGC20, ILUN21]. CSAT [LBHW25]. CSGVD [TTB^{+}23]. CSP [LZP^{+}20]. CSVD [CCC^{+}24]. CSVD-TF [CCC^{+}24]. CTL [ADS^{+}22]. cue [QWHH23]. cues [ZLG25]. Cultural [LCP^{+}24, YRC^{+}25, ZLS22]. culture [BHS24]. Current [SFR23, LLL^{+}22, ECMC20]. Custom [RSL^{+}21]. Custom-tailored [RSL^{+}21]. cuts [GFS21]. CVE [WZC^{+}20]. CVE-assisted [WZC^{+}20]. Cyber [AAG21, FMG^{+}25, SSZ20, vDDPZ24, ACD^{+}21, BRS^{+}22, BGM^{+}21, BCW21, DSM20, FKF^{+}23, GB20, GCLB22, GAB20, LTJ^{+}20, MFF^{+}24, PPL^{+}25, RATS21, SCN^{+}21, dlVRB21, BJB^{+}21, VWSCH23, ZKDP22]. Cyber-Physical [AAG21, FMG^{+}25, SSZ20, ACD^{+}21, BGM^{+}21, BCW21, DSM20, GB20, GCLB22, GAB20, LTJ^{+}20, PPL^{+}25, SCN^{+}21, dlVRB21, BJB^{+}21]. cybersecurity [ZGLPA24]. cycle [RK20, RATS21]. cyclic [LHN20].

D [ZWF^{+}25]. D-kernels [ZWF^{+}25]. DAG [vdG25]. DAG-based [vdG25]. daily [MAC^{+}22, MAC^{+}22]. dark [Mec24]. dashboards [MDAHS25]. Data [FGRF24, GRV^{+}21, LTZ^{+}21, LYZ^{+}22, MBO^{+}22, RJ23, ROL21, SPZW25, VKG^{+}23, YWW22, ZZK^{+}24, AMHB24, BHS24, BAH24, CBZZ24, CVC21, CCR24, CLNQ24, DWG^{+}24, DJS^{+}22, FC20, HGHK24, HOCK22, Hu25, ISKB20, KBW^{+}24, LDH22, LLWL22, LCC^{+}23a, MHH21, MH20, NHA20, NBGC20a, NBGC20b, Pat20, PDDD20, PFC^{+}23, PMDN20, ROGS^{+}24, SHWR22, WZL^{+}25, WYG^{+}24, ZHM^{+}23, ZPSW24, AX21]. Data-driven [RJ23, BHS24, CCR24, HGHK24]. data-related [FGRF24]. Database [Tai24, ALB20, AHP21, BMB20, LMZ^{+}23, RVAC^{+}25, SOMSCT23, Tai20, TGG21]. datacenters [AZR^{+}21]. dataflow [SS23]. dataflows [LtBD24]. datamorphic [ZB22]. dataset [CLZ^{+}24, DFC^{+}23, FGG^{+}20, FVDF21a, FVDF21b, FA24, GDLM21, WUK^{+}21, WZC^{+}20]. days [DL21]. DB [RVAC^{+}25]. DBT [WDF^{+}22]. DDoS [WCZL25]. deal [GPF22]. Dear [Woo25]. Debt [WB23, AAB^{+}22, BGMB20, BVHHO23, FAG^{+}20, FRP^{+}23, GAS25, HTN^{+}25, IZAD21, KRLS^{+}20, LLST20, LBT^{+}21, MVSV25, MCDP24, RMT^{+}22, RPC^{+}24, TKSC20, VKLM21, VCF22, VHB21, WMA^{+}24, ZYZ^{+}23, dTMS21, MFLS22]. debts [OMA^{+}22, YZL^{+}22]. debugging [HH22, HCC22, MHJW22, SXL^{+}22, WTS23, ZMLZ23]. decade [RBS^{+}23]. decades [GMFO^{+}21a, GMFO^{+}21b, SGM23]. December [Ano20c, Ano21c, Ano22c, Ano23c, Ano24c]. Decentralized [ŠMF^{+}23]. decision [AKBN20, BS23, HBK25, LTG^{+}25, MBP23, RPT23, ŠMF^{+}23, SMdS^{+}23, ZXW20]. decision-making [MBP23, RPT23, ŠMF^{+}23]. decisions [KRC23, NBP24, PMDN20, YSB^{+}21, SZ22]. declarative [BBW22]. decoder [ZZY^{+}24]. decompilation [HSVMB20]. decompiler [HSVMB20]. decrease [LLST20]. dedicated [MDAHS25]. deductive [AHP21, LH24a]. Deep [BN23, GDWK24, HBN^{+}25, JST^{+}23, LLL^{+}22, OHM25, TCTL25, ZZK^{+}24, ZXZS24, ZSF^{+}23, AAB^{+}22, CDLN21, DXL^{+}24, FWZZ25, GBK^{+}23, JZW^{+}21, KRP24, LXYL20, MFPB24, MBO^{+}22, NWZ25, QHC^{+}24, SELS21, TTB^{+}23, WLL^{+}24, WJX^{+}24, ZWF^{+}25, ZLS20, ZYZZ21, GTT^{+}20, HHJ^{+}20, YZM^{+}25]. DeepCover [GF24]. DeepFeature [FWZZ25]. DeepKernel [ZWF^{+}25]. defeat [BCOC24]. Defect [ALB20, ASKS20, AKBN20, BS23, CXP^{+}23, DXD25, ET21, GBK^{+}23, GS21a, GS21b, HXJ^{+}20, JSG25, LAL^{+}23, LMVRA^{+}20, MS25, NM25, NQT25, SKHLS24, TLXW23, XLY^{+}21, YZZ^{+}24, ZJXG20, ZJY^{+}24, ZSCD22, ZYZZ21]. Defects [THN20, ALB20, DOG^{+}23, GS21a, GS21b, MS25]. defense [TMV25]. define [ZLS25]. Definition [KOPN22]. definitions [MC20, THG20]. Delay [GS20]. Delivering [PGR24, CCR24]. Delphi [ZH21]. delta [MRS25]. demand [DLCH25, RPT23]. density [HELW20, LMVRA^{+}20]. dependability [LHS^{+}24]. dependable [SJC^{+}22]. Dependence [DLCH25, ZYL^{+}22, LBG^{+}20]. dependencies [ACC^{+}25, HBSV^{+}22, MOP24, Vog20]. dependency [Gu23, HG22, LMGK22, LBF^{+}21a, LBF^{+}21b]. deploy [XDL^{+}22]. deployed [AM23, HH24]. Deployment [AÇCT21, AFJ^{+}20, KRSW22, CAF22, RFSP^{+}25]. depth [AES22]. derivates [MLGBRC24]. derived [SZ22]. Deriving [AvdBG^{+}24]. Description [FZT^{+}22]. Descriptions [RMOGA20]. descriptive [WC20, WC23]. Design [BG24, LHS^{+}24, SBC20, SZ22, WLS^{+}21, ACD^{+}24, ARPDP24, BRSR21, CWGS23, CACHA23, CGM24, FJvdW20, FNB^{+}24, GGB20, IR21, LH24b, LLL^{+}25, LLZ^{+}22, MTA24, NAZ22, NBP24, NNG21, PPMC22, PFW21, PÁngeles MoragaGC25, SJH^{+}25, SZSV22, ZFC^{+}22, CDRV20]. design-pattern [BRSR21]. designs [HQNR^{+}22, MPP24, SZ22]. detect [AA22, KZK22, NDD^{+}24, WC20]. detected [MCZX20]. Detecting [AJFD25, ES24, LDT22, SHWR22, DBM^{+}24, WLLJ24, YML^{+}22, BN23, GDLM21, LGC^{+}25, NMRS22, YZL^{+}22]. Detection [LPS^{+}23, OHM25, QHC^{+}24, SV20, ZZK^{+}24, ASKS20, AS22, AZ21, ABT^{+}22, BKG20, BRSR21, CLZ^{+}23, CCC^{+}24, CLZ^{+}24, CAA^{+}23, CLC^{+}23, CFF^{+}23, CCCW24, CBDK23, CSLN23, DFC^{+}23, FRC24, FA24, HBN^{+}25, HXJ^{+}20, JST^{+}23, LMR^{+}23, LLL^{+}22, LLW^{+}23, LID^{+}22, MCZX20, MMC24, MSRR22, MCRM25, MSB21, NMRS22, NAEAK25, NAZ22, NNV^{+}24, PHLHM24, PDDD20, PMM^{+}25, RSL^{+}21, SK22b, SELS21, SPZW25, SV25, SJC^{+}22, SSP21, TRD^{+}23, TTB^{+}23, TQT^{+}25, TAT^{+}23, VIL24, WDXX23, WTG23, YZM^{+}25, YLG^{+}25, YWLZ23, ZNPR^{+}23, ZFG^{+}23, ZX23, ZGW^{+}20b, ZLD^{+}25, vDDPZ24]. detector [NRRS20, ZYK^{+}25]. determinant [Bat20]. deterministic [RZLC24]. develop [AAW20]. developed [LGKT22]. Developer [AAA24, AES22, Fei23, dSNJ25, AMS23, BGMB20, HATG21, HYF^{+}24, RPS^{+}23, SMdS^{+}23, WLC^{+}20, YWTA25]. developer-informed [RPS^{+}23]. developers [AWA^{+}22, BSDB20, CAC20, EBAR21, FGRF24, GGB^{+}22, MFPB24, MCH22, MRA22b, NDDD20, PFC^{+}23, PCC^{+}20, SGG22, SGG24, WCZW22, ZLS22]. Developing [PKS25, MKB23, NRFS25, PZVN^{+}25, WCZW22]. Development [ALSA21, ASCR23, BFL23, IBP21, KRSW22, LOBD25, PdS23, ZYK^{+}25, ÅKH^{+}22, ACM25, ACSJ23, APB20, AACC21, BCF^{+}21, BCF^{+}22, BCLN21, BRO^{+}22, BBND^{+}20, BCJI22, BCOC24, BM23, BL21a, BL21b, CCS20, CXH^{+}24, CTD^{+}23, CPD20, ECMC20, EKK^{+}24, FvB24, GCSHB20, GM20, GAS25, JH20, Jør24, KKH^{+}21, KKL^{+}21, KR23, KMAB20, LSFE21, LGKT22, LBMF^{+}22, MVGHPT22, MBKL24, Mec24, MMB22, MFLS22, MLJ23, NdSR^{+}21, NF24, NAV^{+}24, NBGC20a, NBGC20b, PPMC22, PGR24, PN21, RPR22, RATS21, SWH^{+}20, SSO^{+}23, SHA21, SFR23, TSP20, UPP^{+}22, VDVC21, VCB24, WDT^{+}24, ZGHG^{+}23, OWGS23, VML21]. device [GW24]. devices [LCT22]. DevOps [CVC21, DPA^{+}24, ESO^{+}24, GGP^{+}24, WLS20, ZZJ25]. DevSecOps [ZCL24]. DHS [RJ23]. diagnostics [DSM20]. diagram [CZLN22]. diagrams [BHHQ^{+}22, SS23, SHWR22, ZXW20]. Dialog [FCW^{+}23]. Diaspora [CID25]. differed [NLTM23]. difference [BRSK24, LHY^{+}25, YCWD23]. Different [FAFJ24, SCPB25, HTG^{+}24, WJX^{+}24]. differential [LXLZ20, YLLY25]. Differentiation [KdJPK^{+}23]. difficulty [RSU^{+}24]. diffusion [CFAT25]. DigBug [KGL^{+}22]. Digging [IR21]. Digital [BNMW23, DJR^{+}22, MH24, PZVN^{+}25]. dimensional [AAG21, LCC^{+}23a]. dimensionality [CWGS23]. dimensions [ZCL24]. direct [SELS21]. direct-learning [SELS21]. directed [RF24b, aSLF^{+}22]. directions [IPB23, MMC24, RFB^{+}22, SCB22]. discipline [KSG^{+}22]. Discovering [ZB22]. discovery [GMS22]. discrete [LHN20]. discuss [WDT^{+}24]. Discussed [THN20, AA24, SHB21]. discussions [FRC24, LSF^{+}25, NSL^{+}21]. Dispersion [LCP^{+}24]. distance [LMVRA^{+}20]. Distributed [ADS^{+}22, RSM^{+}23, AZR^{+}21, ACD^{+}21, ACD^{+}24, BSDB20, Bro24, HKP23, KMS^{+}25, LLWL22, MPRX20, PHLHM24, RAGCSS^{+}20]. Distribution [RSLC24, RK20, WCZL25, dSIdÁ23]. Distribution-aware [RSLC24]. Diverse [ZLW^{+}23]. diversification [SMB^{+}20]. Diversified [CZW20]. Diversity [KA22, HSVMB20, Hyr25, MLBD21, NF24, WKP20]. Diversity-driven [KA22]. divider [OPF24]. DLAP [YZM^{+}25]. Do [BCLN21, PSGD23, TST24, WYFZ24, AKMS23, BEAK21, Fei23]. documentation [BRO^{+}22, Vid22]. DoD [RJ23]. Does [JST^{+}23, LLST20, LNST21, BS23, FEBO22, ZLS25]. Domain [AFJ^{+}20, DJR^{+}22, BVHHO23, BPJ^{+}22, BB22, CVC21, DL22, GKB^{+}21, KBG^{+}24, KL20, QZS^{+}24, SZ22, SJC^{+}22, VWSCH23]. domain-adversarial [QZS^{+}24]. Domain-based [AFJ^{+}20]. domain-oriented [DL22]. domain-specific [BPJ^{+}22, KBG^{+}24, VWSCH23]. Done [KOPN22]. DongTing [DFC^{+}23]. Don’t [WZL^{+}25, AZB25, ACG^{+}21]. dragons [BCOC24]. drawbacks [CRV23]. drift [BAH24, WZZ21]. drift-aware [BAH24, WZZ21]. Driven [ALSA21, DAML23, FBMR20, IBP21, KRSW22, PdS23, AIH20, AZB25, BCF^{+}21, BCF^{+}22, BHS24, BCF23a, BM23, BDIL25, CCR24, CTD^{+}23, CDET22, EBAR21, FJvdW20, GB20, GdOB25, GVBC^{+}24, HGHK24, KA22, KFJA23, LYZ^{+}22, NBGC20a, NBGC20b, PKS25, RVAC^{+}25, RJ23, RHJ25, SSS22, VGS^{+}23, WNZ25]. drivers [MMB22]. driving [ALZ^{+}20]. DSL [MN21a]. Dual [LYZ^{+}22]. duplicate [JST^{+}23]. duration [YZS22]. during [ACSJ23, CXH^{+}24, GBMF22, LTG^{+}25, MAV24, RVAC^{+}25, STM^{+}22, Tok22]. DValidator [MOP24]. Dynamic [HPP24, MCMA21, PCJNP23, ACC^{+}20, AdlBGZ^{+}23, dACO21, CF24, DXD25, ISKB20, MdSKD22, RFB^{+}22, SMKI23, WXL^{+}20, WTS23, ZXW20, dCMM^{+}22, APAF21, GMS22]. dynamically [MDML25]. Dynamics [AAA24, JVZ25, MDAHS25].

E-SC4R [TCA22]. E/E [MW25]. E2E [AMB^{+}25, OLR^{+}25]. EA4Cloud [CNdLL20]. Early [LTJ^{+}20, LL21, SCdPL24, RJ25]. easy [AMHB24]. ECCOLA [VKJ^{+}21]. echo [PSGD23]. Eclipse [KCMD21, NAV^{+}24]. economic [MH20]. ecosystem [CFAT25, DMD23, Gu23, KVP23, LTZ^{+}21, MCSAGB20, MAS23, dlVRB21, SMdS^{+}23]. Ecosystems [ROL21]. Edge [MSC20, MYH^{+}24, BBF^{+}21, CAF22, CFF^{+}23, DHM^{+}21, DTZ^{+}22, KMS^{+}25, MPGB22, XDL^{+}22, AM23]. edge-based [CAF22]. edge-cloud [BBF^{+}21]. Edge-Fog [MSC20]. EdgeWorkflow [XDL^{+}22]. editor [FBMR20]. Editorial [Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano25b, Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h]. edits [JWZM20]. education [BTSC^{+}23, CGM24, CJNDZ21, GRLA20, Hyr25, KI23, NRFS25, OKG25, PNND25, SJH^{+}24, SHG24]. educational [LH24b]. EFACT [ZLW^{+}24]. effect [AWA^{+}22, BIT24, CAC20, LWC20, LST20, VM24a]. Effective [PLP^{+}20, ZCZL23, AHT^{+}21, CLC^{+}23, GGMB^{+}22, HXJ^{+}20, MSMB21, NMT^{+}23, ÖS20, QWHH23, SHLM24, XBS21, YJZZ23, YLW^{+}21, ZWF^{+}25]. Effectiveness [CKY25, DHZ^{+}24, GBMF22, HM24, HYF^{+}24, LWZ21, MH24, TGG21]. Effects [DFB20, JH20, RMT^{+}22, SvdBHV24, Tai20, TSP20, VCB24, VHB21, WLC^{+}20]. efficiency [DCÁ^{+}24, MH24, PSL25, TTL20, XZZ^{+}25]. Efficient [CBZZ24, MSMB21, Mas22, ZFG^{+}23, ATWS25, BDS^{+}24, FR20, HLL21, ISKB20, LCY23, LL25, NMV25, SRTGA^{+}25, SPZW25, TQT^{+}25, YDP22, ZWC^{+}24]. efficiently [EBAR21]. Effort [dACOdS23, DFB20, JH20, RJ25, VCB24, dACO21, LGKT22]. elastic [RFSP^{+}25, VSTK21]. Electricity [AZR^{+}21]. element [CFNA25]. elements [CZLN22, OSM^{+}23, SYB^{+}23]. elicitation [BMHR21, GBMF22]. eliciting [SRM^{+}25]. embedded [CCS20, HLZ^{+}23, MVŞ23]. embedding [ES24, MRVS24, TTB^{+}23, WJX^{+}24, YWLZ23]. embeddings [HPZ^{+}20, ZLSY23]. emergence [PT21]. emergent [RFSP^{+}25]. Emerging [SHG24, BHS24, ROL21]. Emotimonitor [EMCN^{+}22]. emotion [IMTS23]. emotions [EMCN^{+}22, HOCK22]. empathetic [YRC^{+}25]. empathize [LH24b]. Empathy [GGMH23]. Empirical [ANC^{+}23, GW23, LCP^{+}24, MVŞ23, AA24, AKMS23, AM23, BMLH25, BCLN21, BBND^{+}20, BAV20, BTSC^{+}23, CXH^{+}24, CATA21, GMCA21, GBMF22, GFS21, GSK24, GW24, IZAD21, JST^{+}23, JKD25, KH24, LALM23, LST20, LNST21, LABJ23, LSG^{+}24, LFH^{+}22, LMZT22, LQY^{+}22, LCY^{+}24, LDT^{+}25, MCM^{+}24, NSR^{+}23, NEJ24, OZVRD21, OMA^{+}22, OSJB21, PLL^{+}23, PDDD20, PMM^{+}25, RA25, ROL21, SSO^{+}23, SGV22, SXL^{+}22, SBD23, SMVC24, THG20, TKSC20, WCZW22, WCH^{+}23, WJX^{+}24, YWTA25, ZKDP22, ZWY^{+}22, ZAS25, ZRGJ21]. Empowering [GSF24]. enable [ZMLZ23]. enabled [DAP20, HTB21, ISW25, LCY23, NBP24, RPC^{+}24]. Enabling [BB22, KKL^{+}21]. encounter [WCZW22]. Encryption [LYZ^{+}22, FG24]. End [AAZB23, FYH^{+}24, BSCS23, GM20, Jør24, MDWS^{+}23, NAV^{+}24, SSO^{+}23, BFL23]. End-to-end [FYH^{+}24, NAV^{+}24]. End-user [AAZB23, BSCS23, MDWS^{+}23, SSO^{+}23, BFL23]. endpoint [SZ22]. Energy [GS20, CNdLL20, DCÁ^{+}24, LFLS24, PCC^{+}20, PÁngeles MoragaGC25]. energy-aware [CNdLL20]. Energy-Delay [GS20]. enforcement [BRS23]. enforcers [RMM24]. engagement [Tok22]. engine [HIDT21]. engineer [BFM^{+}25, CKP20, OBCR23]. Engineering [BFZC21, Bro24, CID25, DJR^{+}22, DAML23, FBMR20, Gla23, PFdMF21, RBvB^{+}24, WFM^{+}24, ZFGH^{+}22, A– DHM23, AMRS20, AMM25, AACC21, AGP22, BbASP23, BDIL25, BCW21, CDET21, CDP22, CTD^{+}23, CJNDZ21, CDET22, DdM^{+}24, DPD^{+}22, DPGGP23, DBO24, ETD^{+}24, EKHJ^{+}20, EN23, FKGN23, GB20, GCLB22, Gir21, GPPDLF23, GVBC^{+}24, GGMH23, HNM^{+}24, HLS24, Hyr25, IT23, KKH^{+}21, KGS^{+}25, KHEC^{+}23, LSB^{+}22, LFLS24, LGT^{+}23, LK23, LBCG23, LHH^{+}24, MBKL24, MFF^{+}24, MWFK20, MRA22b, MdOT23, MKB23, NRFS25, NEJ24, OKG25, PDGMT20, dPPdJFF21, RA25, SvdBHV24, SM20b, SMVC24, WFR21, WKP20, WMAL21, WGY^{+}20, ZLS23, KZTS22, RFS^{+}21]. engineering-enhancement [PDGMT20]. engineers [SM20a, STM^{+}22]. engines [BQR24, KA24, PPM^{+}21]. enhance [XTF^{+}21]. Enhanced [GIP^{+}24, CCCW24, NAEAK25, NDP^{+}21, VTS22, WLH^{+}25, ZLW^{+}23, ZYZZ21]. enhancement [PDGMT20, ZCZL23]. Enhancing [CJZ^{+}20, CDLN21, GRdAL23, KGM24, NEJ24, PBC^{+}23, ZLD^{+}25]. enrichment [DXL^{+}24]. Ensemble [dACO21, BKG20, GSF24, HBN^{+}25, PPL^{+}25, dACOdS23]. EnseSmells [HBN^{+}25]. Enterprise [SVAGB20]. entities [JZM21, OEW22]. entity [JWZM20, ZLS20]. entropy [YWLZ23]. Environment [MSC20, BBF^{+}21, NMG^{+}20, GMS22]. environments [GTT^{+}20, LCY23, MPGB22, PKB24, RFB^{+}22, RFSP^{+}25, YAaO^{+}21, ZPL21]. equal [ST24]. Equivalence [LZ24, MAK^{+}21]. equivalent [HSW^{+}24, MAK^{+}21]. Era [AAA24, AX21, ZGLPA24]. erosion [WL24]. erroneous [SSP21]. Error [TGG21, BDS^{+}24, GF24, HM24, Liu21]. errors [CKP20]. ERTMS [NMG^{+}20]. ERTMS/ETCS [NMG^{+}20]. ES6 [PZDG21]. EsArCost [WL24]. established [SM21]. establishment [IR21]. estimates [MGSC22]. Estimating [ETY^{+}22, WL24]. estimation [JH20, MCM^{+}24, MR24, RJ23, RJ25, WDO24, dSIdÁ23, dACO21, dACOdS23]. ETCS [NMG^{+}20]. Ethereum [BMP22, HLW^{+}23, PBC^{+}23, YML^{+}22]. ethical [KMS^{+}25, YRC^{+}25]. ethically [VKJ^{+}21]. ethics [Fei23]. ethnography [NdSR^{+}21]. ETL [AZKR20]. ETune [CBZZ24]. EUDability [BFL23]. European [GCF22]. evaluate [AR24, SM24]. EvaluateXAI [AR24]. Evaluating [BHHQ^{+}22, BSDB20, FVDF21a, FVDF21b, FA24, FAFJ24, LBG^{+}20, LWZ21, NMRS22, dORGCG23, HH22, HH25, VKLM21]. Evaluation [PBV24, TMV25, ÅKH^{+}22, AON^{+}24, CND22, CATA21, GGB20, GBMF22, IR21, LALM23, LSG^{+}24, LLK^{+}21, LWSZ23, MH24, MTA24, MBG^{+}25, PLL^{+}23, RK20, RPS^{+}23, SJH^{+}25, SL20, SHLM24, SBM23, SZSV22, SMB^{+}20, SBD23, XCH^{+}20, YWY^{+}21]. Evaluations [LL21]. Event [WTS23, AIH20, CSLN23, OSJB21, LCAC21]. Event-aware [WTS23]. event-driven [AIH20]. events [NEJ24]. evidence [LMR^{+}23, MSKS24, MdOT23, STM^{+}22]. evidence-based [MdOT23]. Evolution [GSK24, LGMP24, MVSV25, MDAHS25, RB23, ARPDP24, DCÁ^{+}24, DSO20, HQNR^{+}22, HBK25, KFJA23, KGS23, LXLZ20, LTZ^{+}21, MAV24, OSJB21, RV22, RFSP^{+}25, SGM23, SOMSCT23, WCH^{+}22, WYG^{+}24, ZHS^{+}24, ZWP^{+}24]. Evolutionary [MLPC20, AA22, BRSR21, BFZC21, DOG^{+}23, JCNS^{+}22, WMS23]. evolve [AKMS23]. evolving [DSB23, IIK21]. Examining [SBF20, VM24a, WLC^{+}20]. example [CJZ^{+}20]. example-based [CJZ^{+}20]. examples [Hor21, SRD^{+}21]. exceptions [MLYL25]. Exchange [ASCR23, PKGA22, ACD^{+}24]. executable [KBG^{+}24, SBD23]. executing [VTS22]. Execution [AZKR20, LWL20, AKL25, CI24, HIDT21, KBBD23, MN23, SH20, YZS^{+}25, YAaO^{+}21]. existing [ZMLZ23]. experience [BB22, FAA22, GKB^{+}21, MBKL24, PBV24, PGR24, PKS25, SMM23, WLCC25, YLC25, ZHM^{+}23]. experiences [GRV^{+}21]. experiment [AC21, DFB20, MDWS^{+}23, PSZ21]. Experimental [FC20, MTA24, MPP24]. experimentation [GAB20, SM21]. experimented [Fei23]. experiments [BCF^{+}22, DPRD21, Fei23, GVBC^{+}24, MPP24, NSL^{+}21, RFS^{+}21, WMS23]. expert [CCC^{+}24, CXP^{+}23, MGSC22, NIJ22, SKP20]. expertise [GBLP25, HTG^{+}24]. Explainable [CACHA23, YZZ^{+}24]. Explaining [CWGS23, MHJW23, TCA22, WCGS23, YZL^{+}22]. explanations [BSCS23]. exploit [YZC^{+}23]. ExploitGen [YZC^{+}23]. Exploiting [YZL^{+}22]. Exploration [HTC^{+}23, AHFP24, SMKI23, SVVD21]. exploratory [BMLH25, CKP20, FPB^{+}24, FSS23, FvB24, GCF22, LXL^{+}23, MSMB21, ZZP21, ZB22]. Exploring [CJNDZ21, DCÁ^{+}24, LMZT22, LQY^{+}22, Pat20, RFSP^{+}25, SS20, VM24b, ZLZ^{+}25, dCMM^{+}22]. exported [DYZ^{+}23]. expression [LL25, LL23]. extended [dACOdS23, CLLJ24, MPGB22, PLP^{+}20]. Extending [ABMV24]. extensible [FvB24, Spa24]. Extension [GBMF22, GKAHMO22]. extensions [CHLT23, GAC20]. extensive [LWYW20, TSLHS21]. Extent [Vog20]. external [LAL^{+}23, ZLW^{+}24]. ExTrA [CWGS23]. extract [SAZN22]. Extracting [DDCC24, WLH^{+}25, FFV^{+}23, SYB^{+}23]. extraction [BCF23a, GF24, LCS^{+}24, SS23, ZPSW24]. eye [BMLH25, HBSV^{+}22, MBG^{+}25]. eye-tracker [BMLH25].

FaaS [ES23, SGG^{+}25]. FaaS-based

[SGG^{+}25]. FaaSten [YSB^{+}21]. Facilitating [AGL21, NAV^{+}24]. facing [SWH^{+}20]. Factors [AF22, Gla23, SMdS^{+}23, Tok22, FGRF24, LBT^{+}21, LMZT22, MVGHPT22, MGSC22, MKB23, NdOdO^{+}22, RCT22, SJH^{+}24, SRS^{+}25, SM20a, ZGW^{+}20b]. factory [DHM^{+}21]. fail [BEAK21]. failure [BS23, CSLN23, JYZ^{+}24, JS22, SXL^{+}22, TTL20]. failures [CDLN21, HSJB23, MS25, MPRX20]. Fairness [PNND25, RSLC24, WZL^{+}25]. familiarity [WLC^{+}20]. families [MMSM^{+}23, SSS^{+}20]. family [DPRD21, RMT^{+}22, SOMSCT23]. far [BS23, DPD^{+}22, YZS^{+}25]. farmers [SSO^{+}23]. Fast [KDB^{+}21, KBP24, WDXX23, LCW^{+}24]. Fault [LLWL22, LWL^{+}23, VHSB22, DSM20, FNB^{+}24, Hu24, Hu25, LWL20, LDH22, LLW^{+}24, LLH^{+}25, MCZX20, MHH21, MHJW23, MRM^{+}22, TPV25, WYFZ24, WDO24, XBS21, XZY^{+}20, YJZZ23, ZXW20, dSFdC^{+}25]. fault-based [WYFZ24]. fault-proneness [TPV25]. Fault-tolerant [LLWL22, FNB^{+}24]. faults [DSM20, Dut24, LST20, OMA^{+}22, RF24a, SV20, SV25].

Faulty [SVV24]. FCCI [SKP20]. feasible [CBMM20]. Feature [BSF^{+}25, DWBA24, MAV24, MCRM25, MPFB23, NAZ22, SBM23, VVBGL^{+}23, Vog20, CLZ^{+}24, CAF22, CKM^{+}20, CCCW24, DMAH24, EFPC21, EM24, HGH^{+}23, LAL^{+}23, MMDL23, MMSM^{+}23, MHR^{+}25, RVAC^{+}25, ROGS^{+}24, TLXW23, WGL^{+}22, ZXH^{+}24, ZAS25, ZYZZ21, ZB22]. Feature-based [NAZ22, MMDL23, ZB22]. feature-driven [RVAC^{+}25]. Feature-oriented [MAV24]. feature-weighted [TLXW23]. featured [SRD^{+}21]. features [EM24, FWZZ25, HYF^{+}24, LFH^{+}22, MHOM22, RSU^{+}24, WGL^{+}22, WHK^{+}23, ZJXG20, ZPSW24]. featuring [CBDK23]. FEBI [HLW^{+}23]. February [Ano20p, Ano21q, Ano22p, Ano23p, Ano24o, Ano25i]. federated [LLZ^{+}22, PKB24, VCT20]. Feedback [DMAH24, CCR24, KDB^{+}21, SHG24]. female [SSO^{+}23]. Few [LYG^{+}24, SPZW25]. Few-shot [LYG^{+}24, SPZW25]. fewer [TST24]. FHIR [RLK^{+}24]. fibre [KBW^{+}24]. Field [GKPS24, ADS^{+}22, CBMM20, WMA^{+}24]. Field-sensitive [GKPS24]. fifty [FKGN23]. file [WWC23]. filter [LCC^{+}23b, LCW^{+}24]. filtering [DLX^{+}23, LXYL20]. filtering-based [DLX^{+}23]. Find [Woo25]. Finding [CKP20, DSM20, TST^{+}21a, TST^{+}21b, YOH^{+}23, RF23]. Findings [BTSC^{+}23]. Fine [CLZ^{+}24, SPZW25, XCZ23]. Fine-grained [CLZ^{+}24, XCZ23]. fine-tuning [SPZW25]. finite [JJS24, Ibi22]. Finland [Hyr25]. first [LYS^{+}23]. fit [Bat20, SS20]. fix [MFBP20, NLS^{+}20, XXC^{+}24]. Fixed [AHT^{+}21]. Fixed-Size-Candidate-Set [AHT^{+}21]. fixes [SK22a, WCH^{+}23]. fixing [ASK^{+}23, EBAR21, LSSZ21, WLC^{+}20]. FL [Hu25, LCC^{+}23b]. flakines [TRD^{+}23]. flaws [KRLS^{+}20]. flexibility [SMH^{+}23]. Flexible [CCR24, Spa24, LCT22, VGS^{+}23].

FlexRay [GAL20]. floating

[MLYL25, YZS^{+}25]. floating-point

[MLYL25, YZS^{+}25]. FLOSS [JVZ25, WSL^{+}20]. flow [CCR24, SHWR22, PBC^{+}23]. Flutter [AA24]. focus [CKP20, MLJ23]. focused [NQT25]. focusing [NBGC20a, NBGC20b]. Fog [MSC20, PKB24, MPGB22]. following [PZVN^{+}25]. forced [ŠMKGH23]. forecast [AAB^{+}22]. forecasting [TKSC20, DCMR20]. Formal [LZP^{+}20, BGM^{+}21, CBDK23, CFP^{+}21, HIDT21, LM21, RHJ25]. formal-model [HIDT21]. Formalization [LZ25]. formatting [OSM^{+}23]. Formulas [MPFB23, MLBD21]. formulation [ÖS20, Tai20]. forum [Pat20, WDT^{+}24, YOH^{+}23]. forums [WDT^{+}24]. Fostering [Gla23]. foundation [LLL^{+}25]. four [SGV22]. fractional [WCZL25]. fragmentation [BDS^{+}24]. fragments [WZFD25]. FragQC} [BDS^{+}24]. Framework [VKB^{+}25, AZB25, Ala21, AZ21, AWP^{+}25, AWHS22, AR24, BAH24, BGC20, BRS23, BBW22, CKM^{+}20, CCCW24, DSB23, ECS23, FNB^{+}24, GMS22, GTT^{+}20, GM20, HDX^{+}23, HGH^{+}23, JRM^{+}22, KRP24, KSG^{+}22, KBG^{+}24, KBPMJ20, LZL^{+}25, MDVB^{+}23, MCH22, MN21a, MRR^{+}20, NAEAK25, NMV25, NAV^{+}24, PKB24, RPS25, RSVW23, SSS21a, SSS21b, SZS25, VWSCH23, VGS^{+}23, XCH^{+}20, YZM^{+}25, YYW^{+}24, YSB^{+}21, ZSF^{+}23]. frameworks [BCLN21, HH24, HKP23, PPM^{+}21, SBF20]. free [NF24, RF24a]. free/libre [NF24]. frequencies [VHSB22]. frequency [CKM^{+}20, MLBD21]. frequency-inverse [CKM^{+}20]. front [Jør24]. front-end [Jør24]. FSECAM [MHR^{+}25]. fumes [ACG^{+}21]. Function [SL20, ZLW^{+}24, HLW^{+}23, LL21, LFFW23, SSP21, VHSB22, YSB^{+}21]. Function-as-a-Service [SL20, YSB^{+}21]. functional [ACD^{+}24, BFHC20, CJZ^{+}20, CBMM20, GGPR24, GBT^{+}20, KRS^{+}21a, KRS^{+}21b, NMG^{+}20, RJ25, GCLB22]. furious [vDDPZ24]. fusing [CCC^{+}24, CXP^{+}23]. fusion [CLLJ24, ZWC^{+}24, ZAS25]. future [ACM25, BLHS23, ECMC20, IPB23, MMC24, MMB22, RFB^{+}22, SCB22, tBS23]. Fuzz [LHY^{+}25, ZLD^{+}25]. Fuzzing [KZK22, KH24, RZLC24, WLL^{+}24, WMH^{+}24]. Fuzzy [LBH^{+}24, SKP20, YB24].

galaxy [OZVRD21]. gambling [CACHA23]. Game [BFZC21, GCSHB20, ASCR23, AA22, CTD^{+}23, PPM^{+}21, ASCR23]. game-specific [ASCR23]. Game-theoretic [GCSHB20]. games [ZGLPA24]. gamification [BTSC^{+}23, dPPdJFF21]. gap [DdM^{+}24, DCMR20]. gaps [HWL^{+}25]. garde [SM21]. gas [ACG^{+}21, DLV^{+}22]. Gated [HHZW23, YZL^{+}22]. gates [ZXW20]. GBSR [LLW^{+}24]. GCC [ZRGJ21]. GDPR [ARPDP24]. GEML [BRSR21]. gender [DdM^{+}24, KGM24, OKG25]. General [BGC20, MVŞ23, ZYZ^{+}23]. generalization [MWD24]. generalized [SV25]. generate [AKBN20, DPRD21, LL23]. generated [CFI^{+}24, GGB20]. Generating [AIH20, MRS25, BFZC21, KBBD23, MFBP20]. Generation [FZT^{+}22, BDMP21, CKY25, DWG^{+}24, DZY^{+}23, FYH^{+}24, GDWK24, Ibi22, KKRT24, KA22, KA24, KBPMJ20, LWP^{+}21, LL25, LXLZ20, LCC^{+}23b, LSSZ21, MCZX20, MMDL23, MFPB24, MRA22b, NAD^{+}20, NHA20, RSLC24, SRTGA^{+}25, aSLF^{+}22, TCTL25, WYG^{+}24, YZC^{+}23, ZWY^{+}22, ZLW^{+}23, ZZY^{+}24, ZYZ^{+}23, ZSF^{+}23, EBSB23]. Generative [RBvB^{+}24, KKE21]. generator [BG24, DSB23]. generators [LH24a]. Generic [DWH23, LCC^{+}23a]. Genetic [GKW25, AA22, Hu25, XBS21]. geo [LLWL22]. geo-distributed [LLWL22]. Geographical [LCP^{+}24]. geographically [AZR^{+}21]. GeQuPI [GKW25]. German [EKK^{+}24, ZH21]. Germany [KBW^{+}24]. girls [SGW^{+}22]. Git [HHK20, DCMR20]. Gitcoin [CPC^{+}23]. GitHub [ASCR23, CXH^{+}24, CDM25, DMN^{+}23, DMD23, GSM^{+}23, GDLM21, LYS^{+}23, LSF^{+}25, WWC23, WCH^{+}22, ZLZ^{+}25]. GitLab [EKHJ^{+}20]. Gitter [SHB21]. Global [CDET21, BBG^{+}25, BCLN21, MHH21, SM20b]. globally [BSDB20]. GloBug [MHH21]. Goal [LLW^{+}23, LCW^{+}24, DDCC24]. Goal-conflict [LCW^{+}24]. goes [FAA22]. Google [Hor21, SWG^{+}20]. governance [LLZ^{+}23]. government [WSL^{+}20]. government-academia [WSL^{+}20]. GPTSniffer [NDD^{+}24]. GPUs [CCS20]. grained [CLZ^{+}24, XCZ23]. Grammar [RF24b, BRSR21, NAD^{+}20, ZHS^{+}24]. Grammar-based [RF24b, BRSR21]. grammars [MRVS24, RF24a]. granularity [BCF23b]. Graph [BSH^{+}20, DLX^{+}23, FFSB23, LLW^{+}24, LFFW23, ZYL^{+}22, AGL21, CLZ^{+}23, CFF^{+}23, CCCW24, CCR24, DHZ^{+}24, KRP24, MRS25, TTB^{+}23, WTG23, WLH^{+}25, YZL^{+}22, YLLY25, ZXH^{+}24, ZLX^{+}22, ZSZ^{+}22, PBC^{+}23, WLH^{+}25]. Graph-based [BSH^{+}20, FFSB23, LLW^{+}24]. graph-transformer [ZXH^{+}24]. Graph4Web [ZLX^{+}22]. graphic [QWHH23]. graphical [AC21, DPRD21]. graphs [AHT^{+}21, RF24b]. gravity [CKM^{+}20]. great [BFM^{+}25]. Greek [Mam23]. Grey [VKG^{+}23, GGP^{+}24, ZHLR23]. ground [GDLM21, VGS^{+}23]. ground-truth [GDLM21]. grounded [DPA^{+}24, DPGGP23]. group [IR21, IT23, KI23]. grow [AZB25]. growth [HCC22]. GRuM [VGS^{+}23]. GT [GJW^{+}22]. GT-SimNet [GJW^{+}22]. guarantees [MH20]. Guest [FBMR20]. GUI [AJFD25, DFT24, RSM^{+}23].

GUI-based [AJFD25]. guidance [MDRdP^{+}24, WR22]. guide [CXY^{+}23, DBO24]. Guided [MLPC20, SMKI23, WLL^{+}24]. guideline [FMF^{+}24]. guidelines [MLS^{+}21, SZSV22, ZFC^{+}22]. Guiding [FWZZ25].

handle [ABMV24]. handling [ASSH22, KCMD21]. Hansie [MN21b]. hard [AMHB24]. hardware [BBND^{+}20, SCN^{+}21]. hardware-in-the-loop [SCN^{+}21]. harming [FAG^{+}20]. Harmonizing [DPA^{+}24]. Harnessing [ACG^{+}25]. heads [ZAS25]. health [AAZB23, LGT^{+}23, LWSZ23, MH24]. heap [ACG^{+}25]. Heat [BIT24]. heaviness [Gu23]. help [BTSC^{+}23, CKP20]. helps [PdS23]. here [SMH^{+}23]. heterogeneous [AZR^{+}21, CLZ^{+}24, RCA^{+}22]. heuristic [GBSO20, YDP22, ZWY^{+}22, ZSF^{+}23]. heuristics [BFHC20, SVVD21, SSP21]. hidden [ZLS25]. Hierarchical [JSTW22, LL25, ZPSW24, ZYL^{+}22, AGL21, APR24, GDWK24, ZSF^{+}23]. hierarchically [LZJ20]. High [SCB22, SH20, WYFZ24]. High-availability [SCB22]. high-order [WYFZ24]. higher [SHG24]. highly [FKF^{+}23, LBHW25, MAV24, MAS23]. highly-configurable [MAV24]. highly-selected [MAS23]. Highway [HHZW23]. Hippo [HIDT21]. histories [BN23]. history [ZWP^{+}24]. HL7 [RLK^{+}24]. HMER [LWL20]. hoc [FR20]. holistic [BL21a, BL21b, TGX^{+}22, YWS23]. Home [ŠMKGH23, CR23, SMH^{+}23, Tok22]. home-based [Tok22]. hospitals [RLK^{+}24]. host [TMV25]. host-based [TMV25]. hot [SH20]. HPC [AHL22, PPL^{+}25]. HSP [MAP^{+}20]. Human [Gla23, LBCG23, MKB23, RBvB^{+}24, GDK24, GVBC^{+}24, HM24, LGT^{+}23, LCY23, MVGHPT22, ZLS23]. Human-Centered [RBvB^{+}24].

Human-centric [LBCG23].

human-oriented [GVBC^{+}24]. HUNTER [TGX^{+}22]. hW [GBSO20]. hW-inference [GBSO20]. Hybrid [FZT^{+}22, LOBD25, SCD^{+}24, Ala21, AMM25, BMB20, DHZ^{+}24, HPP24, LH24b, Liu21, LLH^{+}25, MAP^{+}20, PN21, YZS22, ZCLP21, ZZTC23, ZHLR23, ZGLPA24, ZYZZ21, LWL20, MN21b]. Hybridization [TPGH20]. hypervisor [CDO24].

IADA [MdSKD22]. iCoLa [FvB24]. ideal [KSG^{+}22]. Identification [MFLS22, ASM^{+}21, CDM25, HLW^{+}23, LCW^{+}24, NSL^{+}21, SSS^{+}20, SV20, XCH^{+}20, ZJXG20, dCMM^{+}22]. identifiers [GIP^{+}24, NAD^{+}20, PSGD23]. identify [RPL^{+}21]. Identifying [MSB23, WXL^{+}20, ZCL24, dTMS21, AKL25, AWP^{+}25, BGE^{+}21, SKP20, YB24, ZYZ^{+}23, BSGN21]. identity [LZ24]. IEC [GRV^{+}21, RSL^{+}21]. IFML [PLP^{+}20]. iFogSim [MPGB22]. iFogSim2 [MPGB22]. IIoT [DHM^{+}21]. image [LFH^{+}22]. images [CZLN22]. Imbalance [QWHH23]. Imbalanced [XZY^{+}20]. Impact [ACH25, CLNQ24, PÁngeles MoragaGC25, ZLS22, AMHB24, BS23, CFAT25, CR23, JKD25, KCMD21, KHEC^{+}23, KAA^{+}21, LHG^{+}25, MCY23, MR24, OSM^{+}23, OAH^{+}23, RPC^{+}24, SMVC24, TRD^{+}23, ZGW^{+}20b, ZZP21]. impacts [KdJPK^{+}23, LQY^{+}22]. impaired [CXY^{+}23]. imperative [BBW22]. imperfect [HCC22]. Impermanent [GIP^{+}24]. Implementation [LHN20, MDML25, AC21, BG24, SJH^{+}25]. implemented [MCDP24]. Implementing [Hyr25, VKJ^{+}21, YWTA25]. Implications [ZJY^{+}24, ACSJ23]. Implicit [DMAH24]. importance [DZY^{+}23]. Improve [DXD25, HELW20, JST^{+}23, MHOM22, SSP21, WC20, dCMM^{+}22]. Improved [SRD^{+}21, CLC^{+}23, MCRM25, XZZ^{+}25, YWCX24, ZCZL23]. Improvement [GKW25, HTC^{+}23, MICV23, TTL20]. improves [WMLM22]. Improving [ASSH22, LCAC21, RZLC24, WZL^{+}25, ZLS20, LLW^{+}24, SZSV22, GJW^{+}22]. in-breadth [AES22]. in-context [YLGS24]. in-depth [AES22]. In-situ [VBMB20]. In-the-field [CBMM20]. in-vehicle [GAL20]. inactive [WZL^{+}25]. Incivility [FRC24]. inclusion [Hyr25]. inclusive [CGM24]. incompatibilities [LGC^{+}25]. inconsistent [LGC^{+}25]. incorrect [ASK^{+}23]. Incremental [ZWC^{+}24, HZ24, KDB^{+}21, LCW^{+}24]. Incremental-concurrent [ZWC^{+}24]. independence [AdlBGZ^{+}23]. independent [BCOC24]. indicators [IT23]. individual [KdJPK^{+}23]. individualism [ZLS22]. induced [MS25]. Industrial [GRV^{+}21, BAH24, DWH23, DSM20, FAA22, GAB20, IMTS23, KBPMJ20, KBP24, LCY23, MS25, MAP^{+}20, MDWS^{+}23, NQT25, TST^{+}21a, TST^{+}21b, WLCC25, ZHM^{+}23, ZLC^{+}23, vRMG23, NSL^{+}21]. industry [AF22, CJNDZ21, DAML23, HNM^{+}24, LSFE21, OSJB21, RMT^{+}22, ROL21, SRS^{+}25, ZGLPA24]. inefficient [AHL22, PCC^{+}20]. Inference [LBH^{+}24, GBSO20, RSM^{+}23, XCZ23]. Infinite [VCF22]. Influence [LCP^{+}24, BGMB20, BIT24, MVGHPT22, MLGBRC24]. Influencing [FGRF24]. Information [WGMT25, BDLT21, DFB20, Ibi22, LXYL20, LCS^{+}24, MAP^{+}20, RV22, SHWR22, TSP20, YWCX24]. informed [HBK25, RPS^{+}23]. infrastructure [FPB^{+}24, DDPT20, RAGCSS^{+}20]. infrastructure-agnostic [RAGCSS^{+}20]. infrastructures [CAF22, PPL^{+}25]. Initiatives [dPPdJFF21, RJ25, TH24]. injection [LMGK22]. Inner [ECMC20]. Input [LABJ23, LSG^{+}24]. input-aware [LSG^{+}24]. inquiry [ZLC^{+}23]. insecure [GSM^{+}23]. Insights [LH24b, dSNJ25, WDT^{+}24]. inspection [NIJ22]. inspectors [NIJ22]. Instance [ZWL^{+}24]. Instance-level [ZWL^{+}24]. instances [AM23]. instant [SGG22, SGG24]. instantiation [OBW^{+}25]. institutions [WMAL21]. instructional [MRA22b]. Instrumented [GRdAL23]. instruments [FKGN23]. integer [OPF24]. integrated [PPMC22, RFB20, RO22]. Integrating [APB20, EKHJ^{+}20, LLH^{+}25, Hei20, SK22a]. Integration [DZY^{+}23, BMB20, BB22, DBO24, HPP24, JS22, LWYW20, LTJ^{+}20, LMZT22, THG20, WMLM22, YLHZ20]. integrative [MMB22]. Intel [Ala21]. Intelligence [MMO22, FEBO22, GPF22, LZL^{+}25, SFR23]. intelligent [CXY^{+}23, ESO^{+}24, LTG^{+}25]. intensity [KGS23]. Intensive [KRD^{+}23, BGL^{+}22, VKLM21, ZGHG^{+}23]. Intention [AAA24, SS20]. Intentions [Mam23]. Inter [GS20, AHL22, BSS^{+}25, GSK24, WHK^{+}23, DPGGP23, GPPDLF23]. Inter-Coder [GPPDLF23]. Inter-Process [GS20, AHL22]. inter-rater [BSS^{+}25, DPGGP23]. inter-service [WHK^{+}23]. inter-version [GSK24]. Interaction [CBDK23, BMLH25, BG24, JT20, LCY23, MAC^{+}22]. interactions [dSIdÁ23]. interactive [Bro24, CACHA23]. Interdisciplinary [VHB21, HM25]. interest [dTMS21]. interface [LB24, SWZ^{+}20, WGL^{+}22, ZFC^{+}22]. interfaces [DPRD21, SZSV22]. interference [MdSKD22]. interference-aware [MdSKD22]. interleavings [KZK22]. Intermediate [LZ25, vdG25]. intermittent [SV20, SV25]. internationalisation [LXL^{+}23]. interoperability [BGL^{+}20, HPF23, RLK^{+}24]. Interpretability [ZSCD22]. interpretable [NAEAK25]. interpreters [CF23]. intersection [BFL23, CJNDZ21]. intervention [NRFS25, IR21]. interventions [SJH^{+}25]. intra [GSK24, WHK^{+}23, ZJY^{+}24]. intra- [GSK24]. intra-defect [ZJY^{+}24]. intra-service [WHK^{+}23]. introducing [BDMP21]. Introduction [AMM25, BDLR23, EHB21, HS21, KRD^{+}23, FBMR20, HMR21]. intrusion [FA24]. intrusive [CSLN23]. intuition [GDK24]. invariants [NHA20]. inverse [CKM^{+}20]. Investigating [BbASP23, DFT24, ET21, JZM21, KdJPK^{+}23, SSO^{+}23]. Investigation [LCP^{+}24, TOO^{+}23, GS20, JKD25, KH24, MVŞ23, PMM^{+}25, ROL21, SXL^{+}22]. involved [BHS24]. IoT [CAF22, CPD20, EN23, GPF22, KRSW22, LZB^{+}23, MdOT23, PKB24, POWGH22, RFB20, SBF20, SMB^{+}20, SRM^{+}25, TAF^{+}20, VTS22]. IoT-enhanced [VTS22]. IRFL [HH25]. ISO [GRV^{+}21]. ISO/IEC [GRV^{+}21]. isolate [LWZ21]. isolation [CDO24]. Issue [AX21, EHB21, KRD^{+}23, WDL^{+}25, BDLR23, BCW21, CLNQ24, FRC24, FBMR20, GDLM21, HS21, LCC^{+}23b, LSSZ21, RMOGA20, SHB21]. Issue-commit [WDL^{+}25]. issues [CPC^{+}23, FGRF24, GGB^{+}22, HTC^{+}23, LST20, MMC24, MH24, Mec24, MR24, RMOGA20, ZXZS24]. iStar [GAC20]. IT/SE [KGM24]. item [RF24a]. item-level [RF24a]. Iterative [YYW^{+}24].

January [Ano20q, Ano21r, Ano22q, Ano23q, Ano24p, Ano25j]. Java [CAC20, CA20, DD20, HSVMB20, HG22, HHK20, MREVEA^{+}22, MLBD21, MSB23, SS23, SGV22, THG20]. Java-based [SGV22]. Javadoc [BGE^{+}21]. JavaScript [JZM21, PZDG21, TOO^{+}23]. JITGNN [KRP24]. Job [Bat20, GAS25, KGM24]. Job-work [Bat20]. jobs [ZPL21]. joint [CLZ^{+}23, HPZ^{+}20]. Journal [NBGC20a]. journey [ES23]. judgment [MGSC22]. July [Ano20r, Ano21s, Ano22r, Ano23r, Ano25k]. June [Ano20s, Ano21t, Ano22s, Ano23s, Ano24q, Ano25l]. JUnit [GKAHMO22, WC23]. Jupiter [GKAHMO22]. Just [JSG25, KRP24, LID^{+}22, AAB^{+}22, CXP^{+}23, DXD25, NNV^{+}24, YZZ^{+}24, ZSCD22]. Just-In-Time [KRP24, JSG25, LID^{+}22, AAB^{+}22, CXP^{+}23, DXD25, NNV^{+}24, YZZ^{+}24, ZSCD22].

K-Clique [FC20]. Kaiaulu [PKT24]. Kappa [PDGMT20]. KDM [dSS^{+}22]. KDM-represented [dSS^{+}22]. keep [PdS23]. Kernel [BSS^{+}25, DD20, DFC^{+}23, NEJ24]. kernel-level [NEJ24]. kernels [ZWF^{+}25]. key [MAK^{+}21, MMB22]. keyword [CZW20]. kill [ZZP21]. killing [WYG^{+}24]. KLEE [ZWY^{+}22]. knight [BCOC24]. knowledge [AAZB23, BLTX21, FJvdW20, GFCLB^{+}24, LLL^{+}22, LSFE21, LAL^{+}23, NdSR^{+}21, OWGS23, RPT23, WGL^{+}22, WZFD25, ZLSY23]. knowledge-aware [ZLSY23]. knowledge-based [OWGS23]. Kotlin [CFAT25, MREVEA^{+}22]. Kromaia [BFZC21]. Kubernetes [TBD^{+}23, VSTK21]. Kulla [RAGCSS^{+}20].

L [CCZ^{+}24]. labeled [HOCK22]. labels [SJC^{+}22]. landscape [ZZJ25]. Language [BSF^{+}25, MPFB23, YZM^{+}25, AC25, ADG^{+}20, AMRS20, ES24, ACD^{+}24, BCF23a, BCOC24, BPJ^{+}22, BBW22, CF23, CF24, CLNQ24, DDCC24, DHZ^{+}24, FvB24, HBN^{+}25, KBG^{+}24, LMR^{+}23, LQY^{+}22, NRRS20, OBW^{+}25, PSZ21, PC23, RPL^{+}21, VWSCH23, XXC^{+}24, ZHS^{+}24]. language-parametric [KBG^{+}24]. languages [BRSK24, BB22, ECS23, GSF24, KBG^{+}24, PBV24, RSL^{+}21, YOH^{+}23]. Large [PN21, TPGH20, YZM^{+}25, BHS24, Bat20, BSDB20, CSLN23, DFC^{+}23, IBP21, KKH^{+}21, LST20, LK23, MSMB21, MSB23, OBW^{+}25, PDDD20, PSGD23, PGR24, QWG22, SJC^{+}22, UPP^{+}22, WZC^{+}20]. large-scale [Bat20, BSDB20, CSLN23, DFC^{+}23, IBP21, KKH^{+}21, LST20, MSMB21, PSGD23, SJC^{+}22, UPP^{+}22, WZC^{+}20]. latency [BBF^{+}21, MdSKD22]. latency-sensitive [MdSKD22]. LAURA [TAF^{+}20]. layer [CAF22, NBGC20a, NBGC20b]. layered [BBW22, KL20]. layout [BHHQ^{+}22, CATA21]. layout-based [CATA21]. LCVD [WTG23]. LDA [YOH^{+}23]. lead [TST24]. Leading [WSL^{+}20]. leaks [BF22, PCC^{+}20]. learned [BDMP21]. learner [ZFG^{+}23]. learners [CCR24]. Learning [JST^{+}23, LSG^{+}24, LH24b, MLPC20, PAM^{+}21, QWHH23, YWS23, YZM^{+}25, ZZK^{+}24, ZYL^{+}22, ALZ^{+}20, ES24, AWHS22, BKG20, BRSR21, BK20, BHHQ^{+}22, BN23, CLZ^{+}24, CJR22, CBW^{+}23, CCCW24, CDLN21, DXD25, DL21, DZY^{+}23, DBB20, DHZ^{+}24, GTT^{+}20, Gir21, GBK^{+}23, HGHK24, HTC^{+}23, HHJ^{+}20, ILUN21, ISW25, JZW^{+}21, JSG25, KR23, LLL^{+}22, LC20, LXYL20, LYG^{+}24, LAL^{+}23, LLZ^{+}22, MFPB24, MMB22, MRR^{+}20, MBO^{+}22, MRM^{+}22, NNV^{+}24, NWZ25, PT21, PDDD20, PNND25, PGR24, QWG22, SvdBHV24, SELS21, SKG^{+}24, SBM23, TCTL25, TTB^{+}23, WLL^{+}24, WJX^{+}24, XZY^{+}20, YLHZ20, YYW^{+}24, YLGS24, ZJXG20, ZCLP21, ZXH^{+}24, ZWF^{+}25, ZHLR23, ZGW^{+}20b, ZB22, ZXZS24, ZSF^{+}23]. learning-based [ALZ^{+}20, DL21, DBB20, NNV^{+}24, ZGW^{+}20b]. Legacy [HOAM23, YAaO^{+}21, ASM^{+}21, PZDG21, YLC25]. legal [YRC^{+}25]. legibility [OSM^{+}23]. less [SRTGA^{+}25]. less-than [SRTGA^{+}25]. lessons [BDMP21, OSJB21]. level [BN23, FYH^{+}24, Hyr25, LWYW20, MHR^{+}25, NMG^{+}20, NEJ24, PPB20, RV22, RF24a, XXC^{+}24, ZLSY23, ZWL^{+}24]. levels [HTG^{+}24]. Leveraging [AFJ^{+}20, CF23, HBK25, ZLSY23]. Lexical [CA20, LBG^{+}20]. Liability [DMN^{+}23]. libraries [NDDD20]. library [JZW^{+}21, SWZ^{+}20]. libre [NF24]. licensing [PKGA22]. life [NQT25, RATS21]. lifted [MSDC25]. lifting [ZLW^{+}24]. light [MDRdP^{+}24, Mec24]. light-weight [MDRdP^{+}24]. lightweight [MS25]. likely [NHA20]. limit [PdS23]. Line [KZTS22, ACA^{+}23, GFCLB^{+}24, HPF23, HBK25, KGS23, MRA^{+}22a, MSDC25, NNG21, NZA^{+}25]. linear [CCZ^{+}24, LLNC21, YLW^{+}21]. linear-time [CCZ^{+}24]. Lines [DMAH24, PtBD^{+}25, AGPR20, BMLH25, BEM^{+}23, BCF23a, CBC^{+}24, CF23, CF24, CTD^{+}23, EM24, Ger25, LtBD24, POZ20, tBS23, APAF21]. linguistic [BCF23b, POWGH22]. Link [MLPC20, WDL^{+}25, LMR^{+}23, LSF^{+}25]. linked [LSF^{+}25]. linking [MHR^{+}25, RPR22]. links [WCH^{+}22, ZXZS24]. Linux [BSS^{+}25, DFC^{+}23]. List [BSS^{+}25]. Literature [MYH^{+}24, PFdMF21, SJH^{+}24, VKG^{+}23, ACM25, AÇCT21, AWMW20, BAV20, dACOdS23, DAP20, DN21, DL22, GRLA20, GB20, GAB20, GGP^{+}24, GPF22, GdOB25, HTN^{+}25, HH22, ISW25, Jør24, LMR^{+}23, LHG^{+}25, LBT^{+}21, LZB^{+}23, LLZ^{+}23, MMC24, MH24, MRA^{+}22a, Mec24, MFLS22, MWFK20, MLJ23, MRM^{+}21, OSM^{+}23, PAM^{+}21, PDGMT20, PSAB22, QWGS24, RBS^{+}23, SL20, SHG24, Tai24, VDVC21, VIL24, WFR21, WCL23, ZNPR^{+}23, ZZK^{+}24, ZCL24, ZH22]. Live [AHFP24, CWTL21, KBP24]. lived [EKB^{+}23]. LLM [CKY25]. LLVM [ZRGJ21]. Load [AFJ^{+}20, AA23, DTZ^{+}22, GGMB^{+}22]. local [LCW^{+}24]. Localization [LWL^{+}23, ZWL^{+}24, Hu24, Hu25, KGL^{+}22, LWL20, LCC^{+}23a, LLW^{+}24, LLH^{+}25, MMSM^{+}23, MHH21, RF24a, TSLHS21, TPV25, VHSB22, XBS21, XCZ23, YWY^{+}21, YJZZ23, YWCX24]. Localizing [Dut24]. locate [PCC^{+}20]. Locating [AHL22, WZFD25, JT20, LFFW23]. Location [THB22, CFNA25, EFPC21, EM24]. locators [DFT24]. Log [MFPB24, FYH^{+}24, GLCH25, HDX^{+}23, SHLM24, YLG^{+}25, YCWD23]. log-based [HDX^{+}23]. LogGzip [GLCH25]. logic [BHP^{+}21, SB23, YB24, BJB^{+}21]. logic-based [SB23, YB24, BJB^{+}21]. logical [KSG^{+}22]. logistics [YYT25, ZH21, ZH22]. logo [BCF23a]. LogOW [YLG^{+}25]. Long [WDO24, EKB^{+}23, GBT^{+}20, SAZN22]. long-lived [EKB^{+}23]. Long-term [WDO24]. longevity [GW24]. longitudinal [BHS24]. look [LYS^{+}23, NQT25]. Looking [FGS23]. looming [GSM^{+}23]. Loop [WTG23, SCN^{+}21]. Loop-oriented [WTG23]. lossless [GLCH25]. low [ACM25, BSS^{+}25, YB24]. low-code [ACM25]. LR [RF24b]. LR-graphs [RF24b]. LSTM [LZJ20]. LWS [HDX^{+}23].

Machine [ILUN21, ALZ^{+}20, BKG20, BRSR21, BK20, BHHQ^{+}22, DL21, DBB20, Gir21, GF24, HTC^{+}23, ISW25, MRR^{+}20, MRM^{+}22, NWZ25, PT21, PSZ21, PDDD20, PNND25, QWG22, SvdBHV24, SKG^{+}24, SBM23, TMV25, ZGW^{+}20b, ZB22, DD20, HLW^{+}23, JCNS^{+}22, THB22]. machine-learning [BKG20, PT21]. machine-learning-based [PDDD20]. machine-learning-enabled [ISW25]. Machines [Ibi22]. Maestro2 [HTG^{+}24]. Mailing [BSS^{+}25]. main [SWH^{+}20]. maintainability [CFAT25, JKD25, MVŞ23, MSB23, SBM23, YB24]. maintained [MCH22]. Maintaining [BGL^{+}20]. maintenance [ACH25, AXUO24, DFB20, HELW20, TAT^{+}23]. makes [DDPP23]. making [HBK25, KRC23, LTG^{+}25, MBP23, RPT23, ŠMF^{+}23]. malicious [FR20, SJC^{+}22]. malware [MMC24, dCMM^{+}22]. Management [WB23, AKBN20, AZR^{+}21, FMF^{+}24, GTT^{+}20, GFCLB^{+}24, LWSZ23, LTG^{+}25, MPGB22, MLPC20, MBO^{+}22, NdSR^{+}21, RPC^{+}24, SZS25, TGG21, Tai24, TGX^{+}22, VVBGL^{+}23]. manager [SHA21, WB23]. Managing [BBF^{+}21, FMG^{+}25, MSKS24, TH24, YYT25, VSTK21]. Manifesto [RBvB^{+}24]. Manipulation [EHB21, HS21]. many [DWH23]. many-criteria [DWH23]. Map [SRM^{+}25, CDP22, MGSC22, OEW22]. MAPE [SB23]. MAPE-K [SB23]. Mapping [DJR^{+}22, SGM23, WLS20, AKL25, A– DHM23, AACC21, AGP22, BRS^{+}22, BM23, CJNDZ21, ETD^{+}24, FdSN^{+}20, GMFO^{+}21a, GMFO^{+}21b, GRLA20, HATG21, HWL^{+}25, KI23, KAA^{+}21, LMZ^{+}23, LBMF^{+}22, MAK23, MSB21, NdOdO^{+}22, OBCR23, PPMC22, PC23, PTW22, PNND25, dPPdJFF21, PÁngeles MoragaGC25, RATS21, UPP^{+}22, WUK^{+}21, WLdCM22, POZ20]. March [NBGC20a, Ano20t, Ano21u, Ano22t, Ano23t, Ano24r, Ano25m]. Market [ZPL21]. Market-oriented [ZPL21]. Markov [AKBN20]. MAS4DT [PZVN^{+}25]. masked [XXC^{+}24]. Masking [FC20]. Mastodon [CID25]. match [FG24]. mathematical [MLBD21]. matrix [Hu24, RPR22]. matter [SM20a, TST^{+}21a, TST^{+}21b]. mature [MSS21]. maturity [FAFJ24, WMLM22]. Maven [HBSV^{+}22, KVP23]. Maximizing [MMB22, JS22]. May [Ano20u, Ano21v, Ano22u, Ano23u, Ano24s, Ano25n]. MBSE [GBT^{+}20, KSG^{+}22]. meaning [LB24]. means [BSCS23, KSF20]. measure [LALM23]. measurement [LBMF^{+}22, MFLS22, WFR21, ZNPR^{+}23]. measures [VML21]. Measuring [GAS25, PSL25, ZZL^{+}23]. mechanism [BAH24]. mechanisms [HHK20, HM24, NDP^{+}21, TSP20]. mechatronic [VHB21]. media [CID25]. meetings [SM20b, TC22]. meets [IMTS23].

Membership [LYZ^{+}22]. MeMo [BGE^{+}21]. Memory [HLL21, LWC20, WDF^{+}22]. mental [BAV20, Tok22]. merge [AMS23, FFSB23, SM24]. message [LCT22]. message-based [LCT22]. messages [MRS25, PMDN20, TGG21]. messaging [SGG22, SGG24]. meta [FvB24, HSVMB20, NdSR^{+}21, ZHS^{+}24]. meta-decompilation [HSVMB20]. meta-ethnography [NdSR^{+}21]. meta-language [FvB24]. meta-model-based [ZHS^{+}24]. metaheuristic [ZYZZ21]. metamodel [HPF23]. metamodels [NDP^{+}21]. Metamorphic [AJFD25, BGE^{+}21, CKY25, LLNC21, SG20, aSLF^{+}22, WXZL23, XTF^{+}21, YWY^{+}21, YLW^{+}21]. metaphor [MLGBRC24, MCDP24]. MeTeaM [MSS21]. METFORD [VKB^{+}25]. method [BKG20, CR23, JWZM20, KVP23, KRS^{+}21a, KRS^{+}21b, LWYW20, LXYL20, LM21, MSS21, NdSR^{+}21, PPB20, RJ25, SAZN22, SV25, VKJ^{+}21, WGL^{+}22, WZC^{+}20, XBS21, ZCLP21]. method-level [LWYW20, PPB20]. methodological [ACH25]. Methodology [CNdLL20, AAW20, Bat20, CDRV20, Mec24, WUK^{+}21]. Methods [KRD^{+}23, AIH20, HHK20, HLS24, HWL^{+}25, Mam23, MRM^{+}21, PLL^{+}23, RHJ25, SM20b, ZH21, ZH22]. metric [SHLM24, WUK^{+}21, XZY^{+}20]. metrics [AA23, ACH25, CCC^{+}24, CXP^{+}23, EKHJ^{+}20, HJK^{+}21, LMZT22, MCM^{+}24, MSS21, DDPT20, PT21, PHLHM24, Pat20, WNZ25]. metrics-based [WNZ25]. MGA [DWG^{+}24]. mHealth [AAZB23]. Micro [RRFD24, CDN^{+}22, SGW^{+}22]. Micro-containers [RRFD24]. micro-rejuvenation [CDN^{+}22]. microcontrollers [RRFD24]. MicroFog [PKB24]. MicroIRC [ZWL^{+}24]. Microservice [AFJ^{+}20, LGMP24, WHK^{+}23, ZWL^{+}24, ACC^{+}25, ATWS25, AÇCT21, AKMS23, BBG^{+}25, BSH^{+}20, CAA^{+}23, CDET22, FA24, GGPR24, LHG^{+}25, MPGB22, RFSP^{+}25, SS23, TAT^{+}23, VSTK21, XCZ23, ZLC^{+}23]. microservice-based [CDET22, TAT^{+}23]. Microservices [WLS20, AKMS23, CJR22, DDPP23, GGP^{+}24, HH24, HWL^{+}25, LLST20, LKP^{+}21, MVSV25, MTA24, NSL^{+}21, NSR^{+}23, PKB24, PHLHM24, PSAB22, VTS22, WLS^{+}21, WLH^{+}25, YB24, ZZL^{+}23, dTMS21]. microservices-based [PKB24]. mid [dSNJ25]. mid-sized [dSNJ25]. migrating [LLST20]. migration [HOAM23, HTB21, KBP24, TMV25, YAaO^{+}21]. million [HBSV^{+}22]. MILP [ZZTC23]. Mind [SRM^{+}25]. minimal [Mas22]. minimization [LMVRA^{+}20, ÖS20]. minimum [XCH^{+}20]. Mining [AAG21, BLTX21, FPB^{+}24, GCF22, MLS^{+}21, WLCC25, ZZH^{+}25, CBC^{+}24, GGB^{+}22, MWY^{+}22, ZWP^{+}24, dCMM^{+}22]. misbehaviour [MRM^{+}22]. misexposure [DYZ^{+}23]. Missing [WGL^{+}22]. mission [ACA^{+}23, RCA^{+}22]. mission-specific [ACA^{+}23]. mitigation [WDF^{+}22]. MITU [WZFD25]. mixed [CDO24, SM20b]. mixed-criticality [CDO24]. mixed-methods [SM20b]. mixup [DHZ^{+}24]. mixup-based [DHZ^{+}24]. ML [HGHK24, KMS^{+}25, NBP24, Spa24]. ML-enabled [NBP24]. MLOps [WNZ25]. Mob [SM21]. Mobile [AJFD25, AXUO24, AAZB23, AF22, CXY^{+}23, CDN^{+}22, DTZ^{+}22, GMS22, LSFE21, LFH^{+}22, LDT^{+}25, LCT22, NdOdO^{+}22, NBGC20a, NBGC20b, SSO^{+}23, WLdCM22, ZFC^{+}22, SMdS^{+}23]. mobility [MPGB22]. modal [CWTL21, GJW^{+}22, JSG25, TQT^{+}25]. modality [YWLZ23]. mode [LH24b, RSVW23]. mode-switching [RSVW23]. Model [AMB^{+}25, BDIL25, CDP22, CTD^{+}23, DAML23, GKB^{+}21, Gla23, GdOB25, GAL20, KRSW22, MKB25, POZ20, SVV24, SBM23, YZM^{+}25, BS23, BBW22, CBC^{+}24, CVC21, CLC^{+}23, CLLJ24, CYW21, CDM25, CDET22, DFT24, DLBE22, ESO^{+}24, GB20, GDLM21, HZ24, HIDT21, HLS24, HGH^{+}23, HLZ^{+}23, HCC22, IIK21, KKRT24, KSF20, KRLS^{+}20, LLW^{+}23, LLL^{+}25, NBGC20a, NBGC20b, OZVRD21, PBV24, PPL^{+}25, RPS25, RAGCSS^{+}20, RSM^{+}23, RSVW23, ROGS^{+}24, dORGCG23, RHJ25, SHLM24, SSZ20, SOMSCT23, VGS^{+}23, WNZ25, XXC^{+}24, XZZ^{+}25, YXLH24, YLG^{+}25, YB24, ZZTC23, ZHS^{+}24, ZWF^{+}25, ZSCD22, ZSF^{+}23, FBMR20]. Model-Based [SVV24, CDP22, GKB^{+}21, MKB25, POZ20, ESO^{+}24, HLS24, IIK21, RSVW23]. model-checking [PBV24]. model-connected [RPS25]. Model-Driven [DAML23, KRSW22, BDIL25, CTD^{+}23, GdOB25, CDET22, GB20, NBGC20a, NBGC20b, RHJ25, VGS^{+}23, WNZ25]. model-reduction [SSZ20]. model-to-text [HZ24]. Modelica [SSZ20]. Modeling [APR24, HCB^{+}20, LZJ20, ALZ^{+}20, BRS^{+}22, BPJ^{+}22, CBW^{+}23, GTT^{+}20, LBF^{+}21a, LBF^{+}21b, LHF22, MFF^{+}24, RCA^{+}23, SZS25, WLCC25, CDRV20, DJS^{+}22]. Modelling [SBC20, VTS22, AC21, BSF^{+}25, BGM^{+}21, BCJI22, LK23, SSZ20, ZTK^{+}23]. Models [MLPC20, MPFB23, BFZC21, CJR22, CAF22, CLNQ24, CFP^{+}21, DDCC24, DPRD21, EFPC21, ET21, EBSB23, FCW^{+}23, GBSO20, GGMH23, Hei20, HBN^{+}25, JJS24, JVZ25, KGS^{+}25, LATV22, LSG^{+}24, LDH22, LL25, MICV23, MBO^{+}22, NJF20, NM25, OBW^{+}25, PLP^{+}20, RK20, RJ23, SGG24, SZ22, SFR23, TBD^{+}23, VVBGL^{+}23, WJX^{+}24, XLY^{+}21, ZAS25]. moderating [LWC20]. Moderator [RCT22]. modern [DN21, HM24, THG20, dSNJ25]. modernity [AvdBG^{+}24]. modernization [ASM^{+}21, CF24]. modification [BN23]. modular [HPF23, NMV25, SVVD21]. module [PGR24, YDP22]. modules [DHK^{+}23, OEW22, PZDG21]. moment [CKM^{+}20]. monitor [EMCN^{+}22]. Monitoring [GGP^{+}24, CBMM20, FR20, JRM^{+}22, MN21a, MN23, SZSV22, VWSCH23, VGS^{+}23, WGMT24, WLS^{+}21]. monitors [ACD^{+}21]. monolithic [LLST20, WLH^{+}25]. Monte [HGH^{+}23]. month [BCF^{+}21]. MontiThings [KRSW22]. morale [BGMB20]. motivations [LSFE21, SWG^{+}20]. move [WLdCM22]. moving [TMV25]. MPI [DWG^{+}24]. MPLinker [WDL^{+}25]. MrsP [ZGW^{+}20a]. MSL [AMRS20]. MSTIL [QWHH23]. MT [CNdLL20]. MT-EA4Cloud [CNdLL20]. Multi [AAG21, AM23, AWHS22, BRS^{+}22, DHM^{+}21, GKW25, PZVN^{+}25, QWHH23, SJH^{+}24, WDL^{+}25, CAF22, CXP^{+}23, CYW21, CR23, DOG^{+}23, DLBE22, GGP21, GJW^{+}22, JYZ^{+}24, JWZM20, LTJ^{+}20, LQY^{+}22, LCC^{+}23a, MPRX20, MHR^{+}25, ÖS20, RPL^{+}21, SRD^{+}21, WMA^{+}24, WMS23, WZFD25, YYW^{+}24, YWLZ23, ZCL24, ZLSY23, LFFW23]. Multi-Agent [PZVN^{+}25, DLBE22]. multi-case [GGP21]. multi-concerns [LTJ^{+}20]. multi-criteria [ÖS20]. Multi-cue [QWHH23]. Multi-dimensional [AAG21, LCC^{+}23a]. multi-entity [JWZM20]. Multi-factory [DHM^{+}21]. multi-featured [SRD^{+}21]. multi-language [RPL^{+}21]. multi-layer [CAF22]. multi-level [MHR^{+}25, ZLSY23]. multi-method [CR23]. multi-modal [GJW^{+}22]. multi-modality [YWLZ23]. Multi-Objective [GKW25, AM23, CXP^{+}23, CYW21, DOG^{+}23, WMS23]. Multi-paradigm [BRS^{+}22]. multi-programming-language [LQY^{+}22]. multi-project [WMA^{+}24]. Multi-Relational [LFFW23]. multi-scale [JYZ^{+}24]. multi-source [WZFD25]. multi-task [AWHS22, YYW^{+}24]. Multi-template [WDL^{+}25]. multi-tier [MPRX20]. Multi-triage [AWHS22]. Multi-vocal [SJH^{+}24, ZCL24]. Multilayered [DL21]. Multilevel [DD20, AGL21]. multilingual [ZYK^{+}25]. multiple [BRO^{+}22, CAC20, FKF^{+}23, HTB21, HLS24, LLNC21, SMVC24, XBS21, ZLW^{+}23, ZYZ^{+}23, dTMS21]. multiple-case [HLS24, dTMS21]. multiple-fault [XBS21]. multiple-sampling-enhanced [ZLW^{+}23]. multitasking [RPM^{+}22]. multivariate [ABT^{+}22]. Multivocal [ACC^{+}25, MYH^{+}24, PSAB22, SL20, TRD^{+}23]. Must [RBvB^{+}24]. mutant [ZWF^{+}25]. mutants [MAK^{+}21, WYG^{+}24]. Mutation [LATV22, LWL20, TS22, BMP22, CF23, DWG^{+}24, KDB^{+}21, LLH^{+}25, MAK^{+}21, WYFZ24, YJZZ23, ZLD^{+}25, ZZP21, VKB^{+}25]. Mutation-based [LATV22, DWG^{+}24, LWL20, LLH^{+}25]. mutual [HLZ^{+}23, Ibi22].

Naive [OEW22, TLXW23]. name [SJC^{+}22]. named [PSGD23, ZLS20]. names [PSGD23, WC20, WC23, ZCLP21]. native [NBGC20a, NBGC20b]. natural [DDCC24, PC23, YOH^{+}23]. navigate [MGSC22]. Navigating [ZZJ25]. nearest [LMVRA^{+}20]. need [AES22, LDT22, WYFZ24, WLdCM22]. needs [DAML23, WGMT25]. negative [AMHB24, PPB20]. neighborhood [LMVRA^{+}20]. nervous [WFR21]. Nested [LFFW23, ZGW^{+}20a]. Net [DLCH25]. nets [APR24, SCdPL24, WLLJ24]. Network [FZT^{+}22, MSC20, OHM25, AWA^{+}22, CLC^{+}23, FWZZ25, HATG21, HLZ^{+}23, JYZ^{+}24, KRP24, LATV22, MCSAGB20, QZS^{+}24, RF23, SCPB25, SZSV22, WTG23, YZL^{+}22, YLLY25, ZLX^{+}22, ZLS20, ZYZZ21, LBH^{+}24, LFFW23]. networking [GAL20]. Networks [ZYL^{+}22, AAB^{+}22, CLZ^{+}23, CFF^{+}23, FR20, GJW^{+}22, NDP^{+}21, QHC^{+}24, ZSZ^{+}22]. Neural [HPZ^{+}20, OHM25, CLZ^{+}23, FWZZ25, JSTW22, KRP24, LLH^{+}25, NM25, NDP^{+}21, QZS^{+}24, QHC^{+}24, RF23, WTG23, YZL^{+}22, YLLY25, ZLS20, ZYZZ21, HLW^{+}23]. Neural-FEBI [HLW^{+}23]. next [KA24, dSIdÁ23]. NFV [HTB21]. NHPP [LDH22]. NHPP-based [LDH22]. nine [GCF22]. NLI [SWZ^{+}20]. NLP [FFV^{+}23, LCS^{+}24, SCdPL24]. NLP-based [LCS^{+}24]. no [ACM25]. no-code [ACM25]. node [XXC^{+}24]. node-level [XXC^{+}24]. Noise [BIT24, ASSH22, HTC^{+}23]. Non [WDF^{+}22, BFHC20, CSLN23, WC20, XZZ^{+}25, YU22]. non-autoregressive [XZZ^{+}25]. non-descriptive [WC20]. non-functional [BFHC20]. non-intrusive [CSLN23]. non-security [YU22]. Non-volatile [WDF^{+}22]. NoSQL [ALB20, AON^{+}24, BMB20]. NoSQL-based [AON^{+}24]. novel [CDRV20, DTZ^{+}22, FGG^{+}20, KZTS22, LKP^{+}21, PZVN^{+}25, RJ25]. November [Ano20v, Ano21w, Ano22v, Ano23v, Ano24t]. novice [BIT24, CKP20, CCR24, LWL^{+}23, NIJ22]. npm [ANC^{+}23, MAS23]. nuances [ZLS25]. Numerical [MPFB23].

obfuscation [FC20]. object

[CCZ^{+}24, MCDP24]. object-oriented [CCZ^{+}24, MCDP24]. Objective [GKW25, AM23, CXP^{+}23, CYW21, DOG^{+}23, WMS23, ZPL21, dSIdÁ23]. objectives [BFHC20]. objects [GM20, RA25]. Observability [GRdAL23, LCAC21, ZZP21]. observable [AKBN20]. Observation [LBF^{+}21a, LBF^{+}21b, JS22]. Observation-based [LBF^{+}21a, LBF^{+}21b]. observations [Hei20, LPPG20]. observers [AdlBGZ^{+}23]. occurrence [IZAD21]. occurrent [OAH^{+}23]. October [Ano20w, Ano21x, Ano22w, Ano23w, Ano24u]. off [DCÁ^{+}24, FEBO22]. OMG [CPD20]. on-demand [DLCH25, RPT23]. onboard [SM20a]. Onboarding [AAA24, BSDB20, SS20]. One [XDL^{+}22, SKHLS24, MSC20]. one-class [SKHLS24]. OneSpace [ES24]. Online [EM20, SJC^{+}22, CACHA23, GF24, HM25, LCC^{+}23a, PGR24, SMM23, THB22, WMS23, ZPL21]. Only [LDT22]. onsite [SMM23]. ontology [KL20, LC20]. ontology-based [LC20]. OOD [ZAS25]. OP [CCZ^{+}24]. OP-ART [CCZ^{+}24]. opaque [FC20]. Open [JLL23, JVZ25, NAV^{+}24, SVAGB20, AKMS23, BMHR21, BGL^{+}20, BGL^{+}22, ES23, Fei23, FRC24, JCNS^{+}22, LMZT22, NRRS20, NMV25, NF24, OMA^{+}22, PKGA22, ROGS^{+}24, THG20, TKSC20, VDVC21, WMLM22, YLG^{+}25, ZKDP22, dlVRB21, ROL21]. open-source [AKMS23, ES23, Fei23, JCNS^{+}22, NMV25, OMA^{+}22, THG20, TKSC20, ZKDP22]. open-world [YLG^{+}25]. Operating [A– DHM23]. operation [LM21]. Operational [AFJ^{+}20]. operations [BAH24, LL23]. operator [KGL^{+}22]. Opportunistic [MSC20, WMA^{+}24]. opportunities [BM25, BGC20, GCSHB20, Hyr25, KdJPK^{+}23, MSB23]. opposition [ZHLR23]. opposition-based [ZHLR23]. Optimal [AdlBGZ^{+}23, NJF20, BRO^{+}22]. optimising [CNdLL20]. optimization [BH20, CYW21, EM20, Hu24, KKRT24, OPF24, PSL25, WMH^{+}24, YDP22, ZRGJ21, ZYZZ21]. Optimize [vRMG23, BFHC20]. optimizer [ZHLR23]. Optimizing [YWLZ23, LTG^{+}25, OLR^{+}25]. oracle [MDML25, RPS^{+}23]. oracles [SRTGA^{+}25]. orchestration [KBP24]. order [AdlBGZ^{+}23, DZY^{+}23, WYFZ24]. organisational [KdJPK^{+}23]. organization [BHS24]. Organizational [Gla23, SS20]. organizations [MSKS24, PN21]. oriented [BSH^{+}20, CCZ^{+}24, DL22, ECS23, GVBC^{+}24, MMDL23, MAV24, MCDP24, WTG23, WZL^{+}25, YXLH24, ZPL21]. orthogonal [MPP24, ALB20]. OSCAR [SGG^{+}25]. OSCAR-P [SGG^{+}25]. OSLC [NMG^{+}20]. OSLC-based [NMG^{+}20]. OSS [BDLT21, MSB23]. outcomes [CPC^{+}23, LGKT22]. outdatedness [DMD23]. output [LMVRA^{+}20]. over-reliance [MHJW22]. Overflow [AA24, SZ23, ZLG25, ADG^{+}20, BLTX21, CKP20, CXH^{+}24, GMCA21, MR24, RSU^{+}24, STW23, SZ23, TXW^{+}20, ZLW^{+}23, ZLZ^{+}25, ZLS22, ZLS25, dDLSK23]. overhead [ZGW^{+}20a]. overhead-aware [ZGW^{+}20a]. overloading [Spa24]. oversampling [ZYZ^{+}23]. Overview [JLL23, GAB20]. OWASP [FAFJ24]. own [SM20a, CTD^{+}23]. Owner [KdJPK^{+}23].

P [SGG^{+}25]. PaaS [GGB20]. PaaSArch [GGB20]. package [MCSAGB20, Vid22]. packages [ANC^{+}23, MAS23, MSB23]. PageRank [YJZZ23]. pains [ZLC^{+}23]. pair [DMN^{+}23, ZLZ^{+}25]. pairs [TXW^{+}20]. Pandemic [AAA24, MBKL24, PGR24, STM^{+}22, SMH^{+}23, Tok22]. papers [WCH^{+}22]. paradigm [BRS^{+}22]. paradigms [CTD^{+}23]. Parallel [JYZ^{+}24, PPL^{+}25, RAGCSS^{+}20, SXL^{+}22, SVVD21, ZPL21]. parallelization [OLR^{+}25]. parameter [SPZW25]. parameter-efficient [SPZW25]. parameterized [CWTL21]. parametric [ACG^{+}21, KBG^{+}24]. parser [LWP^{+}21]. Parsing [GLCH25, LL25, LL23, SHLM24, YCWD23]. part [PGR24]. part-time [PGR24]. partial [AdlBGZ^{+}23, SJC^{+}22, ZFS^{+}22]. Partially [AKBN20]. participate [SMdS^{+}23]. participation [NF24]. particle [WMH^{+}24]. particular [MVŞ23]. partitioned [MCMA21]. party [NDDD20]. past [ADS^{+}22, SGM23]. past-CTL [ADS^{+}22]. patches [ASK^{+}23]. patching [IPB23]. Path [aSLF^{+}22, DWG^{+}24, DBO24, WXZL23, ZJXG20, ZWY^{+}22, ZZTC23]. Path-directed [aSLF^{+}22]. paths [BH20]. pathways [HM25]. pattern [AZ21, AMRS20, BRSR21, BCOC24, FJvdW20, HSJB23, LLL^{+}25, MWY^{+}22, NAZ22, SB23, WC20]. pattern-based [HSJB23, WC20]. pattern-driven [FJvdW20]. patterns [ASCR23, AÇCT21, AHL22, BKG20, CAA^{+}23, CID25, GSK24, KFJA23, LMGK22, LLL^{+}25, LLZ^{+}22, MTA24, NAD^{+}20, OBW^{+}25, PFW21, PÁngeles MoragaGC25, SV20, TOO^{+}23, ZFC^{+}22]. pay [LDT22]. payment [FRP^{+}23]. PDFBox [BGL^{+}20]. pedestrian [CXY^{+}23]. People [KRC23]. perceive [SGW^{+}22]. Perceived [RCA^{+}23, STM^{+}22, VML21]. perception [AAZB23, MBG^{+}25]. Perceptions [GW23, MVGHPT22, FAFJ24, dSNJ25]. Performability [AON^{+}24]. Performance [RK20, AMHB24, ABT^{+}22, CJR22, CBW^{+}23, CYW21, CDET22, ECL^{+}22, ET21, ETD^{+}24, JST^{+}23, LATV22, LABJ23, LSG^{+}24, LBH^{+}24, LGKT22, MTA24, MH20, NMRS22, NEJ24, OBCR23, PPB20, RCT22, RZLC24, SGG^{+}25, SL20, Tai24, WLCC25, ZGW^{+}20b, dCMM^{+}22, vRMG23, vDDPZ24]. perils [MHJW22]. periods [GS21a, GS21b]. permission [GGB^{+}22, SLL20, XCH^{+}20]. permission-based [SLL20]. personality [FKGN23, VM24a, VM24b]. personalized [ET21, SZS25]. perspective [EM24, GAB20, Gir21, HOAM23, HNM^{+}24, LGT^{+}23, MKB23, PFC^{+}23, PPM^{+}21, WLS^{+}21, WB23]. perspectives [BbASP23, Pie20]. persuasion [CACHA23]. perturbations [QHC^{+}24]. pervasive [GM20]. Petri [APR24, SCdPL24]. Petri-nets [SCdPL24]. phase [FKF^{+}23, Jør24, vdG25]. phishing [SJH^{+}24, SJH^{+}25]. PHP [RB23]. Physical [AAG21, FMG^{+}25, VWSCH23, ZKDP22, vDDPZ24, ACD^{+}21, BRS^{+}22, BGM^{+}21, BCW21, DSM20, FKF^{+}23, GB20, GCLB22, GAB20, LTJ^{+}20, MFF^{+}24, PPL^{+}25, RATS21, SCN^{+}21, SSZ20, dlVRB21, BJB^{+}21]. pipeline [DJS^{+}22, FGRF24, KBW^{+}24, SFR23]. PL [TS22]. PL/SQL [TS22]. placement [LLWL22, PKB24, THB22]. planning [DHM^{+}21, HTB21, WCGS23, WXZL23, ZZTC23]. platform [BCOC24, BL21a, BL21b, CSLN23, FCW^{+}23, LWSZ23, RLK^{+}24]. platforms [CID25, DHM^{+}21, YSB^{+}21]. platooning [SB23]. pleasingly [ZPL21]. PMTT [JYZ^{+}24]. point [FRP^{+}23, MLYL25, YZS^{+}25]. PointNet [ZCZL23]. points [LL21]. policies [AKBN20, KCMD21, LSB^{+}22, SHWR22, SMH^{+}23]. Policy [LYZ^{+}22, LHF22]. Policy-driven [LYZ^{+}22]. polynomial [vdG25]. polynomial-based [vdG25]. pooling [DHZ^{+}24]. popularity [KVP23, WWC23]. Portfolio [MDAHS25]. positions [HHJ^{+}20]. Post [THN20, PGR24, SMH^{+}23]. post-pandemic [PGR24, SMH^{+}23]. post-processing [KGL^{+}22]. Post-Release [THN20]. posts [CKP20, GMCA21, STW23, SZ23, WDT^{+}24, YOH^{+}23, ZLW^{+}23, dDLSK23]. Potential [MH24, WMH^{+}24, AWP^{+}25, ETY^{+}22, MW25, WDT^{+}24]. Potential-aware [WMH^{+}24]. potentials [SBF20]. Power [LOBD25, DCÁ^{+}24, EMCN^{+}22]. power-up [EMCN^{+}22]. powered [CF24]. PR [GDLM21]. Practical [Ala21, SSP21, PBV24]. Practice [LGMP24, BDMP21, CRV23, FAG^{+}20, FEBO22, GKB^{+}21, HO22, NLTM23, OZVRD21, SM21, SFR23, dSNJ25, FBMR20]. practices [AC25, AACC21, DAML23, GCSHB20, GSM^{+}23, HH25, KKH^{+}21, NSR^{+}23, NBP24, NWZ25, SHG24, SvdBHV24, SGW^{+}22, WKP20, ZLC^{+}23]. practitioner [FRP^{+}23, SJH^{+}25, WLS^{+}21]. practitioners [VML21]. Pragmatic [LMR^{+}23]. Pre [KGL^{+}22, CLNQ24, FCW^{+}23, HOCK22, PGR24, SSS22, ZAS25]. pre- [PGR24]. pre-labeled [HOCK22]. pre-recorded [SSS22]. pre-trained [CLNQ24, FCW^{+}23, ZAS25]. Pre/post [KGL^{+}22]. Pre/post-processing [KGL^{+}22]. Precise [ZYL^{+}22, MC20, PBC^{+}23, WTS23, XCZ23]. precision [LPS^{+}23, SSP21]. Predicting [MS25, MPRX20, PT21, TBD^{+}23, HCB^{+}20, JYZ^{+}24, LMVRA^{+}20, RMOGA20, SGG^{+}25, ZCLP21, ADG^{+}20]. prediction [ASKS20, AMHB24, AMS23, AXUO24, BS23, CXP^{+}23, CLLJ24, CYW21, DOG^{+}23, DXD25, DYZ^{+}23, ET21, FGG^{+}20, GBK^{+}23, GF24, HHJ^{+}20, ILUN21, JSG25, KRP24, LBH^{+}24, LCC^{+}23b, LAL^{+}23, MS25, NM25, NQT25, PPB20, QZS^{+}24, SKHLS24, TXW^{+}20, TLXW23, WDO24, XZY^{+}20, XLY^{+}21, YZZ^{+}24, ZXH^{+}24, ZJY^{+}24, ZSCD22, ZYZZ21]. prediction-based [LCC^{+}23b]. predictions [MHJW23]. predictive [BS23, CLLJ24]. predictor [LALM23]. preemptive [LHF22]. preferences [BEM^{+}23]. preferences-based [BEM^{+}23]. preliminary [GGMH23]. preparation [ZZK^{+}24]. presence [ASK^{+}23]. preserving [Ger25]. Prevalence [RMT^{+}22, RPC^{+}24]. Prevent [Gla23]. Prevent-Model [Gla23]. Preventing [ZFC^{+}22]. prevention [Liu21]. PRHAN [FZT^{+}22]. price [AZR^{+}21]. primary [ZCL24]. prime [BH20]. principal [AAB^{+}22, dTMS21]. principles [ROGS^{+}24, YLC25]. prioritisation [MAP^{+}20, TC22]. prioritization [CQZ^{+}20, GS21a, GS21b, HZT^{+}20, HYF^{+}24, LBT^{+}21, MAV24, MN21b, aSLF^{+}22, ST24, ZFS^{+}22]. prioritize [MICV23]. priority [LCY^{+}24, LHF22]. PRISE [GAC20]. Privacy [BSCS23, GPF22, KMS^{+}25, KRC23, PFC^{+}23, YLLY25, ZZH^{+}25]. privacy-related [KRC23]. Proactive [BBG^{+}25, ATWS25, CLLJ24, TTL20, APAF21]. probability [RK20]. probes [AHFP24]. problem [ACC^{+}20, AWA^{+}22, CF23, FGRF24, KL20, ÖS20, Pie20, Woo25, dSIdÁ23]. problem-solving [AWA^{+}22]. problems [KOPN22, ZLZ^{+}25]. Process [GS20, AGPR20, AKBN20, AHL22, ACSJ23, ASK^{+}23, CBC^{+}24, DOG^{+}23, GAC20, GAS25, JCNS^{+}22, LWC20, MDVB^{+}23, MDRdP^{+}24, MAC^{+}22, PDGMT20, RK20, VVBGL^{+}23, SWH^{+}20]. process-centric [MDVB^{+}23]. Processes [HLS24, PFdMF21, CMP^{+}20, LBT^{+}21, MSC20, MFF^{+}24, NQT25, VTS22, dSNJ25]. Processing [VKG^{+}23, DHZ^{+}24, FGRF24, HH24, KGL^{+}22, KBW^{+}24, LLWL22, PC23]. ProCon [MDVB^{+}23]. ProDSPL [APAF21]. Product [APAF21, HJK^{+}21, KZTS22, LtBD24, NNG21, PtBD^{+}25, ACA^{+}23, BMLH25, BEM^{+}23, BBND^{+}20, BCF23a, CBC^{+}24, CF23, CF24, CTD^{+}23, DMAH24, EM24, Ger25, GFCLB^{+}24, HPF23, HBK25, KGS23, MRA^{+}22a, MSDC25, NZA^{+}25, POZ20, SWH^{+}20, WMLM22, tBS23, KdJPK^{+}23]. Product-line [NNG21]. Production [FMG^{+}25, DHM^{+}21, MDWS^{+}23, MFF^{+}24, MBO^{+}22, vBD21]. Productivity [LCP^{+}24, MVGHPT22, STM^{+}22]. products [MFF^{+}24, VHB21, ZGHG^{+}23]. professional [GBLP25, VM24a]. professionals [SS20, Tok22]. Profiles [AFJ^{+}20, AC21, ZLS22]. Profiling [DLV^{+}22, SGG^{+}25, PHLHM24]. profit [MH20]. prognostics [LWSZ23]. Program [GKW25, YWW22, ABMV24, AHFP24, AGL21, AS22, DWG^{+}24, ETY^{+}22, GKPS24, Ger25, GSF24, GDK24, HLL21, LBG^{+}20, LBF^{+}21a, LBF^{+}21b, LLK^{+}21, NZA^{+}25, SLL20, XXC^{+}24, YWCX24, YMDM21, DLCH25]. ProgrammableWeb [LTZ^{+}21]. programmer [DMN^{+}23, MBG^{+}25]. programmers [BIT24]. Programming [BIT24, AC25, ADG^{+}20, ACD^{+}24, CKP20, CND22, CCR24, DPD^{+}22, ECS23, GDK24, HBN^{+}25, LQY^{+}22, MDWS^{+}23, MBG^{+}25, RSL^{+}21, SGW^{+}22, SRD^{+}21, SM21, ZLZ^{+}25, GKW25]. programs [AIH20, BK20, BAV20, CCZ^{+}24, DDPP23, JZM21, LWL^{+}23, LZJ20, LZ24, MC20, PCJNP23, RF23, SV20, SV25, TS22, XBS21, YLW^{+}21, YZS^{+}25, ZMLZ23]. Project [YLGS24, BS23, BGL^{+}20, CCC^{+}24, DXD25, GS21a, GS21b, Jør24, KSF20, QZS^{+}24, SHA21, SMM23, TLXW23, WMA^{+}24]. project-based [SMM23]. Project-specific [YLGS24]. projects [BCJI22, BDLT21, BSDB20, CND22, CR23, DCMR20, Fei23, HOCK22, ILUN21, IT23, Jør24, JVZ25, KI23, KHEC^{+}23, LWC20, LQY^{+}22, LCY^{+}24, LGKT22, LMVRA^{+}20, MLBD21, NF24, RMOGA20, SHA21, THG20, WMLM22, WCH^{+}23, WWC23, HG22]. Prompt [WDL^{+}25, LYG^{+}24]. Prompt-tuning [WDL^{+}25]. Prompting [YZM^{+}25]. proneness [TPV25]. proof [AC21]. proof-of-concept [AC21]. propagation [LWZ21, ZWP^{+}24]. properties [ACD^{+}21, BGM^{+}21, BRS23, RA25, BJB^{+}21]. Property [AHFP24, DLCH25, PBV24]. proposal [BMB20, GVBC^{+}24, PZVN^{+}25]. proposals [LL21]. Propositional [MPFB23]. protection [YLLY25]. protocol [BCF23a, LKP^{+}21, SH20]. Prototyper [GGB20]. prototypes [GGB20]. prototyping [ZHS^{+}24]. provably [MPP24]. provide [WC23]. provided [MFPB24]. Providentia [BFHC20]. provider [MH20]. providers [AAZB23, LKP^{+}21]. Providing [RPM^{+}22]. pseudocode [XZZ^{+}25]. pseudocode-to-code [XZZ^{+}25]. PSTM [LZP^{+}20]. Psychometric [FKGN23]. public [BHS24, CYW21, WCH^{+}22]. publications [LHH^{+}24]. Pull [FZT^{+}22, LNST21]. Pytester [TCTL25]. Python [LGC^{+}25].

Q&A [AA24, HPZ^{+}20, Pat20]. QExplore [SMKI23]. QMM [FR20]. QMM-VANET [FR20]. QoS [FR20]. qSOA(R) [HPP24]. qualitative [BQR24, GPPDLF23, HLS24, SGG22, VHB21]. Quality [BLTX21, LCC^{+}23b, LBMF^{+}22, MVŞ23, BAH24, BRO^{+}22, BBND^{+}20, BHHQ^{+}22, CND22, CGM24, CLNQ24, EBSB23, FGRF24, FMF^{+}24, GGP21, GRV^{+}21, ISW25, KSF20, KRLS^{+}20, LNST21, LQY^{+}22, MDVB^{+}23, MBO^{+}22, OBCR23, OAH^{+}23, DDPT20, POWGH22, PN21, RBS^{+}23, RFB20, WMLM22, WNZ25, WCGS23, XZZ^{+}25, YB24, NLTM23]. Quantitative [ACPM22, WMLM22, BQR24, CBC^{+}24, CDRV20, ST24]. Quantum [OPF24, AMM25, BDS^{+}24, DPD^{+}22, DCÁ^{+}24, HPP24, KAW^{+}23, LZ24, OMA^{+}22, PCJNP23, SCD^{+}24, ZMLZ23, vdG25, GKW25, LZ25]. quantum/Classical [HPP24]. quasi [SBD23]. quasi-static [SBD23]. Query [YXLH24, Tai20]. Query-oriented [YXLH24]. question [AA24, TXW^{+}20]. question-and-answer [AA24, TXW^{+}20]. questions [ADG^{+}20, MR24, PKGA22, RSU^{+}24]. queueing [LATV22, LHF22]. quick [LL21]. QuixBugs [YMDM21]. quo [FKGN23].

R [MCSAGB20, Vid22]. race [SPZW25]. Raising [ZGLPA24]. Random [AHT^{+}21, ZWY^{+}22, BH20, CCZ^{+}24]. range [ABMV24]. Rank [MLPC20]. ranked [Hor21]. Ranking [CFNA25]. rapid [LBMF^{+}22, ZHS^{+}24]. rate [MN23]. rater [BSS^{+}25, DPGGP23]. ratio [MAK^{+}21]. Rational [SWH^{+}20]. rationale [AES22]. RE [IMTS23]. reaching [MC20]. react [FBV24]. react-based [FBV24]. reactions [BCF^{+}22]. reactive [ATWS25, BBG^{+}25]. readability [MHOM22]. readme [WWC23]. Real [CRX24, HIDT21, KH24, MCMA21, MLBD21, NQT25, RPM^{+}22, WCH^{+}23, YAaO^{+}21]. real-life [NQT25]. Real-Time [CRX24, HIDT21, MCMA21, RPM^{+}22, YAaO^{+}21]. real-world [KH24, MLBD21]. reality [ZLC^{+}23]. realizability [Bro24]. really [JS22]. reasoning [AHP21, GDK24]. reassurance [KRC23]. Recognition [VBMB20, WM24, ZLS20]. recognizing [CZLN22]. Recommendation [MBP20, LXYL20, NMT^{+}23, QWHH23, WZZ21, WGL^{+}22, WZL^{+}25, YYW^{+}24]. recommending [JZM21, KL20, NDDD20]. Reconstruction [AKL25]. recorded [SSS22]. records [Spa24]. recovering [ZXZS24]. Recovery [FMF^{+}24, MLPC20, WHK^{+}23, WDL^{+}25]. Reducing [QWG22]. reduction [ACPM22, AdlBGZ^{+}23, BDS^{+}24, CWGS23, CBZZ24, SSZ20, ZWF^{+}25, LWL20]. ReenRepair [HSW^{+}24]. Reentrancy [OHM25, HSW^{+}24]. Refactoring [AMO21, FBV24, GIP^{+}24, HMR21, HPF23, IZAD21, LPPG20, MRS20, MSB23, PZDG21, PÁngeles MoragaGC25, SAZN22, Vog20, YLC25]. refactorings [OAH^{+}23, PMDN20, PSAB22]. Reference [RA25, GMFO^{+}21a, GMFO^{+}21b]. refinement [DSO20, LLW^{+}24]. refinements [HXJ^{+}20]. Refining [NM25]. reformulation [ZLSY23]. refusals [LHN20]. regional [ZLG25]. Regression [HZT^{+}20, CQZ^{+}20, DWH23, KH24, LLNC21, MAP^{+}20, MBP23, MN21b, SGV22]. regulatory [KMS^{+}25]. reinforcement [DZY^{+}23, NWZ25, TCTL25, WLL^{+}24, YLHZ20, ZHLR23]. rejuvenation [CRX24, CDN^{+}22]. related [ANC^{+}23, FGRF24, GMCA21, KOPN22, KRC23, MLBD21, ZXH^{+}24, dDLSK23]. Relation [CQZ^{+}20, DCÁ^{+}24, KVP23, LGKT22, VM24b, ZLX^{+}22]. relation-aware [ZLX^{+}22]. Relation-based [CQZ^{+}20]. relational [TGG21, LFFW23]. relations [BGE^{+}21, CKY25, LHN20, XTF^{+}21].

Relationship

[THN20, AA23, GFS21, GBLP25]. Release [THN20, dSIdÁ23, KCMD21]. relevance [VML21]. relevant [WZFD25]. Reliability [DPGGP23, GPPDLF23, ZXW20, AR24, HCC22, LDH22, LHF22, NJF20, TTL20, WCZL25, WLdCM22]. Reliable [CLLJ24, KRSW22, HOCK22, ZHLR23]. reliance [MHJW22]. Remodularisation [TCA22]. Remote [GS20]. removal [IZAD21]. removing [LDT22]. rename [PMDN20]. reorganization [ZPSW24]. repair [ABMV24, BF22, ETY^{+}22, GSF24, HSW^{+}24, IIK21, KKE21, LLK^{+}21, NZA^{+}25, WL24, XXC^{+}24, YMDM21]. repairs [JZW^{+}21, KKE21]. REPD [ASKS20]. replacement [SM20a]. replicated [VM24b]. replication [BEAK21, HYF^{+}24, MH20]. RepliComment [BSGN21]. report [BB22, FAA22, GKB^{+}21, JST^{+}23, LYS^{+}23, MBKL24, PKS25, SMM23, WLCC25, WZC^{+}20, YLC25, ZHM^{+}23, ZLSY23]. reported [MR24]. reports [AXUO24, JCNS^{+}22, LFH^{+}22, SHB21, WGMT25]. repositories [AKMS23, TKSC20, WCH^{+}22]. repository [ROGS^{+}24]. representation [JSTW22, JSG25, MRS25, YWS23, ZCLP21, vdG25]. representations [BAV20, LB24, SMVC24, LZ25]. representative [MN23]. represented [dSS^{+}22]. Representing [KSF20]. Reproducibility [MR24, HTN^{+}25]. reproduction [WGMT24]. reprogramming [TQT^{+}25]. Request [FZT^{+}22, LNST21, STW23]. requests [FFSB23]. required [VCB24]. requirement [BRO^{+}22, DDCC24, MKB25, RPR22]. Requirements [AGPR20, FMF^{+}24, KKH^{+}21, MLPC20, BFM^{+}25, BMHR21, BN23, DBB20, FFV^{+}23, GBMF22, KSG^{+}22, KL20, LGT^{+}23, LC20, LK23, LCS^{+}24, MKB23, SJH^{+}25, SCdPL24, SRM^{+}25, ST24, WKP20, YRC^{+}25, dSIdÁ23, GCLB22, MFLS22, FGS23]. Research [FBMR20, GVBC^{+}24, LHH^{+}24, ACM25, A– DHM23, AF22, APB20, AGP22, ECMC20, FKGN23, FMF^{+}24, GW23, GPPDLF23, HATG21, KI23, MMC24, MMB22, NEJ24, POZ20, SGM23, SFR23, UPP^{+}22, NLTM23, Pie20]. researcher [Fei23, RFS^{+}21]. researchers [Woo25]. researches [NdSR^{+}21]. ReSIde [SSS^{+}20]. residence [XZY^{+}20]. residual [CFF^{+}23]. resilience [PLL^{+}23]. Resistant [THB22]. resolutions [AMS23]. resolved [GS21a, GS21b]. resolving [CPC^{+}23]. Resource [AMB^{+}25, ATWS25, AZB25, AZR^{+}21, BF22, GTT^{+}20, Pie20, TGX^{+}22, TBD^{+}23, ZHLR23]. resource-driven [AZB25]. resource-efficient [ATWS25]. Resources [AZB25, MFF^{+}24, OWGS23, ZGW^{+}20a, ZPL21]. response [MFBP20, RK20]. responses [TRD^{+}23]. REST [POWGH22]. restoration [XBS21]. result [PPB20]. Results [FKF^{+}23, NRFS25, ZH21, AHFP24, VM24b, BCF^{+}22, RMT^{+}22]. retainment [BCF^{+}21]. RETORCH* [AMB^{+}25]. retrieval [GDK24, MAP^{+}20, SH20, SRD^{+}21, YWCX24]. retrieval-based [YWCX24]. retrieve [GBSO20]. Retriever [KGS^{+}25]. reusability [Pat20]. Reusable [SSS^{+}20]. Reuse [AX21, AAW20, BCF23b, CXH^{+}24, DBB20, FAG^{+}20, GFS21, GW24, MWY^{+}22, NAV^{+}24, SWZ^{+}20, SBF20, WMA^{+}24]. reusing [HGHK24]. reveal [HBSV^{+}22]. Revealing [UPP^{+}22, ZWP^{+}24]. Reverse [FBMR20, KGS^{+}25, LSB^{+}22]. reversible [Spa24]. Review [MYH^{+}24, PFdMF21, SJH^{+}24, SBC20, VKG^{+}23, VBMB20, AA23, ACM25, AÇCT21, AWMW20, BRS^{+}22, BAV20, dACOdS23, DAP20, DN21, DL22, DL21, FRC24, FFSB23, GGB^{+}22, GB20, GGP^{+}24, GPF22, GdOB25, HOAM23, HTN^{+}25, HH22, HJK^{+}21, HO22, HM24, ISW25, Jør24, KAW^{+}23, LPPG20, LMR^{+}23, LLL^{+}22, LHG^{+}25, LBT^{+}21, LZB^{+}23, LLK^{+}21, LLZ^{+}23, MMC24, MH24, MRA^{+}22a, Mec24, MMB22, MFLS22, MLJ23, MRM^{+}21, OBCR23, OSM^{+}23, PGW^{+}23, PAM^{+}21, PSAB22, QWGS24, RBS^{+}23, RFB^{+}22, SL20, SHG24, TRD^{+}23, Tai24, VDVC21, VIL24, WFR21, WCL23, YSB^{+}21, ZNPR^{+}23, ZZK^{+}24, ZCL24, ZH22, dSFdC^{+}25, dSNJ25, WUK^{+}21]. Reviewer [VBMB20]. reviews [ANC^{+}23, GCF22, MWFK20, NdOdO^{+}22, PDGMT20, TST24, WLCC25, ZZH^{+}25]. Revisiting [ZLC^{+}23]. revolutions [ŠMKGH23]. reward [YLHZ20]. rewriting [SK22a]. rich [DSB23, SS23]. right [DBO24]. risk [GAS25, LWZ21, TTL20, XCH^{+}20]. risks [BCLN21]. RNN [GF24]. road [GBT^{+}20]. roadmap [MdOT23, POZ20]. roadmapping [BPJ^{+}22]. robot [LCY23, A– DHM23]. Robotic [Gla23, MDWS^{+}23, YYT25]. robotics [AGM^{+}25, MLS^{+}21]. robots [RCA^{+}22]. Robust [RFB20, FWZZ25, LCC^{+}23a, RPR22]. robustness [DFT24, GGPR24, LWSZ23]. Role [HQNR^{+}22, GAS25, KdJPK^{+}23, LSB^{+}22, LWC20, OWGS23, PDDD20, SHA21]. role-based [LSB^{+}22]. roles [CBDK23, FSS23, OZVRD21]. roll [KBW^{+}24]. roll-out [KBW^{+}24]. rooms [SHB21]. Root [ZWL^{+}24, BSH^{+}20, FGRF24, LCC^{+}23a, XCZ23]. router [SH20]. Roxygen [Vid22]. RPerf [WLCC25]. RRGcode [GDWK24]. RSFIN [LBH^{+}24]. rule [AR24, CHLT23, RF24a, SV25, LBH^{+}24]. rule- [RF24a]. rule-based [AR24, SV25]. rules [AKL25, HPF23]. Run [CSLN23, SMB^{+}20, ACG^{+}21, ZGW^{+}20a]. Run-time [CSLN23, SMB^{+}20, ZGW^{+}20a]. Runtime [CWTL21, IPB23, JJS24, NQT25, WCZL25, ALZ^{+}20, ADS^{+}22, BRS23, GGB^{+}22, Hei20, MN21a, MN23, RMM24, VGS^{+}23]. runtimes [KBP24].

Safe [Gla23, JS22]. Safety [CID25, GCLB22, WFM^{+}24, APR24, BLHS23, BRS23, CDP22, CDO24, DL21, GdOB25, KRS^{+}21a, KRS^{+}21b, KBBD23, LZ25, MSKS24, MKB25, PGW^{+}23, RPS25]. Safety-critical [WFM^{+}24, APR24, CDO24, KBBD23, MSKS24]. SAMM [FAFJ24]. sample [LXLZ20]. samples [MCH22, MN23]. sampling [MN23, THB22, ZLW^{+}23]. sampling-based [THB22]. sandbox [dCMM^{+}22]. SAP [NQT25]. Satisfaction [KMAB20, GAS25, LGKT22]. satisfiability [LCW^{+}24]. satisficement [BFHC20]. Saudi [AAZB23, AF22]. SBFL [TPV25]. SC4R [TCA22]. Scalability [AFJ^{+}20, ABT^{+}22, BMHR21, HH24]. scalable [PKB24]. Scale [TPGH20, Bat20, BSDB20, CSLN23, DWH23, DFC^{+}23, IBP21, JYZ^{+}24, KKH^{+}21, LST20, MSMB21, PSGD23, PN21, SJC^{+}22, UPP^{+}22, WZC^{+}20]. scaled [ŠMF^{+}23]. Scaling [PPL^{+}25, ATWS25, BBG^{+}25, BCLN21]. scarcity [MPP24]. SCC [ADG^{+}20]. scenario [GBLP25, MMDL23]. scenarios [SH20]. SCGRU [ZYZ^{+}23]. schedulability [ZGW^{+}20a]. scheduling [CRX24, EBAR21, HTB21, ISKB20, LLWL22, MCMA21, MdSKD22, YZS22, ZZTC23, ZHLR23, ZPL21]. scheduling-driven [EBAR21]. schema [OSJB21]. scheme [CND22, DTZ^{+}22]. scholars [WMAL21]. school [SGW^{+}22]. science [HM25, OKG25, ROGS^{+}24]. scientific [AACC21, EKB^{+}23, HGHK24, HTN^{+}25, LLWL22]. Scoping [MDAHS25, DSM20, MRA^{+}22a]. score [SRD^{+}21]. scores [RSVW23]. scoring [BAH24]. Scratch [AC25]. screencasts [MRA22b]. scriptless [RSM^{+}23]. scripts [CATA21, OLR^{+}25]. Scrum [CR23, HO22, KdJPK^{+}23, SWH^{+}20]. SDN [HTB21]. SDN-NFV-enabled [HTB21]. SE [KGM24]. Seamless [RVAC^{+}25]. Search [LBH^{+}24, BFHC20, CJZ^{+}20, CZW20, DOG^{+}23, DXL^{+}24, EFPC21, ETY^{+}22, GSM^{+}23, GDWK24, HGH^{+}23, HPZ^{+}20, HLZ^{+}23, KA24, LCW^{+}24, SMKI23, WGY^{+}20, YXLH24, ZWY^{+}22, ZPSW24]. Search-based [LBH^{+}24, BFHC20, GDWK24]. seas [MGSC22]. SeCNN [LWP^{+}21]. Secondary [FdSN^{+}20, ZLS23, SGW^{+}22]. Section [AMM25]. sector [AF22, BHS24, Mam23]. Secure [Gla23, LCT22, AZ21]. Security [AAG21, BGM^{+}21, ACC^{+}20, ANC^{+}23, AAZB23, AZ21, ARPDP24, BLHS23, CDRV20, CFI^{+}24, FAFJ24, GGP21, GB20, GFS21, GdOB25, HCB^{+}20, JRM^{+}22, KL20, LC20, MREVEA^{+}22, MICV23, MSKS24, NJF20, NSL^{+}21, NSR^{+}23, PGW^{+}23, PSAB22, RCT22, RSVW23, RHJ25, SCD^{+}24, SS23, TC22, TST^{+}21a, TST^{+}21b, VCB24, WLLJ24, WZC^{+}20, YU22, ZCZL23, CDRV20]. Security-by-Design [CDRV20]. security-related [ANC^{+}23]. security-rich [SS23]. seed [LHY^{+}25]. SEET [AZKR20]. select [LMZT22]. selected [MAS23]. selecting [LKP^{+}21]. selection [ASSH22, dACO21, CKM^{+}20, DWH23, JS22, JRM^{+}22, KSG^{+}22, KGL^{+}22, LWYW20, MAP^{+}20, MAV24, PDGMT20, SGV22, SRS^{+}25, ZYZZ21]. selective [KDB^{+}21, SK22a]. Selenium [GKAHMO22]. Selenium-Jupiter [GKAHMO22]. Self [AMO21, GMS22, PtBD^{+}25, YCWD23, AAW20, AGM^{+}25, AMRS20, APAF21, BFHC20, HLZ^{+}23, HHZW23, IZAD21, PGW^{+}23, QWG22, RATS21, WCGS23, YZL^{+}22, ZYZ^{+}23]. self-adaptation [AGM^{+}25, APAF21]. Self-Adaptive [PtBD^{+}25, GMS22, AAW20, AMRS20, BFHC20, PGW^{+}23, QWG22, RATS21, WCGS23]. self-admitted [IZAD21, YZL^{+}22, ZYZ^{+}23]. Self-Affirmed [AMO21]. self-attention [HLZ^{+}23]. Self-supervised [YCWD23, HHZW23]. Semantic [LAL^{+}23, RLK^{+}24, CCC^{+}24, CZLN22, CXP^{+}23, CFF^{+}23, DAP20, DXL^{+}24, HSW^{+}24, JSTW22, LWP^{+}21, MHOM22, OZVRD21, RPR22, DBM^{+}24, WDXX23, YCWD23, ZXH^{+}24]. semantic-aware [JSTW22]. Semantics [ZYL^{+}22, CJZ^{+}20, NAD^{+}20, PSGD23]. semi [YLG^{+}25, ZXZS24]. semi-supervised [YLG^{+}25, ZXZS24]. semisupervised [CBW^{+}23]. sensitive [AdlBGZ^{+}23, GKPS24, HLL21, MdSKD22]. sensitivity [LABJ23]. sensor [SCPB25]. Sensorina [SRM^{+}25]. Sentiment [SZ23, HOCK22]. separate [FG24]. separation [ZX23]. September [Ano20x, Ano21y, Ano21z, Ano22x, Ano23x, Ano24v]. Sequence [JH20, CWTL21, LL25, TTB^{+}23]. Sequences [SVV24, CDM25, LL23]. series [MDAHS25]. serious [ZGLPA24]. server [BCF23a]. serverless [ECL^{+}22]. Service [RFB^{+}22, SL20, SSS21a, SSS21b, SBC20, ASM^{+}21, AM23, BS23, BSH^{+}20, CZW20, GMS22, LCC^{+}23a, LTZ^{+}21, NF24, NAV^{+}24, RK20, RFB20, dORGCG23, SSS^{+}20, WZZ21, WHK^{+}23, WGY^{+}20, YYW^{+}24, YSB^{+}21, ZLX^{+}22, ZPL21]. Service-Based [SSS21a, SSS21b, dORGCG23, WGY^{+}20]. service-oriented [BSH^{+}20]. ServiceAnomaly [PHLHM24]. services [EM20, EN23, LLW^{+}23, ZGHG^{+}23]. session [HDX^{+}23]. session-based [HDX^{+}23]. sessions [EKHJ^{+}20]. Set [WYG^{+}24, EM24, RO22, XCH^{+}20, AHT^{+}21]. sets [HOCK22]. setting [Tok22, YLG^{+}25]. seven [GGP21]. severity [AXUO24, MCM^{+}24, RPC^{+}24, TXW^{+}20]. SEXTAMT [MGSC22]. SGT [ZXH^{+}24]. SGX [Ala21]. Shape [QWHH23, ZLG25]. Shape-aware [QWHH23]. shared [LWC20, SV20]. Sharing [GGB^{+}22, LYZ^{+}22]. Shedding [Mec24]. ship [ZZTC23]. short [SGG24]. shot [LYG^{+}24, SPZW25]. should [EBSB23]. shrinkage [WDO24]. side [Mec24]. Signal [BJB^{+}21]. Signal-Based [BJB^{+}21]. signatures [AvdBG^{+}24, YML^{+}22]. significant [LST20]. similar [DDPP23, GGB^{+}22, MWY^{+}22, WGL^{+}22]. similar-app [GGB^{+}22]. Similarity [AWMW20, CFNA25, GJW^{+}22, NRRS20, RPR22, WGY^{+}20, ZNPR^{+}23]. Similarity-based [AWMW20, CFNA25]. SimNet [GJW^{+}22]. Simple [GSM^{+}23]. simpler [TAF^{+}20]. simplification [YWCX24]. Simulated [MSC20]. Simulation [SBC20, GTT^{+}20, HDX^{+}23, HTG^{+}24, RK20]. simulations [PPL^{+}25]. simulator [MPGB22, MSC20]. sine [ZHLR23]. single [AM23, MC20, OAH^{+}23]. single-tenant [AM23]. sites [HPZ^{+}20, LXYL20, PKGA22]. situ [VBMB20]. situation [TAF^{+}20]. situation-aware [TAF^{+}20]. six [BCF^{+}21, LPS^{+}23]. six-month [BCF^{+}21]. Size [AHT^{+}21]. sized [dSNJ25]. skills [HNM^{+}24, LGKT22]. skip [JS22]. SLA [HTB21]. SLA-aware [HTB21]. Slack [SM20b]. SLAs [CDRV20]. slice [WL24]. sliced [CLZ^{+}23]. slices [SBD23]. slicing [AS22, DLCH25, GKPS24, Ger25, LBF^{+}21a, LBF^{+}21b, SK22b, SB23, WTS23]. slow [vDDPZ24]. small [APB20, AHT^{+}21, LST20, ACA^{+}23]. Smart [CFF^{+}23, OHM25, ACG^{+}21, BHP^{+}21, CLZ^{+}23, CLZ^{+}24, DLV^{+}22, GM20, HCB^{+}20, HSW^{+}24, LFFW23, PBC^{+}23, RLK^{+}24, SCD^{+}24, VDVC21, VIL24, VCT20, WCH^{+}23, YML^{+}22, YWLZ23, dSFdC^{+}25]. smart-contracts [PBC^{+}23]. SmartCLIDE [NAV^{+}24]. Smell [ZZK^{+}24, HYF^{+}24, PDDD20, PMM^{+}25, SAZN22, SELS21, TSLHS21]. smell-aware [TSLHS21]. Smells [PSAB22, CAA^{+}23, GSK24, HMR21, HBN^{+}25, JKD25, LPPG20, MSB21, PT21, PFW21, PÁngeles MoragaGC25, RB23, TST24, WCL23]. SMEs [GGP21]. Snapshot [CRV23]. snippets [ADG^{+}20]. SOA [CPD20]. SOA-based [CPD20]. social [CID25, GCLB22, HATG21, MVGHPT22, YRC^{+}25]. society [VCT20, BNMW23]. socio [PT21]. socio-technical [PT21]. Software [AGM^{+}25, A– DHM23, AX21, AACC21, APAF21, BCF23a, BDIL25, CF24, CDN^{+}22, DJR^{+}22, DAML23, DPD^{+}22, DMAH24, EM24, FRP^{+}23, GRLA20, GFS21, GBLP25, KRD^{+}23, KAW^{+}23, LCP^{+}24, LOBD25, MMO22, MRA^{+}22a, MN23, NRFS25, NBGC20a, OWGS23, PFdMF21, PtBD^{+}25, RBvB^{+}24, SVV24, SvdBHV24, SVAGB20, SMdS^{+}23, TCA22, Tok22, VML21, ZFGH^{+}22, ZYZZ21, AAW20, ASM^{+}21, AGPR20, AWA^{+}22, AKBN20, AMS23, AZ21, APB20, AMM25, AGP22, AWMW20, ABT^{+}22, AR24, ARPDP24, BS23, BMLH25, BCLN21, BRO^{+}22, BCJI22, BGMB20, BL21a, BL21b, BTSC^{+}23, BFZC21, BbASP23, BSDB20, Bro24, BCW21, BGL^{+}20, BGL^{+}22, CCC^{+}24, CDET21, CBW^{+}23, CBZZ24, CAC20, CA20, CRX24, CKP20, CKM^{+}20, CLC^{+}23, CXP^{+}23, CXH^{+}24, CGM24, CTD^{+}23, CJNDZ21, CDLN21, DdM^{+}24, DAP20, DSB23, DPGGP23, DLCH25, EFPC21, ECMC20, ET21, ETD^{+}24, EKHJ^{+}20, EKB^{+}23, FMG^{+}25]. software [FCW^{+}23, FJvdW20, FKGN23, FVDF21a, FVDF21b, FEBO22, FSS23, FAFJ24, GW23, GMFO^{+}21a, GMFO^{+}21b, Ger25, Gir21, GBK^{+}23, GSK24, GPPDLF23, GVBC^{+}24, GGMH23, GAS25, GFCLB^{+}24, HTN^{+}25, HNM^{+}24, HOCK22, HPP24, HQNR^{+}22, HPF23, HBK25, HPZ^{+}20, HCC22, Hyr25, ILUN21, IBP21, IPB23, IT23, JCNS^{+}22, JYZ^{+}24, JSG25, JKD25, JH20, Jør24, KSF20, KDB^{+}21, KGS^{+}25, KR23, KHEC^{+}23, KGS23, KMAB20, KBBD23, LMGK22, LWC20, LFLS24, LXYL20, LDH22, LMZT22, LQY^{+}22, LBH^{+}24, LBHW25, LL25, LGKT22, LHF22, LWSZ23, LBCG23, LAL^{+}23, LHH^{+}24, LZL^{+}25, LID^{+}22, LTG^{+}25, LMVRA^{+}20, LBMF^{+}22, MSMB21, MVGHPT22, MS25, MLS^{+}21, MAK23, MMDL23, MGSC22, MBKL24, Mec24, MSS21, MMB22, MTA24, MFLS22, MWFK20, MN21a, MLJ23, MSRR22, MHR^{+}25, MLGBRC24, MCDP24, MRA22b, MCRM25, MdOT23, MHJW23, MSDC25, NRRS20, NJF20, NdSR^{+}21, NM25]. software [NAZ22, NF24, NDDD20, NZA^{+}25, NEJ24, OBCR23, OKG25, OMA^{+}22, DDPT20, PLL^{+}23, PKGA22, PPMC22, Pat20, PC23, PFC^{+}23, PAM^{+}21, PDGMT20, POZ20, PNND25, PdS23, PGR24, PPM^{+}21, PKS25, dPPdJFF21, RV22, RPR22, RMOGA20, RCT22, RJ23, RHJ25, SGM23, SC22, SS20, SHA21, SSS^{+}20, SM24, SRS^{+}25, SGG22, SGG24, STM^{+}22, SV20, SV25, SM20b, SYB^{+}23, SMVC24, SZ23, TH24, TC22, VCB24, VKLM21, VM24a, Vog20, WDT^{+}24, WL24, WFR21, WMAL21, WXZL23, WDO24, YZM^{+}25, YAaO^{+}21, YDP22, ZGHG^{+}23, ZKDP22, ZZTC23, ZCZL23, ZSCD22, ZLS20, ZLS22, ZLS23, tBS23, KZTS22, NLTM23, RFS^{+}21]. software-architecture [GSK24]. Software-Intensive [KRD^{+}23, BGL^{+}22, VKLM21, ZGHG^{+}23]. Software-testing [GRLA20]. SOLID [YLC25]. solidity [DLV^{+}22, WCH^{+}23]. solution [BPJ^{+}22]. Solutions [SVAGB20, BHS24, CDET21, HGHK24, HWL^{+}25, MBO^{+}22, PKS25, SRD^{+}21, ZLZ^{+}25]. solve [dSIdÁ23]. solving [AWA^{+}22, CHLT23]. Some [LST20]. SonarCloud [NRFS25]. SonarQube [LST20]. Source [EHB21, JVZ25, SVAGB20, ZYL^{+}22, AA23, AZR^{+}21, AKMS23, BGL^{+}20, BGL^{+}22, ECMC20, ES23, Fei23, FRC24, GIP^{+}24, HS21, HELW20, HHJ^{+}20, HHZW23, JCNS^{+}22, LMZT22, LAL^{+}23, MCM^{+}24, NRRS20, NMV25, NAD^{+}20, NF24, NDD^{+}24, OEW22, OMA^{+}22, PKGA22, SKG^{+}24, aSLF^{+}22, SMVC24, TTB^{+}23, THG20, TKSC20, WMLM22, WZFD25, ZNPR^{+}23, ZKDP22, ZSZ^{+}22, dlVRB21, ASKS20]. source-code [AA23]. sourced [OSJB21]. Sourcing [LCAC21]. Space [SBC20, ES24, CBZZ24, LTJ^{+}20]. spacecraft [GBT^{+}20]. spaces [BPJ^{+}22, CWGS23, ETY^{+}22, PAM^{+}21, QWG22]. span [BCF^{+}21]. Spanish [GGP21]. spare [ZXW20]. spares [CRX24]. Spark [CYW21, ISKB20, WCZW22]. spatial [ACD^{+}21]. speaker [WM24]. Special [AX21, AMM25, BCW21, LSSZ21, BDLR23, FBMR20, HS21, EHB21, KRD^{+}23]. specific [ASCR23, ACA^{+}23, BPJ^{+}22, CVC21, KSG^{+}22, KBG^{+}24, VWSCH23, WCL23, YLGS24, ZLS20]. specific-model [CVC21]. Specification [AAG21, YRC^{+}25, DL22, LM21, Liu21, PBV24, RHJ25, SCN^{+}21, TAT^{+}23, WYFZ24]. specifications [DDCC24, KRLS^{+}20, LB24, LGC^{+}25, LCS^{+}24, MKB25, PSZ21, RO22, SLL20]. Spectrum [MMSM^{+}23, RF24a, HBSV^{+}22]. Spectrum-based [MMSM^{+}23, RF24a]. speech [LCY23]. speech-enabled [LCY23]. SPELLing [PCC^{+}20]. SPL [MDAHS25, RVAC^{+}25]. SPL-DB-Sync [RVAC^{+}25]. split [CHLT23]. SpongeBugs [MFBP20]. Spotify [ŠMF^{+}23]. Spreadsheet [MHJW22, HXJ^{+}20, MHJW23]. spreadsheets [HJK^{+}21]. Spring [MCH22]. SQL [BMB20, Tai20, TS22]. SQL/NoSQL [BMB20]. srcClone [AS22]. SSL [WXL^{+}20]. SSL/TLS [WXL^{+}20]. Stability [YWY^{+}21, CKY25, ECL^{+}22, YLW^{+}21]. Stack [ASCR23, LZJ20, SZ23, ZLG25, AA24, ADG^{+}20, BLTX21, CKP20, CXH^{+}24, GMCA21, MR24, PKGA22, RSU^{+}24, STW23, SZ23, TXW^{+}20, ZLW^{+}23, ZLZ^{+}25, ZLS22, ZLS25, dDLSK23]. stack-augmented [LZJ20]. StackOverflow [FG24]. StaDART [ACC^{+}20]. stage [YXLH24]. stakeholders [BEM^{+}23, FAFJ24]. standard [WGL^{+}22]. standards [CPD20]. standby [SH20]. start [ZH22]. start-ups [ZH22]. startup [MBKL24]. startups [BBND^{+}20, ZH21, GBLP25]. State [BLHS23, Ibi22, LHS^{+}24, AGP22, GF24, JJS24, KBP24, PSZ21, PTW22, RSM^{+}23, SFR23, UPP^{+}22, Gir21]. State-of-the-art [BLHS23]. stateful [VSTK21]. statement [FYH^{+}24]. statements [MFPB24]. states [OBCR23]. Static [ZX23, AvdBG^{+}24, LPS^{+}23, MFBP20, MCM^{+}24, MC20, PBC^{+}23, SVVD21, SBD23, WXL^{+}20, ZLW^{+}24, ZLD^{+}25, dCMM^{+}22]. statistic [PDGMT20]. statistical [CBC^{+}24, MVGHPT22, PPL^{+}25]. Status [FKGN23, IT23, LHH^{+}24]. stay [SMH^{+}23]. step [Liu21, Woo25, MSC20]. STEP-ONE [MSC20]. stereotypes [DPRD21, HQNR^{+}22]. STILE [OLR^{+}25]. still [THG20]. stochastic [CLLJ24, YZS22]. storage [AON^{+}24, EM20]. store [NdOdO^{+}22]. stories [AWP^{+}25, APB20]. story [BDMP21]. strategic [DBO24, MRVS24]. Strategies [LBT^{+}21, LGMP24, EFPC21, FVDF21a, FVDF21b, RPC^{+}24, SRS^{+}25, ZWY^{+}22]. strategizing [BSDB20]. strategy [DSO20, MH20, SMKI23, THB22]. Stream [VKG^{+}23, HH24]. Streaming [KR23]. strength [DYZ^{+}23]. strengthen [ZLW^{+}24]. strong [Mas22]. structural [CAC20, MWD24, MHOM22, MSDC25]. Structure [ZYL^{+}22, GVBC^{+}24, HPZ^{+}20, LBHW25, NAD^{+}20, PSZ21, SGM23]. structure-aware [LBHW25]. stubborn [WYG^{+}24]. Student [SWG^{+}20, CND22, KDB^{+}21, KHEC^{+}23, LGKT22, PKS25, YU22]. student-driven [PKS25]. studies [BAV20, DPGGP23, FdSN^{+}20, FA24, LH24b, TST^{+}21a, TST^{+}21b, WUK^{+}21, ZMLZ23, ZLS23].

Study [DJR^{+}22, LGMP24, VBMB20, WWC23, WLS20, ACC^{+}25, AIH20, AC25, AES22, AA24, A– DHM23, AAZB23, AF22, APB20, AACC21, AGP22, BMLH25, BHS24, BEAK21, BCLN21, BRO^{+}22, BBND^{+}20, BM23, BTSC^{+}23, BMHR21, BGL^{+}20, CAC20, CA20, CAA^{+}23, CKP20, CXH^{+}24, CDO24, CID25, CR23, DPA^{+}24, DSM20, ECL^{+}22, EKK^{+}24, ETD^{+}24, FPB^{+}24, FEBO22, FFV^{+}23, FKF^{+}23, FSS23, GGP21, GMFO^{+}21a, GMFO^{+}21b, GMCA21, GSK24, GW24, HG22, HATG21, HLS24, HYF^{+}24, HWL^{+}25, IZAD21, IMTS23, JST^{+}23, KCMD21, KI23, KOPN22, KAA^{+}21, LST20, LNST21, LABJ23, LWYW20, LFH^{+}22, LMZT22, LQY^{+}22, LMZ^{+}23, LCY^{+}24, LDT^{+}25, LSF^{+}25, LXL^{+}23, MCZX20, MAP^{+}20, MAK23, MCM^{+}24, MAS23, MSB21, NdOdO^{+}22, NSR^{+}23, OZVRD21, OMA^{+}22, PLL^{+}23, PC23, PTW22, PDGMT20, PSGD23, POZ20, PNND25, PFW21, PPM^{+}21, PÁngeles MoragaGC25, PN21, RSU^{+}24, RA25, RATS21]. study [SSO^{+}23, SH20, STW23, SMB^{+}20, SB23, SM20b, SMVC24, TSLHS21, THG20, TKSC20, UPP^{+}22, VM24b, VHB21, WUK^{+}21, WCZW22, WMLM22, WCH^{+}23, WLdCM22, WR22, WMA^{+}24, XLY^{+}21, YLHZ20, YOH^{+}23, YWTA25, YMDM21, YU22, ZWY^{+}22, ZGW^{+}20b, ZRGJ21, ZLZ^{+}25, ZZP21, ZH21, ZLS22, ZLS23, dTMS21, dSNJ25, vBD21, vRMG23]. Studying [BCF^{+}21, CPC^{+}23, THN20]. stupid [GSM^{+}23]. sub [CMP^{+}20]. sub-processes [CMP^{+}20]. subjectivity [HOCK22]. subjects [MPP24]. Substructure [WGY^{+}20]. success [BDMP21, CR23, ILUN21, Jør24, RMOGA20, SJH^{+}24, SS20]. successful [WSL^{+}20]. suggestion [HHZW23, JWZM20]. suggestions [MFBP20]. suite [Ibi22, ÖS20, RF24b]. suites [IIK21]. summaries [AIH20]. summarization [FCW^{+}23, GJW^{+}22, YWS23, YLGS24, ZSZ^{+}22]. Summer [SWG^{+}20]. SuMo [BMP22]. superoptimized [ACG^{+}25]. supervised [HHZW23, YLG^{+}25, YCWD23, ZXZS24]. Support [GBLP25, ETD^{+}24, FvB24, GAC20, MFPB24, MBP23, MBG^{+}25, RV22, TAT^{+}23, WNZ25]. Supporting [CAF22, ZHS^{+}24, LYZ^{+}22, NDDD20]. Survey [LH24a, BCF^{+}22, BLHS23, BGC20, DAML23, FdSN^{+}20, IPB23, KOPN22, LFLS24, MRS20, MSRR22, MRA22b, SLL20, SKG^{+}24, SCB22, VML21]. surveys [NSL^{+}21, RMT^{+}22]. survivability [BGM^{+}21]. survival [RB23]. suspicious [SV20]. suspiciousness [LWL^{+}23, LLW^{+}24]. Sustainability [HNM^{+}24]. sustainable [RATS21, TGX^{+}22, BNMW23]. sustainable-development [RATS21]. Sustaining [LGT^{+}23, PKS25]. SUT [HDX^{+}23]. swapping [SSP21]. swarm [WMH^{+}24]. SWFC [AHT^{+}21]. SWFC-ART [AHT^{+}21]. switching [RSVW23, SM20a]. Symbolic [MLYL25, YZS^{+}25, CHLT23, KBBD23, AZKR20]. symmetric [CKY25]. symptoms [JZW^{+}21]. Sync [RVAC^{+}25]. synchronous [ZZY^{+}24]. Syntax [Ger25, BN23]. Syntax-preserving [Ger25]. synthesis [ACG^{+}25, GDK24]. Synthesizing [NdSR^{+}21]. SynthoMinds [GDK24]. SYNTONY [WMH^{+}24]. System [AAG21, GGP21, CXY^{+}23, DLCH25, HSJB23, KKH^{+}21, KKL^{+}21, LLST20, LWSZ23, LTG^{+}25, MBKL24, NMG^{+}20, NEJ24, PZVN^{+}25, RA25, SKP20, SH20, SB23, TSP20, Tai24, WNZ25, WFR21, YYT25, ZZTC23, A– DHM23, RATS21]. system-level [NMG^{+}20]. system-of-systems [HSJB23]. Systematic [BAV20, DJR^{+}22, DL22, MRM^{+}21, PFdMF21, PDGMT20, WLS20, AA23, AAW20, ACM25, AÇCT21, A– DHM23, AACC21, AGP22, AWMW20, BRS^{+}22, BM23, dACOdS23, CJNDZ21, DAP20, DAML23, DN21, ETD^{+}24, FdSN^{+}20, GMFO^{+}21a, GMFO^{+}21b, GRLA20, GB20, GGP^{+}24, GPF22, GdOB25, HTN^{+}25, HATG21, HH22, HJK^{+}21, HO22, HWL^{+}25, ISW25, Jør24, KAW^{+}23, KI23, KAA^{+}21, LPPG20, LMR^{+}23, LHG^{+}25, LBT^{+}21, LMZ^{+}23, LLZ^{+}23, LBMF^{+}22, MMC24, MH24, MRA^{+}22a, MAK23, MGSC22, Mec24, MMB22, MFLS22, MWFK20, MLJ23, MSB21, NdOdO^{+}22, OBCR23, OSM^{+}23, PPMC22, PC23, PTW22, PGW^{+}23, PAM^{+}21, PNND25, dPPdJFF21, PÁngeles MoragaGC25, QWGS24, RBS^{+}23, RFB^{+}22, RATS21, SHG24, Tai24, UPP^{+}22, VDVC21, VIL24, WUK^{+}21, WFR21, WLdCM22, WCL23, YLHZ20, ZNPR^{+}23, ZZK^{+}24, ZH22, dSFdC^{+}25]. Systems [AX21, FMG^{+}25, Gla23, NBGC20a, PtBD^{+}25, SVAGB20, ZWL^{+}24, tBS23, vDDPZ24, AAW20, ASM^{+}21, AHL22, APR24, AON^{+}24, AMRS20, AMM25, ACD^{+}21, ACD^{+}24, BRS^{+}22, BGM^{+}21, BFHC20, BMHR21, Bro24, BCW21, CJR22, CCS20, CNdLL20, CBW^{+}23, CBZZ24, CRX24, CWTL21, CBDK23, CDO24, CDET22, CPD20, CDLN21, CDN^{+}22, DL21, DSO20, DSM20, DLBE22, EN23, FWZZ25, FVDF21a, FVDF21b, FNB^{+}24, FKF^{+}23, GB20, GCLB22, GAB20, GPF22, Gir21, GAL20, HCB^{+}20, HOAM23, HPP24, HKP23, HIDT21, HSJB23, ISW25, IBP21, JCNS^{+}22, JYZ^{+}24, KAW^{+}23, KBP24, LMGK22, LWC20, LHN20, LHG^{+}25, LABJ23, LSG^{+}24, LTJ^{+}20, LCC^{+}23a, LBH^{+}24, LBHW25, LK23, LCAC21, LSBG21, LLK^{+}21, LLZ^{+}22, LLNC21, MSMB21, MPRX20, MMDL23, MDML25, MFF^{+}24, MAV24, MN21a, MMSM^{+}23, MVŞ23, MdOT23, MRM^{+}21, MRM^{+}22, NSL^{+}21, NSR^{+}23, NBP24, NZA^{+}25, NQT25, NWZ25, OMA^{+}22]. systems [OSJB21, PTW22, PGW^{+}23, PPL^{+}25, QWG22, RPT23, RPC^{+}24, SWH^{+}20, SCN^{+}21, SB23, SJC^{+}22, TGG21, TAT^{+}23, TBD^{+}23, VKJ^{+}21, VKLM21, Vog20, WLL^{+}24, WZZ21, WM24, WCZL25, WLS^{+}21, WCGS23, WGY^{+}20, WDF^{+}22, YWY^{+}21, dSS^{+}22, dlVRB21, ACA^{+}23, BJB^{+}21, KRD^{+}23, VWSCH23, WFM^{+}24, ZKDP22]. systems-of-systems [FNB^{+}24, HCB^{+}20]. SZZ [RPS^{+}23].

T [FVDF21a, FVDF21b]. T-wise

[FVDF21a, FVDF21b]. T5APR [GSF24]. Tactics [BLTX21, ISW25, MAK23]. tag [LXYL20]. TagDC [LXYL20]. tailored [RSL^{+}21]. tailoring [LWC20]. take [FEBO22]. target [TMV25]. targeted [QHC^{+}24]. Targeting [BHP^{+}21]. task [AWHS22, LWC20, LYG^{+}24, YZS22, YYW^{+}24, ZAS25]. task-adapted [LYG^{+}24]. tasks [AR24, EBAR21, MCMA21, SMVC24, YZS22, ZAS25]. taxonomies [DPA^{+}24, SC22]. Taxonomy [IPB23, MREVEA^{+}22, ASM^{+}21, DN21, GGMH23, MSDC25, SCB22, ZGHG^{+}23, BJB^{+}21]. team [KHEC^{+}23, SvdBHV24, VM24a, VM24b, Woo25, ZZTC23]. Teams [LOBD25, EMCN^{+}22, KdJPK^{+}23, MVGHPT22, MSS21, MMB22, PdS23]. Technical [MFLS22, OMA^{+}22, RPC^{+}24, TKSC20, WB23, AAB^{+}22, BGMB20, BVHHO23, FAG^{+}20, FRP^{+}23, IZAD21, KRLS^{+}20, LLST20, LBT^{+}21, MVSV25, MCDP24, PT21, RMT^{+}22, VKLM21, VCF22, VHB21, YZL^{+}22, ZYZ^{+}23, dTMS21]. technicalities [KRC23]. Technique [AJFD25, BDS^{+}24, Dut24, YWCX24]. Techniques [MMC24, SRM^{+}25, ZNPR^{+}23, AR24, BMLH25, BGC20, CAA^{+}23, DL22, JRM^{+}22, KH24, LDT^{+}25, MSRR22, PÁngeles MoragaGC25, RZLC24, SHLM24, SKG^{+}24, SGV22, SBM23, SSZ20, VIL24, dSFdC^{+}25, VDVC21]. technologies [GS20, GMCA21, LBCG23, SHG24]. technology [BPJ^{+}22, HTC^{+}23, WL24, YSB^{+}21]. technostress [SRS^{+}25]. TEE [LDT22]. TEE-based [LDT22]. telecom [VM24b]. telework [ŠMKGH23]. Template [YZC^{+}23, DFT24, HZ24, WDL^{+}25]. Template-augmented [YZC^{+}23]. template-based [DFT24, HZ24]. templates [LYS^{+}23, MWD24, XXC^{+}24]. Temporal [CDO24, AAB^{+}22, CLC^{+}23, DLBE22, JYZ^{+}24, LWC20, WZZ21]. tenant [AM23, MH20]. tensions [BHS24]. term [CKM^{+}20, WDO24]. tertiary [AA23, CAA^{+}23, LPPG20, ZLS23]. Test [DWG^{+}24, LCS^{+}24, MMDL23, SVV24, TRD^{+}23, WMLM22, ZFS^{+}22, ASSH22, BCF^{+}21, BCF^{+}22, BDLR23, BGE^{+}21, BDMP21, CJR22, CQZ^{+}20, CATA21, DDPP23, DZY^{+}23, DWH23, GF24, HZT^{+}20, Ibi22, IIK21, KKRT24, KA22, KBG^{+}24, KBBD23, LWYW20, LFH^{+}22, LL25, LL23, LTG^{+}25, MCZX20, MS25, MAP^{+}20, MDML25, MAV24, MN21b, NJF20, NHA20, OLR^{+}25, ÖS20, PMM^{+}25, RSLC24, RF24b, SKP20, SGV22, aSLF^{+}22, SSS22, THG20, WYG^{+}24, WC20, WC23, XBS21, XDL^{+}22, ZWY^{+}22, vBD21]. test-based [CJR22]. test-driven [BCF^{+}21, BCF^{+}22]. Test4Enforcers [RMM24]. testbed [MSC20]. testcase [TCTL25]. testers [FSS23]. Testing [AJFD25, AMB^{+}25, CI24, FKF^{+}23, LLNC21, PTW22, SK22a, VKB^{+}25, ABT^{+}22, BMLH25, BM25, BMP22, BK20, BH20, BTSC^{+}23, BG24, CNdLL20, CF23, CCZ^{+}24, CQZ^{+}20, CRV23, DAP20, DFT24, DYZ^{+}23, FWZZ25, FVDF21a, FVDF21b, GRLA20, GKB^{+}21, GGPR24, GGMB^{+}22, GBSO20, HWL^{+}25, JT20, KKRT24, LHY^{+}25, LHN20, LZ24, LTG^{+}25, MLYL25, MSMB21, MDML25, MAC^{+}22, MBP23, NMG^{+}20, PLP^{+}20, POZ20, Pie20, QWGS24, RSM^{+}23, RMM24, SGM23, SCN^{+}21, SCPB25, SG20, aSLF^{+}22, SSS22, SZ23, TS22, VKG^{+}23, VCT20, WYFZ24, WLCC25, WLS^{+}21, WLdCM22, WXZL23, YWY^{+}21, YLHZ20, YLW^{+}21, ZWF^{+}25, ZMLZ23, ZZP21, ZB22, ZSF^{+}23, AHT^{+}21]. testing-based [MAC^{+}22]. Tests [AFJ^{+}20, ECL^{+}22, FFV^{+}23, HG22, KDB^{+}21, DBM^{+}24]. text [HZ24, SGG24, TCTL25, YWY^{+}21, ZYZ^{+}23]. text-to-testcase [TCTL25]. textual [AC21, BRSK24]. TF [CCC^{+}24]. their [GBMF22, HQNR^{+}22, JCNS^{+}22, Jør24, LDH22, LQY^{+}22, LGKT22, OBCR23, OSJB21, SRS^{+}25, SM20a, Tok22, WWC23, ZLZ^{+}25, dDLSK23]. them [ZZP21]. thematic [MHR^{+}25]. themes [SGG22, WMAL21]. theoretic [GCSHB20]. Theoretical [BbASP23]. Theory [IMTS23, DPA^{+}24, EM24, JVZ25, VKLM21, DPGGP23]. there [BRSK24, LID^{+}22]. thermal [GTT^{+}20]. thermal-aware [GTT^{+}20]. ThermoSim [GTT^{+}20]. thinking [ECMC20, LH24b, PPMC22, BFL23, SGW^{+}22]. third [NDDD20]. third-party [NDDD20]. Thoughts [Win24]. threat [BCJI22, GSM^{+}23]. threats [TST^{+}21a, TST^{+}21b]. Three [GMFO^{+}21a, GMFO^{+}21b, SVV24, BCF^{+}22, KL20, Liu21, PPM^{+}21, SGM23]. three-layered [KL20]. three-perspective [PPM^{+}21]. three-step [Liu21]. Three-Tier [SVV24]. Thriving [ZGLPA24]. Tier [SVV24, MPRX20]. tiered [EM20]. Tigris [MN21a]. Time [Bro24, CRX24, FG24, KRP24, AAB^{+}22, BS23, BCF^{+}21, CXP^{+}23, CCZ^{+}24, CSLN23, DXD25, HIDT21, JYZ^{+}24, JSG25, LHN20, LID^{+}22, MCMA21, MDAHS25, NLTM23, NNV^{+}24, PGR24, RK20, RPM^{+}22, SKHLS24, SMB^{+}20, SB23, TMV25, YZZ^{+}24, YAaO^{+}21, ZGW^{+}20a, ZSCD22]. time-based [TMV25]. time-series [MDAHS25]. Timed [HSJB23]. times [ST24]. timing [YAaO^{+}21]. title [LCC^{+}23b, ZLW^{+}23]. TitleGen [LCC^{+}23b]. TitleGen-FL [LCC^{+}23b]. TLS [WXL^{+}20]. token [WDXX23, ZFG^{+}23]. tolerance [MRM^{+}22, dSFdC^{+}25]. tolerant [FNB^{+}24, LLWL22]. tool [BMP22, CDM25, GGB20, JCNS^{+}22, KSG^{+}22, MHJW22, OLR^{+}25, RO22, ZLW^{+}24]. tools [BG24, GGP^{+}24, HLS24, JLL23, LBT^{+}21, LPS^{+}23, MBG^{+}25, MYH^{+}24, NMRS22, RCA^{+}23, SM24, SZSV22, VDVC21]. top [Hor21, ZLG25]. topic [SGG24, WLCC25]. topics [ZZH^{+}25, GMCA21]. TOSCA [DJS^{+}22]. TOSCAdata [DJS^{+}22]. Tower [PKT24]. Trace [Hu24]. Traceability [MLPC20, FAA22, PC23, WCH^{+}22, ZXZS24]. traces [AHL22, DD20, MN23, PHLHM24, SH20]. tracing [GCF22, JLL23, NEJ24]. tracked [DMAH24]. tracker [BMLH25]. tracking [HHK20, MBG^{+}25, MRS20]. TrAdaBoost [CCC^{+}24]. trade [DCÁ^{+}24]. trade-off [DCÁ^{+}24]. tradeoff [CWGS23]. tradeoffs [WCGS23]. traditional [ZH21, ZH22]. Train [MLPC20, CWTL21, CDP22]. trained [CLNQ24, FCW^{+}23, ZAS25]. training [AMHB24, NRFS25, SJH^{+}24, WDL^{+}25]. trainings [ZGLPA24]. traits [VM24b]. transactional [EN23]. transactive [LWC20]. transfer [DXD25, KBP24, SELS21, TLXW23, ZJXG20]. transfer-learning [SELS21]. Transferable [QWHH23]. Transformation [DLBE22, YWW22, HZ24, KKRT24, KRLS^{+}20, MCRM25, PN21, RVAC^{+}25, ZHS^{+}24]. Transformation-based [DLBE22]. transformations [BBW22, SOMSCT23, AZKR20]. Transformed [LMVRA^{+}20]. transformer [JYZ^{+}24, NAEAK25, ZLW^{+}23, ZFG^{+}23, ZXH^{+}24, HHZW23, WLH^{+}25]. transformer-based [NAEAK25]. Transforming [KKE21, MPFB23]. transition [MDML25]. Transitioning [SMM23]. Translating [LB24]. Translation [CATA21, LYG^{+}24]. tree [BN23, HGH^{+}23, LL25]. tree-based [LL25]. trees [AZB25, ZXW20]. Trello [EMCN^{+}22]. Trends [BDLR23, CJNDZ21, LHS^{+}24, LHH^{+}24, Pie20]. triage [AWHS22, JCNS^{+}22]. triaging [DLX^{+}23]. Triggers [GBMF22]. triple [TLXW23]. Trust [CID25, BSCS23, DLBE22, HG22]. Trustworthiness [MICV23]. trustworthy [BCW21]. truth [GDLM21]. Tuning [CYW21, CBZZ24, FCW^{+}23, LBHW25, SPZW25, WDL^{+}25]. Turkey [Tok22]. Turnover [AAA24, SS20]. tutorial [WZFD25]. twice [VM24b]. Twins [DJR^{+}22, PZVN^{+}25]. Two [ŠMKGH23, TST^{+}21a, AAZB23, BDLT21, CTD^{+}23, FKF^{+}23, LH24b, LDH22, YXLH24, TST^{+}21b]. two-phase [FKF^{+}23]. two-stage [YXLH24]. type [LDH22]. types [PMDN20, RPL^{+}21].

UAV [WXZL23, ZZTC23]. UAV-ship [ZZTC23]. UML [AC21, BHHQ^{+}22, CZLN22]. Uncertainty [ALZ^{+}20, SCN^{+}21, TSP20, ACPM22, BHP^{+}21, HBK25, RFB20]. Uncertainty-aware [SCN^{+}21]. Uncovering [DdM^{+}24]. underlying [SRS^{+}25]. understandability [LALM23, NWZ25, PSZ21]. Understanding [AC25, AAA24, BSS^{+}25, KL20, LGKT22, NF24, SJH^{+}25, SM20b, Vid22, VML21, YU22, KGM24, NM25, TH24, ZJY^{+}24]. unequal [KHEC^{+}23]. Unified [SWH^{+}20, LCS^{+}24]. uniform [BMB20]. uniqueness [WC23]. uniqueness-based [WC23]. unit [KA22, DBM^{+}24, THG20]. unitarity [LZ24]. units [PdS23]. Universal [BSF^{+}25, MPFB23, NRRS20, QHC^{+}24]. university [Hyr25]. Unlocking [CGM24]. Unmanned [ACA^{+}23]. unnecessary [LDT22]. unsupervised [XLY^{+}21]. Unveiling [HWL^{+}25, MDAHS25, SVV24]. Update [SK22a, MWFK20, STW23]. updated [dACOdS23]. updates [ACC^{+}20, HG22, KBP24, SK22a]. uphill [ES23]. ups [ZH22]. Usability [NIJ22, MH24, RCA^{+}23, SZSV22]. Usage [THN20, DMAH24, KFJA23, MWD24]. usage-driven [KFJA23]. usages [HBSV^{+}22]. use [BMB20, GPF22, GBK^{+}23, LBF^{+}21a, LBF^{+}21b, Mec24, SLL20, SM20b, dCMM^{+}22]. used [ACSJ23, MCH22, OKG25, PPMC22, PSGD23, ZH22]. usefulness [TGG21, dDLSK23]. User [MWY^{+}22, SVV24, AAZB23, AWP^{+}25, APB20, BSCS23, DPRD21, GCF22, GM20, MDWS^{+}23, MRA22b, SSO^{+}23, TGG21, WGL^{+}22, WZL^{+}25, WLCC25, ZFC^{+}22, ZZH^{+}25, ZLS25, BFL23]. user-oriented [WZL^{+}25]. users [LH24b, MHJW23, WZL^{+}25]. userspace [DD20]. Using [ALB20, AAB^{+}22, HELW20, Ibi22, LL23, NHA20, XTF^{+}21, AHP21, AHL22, APR24, ABT^{+}22, BDS^{+}24, BAH24, BHHQ^{+}22, dACO21, CAF22, CKM^{+}20, CLNQ24, DOG^{+}23, DWG^{+}24, DFT24, DHM^{+}21, Dut24, FWZZ25, FAFJ24, GGB20, GF24, GRV^{+}21, Hu25, KDB^{+}21, KOPN22, LCS^{+}24, LZP^{+}20, Mam23, MRS25, MLPC20, MCM^{+}24, MC20, MHOM22, MHH21, NdSR^{+}21, OBW^{+}25, PT21, PLP^{+}20, PHLHM24, PDGMT20, PMDN20, dPPdJFF21, QZS^{+}24, QHC^{+}24, QWG22, RLK^{+}24, ROGS^{+}24, RF24b, SCdPL24, SKHLS24, SMKI23, SRD^{+}21, SBM23, SOMSCT23, TXW^{+}20, TBD^{+}23, VHSB22, WMLM22, WL24, WLCC25, WLH^{+}25, WSL^{+}20, YWCX24, YCWD23, ZFC^{+}22, ZFS^{+}22, ZCZL23, ZHLR23, ZGLPA24, ZXW20, vdG25, BFHC20]. utilizing [RSU^{+}24]. UVL [BSF^{+}25, ROGS^{+}24]. UVLHub [ROGS^{+}24]. UX [APB20, GBLP25, NdOdO^{+}22].

V4rdiac [FMG^{+}25]. valid [THG20]. validating [CKY25, MOP24]. validation [BG24, CBC^{+}24, FGG^{+}20, LTJ^{+}20, LM21, MRR^{+}20, MRM^{+}21, YRC^{+}25, ZHM^{+}23]. validity [HXJ^{+}20]. validity-based [HXJ^{+}20]. Valkyrie [RZLC24]. values [KKE21, RFB20, SSO^{+}23, WSL^{+}20, ZB22]. VANET [FR20]. Variability [BSF^{+}25, DHK^{+}23, GFCLB^{+}24, MFF^{+}24, MPFB23, WMA^{+}24, FMG^{+}25, KGS23, MCDP24, NZA^{+}25, YYT25, FGS23]. Variable [DYZ^{+}23, LWL^{+}23, SV20, ZPL21]. Variable-strength [DYZ^{+}23]. Variable-suspiciousness-based [LWL^{+}23]. variant [DSB23]. variant-rich [DSB23]. variants [DMAH24, RPS^{+}23]. variation [LSBG21]. variations [YZS22]. various [GS21a, GS21b]. vehicle [GAL20]. vehicles [ALZ^{+}20, FR20, MKB23, VGS^{+}23]. vehicular [FR20]. Vendor [SM20a]. verification [AHP21, ACPM22, ALZ^{+}20, ADS^{+}22, BGM^{+}21, CJR22, CWTL21, DLCH25, Dut24, JJS24, LZP^{+}20, LCW^{+}24, MBKL24, RCT22, RHJ25, SLL20, SB23, WXL^{+}20, YRC^{+}25, ZTK^{+}23]. verifiers [LH24a]. verify [HIDT21, XTF^{+}21]. verifying [RO22]. Version [WM24, GSK24, LGC^{+}25]. versioning [OZVRD21]. versus [NIJ22, SWH^{+}20]. very [LK23]. very-large [LK23]. via [ARPDP24, CWGS23, CBW^{+}23, CBZZ24, CLLJ24, CSLN23, DBB20, FCW^{+}23, GJW^{+}22, JSG25, KBBD23, LTJ^{+}20, LYG^{+}24, MCZX20, MWD24, MFPB24, PPL^{+}25, SPZW25, TTL20, WTG23, WZL^{+}25, YWY^{+}21, YML^{+}22, YWLZ23, ZJXG20, ZXH^{+}24, ZYL^{+}22, ZSF^{+}23]. VIBE [FGS23]. vicious [FEBO22]. video [AA22, KSF20]. videos [KSF20]. view [AKL25, FRP^{+}23, KGS^{+}25, KKL^{+}21]. view-based [KGS^{+}25, KKL^{+}21]. viewpoints [NLTM23]. views [Fei23, NLTM23]. vintage [GW24]. violations [SHWR22]. Virtual [AAA24, DD20, HLW^{+}23, THB22, LCY23, TMV25]. virtualized [CRX24, CDO24]. vision [KSF20, RPT23]. visual [AC25, BRSK24, CATA21, MHOM22, PBV24]. Visualization [MCDP24, RV22, BMLH25, FFSB23, KBW^{+}24, MLGBRC24]. visually [CXY^{+}23]. ViT [ZAS25]. Vitruvius [KKL^{+}21]. vocal [SJH^{+}24, ZCL24]. voice [STM^{+}22]. volatile [WDF^{+}22]. volume [NBGC20a]. voluntary [ŠMKGH23]. vs [LOBD25]. VsusFL [LWL^{+}23]. Vulnerabilities [AJFD25, GFS21, RHJ25, WXL^{+}20, WLLJ24, YU22]. Vulnerability [OHM25, SK22b, VIL24, AMHB24, ASK^{+}23, CLZ^{+}23, CCC^{+}24, CLZ^{+}24, CKM^{+}20, CLC^{+}23, CFF^{+}23, CCCW24, LMZ^{+}23, LID^{+}22, NNV^{+}24, RSVW23, SYB^{+}23, TTB^{+}23, TQT^{+}25, VVBGL^{+}23, WTG23, WLLJ24, YZM^{+}25, YML^{+}22, YWLZ23, ZX23, ZGW^{+}20b, ZLD^{+}25]. Vulnerable [LFFW23, YML^{+}22]. Vulpedia [YML^{+}22]. VulSlicer [SK22b].

want [KRC23]. WARDER [HXJ^{+}20]. Warn [EKK^{+}24]. warnings [MFBP20]. WASMICO [RRFD24]. Waterfall [LOBD25]. wavelet [WDO24]. way [TAF^{+}20, vRMG23]. Wayback [JCNS^{+}22]. ways [GFS21]. WDBT [WDF^{+}22]. weak [Mas22]. weaknesses [MREVEA^{+}22]. wear [WDF^{+}22]. Weaving [ARPDP24]. web [CZW20, CFNA25, DAP20, GMS22, GKB^{+}21, IIK21, LSB^{+}22, NRFS25, OLR^{+}25, RK20, RB23, SMKI23, WGMT24, WGMT25, ZLX^{+}22, vRMG23, BM25, DFT24, FBV24, GRdAL23, KFJA23, SSS22, TPGH20]. web-based [WGMT24]. WebAssembly [RRFD24]. WebDriver [GKAHMO22]. WebRTC [GGMB^{+}22]. website [AA24]. websites [ACSJ23]. weeks [GGP21]. weight [MDRdP^{+}24]. weighted [TLXW23]. well [Tok22]. well-being [Tok22]. Which [JS22]. White [CBC^{+}24]. White-box [CBC^{+}24]. who [BDLT21]. Wide [OHM25, FVDF21a, FVDF21b, MGSC22]. Widening [MFPB24]. winding [GBT^{+}20]. Windows [MMC24]. wise [FVDF21a, FVDF21b]. within [DdM^{+}24, LCT22, MH24]. wolf [ZHLR23]. word [CHLT23, WJX^{+}24]. Words [LOBD25]. Work [SMH^{+}23, APB20, BLHS23, Bat20, GBLP25, KI23, Tok22, ZGLPA24]. Work-from-home [SMH^{+}23]. workflow [LLWL22, XDL^{+}22]. workflows [DMD23, YWTA25]. Working [ŠMKGH23, CR23]. working-from-anywhere [ŠMKGH23]. Working-From-Home [ŠMKGH23]. worklist [SVVD21]. workload [HDX^{+}23]. workloads [MdSKD22]. world [AHT^{+}21, KH24, MLBD21, YLG^{+}25]. worry [WZL^{+}25]. Written [LOBD25, NDD^{+}24, SS23].

XAI [AR24]. XC [ACD^{+}24]. Xen [CDO24]. XOR [FC20]. XOR-Masking [FC20].

years [FKGN23].

zip [ZLS25]. Zipper [MRVS24]. Zipperbased [MRVS24].

References

Albaghajati:2022:CEG

[AA22] Aghyad Albaghajati and Moataz Ahmed. A co-evolutionary genetic algorithms approach to detect video game bugs. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                                             URL http://www.
                                                             sciencedirect.com/

science/article/
pii/S0164121222000292.

Abbad-Andaloussi:2023:RBS

[AA23] Amine Abbad-Andaloussi.

On the relationship between sourcecode metrics and cognitive load: a systematic tertiary review. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223000146.

Alanazi:2024:WDA

[AA24] Afit Alanazi and Reem Alfayez. What is discussed about Flutter on Stack Overflow (SO) questionand-answer (Q&A) website: an empirical study. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001341.

Akdur:2024:UVO

[AAA24] Gorkem Akdur, Mehmet N. Aydin, and Gizdem Akdur. Understanding virtual onboarding dynamics and developer turnover intention in the era of pandemic. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122400181X.

Ardimento:2022:UDT

[AAB^{+}22] Pasquale Ardimento, Lerina Aversano, Mario Luca Bernardi, Marta Cimitile, and Martina Iammarino. Using deep temporal convolutional networks to just-in-time forecast technical debt principal. The Journal of Systems and Software, 194(??):??, December 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222001649.

Arvanitou:2021:SEP

[AACC21] Elvira-Maria Arvanitou, Apostolos Ampatzoglou, Alexander Chatzigeorgiou, and Jeffrey C. Carver. Software engineering practices for scientific software development: a systematic mapping study. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302387.

Aliabadi:2021:AMD

[AAG21] Maryam Raiyat Aliabadi, Mojtaba Vahidi Asl, and Ramak Ghavamizadeh. ARTI-NALI++: Multidimensional specification mining for complex cyber-physical system security. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001138.

Abbas:2020:AMD

[AAW20] Nadeem Abbas, Jesper Andersson, and Danny Weyns. AS-PLe: a methodology to develop selfadaptive software systems with systematic reuse. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301047.

Aljedaani:2023:EUK

[AAZB23] Bakheet Aljedaani, Aakash Ahmad, Mansooreh Zahedi, and M. Ali Babar. End-users’ knowledge and perception about security of clinical mobile health apps: a case study with two Saudi Arabian mHealth providers. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001959.

Al-Bataineh:2024:ERB

[ABMV24] Omar I. Al-Bataineh, Leon Moonen, and Linas Vidziunas. Extending the range of bugs that automated program repair can handle. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003138.

Avritzer:2022:STA

[ABT^{+}22] Alberto Avritzer, Ricardo Britto, Catia Trubiani, Matteo Camilli, Andrea Janes, Barbara Russo, André van Hoorn, Robert Heinrich, Martina Rapp, Jörg Henß, and Ram Kishan Chalawadi. Scalability testing automation using multivariate characterization and detection of software performance antipatterns. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122200142X.

Addazi:2021:BGT

[AC21] Lorenzo Addazi and Federico Ciccozzi. Blended graphical and textual modelling for UML profiles: a proofof-concept implementation and experiment. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000091.

Akanda:2025:UCP

[AC25] Wahiduzzaman Akanda and James Clause. Understanding comment practices in Scratch: a study of comments in a block-based visual programming language. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S016412122400373X.

AlIslam:2023:CMS

[ACA^{+}23] Md Nafee Al Islam, Muhammed Tawfiq Chowdhury, Ankit Agrawal, Michael Murphy, Raj Mehta, Daria Kudriavtseva, Jane Cleland-Huang, Michael Vierhauser, and Marsha Chechik. Configuring missionspecific behavior in a product line of collaborating Small Unmanned Aerial Systems. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/

                                        pii/S0164121222002199.

Ahmad:2020:SAP

[ACC^{+}20] Maqsood Ahmad, Valerio Costamagna, Bruno Crispo, Francesco Bergadano, and Yury Zhauniarovich. StaDART: Addressing the problem of dynamic code updates in the security analysis of Android applications. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219301530.

Abdelfattah:2025:MSM

[ACC^{+}25] Amr S. Abdelfattah, Tomas Cerny, Md Showkat Hossain Chy, Md Arfan Uddin, Samantha Perry, Cameron Brown, Lauren Goodrich, Miguel Hurtado, Muhid Hassan, Yuanfang Cai, and Rick Kazman. Multivocal study on microservice dependencies. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000020.

Aksakalli:2021:DCP

[AÇCT21] Işıl Karabey Aksakalli, Turgay Çelik, Ahmet Burak Can, and Bedir Teki̇nerdoğan. Deployment and communication patterns in microservice architectures: a systematic literature review. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001114.

Audrito:2021:ADM

[ACD^{+}21] Giorgio Audrito, Roberto Casadei, Ferruccio Damiani, Volker Stolz, and Mirko Viroli. Adaptive distributed monitors of spatial properties for cyberphysical systems.

The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000054.

Audrito:2024:ECX

[ACD^{+}24] Giorgio Audrito, Roberto Casadei, Ferruccio Damiani, Guido Salvaneschi, and Mirko Viroli. The eXchange Calculus (XC): a functional programming language design for distributed collective systems. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000190.

Albert:2021:DRF

[ACG^{+}21] Elvira Albert, Jesús Correas, Pablo Gordillo, Guillermo Román-Díez, and Albert Rubio. Don’t run on fumes — parametric gas bounds for smart contracts. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000200.

Albert:2025:HHA

[ACG^{+}25] Elvira Albert, Jesús Correas, Pablo Gordillo, Guillermo Román-Díez, and Albert Rubio. Harnessing heap analysis for the synthesis of superoptimized bytecode. The Journal of Systems and Software, 221(??):??, March 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003285.

Ahmad:2025:IMC

[ACH25] Syed Ishtiaque Ahmad, Shaiful Chowdhury, and Reid Holmes. Impact of methodological choices on the analysis of code metrics and maintenance. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003078.

Ajimati:2025:ALC

[ACM25] Matthew Oladeji Ajimati, Noel Carroll, and Mary Maher. Adoption of low-code and no-code development: a systematic literature review and future research agenda. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003443.

Alasmari:2022:QVA

[ACPM22] Naif Alasmari, Radu Calinescu, Colin Paterson, and Raffaela Mirandola. Quantitative verification with adaptive uncertainty reduction. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200036X.

Alghamdi:2023:HWU

[ACSJ23] Omar Alghamdi, Sarah Clinch, Rigina Skeva, and Caroline Jay. How are websites used during development and what are the implications for the coding process? The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300198X.

Alrashedy:2020:SPP

[ADG^{+}20] Kamel Alrashedy, Dhanush Dharmaretnam, Daniel M. German, Venkatesh Srinivasan, and T. Aaron Gulliver. SCC++:

Predicting the programming language of questions and snippets of Stack Overflow. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302791.

Albert:2023:ODP

[AdlBGZ^{+}23] Elvira Albert, Maria Garcia de la Banda, Miguel Gómez-Zamalloa, Miguel Isabel, and Peter Stuckey. Optimal dynamic partial order reduction with contextsensitive independence and observers. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001255.

Audrito:2022:DRV

[ADS^{+}22] Giorgio Audrito, Ferruccio Damiani, Volker Stolz, Gianluca Torta, and Mirko Viroli. Distributed runtime verification by past-CTL and the field calculus. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000243.

AlSafwan:2022:DNR

[AES22] Khadijah Al Safwan, Mohammed Elarnaoty, and Francisco Servant. Developers’ need for the rationale of code commits: an inbreadth and indepth study. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000668.

Altuwaijri:2022:FAA

[AF22] Fahad S. Altuwaijri and Maria Angela Ferrario. Factors affecting Agile adoption: an industry research study of the mobile app sector in Saudi Arabia. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200084X.

Avritzer:2020:SAM

[AFJ^{+}20] Alberto Avritzer, Vincenzo Ferme, Andrea Janes, Barbara Russo, André van Hoorn, Henning Schulz, Daniel Menasché, and Vilc Rufino. Scalability assessment of microservice architecture deployment configurations: a domain-based approach leveraging operational profiles and load tests. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030042X.

Alanazi:2021:FPC

[AGL21] Rakan Alanazi, Gharib Gharibi, and Yugyung Lee. Facilitating program comprehension with call graph multilevel hierarchical abstractions. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100042X.

Alberts:2025:SAB

[AGM^{+}25] Elvin Alberts, Ilias Gerostathopoulos, Ivano Malavolta, Carlos Hernández Corbato, and Patricia Lago. Software architecture-based self-adaptation in robotics. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003029.

Assyne:2022:SRS

[AGP22] Nana Assyne, Hadi Ghanbari, and Mirja Pulkkinen. The state of research on software engineering competencies: a systematic mapping study. The Journal of Systems and Software, 185(??):??, March 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002648.

Agh:2020:RAS

[AGPR20] Halimeh Agh, Félix Garcia, Mario Piattini, and Raman Ramsin. Requirements for adopting software process lines. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121220300285.

Alakula:2024:PPL

[AHFP24] Anton Risberg Alaküla, Görel Hedin, Niklas Fors, and Adrian Pop. Property probes: Live exploration of program analysis results. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000232.

Alawneh:2022:LCI

[AHL22] Luay Alawneh and Abdelwahab Hamou-Lhadj. Locating and categorizing inefficient communication patterns in HPC systems using interprocess communication traces. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121222001704.

Albonico:2023:SER

[A– DHM23] Michel Albonico, Milica – Dor– dević, Engel Hamer, and Ivano Malavolta. Software engineering research on the Robot Operating System: a systematic mapping study. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002503.

Alam:2021:DRA

[AHP21] Md. Imran Alam, Raju Halder, and Jorge Sousa Pinto. A deductive reasoning approach for database applications using verification conditions. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121220302934.

Ashfaq:2021:SAC

[AHT^{+}21] Muhammad Ashfaq, Rubing Huang, Dave Towey, Michael Omari, Dmitry Yashunin, Patrick Kwaku Kudjo, and Tao Zhang. SWFC-ART: a costeffective approach for Fixed-Size-Candidate-Set Adaptive Random Testing through small world graphs. The Journal of Systems and Software, 180(??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001059.

Aghamohammadi:2020:GSM

[AIH20] Alireza Aghamohammadi, Maliheh Izadi, and Abbas Heydarnoori. Generating summaries for methods of event-driven programs: an Android case study. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302065.

Amalfitano:2025:GBM

[AJFD25] Domenico Amalfitano, Misael Júnior, Anna Rita Fasolino, and Marcio Delamaro. A GUIbased metamorphic testing technique for detecting authentication vulnerabilities in Android mobile apps. The Journal of Systems and Software, 224(??):??, June 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000329.

Akbarinasaji:2020:POM

[AKBN20] Shirin Akbarinasaji, Can Kavaklioglu, Ayse Basar, and Adam Neal. Partially observable Markov decision process to generate policies in software defect management. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300017.

Aagren:2022:AEC

[ÅKH^{+}22] S. Magnus Ågren, Eric Knauss, Rogardt Heldal, Patrizio Pelliccione, Anders Alminger, Magnus Antonsson, Thomas Karlkvist, and Anders Lindeborg. Architecture evaluation in continuous development. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002089.

Ahn:2025:REA

[AKL25] Hwi Ahn, Sungwon Kang, and Seonah Lee. Reconstruction of an execution architecture view by identifying mapping rules for connectors. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003121.

Assuncao:2023:HDM

[AKMS23] Wesley K. G. Assunção, Jacob Krüger, Sébastien Mosser, and Sofiane Selaoui. How do microservices evolve? An empirical analysis of changes in opensource microservice repositories. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001838.

Alabdulatif:2021:PHC

[Ala21] Abdulatif Alabdulatif. Practical hybrid confidentialitybased analytics framework with Intel SGX. The Journal of Systems and Software, 181(??): ??, November 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001424.

Agnelo:2020:UOD

[ALB20] João Agnelo, Nuno Laranjeiro, and Jorge Bernardino. Using Orthogonal Defect Classification to characterize NoSQL database defects. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302250.

Aghayi:2021:CBD

[ALSA21] Emad Aghayi, Thomas D.

LaToza, Paurav Surendra, and Seyedmeysam Abolghasemi. Crowdsourced behaviordriven development. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302302.

An:2020:UMR

[ALZ^{+}20] Dongdong An, Jing Liu, Min Zhang, Xiaohong Chen, Mingsong Chen, and Haiying Sun. Uncertainty modeling and runtime verification for autonomous vehicles driving control: a machine learningbased approach. The Journal of Systems and Software, 167(??):??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220300959.

Athanasopoulos:2023:MOE

[AM23] Dionysis Athanasopoulos and Mitchell McEwen. Multiobjective empirical computational complexity of singletenant service instances deployed at the Edge. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000602.

Augusto:2025:RCR

[AMB^{+}25] Cristian Augusto, Jesús Morán, Antonia Bertolino, Claudio de la Riva, and Javier Tuya. RE-TORCH*: a cost and resource aware model for E2E testing in the cloud. The Journal of Systems and Software, 221(??):??, March 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224002814.

AlDebeyan:2024:IHE

[AMHB24] Fahad Al Debeyan, Lech Madeyski, Tracy Hall, and David Bowes. The impact of hard and easy negative training data on vulnerability prediction performance. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000463.

Arcaini:2025:ISS

[AMM25] Paolo Arcaini, Andriy Miranskyy, and Hausi Müller. Introduction to the special section on software engineering for hybrid quantum computing systems. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000305.

AlOmar:2021:TAC

[AMO21]  Eman  Abdullah

AlOmar, Mohamed Wiem Mkaouer, and Ali Ouni. Toward the automatic classification of self-affirmed refactoring. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030217X.

Arcaini:2020:MPL

[AMRS20] Paolo Arcaini, Raffaela Mirandola, Elvinia Riccobene, and Patrizia Scandurra. MSL: a pattern language for engineering selfadaptive systems. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300406.

Aldndni:2023:APD

[AMS23] Waad Aldndni, Na Meng, and Francisco Servant. Automatic prediction of developers’ resolutions for software merge conflicts. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002315.

Alfadel:2023:EAS

[ANC^{+}23] Mahmoud Alfadel, Nicholas Alexandre Nagy, Diego Elias Costa, Rabe Abdalkareem, and Emad Shihab. Empirical analysis of securityrelated code reviews in npm packages. The Journal of Systems and Software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223001474.

Anonymous:2020:Aa

[Ano20a] Anonymous. April 2020. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. August 2020. The Journal of Systems and Software, 166(??):??, August 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:D

[Ano20c] Anonymous. December 2020. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:EBa

[Ano20d] Anonymous. Editorial Board. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302432.

Anonymous:2020:EBb

[Ano20e] Anonymous. Editorial Board. The Journal of Systems and Software, 160 (??):??, February 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302493.

Anonymous:2020:EBc

[Ano20f] Anonymous. Editorial Board. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302754.

Anonymous:2020:EBd

[Ano20g] Anonymous. Editorial Board. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300133.

Anonymous:2020:EBe

[Ano20h] Anonymous. Editorial Board. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300364.

Anonymous:2020:EBf

[Ano20i] Anonymous. Editorial Board. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300625.

Anonymous:2020:EBg

[Ano20j] Anonymous. Editorial Board. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300820.

Anonymous:2020:EBh

[Ano20k] Anonymous. Editorial Board. The Journal of Systems and Software, 166(??):??, August 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122030100X.

Anonymous:2020:EBi

[Ano20l] Anonymous. Editorial Board. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301345.

Anonymous:2020:EBj

[Ano20m] Anonymous. Editorial Board. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301795.

Anonymous:2020:EBk

[Ano20n] Anonymous. Editorial Board. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302016.

Anonymous:2020:EBl

[Ano20o] Anonymous. Editorial Board. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302247.

Anonymous:2020:F

[Ano20p] Anonymous. February 2020. The Journal of Systems and Software, 160 (??):??, February 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:Ja

[Ano20q] Anonymous. January 2020. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:Jc

[Ano20r] Anonymous. July 2020. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:Jb

[Ano20s] Anonymous. June 2020. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:Ma

[Ano20t] Anonymous. March 2020. The Journal of Systems and Software, 161(??): ??, March 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:Mb

[Ano20u] Anonymous. May 2020. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:N

[Ano20v] Anonymous. November 2020. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:O

[Ano20w] Anonymous. October 2020. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2020:S

[Ano20x] Anonymous. September 2020. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Ab

[Ano21b] Anonymous. August 2021. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:EBa

[Ano21d] Anonymous. Editorial Board. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302478.

Anonymous:2021:EBb

[Ano21e] Anonymous. Editorial Board. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302673.

Anonymous:2021:EBc

[Ano21f] Anonymous. Editorial Board. The Journal of Systems and Software, 173(??):??, March 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302867.

Anonymous:2021:EBd

[Ano21g] Anonymous. Editorial Board. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000133.

Anonymous:2021:EBe

[Ano21h] Anonymous. Editorial Board. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000261.

Anonymous:2021:EBf

[Ano21i] Anonymous. Editorial Board. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000534.

Anonymous:2021:EBg

[Ano21j] Anonymous. Editorial Board. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000777.

Anonymous:2021:EBh

[Ano21k] Anonymous. Editorial Board. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000947.

Anonymous:2021:EBi

[Ano21l] Anonymous. Editorial Board. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001199.

Anonymous:2021:EBj

[Ano21m] Anonymous. Editorial Board. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001199.

Anonymous:2021:EBk

[Ano21n] Anonymous. Editorial Board. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001527.

Anonymous:2021:EBl

[Ano21o] Anonymous. Editorial Board. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001722.

Anonymous:2021:EBm

[Ano21p] Anonymous. Editorial Board. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001990.

Anonymous:2021:F

[Ano21q] Anonymous. February 2021. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Ja

[Ano21r] Anonymous. January 2021. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Jc

[Ano21s] Anonymous. July 2021. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Jb

[Ano21t] Anonymous. June 2021. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Ma

[Ano21u] Anonymous. March 2021. The Journal of Systems and Software, 173(??): ??, March 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Mb

[Ano21v] Anonymous. May 2021. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:N

[Ano21w] Anonymous. November 2021. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:O

[Ano21x] Anonymous. October 2021. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Sa

[Ano21y] Anonymous. September 2021. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2021:Sb

[Ano21z] Anonymous. September 2021. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002272.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002557.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. The Journal of Systems and Software, 185(??):??, March 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002764.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121222000152.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000401.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000620.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000747.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000991.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001194.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001546.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001753.

Anonymous:2022:EBl

[Ano22o] Anonymous. Editorial Board. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002059.

Anonymous:2022:F

[Ano22p] Anonymous. February 2022. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:Ja

[Ano22q] Anonymous. January 2022. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:Jc

[Ano22r] Anonymous. July 2022. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:Jb

[Ano22s] Anonymous. June 2022. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:Ma

[Ano22t] Anonymous. March 2022. The Journal of Systems and Software, 185(??):

??, March 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:Mb

[Ano22u] Anonymous. May 2022. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:N

[Ano22v] Anonymous. November 2022. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:O

[Ano22w] Anonymous. October 2022. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2022:S

[Ano22x] Anonymous. September 2022. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002436.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. The Journal of Systems and Software, 196 (??):??, February 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200259X.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002746.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300033X.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000535.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121223000857.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001085.

Anonymous:2023:EBh

[Ano23k] Anonymous. Editorial Board. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001541.

Anonymous:2023:EBi

[Ano23l] Anonymous. Editorial Board. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001747.

Anonymous:2023:EBj

[Ano23m] Anonymous. Editorial Board. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002091.

Anonymous:2023:EBk

[Ano23n] Anonymous. Editorial Board. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002406.

Anonymous:2023:EBl

[Ano23o] Anonymous. Editorial Board. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002595.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023. The Journal of Systems and Software, 196 (??):??, February 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:Jc

[Ano23r] Anonymous. July 2023. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:Jb

[Ano23s] Anonymous. June 2023. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:Ma

[Ano23t] Anonymous. March 2023. The Journal of Systems and Software, 197(??): ??, March 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:Mb

[Ano23u] Anonymous. May 2023. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:N

[Ano23v] Anonymous. November 2023. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2023:S

[Ano23x] Anonymous. September 2023. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:D

[Ano24c] Anonymous. December 2024. The Journal of Systems and Software, 218 (??):??, December 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:EBa

[Ano24d] Anonymous. Editorial Board. The Journal of Systems and Software, 207 (??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002868.

Anonymous:2024:EBb

[Ano24e] Anonymous. Editorial Board. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003205.

Anonymous:2024:EBc

[Ano24f] Anonymous. Editorial Board. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003497.

Anonymous:2024:EBd

[Ano24g] Anonymous. Editorial Board. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000384.

Anonymous:2024:EBe

[Ano24h] Anonymous. Editorial Board. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121224000670.

Anonymous:2024:EBf

[Ano24i] Anonymous. Editorial Board. The Journal of Systems and Software, 212(??):??, June 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000931.

Anonymous:2024:EBg

[Ano24j] Anonymous. Editorial Board. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001432.

Anonymous:2024:EBh

[Ano24k] Anonymous. Editorial Board. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001729.

Anonymous:2024:EBi

[Ano24l] Anonymous. Editorial Board. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400219X.

Anonymous:2024:EBj

[Ano24m] Anonymous. Editorial Board. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002371.

Anonymous:2024:EBk

[Ano24n] Anonymous. Editorial Board. The Journal of Systems and Software, 218 (??):??, December 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002553.

Anonymous:2024:F

[Ano24o] Anonymous. February 2024. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:Ja

[Ano24p] Anonymous. January 2024. The Journal of Systems and Software, 207 (??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:J

[Ano24q] Anonymous. June 2024. The Journal of Systems and Software, 212(??):??, June 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:Ma

[Ano24r] Anonymous. March 2024. The Journal of Systems and Software, 209(??): ??, March 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:Mb

[Ano24s] Anonymous. May 2024. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:N

[Ano24t] Anonymous. November 2024. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:O

[Ano24u] Anonymous. October 2024. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2024:S

[Ano24v] Anonymous. September 2024. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2025:Aa

[Ano25a] Anonymous. April 2025. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2025:EBa

[Ano25b] Anonymous. Editorial Board. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003169.

Anonymous:2025:EBb

[Ano25c] Anonymous. Editorial Board. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003376.

Anonymous:2025:EBc

[Ano25d] Anonymous. Editorial Board. The Journal of Systems and Software, 221(??):??, March 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S016412122400356X.

Anonymous:2025:EBd

[Ano25e] Anonymous. Editorial Board. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000378.

Anonymous:2025:EBe

[Ano25f] Anonymous. Editorial Board. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000640.

Anonymous:2025:EBf

[Ano25g] Anonymous. Editorial Board. The Journal of Systems and Software, 224(??):??, June 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000810.

Anonymous:2025:EBg

[Ano25h] Anonymous. Editorial Board. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000913.

Anonymous:2025:F

[Ano25i] Anonymous. February 2025. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2025:Ja

[Ano25j] Anonymous. January 2025. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2025:Jc

[Ano25k] Anonymous. July 2025. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2025:Jb

[Ano25l] Anonymous. June 2025. The Journal of Systems and Software, 224(??):??, June 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2025:Ma

[Ano25m] Anonymous. March 2025. The Journal of Systems and Software, 221(??): ??, March 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2025:M

[Ano25n] Anonymous. May 2025. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Araujo:2024:PEN

[AON^{+}24] Carlos Araújo, Meuse Oliveira, Bruno Nogueira, Paulo Maciel, and Eduardo Tavares. Performability evaluation of NoSQLbased storage systems. The Journal of Systems and Software, 208(??): ??, February 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223002807.

Ayala:2021:PPS

[APAF21] Inmaculada Ayala, Alessandro V. Papadopoulos, Mercedes Amor, and Lidia Fuentes. ProD-SPL: Proactive selfadaptation based on Dynamic Software Product Lines. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000066.

Ananjeva:2020:IUW

[APB20] Alisa Ananjeva, John Stouby Persson, and Anders Bruun. Integrating UX work with agile development through user stories: an action research study in a small software company. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121220301953.

Ali:2024:MSA

[APR24] Nazakat Ali, Sasikumar Punnekkat, and Abdul Rauf. Modeling and safety analysis for collaborative safetycritical systems using hierarchical colored Petri nets. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000013.

Awal:2024:EFE

[AR24] Md. Abdul Awal and Chanchal K. Roy. EvaluateXAI: a framework to evaluate the reliability and consistency of rulebased XAI techniques for software analytics tasks. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002048.

Ayala-Rivera:2024:GCS

[ARPDP24] Vanessa Ayala-Rivera, A. Omar Portillo-Dominguez, and Liliana Pasquale. GDPR compliance via software evolution: Weaving security controls in software design. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001894.

Alomari:2022:CDT

[AS22] Hakam W. Alomari and Matthew Stephan. Clone detection through srcClone: a program slicing based approach. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121221002120.

Agrahari:2023:CGS

[ASCR23] Vartika Agrahari, Shriram Shanbhag, Sridhar Chimalakonda, and A. Eashaan Rao. A catalogue of game-specific antipatterns based on GitHub and Game Development Stack Exchange. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300184X.

Anjum:2023:AVF

[ASK^{+}23] Misbah Anjum, Shakshi Singhal, P. K. Kapur, Sunil Kumar Khatri, and Saurabh Panwar. Analysis of vulnerability fixing process in the presence of incorrect patches. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002011.

Afric:2020:RSC

[ASKS20] Petar Afric, Lucija Sikic, Adrian Satja Kurdija, and Marin Silic. REPD: Source code defect prediction as anomaly detection. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301138.

Sun:2022:PDS

[aSLF^{+}22] Chang ai Sun, Baoli Liu, An Fu, Yiqiang Liu, and Huai Liu. Path-directed source test case generation and prioritization in metamorphic testing. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121221001886.

Abdellatif:2021:TSI

[ASM^{+}21] Manel Abdellatif, Anas Shatnawi, Hafedh Mili, Naouel Moha, Ghizlane El Boussaidi, Geoffrey Hecht, Jean Privat, and Yann-Gaël Guéhéneuc. A taxonomy of service identification approaches for legacy software systems modernization. The Journal of Systems and Software, 173(??): ??, March 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220302582.

Al-Sabbagh:2022:ITC

[ASSH22] Khaled Walid Al-Sabbagh, Miroslaw Staron, and Regina Hebig. Improving test case selection by handling class and attribute noise. The Journal of Systems and Software, 183(??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001904.

Ahmad:2025:TRE

[ATWS25] Hussain Ahmad, Christoph Treude, Markus Wagner, and Claudia Szabo. Towards resourceefficient reactive and proactive autoscaling for microservice architectures. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000585.

Admiraal:2024:DMS

[AvdBG^{+}24]  Chris  Admiraal,

Wouter van den Brink, Marcus Gerhold, Vadim Zaytsev, and Cristian Zubcu. Deriving modernity signatures of codebases with static analysis. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000165.

Ajimati:2022:EAN

[AWA^{+}22] Matthew O. Ajimati, Eoin Whelan, Rick Aalbers, Lorraine Morgan, and Hans van Kranenburg. The effect of advice network connectedness on problemsolving competence among software developers. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001686.

Aung:2022:MTM

[AWHS22] Thazin Win Win Aung, Yao Wan, Huan Huo, and Yulei Sui. Multitriage: a multi-task learning framework for bug triage. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002302.

Auch:2020:SBA

[AWMW20] Maximilian Auch, Manuel Weber, Peter Mandl, and Christian Wolff. Similarity-based analyses on software applications: a systematic literature review. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301278.

Amna:2025:AFI

[AWP^{+}25] Anis R. Amna, Yves Wautelet, Stephan Poelmans, Samedi Heng, and Geert Poels. The AmbiTRUS framework for identifying potential ambiguity in user stories. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000251.

Ampatzoglou:2021:SIS

[AX21] Apostolos Ampatzoglou and Peng Xin. Special issue on software and systems reuse in the Big Data era. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000182.

Ali:2024:BBS

[AXUO24] Asif Ali, Yuanqing Xia, Qasim Umer, and Mohamed Osman. BERT based severity prediction of bug reports for the maintenance of mobile applications. The Journal of Systems and Software, 208(??):??, February 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223002935.

Alvi:2021:SPD

[AZ21] Aleem Khalid Alvi and Mohammad Zulkernine. A security pattern detection framework for building more secure software. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302296.

Akiki:2025:RDG

[AZB25] Paul A. Akiki, Andrea Zisman, and Amel Bennaceur. Resources don’t grow on trees: a framework for resource-driven adaptation. The Journal of Systems and Software, 220(??): ??, February 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412122400311X.

Azizi:2020:SSE

[AZKR20] Banafsheh Azizi, Bahman Zamani, and Shekoufeh Kolahdouz-Rahimi. SEET: Symbolic Execution of ETL Transformations. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301308.

Ali:2021:EPE

[AZR^{+}21] Hashim Ali, Muhammad Zakarya, Izaz Ur Rahman, Ayaz Ali Khan, and Rajkumar Buyya. Electricity price and source aware resource management in geographically distributed heterogeneous datacenters. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221000042.

Bayram:2024:ADQ

[BAH24] Firas Bayram, Bestoun S. Ahmed, and Erik Hallin. Adaptive data quality scoring operations framework using drift-aware mechanism for industrial applications. The Journal of Systems and Software, 217(??): ??, November 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224002280.

Batra:2020:JWF

[Bat20] Dinesh Batra. Jobwork fit as a determinant of the acceptance of large-scale agile methodology. The Journal of Systems and Software, 168(??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121220300583.

Bidlake:2020:SLR

[BAV20] Leah Bidlake, Eric Aubanel, and Daniel Voyer. Systematic literature review of empirical studies on mental representations of programs. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300467.

Bucaioni:2022:EAI

[BB22] Alessio Bucaioni and Matthias Becker. Enabling automated integration of architectural languages: an experience report from the automotive domain. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100203X.

Borstler:2023:IAB

[BbASP23]  Jürgen  Börstler,

Nauman bin Ali, Martin Svensson, and Kai Petersen. Investigating acceptance behavior in software engineering — theoretical perspectives. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002680.

Bulej:2021:MLE

[BBF^{+}21] Lubomír Bulej, Tomáš Bureš, Adam Filandr, Petr Hnětynka, Iveta Hnětynková, Jan Pacovský, Gabor Sandor, and Ilias Gerostathopoulos. Managing latency in edge-cloud environment. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121220302624.

Bacchiani:2025:PRM

[BBG^{+}25]  Lorenzo  Bacchi-

ani, Mario Bravetti, Saverio Giallorenzo, Maurizio Gabbrielli, Gianluigi Zavattaro, and Stefano Pio Zingaro. Proactive– reactive microservice architecture global scaling. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003066.

Berg:2020:AAQ

[BBND^{+}20] Vebjørn Berg, Jørgen Birkeland, Anh Nguyen-Duc, Ilias O. Pappas, and Letizia Jaccheri. Achieving agility and quality in product development — an empirical study of hardware startups. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300777.

Buchmann:2022:BLF

[BBW22] Thomas Buchmann, Matthias Bank, and Bernhard Westfechtel. BXtendDSL: a layered framework for bidirectional model transformations combining a declarative and an imperative language. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000462.

Baldassarre:2021:STD

[BCF^{+}21] Maria Teresa Baldassarre, Danilo Caivano, Davide Fucci, Natalia Juristo, Simone Romano, Giuseppe Scanniello, and Burak Turhan. Studying test-driven development and its retainment over a six-month time span.

The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000340.

Baldassarre:2022:ART

[BCF^{+}22] Maria Teresa Baldassarre, Danilo Caivano, Davide Fucci, Simone Romano, and Giuseppe Scanniello. Affective reactions and test-driven development: Results from three experiments and a survey. The Journal of Systems and Software, 185(??):??, March 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002454.

Bertolotti:2023:SPL

[BCF23a] Francesco Bertolotti, Walter Cazzola, and Luca Favalli. [logo]: Software product lines extraction driven by language server protocol. The Journal of Systems and Software, 205(??): ??, November 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223002042.

Bertolotti:2023:GLR

[BCF23b] Francesco Bertolotti, Walter Cazzola, and Luca Favalli. On the granularity of linguistic reuse. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000997.

Bernsmed:2022:ATM

[BCJI22] Karin Bernsmed, Daniela Soares Cruzes, Martin Gilje Jaatun, and Monica Iovan. Adopting threat modelling in agile software development projects. The Journal of Systems and Software, 183(??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001874.

Beecham:2021:DSA

[BCLN21]  Sarah  Beecham,

Tony Clear, Ramesh Lal, and John Noll. Do scaling agile frameworks address global software development risks? An empirical study. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302181.

Bertolotti:2024:WDD

[BCOC24] Francesco Bertolotti, Walter Cazzola, Dario Ostuni, and Carlo Castoldi. When the dragons defeat the knight: Basilisk an architectural pattern for platform and language independent development. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400133X.

Bures:2021:SIS

[BCW21] Tomas Bures, Radu Calinescu, and Danny Weyns. Special issue on software engineering for trustworthy cyber-physical systems. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000698.

Bucaioni:2025:MDE

[BDIL25] Alessio Bucaioni, Amleto Di Salle, Ludovico Iovino, and Peng Liang. Model-driven engineering for software architecture. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003650.

Bertolino:2023:ISI

[BDLR23] Antonia Bertolino, Guglielmo De Angelis, Maurizio Leotta, and Filippo Ricca. Introduction to the special issue on test automation: Trends, benefits, and costs. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001462.

Bi:2021:AIC

[BDLT21] Tingting Bi, Wei Ding, Peng Liang, and Antony Tang. Architecture information communication in two OSS projects: the why, who, when, and what. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001321.

Brunetto:2021:IAT

[BDMP21] Matteo Brunetto, Giovanni Denaro, Leonardo Mariani, and Mauro Pezzè. On introducing automatic test case generation in practice: a success story and lessons learned. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000303.

Basu:2024:FEQ

[BDS^{+}24] Saikat Basu, Arnav Das, Amit Saha, Amlan Chakrabarti, and Susmita Sur-Kolay. FragQC : an efficient quantum error reduction technique using quantum circuit fragmentation. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001304.

Barrak:2021:WDB

[BEAK21] Amine Barrak, Ellis E. Eghan, Bram Adams, and Foutse Khomh. Why do builds fail? — a conceptual replication study. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000364.

BenSassi:2023:CCS

[BEM^{+}23] Sihem Ben Sassi, Sabrine Edded, Raúl Mazo, Henda Ben Ghezala, and Camille Salinesi. Colla-Config: a stakeholders preferencesbased approach for product lines collaborative configuration. The Journal of Systems and Software, 197(??): ??, March 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412122200262X.

Bhatt:2022:ARR

[BF22] Bhargav Nagaraja Bhatt and Carlo A. Furia. Automated repair of resource leaks in Android applications. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001273.

Bowers:2020:PUS

[BFHC20] Kate M. Bowers, Erik M. Fredericks, Reihaneh H. Hariri, and Betty H. C. Cheng. Providentia: Using searchbased heuristics to optimize satisficement and competing concerns between functional and non-functional objectives in selfadaptive systems. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302717.

Barricelli:2023:ENC

[BFL23] Barbara Rita Barricelli, Daniela Fogli, and Angela Locoro. EUDability: a new construct at the intersection of End-User Development and Computational Thinking. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001923.

Barbosa:2025:AGR

[BFM^{+}25] Larissa Barbosa, Sávio Freire, Rita S. P. Maciel, Manoel Mendonça, Marcos Kalinowski, Zadia Codabux, and Rodrigo Spínola. Attributes of a great requirements engineer. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002449.

Blasco:2021:EAG

[BFZC21] Daniel Blasco, Jaime Font, Mar Zamorano, and Carlos Cetina. An evolutionary approach for generating software models: the case of Kromaia in game software engineering. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302089.

Bombarda:2024:DIV

[BG24] Andrea Bombarda and Angelo Gargantini. Design, implementation, and validation of a benchmark generator for combinatorial interaction testing tools. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003151.

Bhatti:2020:GFO

[BGC20]  Shahzad  Sarwar

Bhatti, Xiaofeng Gao, and Guihai Chen. General framework, opportunities and challenges for crowdsourcing techniques: a comprehensive survey. The Journal of Systems and Software, 167(??): ??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220300893.

Blasi:2021:MAI

[BGE^{+}21] Arianna Blasi, Alessandra Gorla, Michael D. Ernst, Mauro Pezzè, and Antonio Carzaniga.

MeMo: Automatically identifying metamorphic relations in Javadoc comments for test automation. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001382.

Butler:2020:MIO

[BGL^{+}20] Simon Butler, Jonas Gamalielsson, Björn Lundell, Christoffer Brax, Anders Mattsson, Tomas Gustavsson, Jonas Feist, and Erik Lönroth. Maintaining interoperability in open source software: a case study of the Apache PDFBox project. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302262.

Butler:2022:CCA

[BGL^{+}22] Simon Butler, Jonas Gamalielsson, Björn Lundell, Christoffer Brax, Anders Mattsson, Tomas Gustavsson, Jonas Feist, Bengt Kvarnström, and Erik Lönroth. Considerations and challenges for the adoption of open source components in softwareintensive businesses. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002442.

Bernardi:2021:SMF

[BGM^{+}21] S. Bernardi, U. Gentile, S. Marrone, J. Merseguer, and R. Nardone. Security modelling and formal verification of survivability properties: Application to cyberphysical systems. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301710.

Besker:2020:ITD

[BGMB20] Terese Besker, Hadi Ghanbari, Antonio Martini, and Jan Bosch. The influence of technical debt on software developer morale. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300674.

Bidgoli:2020:AAC

[BH20] Atieh Monemi Bidgoli and Hassan Haghighi. Augmenting ant colony optimization with adaptive random testing to cover prime paths. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302699.

Bergstrom:2022:ELQ

[BHHQ^{+}22] Gustav Bergström, Fadhl Hujainah, Truong Ho-Quang, Rodi Jolak, Satrio Adi Rukmono, Arif Nurwidyantoro, and Michel R. V. Chaudron. Evaluating the layout quality of UML class diagrams using machine learning. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200125X.

Bures:2021:TUS

[BHP^{+}21] Tomáš Bureš, Petr Hnětynka, František Plášil, Dominik Škoda, Jan Kofroň, Rima Al Ali, and Ilias Gerostathopoulos. Targeting uncertainty in smart CPS by confidencebased logic. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100162X.

Barbala:2024:TCD

[BHS24] Astri Moksnes Barbala, Geir Kjetil Hanssen, and Tor Sporsem. Towards a common datadriven culture: a longitudinal study of the tensions and emerging solutions involved in becoming data-driven in a large public sector organization. The Journal of Systems and Software, 218 (??):??, December 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002292.

Brabrand:2024:PUI

[BIT24] Claus Brabrand, Nanna Inie, and Paolo Tell. Programming under the influence: On the effect of heat, noise, and alcohol on novice programmers. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002820.

Boufaied:2021:SBP

[BJB^{+}21] Chaima Boufaied, Maris Jukss, Domenico Bianculli, Lionel Claude Briand, and Yago Isasi Parache. Signalbased Properties of Cyber-Physical Systems: Taxonomy and Logicbased Characterization. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220302715.

BenBraiek:2020:TML

[BK20] Houssem Ben Braiek and Foutse Khomh. On testing machine learning programs.

The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300248.

Barbez:2020:MLB

[BKG20] Antoine Barbez, Foutse Khomh, and Yann-Gaël Guéhéneuc. A machine-learning based ensemble method for anti-patterns detection. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302602.

Blanco:2021:HACa

[BL21a] J. Z. Blanco and D. Lucrédio. A holistic approach for cross-platform software development. The Journal of Systems and Software, 179(??):??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221000820.

Blanco:2021:HACb

[BL21b] J. Z. Blanco and D. Lucrédio. A holistic approach for cross-platform software development. The Journal of Systems and Software, 179(??):??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221000820.

Bao:2023:SBC

[BLHS23] Qihao Bao, Bixin Li, Tianyuan Hu, and Xueyong Sun. A survey of blockchain consensus safety and security: Stateof-the-art, challenges, and future work. The Journal of Systems and Software, 196 (??):??, February 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122200231X.

Bi:2021:MAT

[BLTX21] Tingting Bi, Peng Liang, Antony Tang, and Xin Xia. Mining architecture tactics and quality attributes knowledge in Stack Overflow. The Journal of Systems and Software, 180(??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001023.

Binamungu:2023:BDD

[BM23] Leonard Peter Binamungu and Salome Maro. Behaviour driven development: a systematic mapping study. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001449.

Balsam:2025:WAT

[BM25] Sebastian Balsam and Deepti Mishra. Web application testing — challenges and opportunities. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002309.

Bjeladinovic:2020:PAI

[BMB20] Srdja Bjeladinovic, Zoran Marjanovic, and Sladjan Babarogic. A proposal of architecture for integration and uniform use of hybrid SQL/NoSQL database components. The Journal of Systems and Software, 168(??): ??, October 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220301060.

Brataas:2021:AES

[BMHR21] Gunnar Brataas, Antonio Martini, Geir Kjetil Hanssen, and Georg Ræder. Agile elicitation of scalability requirements for open systems: a case study. The Journal of Systems and Software, 182(??):??, December 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001618.

Balouchi:2025:EEE

[BMLH25]  Kambiz  Nezami

Balouchi, Julien Mercier, and Roberto E. Lopez-Herrejon. An exploratory empirical eye-tracker study of visualization techniques for coverage of combinatorial interaction testing in software product lines. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003054.

Barboni:2022:SMT

[BMP22] Morena Barboni, Andrea Morichetta, and Andrea Polini. SuMo: a mutation testing approach and tool for the Ethereum blockchain. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001418.

Buyuk:2023:DLC

[BN23] O. O. Büyük and A. Nizam. Deep learning with classlevel abstract syntax tree and code histories for detecting code modification requirements. The Journal of Systems and Software, 206(??):??, December 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223002467.

Biffl:2023:ASD

[BNMW23] Stefan Biffl, Elena Navarro, Raffaela Mirandola, and Danny Weyns. Architecting for a Sustainable Digital Society. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000638.

Breckel:2022:DSL

[BPJ^{+}22] Alexander Breckel, Jakob Pietron, Katharina Juhnke, Florian Sihler, and Matthias Tichy. A domain-specific language for modeling and analyzing solution spaces for technology roadmapping. The Journal of Systems and Software, 184(??): ??, February 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221001916.

Baresi:2024:QQA

[BQR24] Luciano Baresi, Giovanni Quattrocchi, and Nicholas Rasi. A qualitative and quantitative analysis of container engines. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000086.

Behutiye:2022:TOQ

[BRO^{+}22] Woubshet Behutiye, Pilar Rodríguez, Markku Oivo, Sanja Aaramaa, Jari Partanen, and Antonin Abhervé. Towards optimal quality requirement documentation in agile software development: a multiple case study. The Journal of Systems and Software, 183(??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121221002090.

Broy:2024:TCR

[Bro24] Manfred Broy. Time, causality, and realizability: Engineering interactive, distributed software systems. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003357.

Barisic:2022:MPM

[BRS^{+}22] Ankica Barišić, Ivan Ruchkin, Dušan Savić, Mustafa Abshir Mohamed, Rima Al-Ali, Letitia W. Li, Hana Mkaouar, Raheleh Eslampanah, Moharram Challenger, Dominique Blouin, Oksana Nikiforova, and Antonio Cicchetti. Multiparadigm modeling for cyber–physical systems: a systematic mapping review. The Journal of Systems and Software, 183(??): ??, January 2022.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001783.

Bonfanti:2023:CFR

[BRS23] Silvia Bonfanti, Elvinia Riccobene, and Patrizia Scandurra. A component framework for the runtime enforcement of safety properties. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002813.

Boll:2024:BCT

[BRSK24] Alexander Boll, Pooja Rani, Alexander Schultheiß, and Timo Kehrer. Beyond code: Is there a difference between comments in visual and textual languages? The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001328.

Barbudo:2021:GGB

[BRSR21] Rafael Barbudo, Aurora Ramírez, Francisco Servant, and José Raúl Romero. GEML: a grammarbased evolutionary machine learning approach for designpattern detection. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000169.

B:2023:HFD

[BS23] Umamaheswara Sharma B. and Ravichandra Sadam. How far does the predictive decision impact the software project? The cost, service time, and failure analysis from a cross-project defect prediction model. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001984.

Brunotte:2023:PEM

[BSCS23] Wasja Brunotte, Alexander Specht, Larissa Chazette, and Kurt Schneider. Privacy explanations — a means to end-user trust. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002217.

Britto:2020:ESO

[BSDB20] Ricardo Britto, Darja Smite, Lars-Ola Damm, and Jürgen Börstler. Evaluating and strategizing the onboarding of software developers in large-scale globally distributed projects. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301473.

Benavides:2025:UFM

[BSF^{+}25] David Benavides, Chico Sundermann, Kevin Feichtinger, José A. Galindo, Rick Rabiser, and Thomas Thüm. UVL: Feature modelling with the universal variability language. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224003704.

Blasi:2021:RIC

[BSGN21] Arianna Blasi, Nataliia Stulova, Alessandra Gorla, and Oscar Nierstrasz. RepliComment: Identifying clones in code comments. The Journal of Systems and Software, 182(??):??, December 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001667.

Brandon:2020:GBR

[BSH^{+}20] Álvaro Brandón, Marc Solé, Alberto Huélamo, David Solans, María S. Pérez, and Victor Muntés-Mulero. Graph-based root cause analysis for service-oriented and microservice architectures. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302067.

Bock:2025:ULI

[BSS^{+}25] Thomas Bock, Niklas Schneider, Angelika Schmid, Sven Apel, and Janet Siegmund. Understanding the low inter-rater agreement on aggressiveness on the Linux Kernel Mailing List. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122500007X.

Blanco:2023:CGH

[BTSC^{+}23] Raquel Blanco, Manuel Trinidad, María José Suárez-Cabal, Alejandro Calderón, Mercedes Ruiz, and Javier Tuya. Can gamification help in software testing education? findings from an empirical study. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000420.

Bi:2023:CCC

[BVHHO23] Fandi Bi, Birgit Vogel-Heuser, Ziyi Huang, and Felix Ocker. Characteristics, causes, and consequences of technical debt in the automation domain. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001206.

Capiluppi:2020:LCC

[CA20] Andrea Capiluppi and Nemitari Ajienka. Lexical content as a cooperation aide: a study based on Java software. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030025X.

Cerny:2023:CDT

[CAA^{+}23] Tomas Cerny, Amr S.

Abdelfattah, Abdullah Al Maruf, Andrea Janes, and Davide Taibi. Catalog and detection techniques of microservice antipatterns and bad smells: a tertiary study. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002248.

Capiluppi:2020:EMD

[CAC20] Andrea Capiluppi, Nemitari Ajienka, and Steve Counsell. The effect of multiple developers on structural attributes: a study based on Java software. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122030073X.

Cemiloglu:2023:EPI

[CACHA23] Deniz Cemiloglu, Emily Arden-Close, Sarah E. Hodge, and Raian Ali. Explainable persuasion for interactive design: the case of online gambling. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001935.

Canete:2022:SIA

[CAF22] Angel Cañete, Mercedes Amor, and Lidia Fuentes. Supporting IoT applications deployment on edge-based infrastructures using multi-layer feature models. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001837.

Coppola:2021:TLB

[CATA21] Riccardo Coppola, Luca Ardito, Marco Torchiano, and Emil Alégroth. Translation from layoutbased to visual Android test scripts: an empirical evaluation. The Journal of Systems and Software, 171(??): ??, January 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220302351.

Casaluce:2024:WBV

[CBC^{+}24] Roberto Casaluce, Andrea Burattin, Francesca Chiaromonte, Alberto Lluch Lafuente, and Andrea Vandin. Whitebox validation of quantitative product lines by statistical model checking and process mining. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000268.

Chrszon:2023:IDC

[CBDK23] Philipp Chrszon, Christel Baier, Clemens Dubslaff, and Sascha Klüppelholz. Interaction detection in configurable systems — a formal approach featuring roles. The Journal of Systems and Software, 196 (??):??, February 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002321.

Cornejo:2020:FMF

[CBMM20] Oscar Cornejo, Daniela Briola, Daniela Micucci, and Leonardo Mariani. In-the-field monitoring of functional calls: Is it feasible? The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300066.

Cao:2023:CCC

[CBW^{+}23] Rong Cao, Liang Bao, Chase Wu, Panpan Zhangsun, Yufei Li, and Zhe Zhang. CM-CASL: Comparison-based performance modeling of software systems via collaborative active and semisupervised learning. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300081X.

Cao:2024:EEC

[CBZZ24] Rong Cao, Liang Bao, Kaibi Zhao, and Panpan Zhangsun. ETune: Efficient configuration tuning for big-data software systems via configuration space reduction. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300331X.

Cai:2024:CTC

[CCC^{+}24] Zhilong Cai, Yongwei Cai, Xiang Chen, Guilong Lu, Wenlong Pei, and Junjie Zhao. CSVD-TF: Cross-project software vulnerability detection with TrAdaBoost by fusing expert metrics and semantic metrics. The Journal of Systems and Software, 213(??):??, July 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000815.

Cheng:2024:VDF

[CCCW24] Jianxin Cheng, Yizhou Chen, Yongzhi Cao, and Hanpin Wang. A vulnerability detection framework with enhanced graph feature learning. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121224001638.

Chowdhury:2024:FCF

[CCR24] Md Towhidul Absar Chowdhury, Maheen Riaz Contractor, and Carlos R. Rivero. Flexible control flow graph alignment for delivering datadriven feedback to novice programming learners. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000037.

Campeanu:2020:CBD

[CCS20] Gabriel Campeanu, Jan Carlson, and Séverine Sentilles. Component-based development of embedded systems with GPUs. The Journal of Systems and Software, 161(??): ??, March 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121219302626.

Chen:2024:OAL

[CCZ^{+}24] Jinfu Chen, Jingyi Chen, Lili Zhu, Chengying Mao, Qihao Bao, and Rubing Huang. L′OP-ART: a lineartime adaptive random testing algorithm for objectoriented programs. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122400013X.

Calefato:2021:GSE

[CDET21] Fabio Calefato, Alpana Dubey, Christof Ebert, and Paolo Tell. Global software engineering: challenges and solutions. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121220302776.

Cortellessa:2022:MDA

[CDET22] Vittorio Cortellessa, Daniele Di Pompeo, Romina Eramo, and Michele Tucci. A modeldriven approach for continuous performance engineering in microservicebased systems. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001813.

Cotroneo:2021:EAS

[CDLN21] Domenico Cotroneo, Luigi De Simone, Pietro Liguori, and Roberto Natella. Enhancing the analysis of software failures in cloud computing systems with deep learning. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001400.

Chidambaram:2025:BIM

[CDM25] Natarajan Chidambaram, Alexandre Decan, and Tom Mens. A bot identification model and tool based on GitHub activity sequences. The Journal of Systems and Software, 221(??):??, March 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003315.

Cotroneo:2022:SMR

[CDN^{+}22] Domenico Cotroneo, Luigi De Simone, Roberto Natella, Roberto Pietrantuono, and Stefano Russo. Software micro-rejuvenation for Android mobile systems. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002636.

Cinque:2024:TIA

[CDO24] Marcello Cinque, Luigi De Simone, and Daniele Ottaviano. Temporal isolation assessment in virtualized safety-critical mixed-criticality systems: a case study on Xen hypervisor. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001924.

Chouchani:2022:MBS

[CDP22] Nadia Chouchani, Sana Debbech, and Matthieu Perin. Model-based safety engineering for autonomous train map. The Journal of Systems and Software, 183(??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121221001795.

Casola:2020:NSD

[CDRV20] Valentina Casola, Alessandra De Benedictis, Massimiliano Rak, and Umberto Villano. A novel Security-by-Design methodology: Modeling and assessing security by SLAs with a quantitative approach. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300042.

Cazzola:2023:LMP

[CF23] Walter Cazzola and Luca Favalli. The language mutation problem: Leveraging language product lines for mutation testing of interpreters. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002096.

Cazzola:2024:SMP

[CF24] Walter Cazzola and Luca Favalli. Software modernization powered by dynamic language product lines. The Journal of Systems and Software, 218 (??):??, December 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002322.

Coppola:2025:KAA

[CFAT25] Riccardo Coppola, Tommaso Fulcini, Luca Ardito, and Marco Torchiano. Kotlin assimilating the Android ecosystem: an appraisal of diffusion and impact on maintainability. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.

sciencedirect.com/
science/article/
pii/S0164121225000147.

Chen:2023:SCV

[CFF^{+}23] Da Chen, Lin Feng, Yuqi Fan, Siyuan Shang, and Zhenchun Wei. Smart contract vulnerability detection based on semantic graph and residual graph convolutional networks with edge attention. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001000.

Cotroneo:2024:ACA

[CFI^{+}24] Domenico Cotroneo, Alessio Foggia, Cristina Improta, Pietro Liguori, and Roberto Natella. Automating the correctness assessment of AI-generated code for security contexts. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001584.

Coppola:2025:RAS

[CFNA25] Riccardo Coppola, Robert Feldt, Michel Nass, and Emil Alégroth. Ranking approaches for similarity-based web element location. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003303.

Corradini:2021:FAA

[CFP^{+}21] Flavio Corradini, Fabrizio Fornari, Andrea Polini, Barbara Re, Francesco Tiezzi, and Andrea Vandin. A formal approach for the analysis of BPMN collaboration models. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001047.

Chinchay:2024:UIE

[CGM24] Yussy Chinchay, Javier Gomez, and Germán Montoro. Unlocking inclusive education: a quality assessment of software design in applications for children with autism. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002097.

Chen:2023:SAC

[CHLT23] Yu-Fang Chen, Vojtěch Havlena, Ondřej Lengál, and Andrea Turrini. A symbolic algorithm for the case-split rule in solving word constraints with extensions. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000687.

Coppa:2024:TCE

[CI24] Emilio Coppa and Alessio Izzillo. Testing concolic execution through consistency checks. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400044X.

Cramer:2025:EPT

[CID25] Geoffrey Cramer, William P. Maxam III, and James C. Davis. Engineering patterns for trust and safety on social media platforms: a case study of Mastodon and Diaspora. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003662.

Cico:2021:EIB

[CJNDZ21] Orges Cico, Letizia Jaccheri, Anh Nguyen-Duc, and He Zhang. Exploring the intersection between software industry and software engineering education — a systematic mapping of software engineering trends. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301667.

Camilli:2022:ATB

[CJR22] Matteo Camilli, Andrea Janes, and Barbara Russo. Automated test-based learning and verification of performance models for microservices systems. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000061.

Chen:2020:EEB

[CJZ^{+}20] Zhengzhao Chen, Renhe Jiang, Zejun Zhang, Yu Pei, Minxue Pan, Tian Zhang, and Xuandong Li. Enhancing example-based code search with functional semantics. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300492.

Chen:2020:ASV

[CKM^{+}20] Jinfu Chen, Patrick Kwaku Kudjo, Solomon Mensah, Selasie Aformaley Brown, and George Akorfu. An automatic software vulnerability classification framework using term frequency-inverse gravity moment and feature selection. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300947.

Chatterjee:2020:FHP

[CKP20] Preetha Chatterjee, Minji Kong, and Lori Pollock. Finding help with programming errors: an exploratory study of novice software engineers’ focus in Stack Overflow posts. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302286.

Chan:2025:ESM

[CKY25] Pak Yuen Patrick Chan, Jacky Keung, and Zhen Yang. Effectiveness of symmetric metamorphic relations on validating the stability of code generation LLM. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003741.

Chen:2023:BES

[CLC^{+}23] Jinfu Chen, Wei Lin, Saihua Cai, Yemin Yin, Haibo Chen, and Dave Towey. BiTCN DRSN: an effective software vulnerability detection model based on an improved temporal convolutional network. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122300167X.

Chen:2024:RPA

[CLLJ24] Zhengyin Chen, Jialong Li, Nianyu Li, and Wenpin Jiao. Reliable proactive adaptation via prediction fusion and extended stochastic model predictive control. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002115.

Colavito:2024:IDQ

[CLNQ24] Giuseppe Colavito, Filippo Lanubile, Nicole Novielli, and Luigi Quaranta. Impact of data quality for automatic issue classification using pre-trained language models. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002339.

Cai:2023:CSJ

[CLZ^{+}23] Jie Cai, Bin Li, Jiale Zhang, Xiaobing Sun, and Bing Chen. Combine sliced joint graph with graph neural networks for smart contract vulnerability detection. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002266.

Cai:2024:FGS

[CLZ^{+}24] Jie Cai, Bin Li, Tao Zhang, Jiale Zhang, and Xiaobing Sun. Fine-grained smart contract vulnerability detection by heterogeneous code feature learning and automated dataset construction. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S016412122300314X.

Corradini:2020:CCB

[CMP^{+}20] Flavio Corradini, Andrea Morichetta, Andrea Polini, Barbara Re, Lorenzo Rossi, and Francesco Tiezzi. Correctness checking for BPMN collaborations with sub-processes. The Journal of Systems and Software, 166(??):??, August 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300716.

Chen:2022:CQE

[CND22] Hsi-Min Chen, Bao-An Nguyen, and Chyi-Ren Dow. Codequality evaluation scheme for assessment of student contributions to programming projects. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121222000358.

Canizares:2020:MEM

[CNdLL20] Pablo C. Cañizares, Alberto Núñez, Juan de Lara, and Luis Llana. MT-EA4Cloud: a methodology for testing and optimising energy-aware cloud systems. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300054.

Choetkiertikul:2023:SAB

[CPC^{+}23] Morakot Choetkiertikul, Arada Puengmongkolchaikit, Pandaree Chandra, Chaiyong Ragkhitwetsagul, Rungroj Maipradit, Hideaki Hata, Thanwadee Sunetnanta, and Kenichi Matsumoto. Studying the association between Gitcoin’s issues and resolving outcomes. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002303.

Costa:2020:TAO

[CPD20] Bruno Costa, Paulo F.

Pires, and Flávia C. Delicato. Towards the adoption of OMG standards in the development of SOA-based IoT systems. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301588.

Chi:2020:RBT

[CQZ^{+}20] Jianlei Chi, Yu Qu, Qinghua Zheng, Zijiang Yang, Wuxia Jin, Di Cui, and Ting Liu. Relationbased test case prioritization for regression testing. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300212.

Cucolas:2023:IWH

[CR23] Adrian-Alexandru Cucolaş and Daniel Russo. The impact of working from home on the success of Scrum projects: a multimethod study. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002382.

Cruz:2023:STP

[CRV23] Victor Pezzi Gazzinelli Cruz, Henrique Rocha, and Marco Tulio Valente. Snapshot testing in practice: Benefits and drawbacks. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001929.

Carberry:2024:RTR

[CRX24] Joshua R. Carberry, John Rahme, and Haiping Xu. Realtime rejuvenation scheduling for cloud systems with virtualized software spares. The Journal of Systems and Software, 217(??): ??, November 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224002139.

Cotroneo:2023:RTF

[CSLN23] Domenico Cotroneo, Luigi De Simone, Pietro Liguori, and Roberto Natella. Run-time failure detection via nonintrusive event analysis in a largescale cloud computing platform. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000067.

Chueca:2023:CSP

[CTD^{+}23] Jorge Chueca, Jose Ignacio Trasobares, África Domingo, Lorena Arcega, Carlos Cetina, and Jaime Font. Comparing software product lines and Clone and Own for game software engineering under two paradigms: Model-driven development and codedriven development. The Journal of Systems and Software, 205(??):??, November 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223002194.

Castellanos:2021:ADS

[CVC21] Camilo Castellanos, Carlos A. Varela, and Dario Correal. ACCORDANT: a domain specificmodel and DevOps approach for big data analytics architectures. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302594.

Camara:2023:EEA

[CWGS23] Javier Cámara, Rebekka Wohlrab, David Garlan, and Bradley Schmerl. ExTrA: Explaining architectural design tradeoff spaces via dimensionality reduction. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002540.

Chai:2021:RVT

[CWTL21] Ming Chai, Haifeng Wang, Tao Tang, and Hongjie Liu. Runtime verification of train control systems with parameterized modal live sequence charts. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000595.

Chen:2024:ESC

[CXH^{+}24] Xiangping Chen, Furen Xu, Yuan Huang, Xiaocong Zhou, and Zibin Zheng. An empirical study of code reuse between GitHub and Stack Overflow during software development. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000074.

Chen:2023:BMO

[CXP^{+}23] Xiang Chen, Hongling Xia, Wenlong Pei, Chao Ni, and Ke Liu. Boosting multi-objective just-in-time software defect prediction by fusing expert metrics and semantic metrics. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002480.

Chen:2023:MIG

[CXY^{+}23] Wenjie Chen, Zimiao Xie, Pengxin Yuan, Ruolin Wang, Hongwei Chen, and Bo Xiao. A mobile intelligent guide system for visually impaired pedestrian. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002229.

Cheng:2021:TCA

[CYW21] Guoli Cheng, Shi Ying, and Bingming Wang. Tuning configuration of Apache Spark on public clouds by combining multi-objective optimization and performance prediction model. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001254.

Chen:2022:ARS

[CZLN22] Fangwei Chen, Li Zhang, Xiaoli Lian, and Nan Niu. Automatically recognizing the semantic elements from UML class diagram images. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001340.

Cheng:2020:DKS

[CZW20] Huanyu Cheng, Ming Zhong, and Jian Wang. Diversified keyword search based web service composition. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300224.

Cabral:2021:EEE

[dACO21] Jose Thiago H.

de A. Cabral and Adriano L. I. Oliveira. Ensemble Effort Estimation using dynamic selection. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000017.

Cabral:2023:EEE

[dACOdS23] José Thiago H.

de A. Cabral, Adriano L. I. Oliveira, and Fabio Q. B. da Silva. Ensemble Effort Estimation: an updated and extended systematic literature review. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002187.

David:2023:CMD

[DAML23] Istvan David, Kousar Aslam, Ivano Malavolta, and Patricia Lago. Collaborative model-driven software engineering — a systematic survey of practices and needs in industry. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000213.

Dadkhah:2020:SLR

[DAP20] Mahboubeh Dadkhah, Saeed Araban, and Samad Paydar. A systematic literature review on semantic web enabled software testing. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302596.

Do:2020:CCR

[DBB20] Quoc Anh Do, Tanmay Bhowmik, and Gary L. Bradshaw. Capturing creative requirements via requirements reuse: a machine learningbased approach. The Journal of Systems and Software, 170(??):??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220301631.

Silva:2024:DSC

[DBM^{+}24] Léuson Da Silva, Paulo Borba, Toni Maciel, Wardah Mahmood, Thorsten Berger, João Moisakis, Aldiberg Gomes, and Vinícius Leite. Detecting semantic conflicts with unit tests. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001158.

Dzhusupova:2024:CRP

[DBO24] Rimma Dzhusupova, Jan Bosch, and Helena Holmström Olsson. Choosing the right path for AI integration in engineering companies: a strategic guide. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003400.

Desdentado:2024:ETB

[DCÁ^{+}24] Elena Desdentado, Coral Calero, Ma. Ángeles Moraga, Manuel Serrano, and Félix García. Exploring the tradeoff between computational power and energy efficiency: an analysis of the evolution of quantum computing and its relation to classical computing. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002103.

daCosta:2022:EUS

[dCMM^{+}22] Francisco Handrick da Costa, Ismael Medeiros, Thales Menezes, João Victor da Silva, Ingrid Lorraine da Silva, Rodrigo Bonifácio, Krishna Narasimhan, and Márcio Ribeiro. Exploring the use of static and dynamic analysis to improve the performance of the mining sandbox approach for Android malware identification. The Journal of Systems and Software, 183(??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121221001898.

Decan:2020:GFC

[DCMR20] Alexandre Decan, Eleni Constantinou, Tom Mens, and Henrique Rocha. GAP: Forecasting commit activity in git projects. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300546.

Daoud:2020:MAJ

[DD20]  Houssem  Daoud

and Michel Dagenais. Multilevel analysis of the Java Virtual Machine based on kernel and userspace traces. The Journal of Systems and Software, 167(??):??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220300698.

Das:2024:EGM

[DDCC24] Souvick Das, Novarun Deb, Agostino Cortesi, and Nabendu Chaki. Extracting goal models from natural language requirement specifications. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000244.

deDieu:2023:CAR

[dDLSK23] Musengamana Jean de Dieu, Peng Liang, Mojtaba Shahin, and Arif Ali Khan. Characterizing architecture related posts and their usefulness in Stack Overflow. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000031.

DAngelo:2024:UGG

[DdM^{+}24] Andrea D’Angelo, Giordano d’Aloisio, Francesca Marzi, Antinisca Di Marco, and Giovanni Stilo. Uncovering gender gap in academia: a comprehensive analysis within the software engineering community. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002073.

DeAngelis:2023:WMT

[DDPP23] Emanuele De Angelis, Guglielmo De Angelis, Alessandro Pellegrini, and Maurizio Proietti. What makes test programs similar in microservices applications? The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121223000699.

Palma:2020:TCS

[DDPT20] Stefano Dalla Palma, Dario Di Nucci, Fabio Palomba, and Damian Andrew Tamburri. Toward a catalog of software quality metrics for infrastructure code. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301618.

DAvila:2020:ECI

[DFB20] Leandro Ferreira D’Avila, Kleinner Farias, and Jorge Luis Victória Barbosa. Effects of contextual information on maintenance effort: a controlled experiment. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302171.

Duan:2023:DLS

[DFC^{+}23] Guoyun Duan, Yuanzhi Fu, Minjie Cai, Hao Chen, and Jianhua Sun. DongTing: a large-scale dataset for anomaly detection of the Linux kernel. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001401.

DeLuca:2024:IRL

[DFT24] Marco De Luca, Anna Rita Fasolino, and Porfirio Tramontana. Investigating the robustness of locators in templatebased Web application testing using a GUI change classification model. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003278.

Damiani:2023:VM

[DHK^{+}23] Ferruccio Damiani, Reiner Hähnle, Eduard Kamburjan, Michael Lienhardt, and Luca Paolini. Variability modules. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001868.

Dhungana:2021:MFP

[DHM^{+}21] Deepak Dhungana, Alois Haselböck, Sebastian Meixner, Daniel Schall, Johannes Schmid, Stefan Trabesinger, and Stefan Wallner. Multi-factory production planning using edge computing and IIoT platforms. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001801.

Dong:2024:EHP

[DHZ^{+}24] Zeming Dong, Qiang Hu, Zhenya Zhang, Yuejun Guo, Maxime Cordy, Mike Papadakis, Yves Le Traon, and Jianjun Zhao. On the effectiveness of hybrid pooling in mixup-based graph learning for language processing. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001845.

Dalibor:2022:CDS

[DJR^{+}22] Manuela Dalibor, Nico Jansen, Bernhard Rumpe, David Schmalzing, Louis Wachtmeister, Manuel Wimmer, and Andreas Wortmann. A cross-domain systematic mapping study on software engineering for digital twins. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000917.

Dehury:2022:TMD

[DJS^{+}22] Chinmaya Kumar Dehury, Pelle Jakovits, Satish Narayana Srirama, Giorgos Giotis, and Gaurav Garg. TOSCAdata: Modeling data pipeline applications in TOSCA. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002508.

Dey:2021:MRS

[DL21] Sangeeta Dey and Seok-Won Lee. Multilayered review of safety approaches for machine learningbased systems in the days of AI. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000388.

Deckers:2022:SLR

[DL22] Robert Deckers and Patricia Lago. Systematic literature review of domainoriented specification techniques. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001261.

Drawel:2022:TBM

[DLBE22] Nagat Drawel, Amine Laarej, Jamal Bentahar, and Mohamed El Menshawy. Transformationbased model checking temporal trust in multi-agent systems. The Journal of Systems and Software, 192(??): ??, October 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222001066.

Ding:2025:PDN

[DLCH25] Zhijun Ding, Shuo Li, Cheng Chen, and Cong He. Program Dependence Net and on-demand slicing for property verification of concurrent system and software. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002656.

DiSorbo:2022:PGC

[DLV^{+}22] Andrea Di Sorbo, Sonia Laudanna, Anna Vacca, Corrado A. Visaggio, and Gerardo Canfora. Profiling gas consumption in solidity smart contracts. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002697.

delaVara:2021:ACC

[dlVRB21] Jose Luis de la Vara, Alejandra Ruiz, and Gaël Blondelle. Assurance and certification of cyberphysical systems: the AMASS open source ecosystem. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302120.

Dai:2023:GCF

[DLX^{+}23] Jie Dai, Qingshan Li, Hui Xue, Zhao Luo, Yinglin Wang, and Siyuan Zhan. Graph collaborative filtering-based bug triaging. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000626.

Diaz:2024:HCF

[DMAH24] Oscar Díaz, Raul Medeiros, and Mustafa Al-Hajjaji. How can feature usage be tracked across product variants? Implicit feedback in software product lines. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000566.

Decan:2023:OWG

[DMD23] Alexandre Decan, Tom Mens, and Hassan Onsori Delicheh. On the outdatedness of workflows in the GitHub Actions ecosystem. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002224.

Dakhel:2023:GCA

[DMN^{+}23] Arghavan Moradi Dakhel, Vahid Majdinasab, Amin Nikanjam, Foutse Khomh, Michel C. Desmarais, and Zhen Ming (Jack) Jiang. GitHub Copilot AI pair programmer: Asset or liability? The Journal of Systems and Software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223001292.

Davila:2021:SLR

[DN21] Nicole Davila and Ingrid Nunes. A systematic literature review and taxonomy of modern code review. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000480.

Daaji:2023:BPD

[DOG^{+}23] Marwa Daaji, Ali Ouni, Mohamed Mohsen Gammoudi, Salah Bouktif, and Mohamed Wiem Mkaouer. BPEL process defects prediction using multi-objective evolutionary search. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001620.

Rosa:2023:CEM

[dORGCG23] Thatiane de Oliveira Rosa, Eduardo Martins Guerra, Filipe Figueiredo Correia, and Alfredo Goldman. CharM — evaluating a model for characterizing service-based architectures. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002212.

Diaz:2024:HDT

[DPA^{+}24] Jessica Díaz, Jorge Pérez, Isaque Alves, Fabio Kon, Leonardo Leite, Paulo Meirelles, and Carla Rocha. Harmonizing DevOps taxonomies

— a grounded theory study. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003035.

DeStefano:2022:SEQ

[DPD^{+}22] Manuel De Stefano, Fabiano Pecorelli, Dario Di Nucci, Fabio Palomba, and Andrea De Lucia. Software engineering for quantum programming: How far are we? The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000693.

Diaz:2023:AIR

[DPGGP23] Jessica Díaz, Jorge Pérez, Carolina Gallardo, and Ángel González-Prieto. Applying Inter-Rater Reliability and Agreement in collaborative Grounded Theory studies in software engineering. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001960.

Porto:2021:ICU

[dPPdJFF21] Daniel de Paula Porto, Gabriela Martins de Jesus, Fabiano Cutigi Ferrari, and Sandra Camargo Pinto Ferraz Fabbri. Initiatives and challenges of using gamification in software engineering: a systematic mapping. The Journal of Systems and Software, 173(??):??, March 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302600.

Diaz:2021:FEG

[DPRD21] Eduardo Díaz, José Ignacio Panach, Silvia Rueda, and Damiano Distante. A family of experiments to generate graphical user interfaces from BPMN models with stereotypes. The Journal of Systems and Software, 173(??): ??, March 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220302739.

Derks:2023:BGF

[DSB23] Christoph Derks, Daniel Strüber, and Thorsten Berger. A benchmark generator framework for evolving variantrich software. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001310.

deSouza:2025:SRF

[dSFdC^{+}25] Kathiani Elisa de Souza, Fabiano Cutigi Ferrari, Valter Vieira de Camargo, Márcio Ribeiro, and Jeff Offutt. A systematic review of fault tolerance techniques for smart city applications. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002930.

delSagrado:2023:EDA

[dSIdÁ23] José del Sagrado, José Antonio Sierra Ibañez, and Isabel M. del Águila. An estimation of distribution algorithm based on interactions between requirements to solve the bi-objective Next Release Problem. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000274.

Dowdeswell:2020:FFS

[DSM20] Barry Dowdeswell, Roopak Sinha, and Stephen G. Mac-Donell. Finding faults: a scoping study of fault diagnostics for industrial cyber-physical systems. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301114.

dosSantos:2025:DPM

[dSNJ25]  Eduardo  Witter

dos Santos, Ingrid Nunes, and Dietmar Jannach. Developer perceptions of modern code review processes in practice: Insights from a case study in a mid-sized company. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003327.

Dihego:2020:RCB

[DSO20] José Dihego, Augusto Sampaio, and Marcel Oliveira. A refinement checking based strategy for componentbased systems evolution. The Journal of Systems and Software, 167(??): ??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220300765.

deSLandi:2022:ACC

[dSS^{+}22] André de S.Landi, Daniel San Martín, Bruno M. Santos, Warteruzannan S. Cunha, Rafael S. Durelli, and Valter V. Camargo. Architectural conformance checking for KDM-represented systems. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002132.

deToledo:2021:IAT

[dTMS21] Saulo S. de Toledo, Antonio Martini, and Dag I. K. Sjøberg. Identifying architectural technical debt, principal, and interest in microservices: a multiple-case study. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000650.

Duan:2022:NLB

[DTZ^{+}22] Zhenhua Duan, Cong Tian, Nan Zhang, Mengchu Zhou, Bin Yu, Xiaobing Wang, Jiangen Guo, and Ying Wu. A novel load balancing scheme for mobile edge computing. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002703.

Dutta:2024:LFU

[Dut24] Sudakshina Dutta.

Localizing faults using verification technique. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002923.

Dubslaff:2024:FC

[DWBA24] Clemens Dubslaff, Kallistos Weis, Christel Baier, and Sven Apel. Feature causality. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003102.

Dang:2024:TDG

[DWG^{+}24] Xiangying Dang, Jinyong Wang, Dunwei Gong, Xiangjuan Yao, Changqing Wei, and Biao Xu. Test data generation for covering mutation-based path using MGA for MPI program. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000050.

Dobslaw:2023:GIS

[DWH23] Felix Dobslaw, Ruiyuan Wan, and Yuechan Hao. Generic and industrial scale many-criteria regression test selection. The Journal of Systems and Software, 205(??): ??, November 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223001978.

Dai:2025:ICP

[DXD25] Hongming Dai, Jianqing Xi, and Hong-Liang Dai. Improve cross-project just-in-time defect prediction with dynamic transfer learning. The Journal of Systems and Software, 219(??): ??, January 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224002589.

Deng:2024:CSE

[DXL^{+}24] Zhongyang Deng, Ling Xu, Chao Liu, Luwen Huangfu, and Meng Yan. Code semantic enrichment for deep code search. The Journal of Systems and Software, 207 (??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002510.

Deng:2023:VSC

[DYZ^{+}23] Xi Deng, Jiwei Yan, Shixin Zhang, Jun Yan, and Jian Zhang. Variablestrength combinatorial testing of exported activities based on misexposure prediction. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001681.

Ding:2023:ITO

[DZY^{+}23] Yanru Ding, Yanmei Zhang, Guan Yuan, Shujuan Jiang, Wei Dai, and Yinghui Zhang. Integration test order generation based on reinforcement learning considering class importance. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002182.

Etemadi:2021:SDA

[EBAR21] Vahid Etemadi, Omid Bushehrian, Reza Akbari, and Gregorio Robles. A scheduling-driven approach to efficiently assign bug fixing tasks to developers. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000649.

Evtikhiev:2023:BHS

[EBSB23] Mikhail Evtikhiev, Egor Bogomolov, Yaroslav Sokolov, and Timofey Bryksin. Out of the BLEU: How should we assess quality of the Code Generation models? The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122300136X.

Eismann:2022:CSS

[ECL^{+}22] Simon Eismann, Diego Elias Costa, Lizhi Liao, Cor-Paul Bezemer, Weiyi Shang, André van Hoorn, and Samuel Kounev. A case study on the stability of performance tests for serverless applications. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000498.

Edison:2020:ISS

[ECMC20] Henry Edison, Noel Carroll, Lorraine Morgan, and Kieran Conboy. Inner source software development: Current thinking and an agenda for future research. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300030.

Elyasaf:2023:FAC

[ECS23] Achiya Elyasaf, Nicolás Cardozo, and Arnon Sturm. A framework for analyzing context-oriented programming languages. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223000092.

Echeverria:2021:CSS

[EFPC21] Jorge Echeverría, Jaime Font, Francisca Pérez, and Carlos Cetina. Comparison of search strategies for feature location in software models. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001345.

Ernst:2021:ISI

[EHB21] Neil A. Ernst, Mark Hills, and Árpád Beszédes. Introduction to the Special Issue on Source Code Analysis and Manipulation 2018. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301497.

Ernst:2023:ACL

[EKB^{+}23] Neil A. Ernst, John Klein, Marco Bartolini, Jeremy Coles, and Nick Rees. Architecting complex, long-lived scientific software. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001279.

Eraslan:2020:IGM

[EKHJ^{+}20] Sukru Eraslan, Kamilla Kopec-Harding, Caroline Jay, Suzanne M. Embury, Robert Haines, Julio César Cortés Ríos, and Peter Crowther. Integrating GitLab metrics into coursework consultation sessions in a software engineering course. The Journal of Systems and Software, 167(??):??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220300911.

Enaya:2024:CSD

[EKK^{+}24] Mohamad Fawaz Enaya, Thomas Klingbeil, Jacob Krüger, David Broneske, Frank Feinbube, and Gunter Saake. A case study on the development of the German Corona-Warn-App. The Journal of Systems and Software, 213(??):??, July 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000633.

Erradi:2020:OCO

[EM20] Abdelkarim Erradi and Yaser Mansouri. Online cost optimization algorithms for tiered cloud storage services. The Journal of Systems and Software, 160(??): ??, February 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219302316.

Eisenecker:2024:SPL

[EM24] Ulrich Eisenecker and Richard Müller. Software product lines and features from the perspective of set theory with an application to feature location.

The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003333.

El-Migid:2022:ETP

[EMCN^{+}22] Mohammed-Amr Abd El-Migid, Damon Cai, Thomas Niven, Jeffrey Vo, Kashumi Madampe, John Grundy, and Rashina Hoda. Emotimonitor: a Trello power-up to capture and monitor emotions of Agile teams. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100279X.

Ettazi:2023:TCE

[EN23] Widad Ettazi and Mahmoud Nassar. Towards a cognitive engineering of transactional services in IoT based systems. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000298.

Eskandani:2023:UJF

[ES23] Nafise Eskandani and Guido Salvaneschi. The uphill journey of FaaS in the opensource community. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002655.

Arnaoty:2024:ODC

[ES24] Mohammed El Arnaoty and Francisco Servant. OneSpace : Detecting cross-language clones by learning a common embedding space. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003060.

Eramo:2024:AMB

[ESO^{+}24] Romina Eramo, Bilal Said, Marc Oriol, Hugo Bruneliere, and Sergio Morales. An architecture for model-based and intelligent automation in DevOps. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002255.

Eken:2021:IPP

[ET21] Beyza Eken and Ayse Tosun. Investigating the performance of personalized models for software defect prediction. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001357.

Eramo:2024:ASS

[ETD^{+}24] Romina Eramo, Michele Tucci, Daniele Di Pompeo, Vittorio Cortellessa, Antinisca Di Marco, and Davide Taibi. Architectural support for software performance in continuous software engineering: a systematic mapping study. The Journal of Systems and Software, 207(??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002285.

Etemadi:2022:EPP

[ETY^{+}22] Khashayar Etemadi, Niloofar Tarighat, Siddharth Yadav, Matias Martinez, and Martin Monperrus. Estimating the potential of program repair search spaces with commit analysis. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000309.

Flora:2024:EID

[FA24] José Flora and Nuno Antunes. Evaluating intrusion detection for microservice applications: Benchmark, dataset, and case studies. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001870.

Fucci:2022:WTG

[FAA22] Davide Fucci, Emil Alégroth, and Thomas Axelsson. When traceability goes awry: an industrial experience report. The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001091.

Fucci:2024:ESS

[FAFJ24] Davide Fucci, Emil Alégroth, Michael Felderer, and Christoffer Johannesson. Evaluating software security maturity using OWASP SAMM: Different approaches and stakeholders perceptions. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001079.

Feitosa:2020:CRP

[FAG^{+}20] Daniel Feitosa, Apostolos Ampatzoglou, Antonios Gkortzis, Stamatia Bibi, and Alexander Chatzigeorgiou. CODE reuse in practice: Benefiting or harming technical debt.

The Journal of Systems and Software, 167(??):??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220300960.

Fontana:2020:GEI

[FBMR20] Francesca Arcelli Fontana, Hugo Bruneliere, Hausi Müller, and Claudia Raibulet. Guest editors’ introduction to the special issue on Model Driven Engineering and Reverse Engineering: Research and Practice. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302201.

Ferreira:2024:RRB

[FBV24] Fabio Ferreira, Hudson Silva Borges, and Marco Tulio Valente. Refactoring react-based Web apps. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400150X.

Fellin:2020:EAX

[FC20] Roberto Fellin and Mariano Ceccato. Experimental assessment of XOR-Masking data obfuscation based on K-Clique opaque constants. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302663.

Fan:2023:DSS

[FCW^{+}23] Guodong Fan, Shizhan Chen, Hongyue Wu, Cuiyun Gao, Jianmao Xiao, Xiao Xue, and Zhiyong Feng. Dialog summarization for software collaborative platform via tuning pre-trained models. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001589.

Felizardo:2020:SSA

[FdSN^{+}20] Katia Romero Felizardo, Érica Ferreira de Souza, Bianca Minetto Napoleão, Nandamudi Lankalapalli Vijaykumar, and Maria Teresa Baldassarre. Secondary studies in the academic context: a systematic mapping and survey. The Journal of Systems and Software, 170(??): ??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220301655.

Figalist:2022:BVC

[FEBO22] Iris Figalist, Christoph Elsner, Jan Bosch, and Helena Holmström Olsson. Breaking the vicious circle: a case study on why AI for software analytics and business intelligence does not take off in practice. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002326.

Feitelson:2023:WDA

[Fei23] Dror G. Feitelson. “We do not appreciate being experimented on”: Developer and researcher views on the ethics of experiments on opensource projects. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001693.

Fregnan:2023:GBV

[FFSB23] Enrico Fregnan, Josua Fröhlich, Davide Spadini, and Alberto Bacchelli. Graph-based visualization of merge requests for code review. The Journal of Systems and Software, 195(??): ??, January 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222001820.

Fischbach:2023:ACA

[FFV^{+}23] Jannik Fischbach, Julian Frattini, Andreas Vogelsang, Daniel Mendez, Michael Unterkalmsteiner, Andreas Wehrle, Pablo Restrepo Henao, Parisa Yousefi, Tedi Juricic, Jeannette Radduenz, and Carsten Wiecher. Automatic creation of acceptance tests by extracting conditionals from requirements: NLP approach and case study. The Journal of Systems and Software, 197(??): ??, March 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121222002254.

Firouzi:2024:TSS

[FG24] Ehsan Firouzi and Mohammad Ghafari. Time to separate from Stack-Overflow and match with ChatGPT for encryption. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001808.

Ferenc:2020:ACN

[FGG^{+}20] Rudolf Ferenc, Péter Gyimesi, Gábor Gyimesi, Zoltán Tóth, and Tibor Gyimóthy. An automatically created novel bug dataset and its validation in bug prediction. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121220301436.

Foidl:2024:DPQ

[FGRF24] Harald Foidl, Valentina Golendukhina, Rudolf Ramler, and Michael Felderer. Data pipeline quality: Influencing factors, root causes of datarelated issues, and processing problem areas for developers. The Journal of Systems and Software, 207(??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002509.

Fantechi:2023:VLV

[FGS23]  Alessandro  Fan-

techi, Stefania Gnesi, and Laura Semini. VIBE: Looking for Variability In amBiguous rEquirements. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002163.

Farshidi:2020:CSA

[FJvdW20] Siamak Farshidi, Slinger Jansen, and Jan Martijn van der Werf. Capturing software architecture knowledge for patterndriven design. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301552.

Fischer:2023:THC

[FKF^{+}23] Stefan Fischer, Claus Klammer, Antonio Manuel Gutiérrez Fernández, Rick Rabiser, and Rudolf Ramler. Testing of highly configurable cyber–physical systems — results from a two-phase multiple case study. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000195.

Felipe:2023:PIS

[FKGN23] Danilo Almeida Felipe, Marcos Kalinowski, Daniel Graziotin, and Jean Carlos Natividade. Psychometric instruments in software engineering research on personality: Status quo after fifty years. The Journal of Systems and Software, 203(??): ??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223001358.

Frattini:2024:RQR

[FMF^{+}24] Julian Frattini, Lloyd Montgomery, Davide Fucci, Michael Unterkalmsteiner, Daniel Mendez, and Jannik Fischbach. Requirements quality research artifacts: Recovery, analysis, and management guideline. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001651.

Fadhlillah:2025:MCS

[FMG^{+}25] Hafiyyan Sayyid Fadhlillah, Kristof Meixner, Sandra Greiner, Antonio Manuel Gutiérrez-Fernández, and Rick Rabiser. Managing control software variability in cyber-physical production systems: the V4rdiac approach. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224003698.

Ferreira:2024:FDF

[FNB^{+}24] Francisco Henrique Cerdeira Ferreira, Elisa Yumi Nakagawa, Antonia Bertolino, Francesca Lonetti, Vânia de Oliveira Neves, and Rodrigo Pereira dos Santos. A framework for the design of faulttolerant systemsof-systems. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000530.

Feitosa:2024:MCA

[FPB^{+}24] Daniel Feitosa, Matei-Tudor Penca, Massimiliano Berardi, Rares-Dorian Boza, and Vasilios Andrikopoulos. Mining for cost awareness in the infrastructure as code artifacts of cloud-based applications: an exploratory study. The Journal of Systems and Software, 215(??):??, September 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224001572.

Fatemidokht:2020:QVE

[FR20] Hamideh Fatemidokht and Marjan Kuchaki Rafsanjani. QMM-VANET: an efficient clustering algorithm based on QoS and monitoring of malicious vehicles in vehicular ad hoc networks. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300431.

Ferreira:2024:IDO

[FRC24] Isabella Ferreira, Ahlaam Rafiq, and Jinghui Cheng. Incivility detection in open source code review and issue discussions. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003308.

Freire:2023:SPP

[FRP^{+}23] Sávio Freire, Nicolli Rios, Boris Pérez, Camilo Castellanos, Darío Correal, Robert Ramač, Vladimir Mandić, Nebojša Taušan, Gustavo López, Alexia Pacheco, Manoel Mendonça, Davide Falessi, Clemente Izurieta, Carolyn Seaman, and Rodrigo Spínola. Software practitioners’ point of view on technical debt payment. The Journal of Systems and Software, 196 (??):??, February 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002308.

Florea:2023:RST

[FSS23] Raluca Florea, Viktoria Stray, and Dag I. K. Sjøberg. On the roles of software testers: an exploratory study. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001371.

Frolich:2024:EML

[FvB24] Damian Frölich and L. Thomas van Binsbergen. iCoLa^{+}: an extensible metalanguage with support for exploratory language development. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224000220.

Ferreira:2021:EWTa

[FVDF21a] Fischer Ferreira, Gustavo Vale, João P. Diniz, and Eduardo Figueiredo. Evaluating T-wise testing strategies in a communitywide dataset of configurable software systems. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S016412122100087X.

Ferreira:2021:EWTb

[FVDF21b] Fischer Ferreira, Gustavo Vale, João P. Diniz, and Eduardo Figueiredo. Evaluating T-wise testing strategies in a communitywide dataset of configurable software systems. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100087X.

Feng:2025:DGA

[FWZZ25] Lichao Feng, Xingya Wang, Shiyu Zhang, and Zhihong Zhao. DeepFeature: Guiding adversarial testing for deep neural network systems using robust features. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002450.

Fu:2024:EEL

[FYH^{+}24] Ying Fu, Meng Yan, Pinjia He, Chao Liu, Xiaohong Zhang, and Dan Yang. End-to-end log statement generation at block-level. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001912.

Fang:2022:PAP

[FZT^{+}22] Sen Fang, Tao Zhang, You-Shuai Tan, Zhou Xu, Zhi-Xin Yuan, and Ling-Ze Meng. PRHAN: Automated pull request description generation based on hybrid attention network. The Journal of Systems and Software, 185(??):??, March 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100248X.

Giaimo:2020:CEC

[GAB20] Federico Giaimo, Hugo Andrade, and Christian Berger. Continuous experimentation and the cyber-physical systems challenge: an overview of the literature and the industrial perspective. The Journal of Systems and Software, 170(??):??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412122030193X.

Goncalves:2020:PPS

[GAC20] Enyo Gonçalves, João Araujo, and Jaelson Castro. PRISE: a process to support iStar extensions. The Journal of Systems and Software, 168(??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121220301175.

Guo:2020:MCV

[GAL20] Xiaoyun Guo, Toshiaki Aoki, and Hsin-Hung Lin. Model checking of invehicle networking systems with CAN and FlexRay. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302353.

Gustavsson:2025:JSR

[GAS25] Tomas Gustavsson, Muhammad Ovais Ahmad, and Hina Saeeda. Job satisfaction at risk: Measuring the role of process debt in agile software development. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/

science/article/
pii/S0164121225000184.

Geismann:2020:SLR

[GB20] Johannes Geismann and Eric Bodden. A systematic literature review of model-driven security engineering for cyber-physical systems. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301461.

Giray:2023:UDL

[GBK^{+}23] Görkem Giray, Kwabena Ebo Bennin, Ömer Köksal, Önder Babur, and Bedir Tekinerdogan. On the use of deep learning in software defect prediction. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121222002138.

Guerino:2025:TUS

[GBLP25] Guilherme Corredato Guerino, Renato Balancieri, Gislaine Camila Lapasini Leal, and Rafael Prikladnicki. Towards the UX Support to Software Startups: On the relationship of professional expertise and UX work in the Brazilian scenario. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002905.

Giunta:2022:CTE

[GBMF22] Benito Giunta, Corentin Burnay, Neil Maiden, and Stéphane Faulkner. Creativity Triggers: Extension and empirical evaluation of their effectiveness during requirements elicitation. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000930.

Groz:2020:HIH

[GBSO20] Roland Groz, Nicolas Bremond, Adenilso Simao, and Catherine Oriat. hW-inference: a heuristic approach to retrieve models through black box testing. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302006.

Gregory:2020:LWR

[GBT^{+}20] Joe Gregory, Lucy Berthoud, Theo Tryfonas, Alain Rossignol, and Ludovic Faure. The long and winding road: MBSE adoption for functional avionics of spacecraft. The Journal of Systems and Software, 160 (??):??, February 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302274.

Garousi:2022:MUR

[GCF22] Vahid Garousi, David Cutting, and Michael Felderer. Mining user reviews of COVID contacttracing apps: an exploratory analysis of nine European apps. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002338.

Gharib:2022:CPS

[GCLB22] Mohamad Gharib, Andrea Ceccarelli, Paolo Lollini, and Andrea Bondavalli. A cyber-physical– social approach for engineering Functional Safety Requirements for automotive systems.

The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000577.

Gavidia-Calderon:2020:GTA

[GCSHB20] Carlos Gavidia-Calderon, Federica Sarro, Mark Harman, and Earl T. Barr. Game-theoretic analysis of development practices: Challenges and opportunities. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219301980.

Gou:2024:SBH

[GDK24] Qianwen Gou, Yunwei Dong, and Qiao Ke. Syntho-Minds: Bridging human programming intuition with retrieval, analogy, and reasoning in program synthesis.

The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001857.

Golzadeh:2021:GTD

[GDLM21] Mehdi Golzadeh, Alexandre Decan, Damien Legay, and Tom Mens. A ground-truth dataset and classification model for detecting bots in GitHub issue and PR comments. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100008X.

Grechi:2025:MDS

[GdOB25] Victor Luiz Grechi, André Luiz de Oliveira, and Rosana T. Vaccare Braga. Modeldriven safety and security co-analysis: a systematic literature review. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002954.

Gou:2024:RDH

[GDWK24] Qianwen Gou, Yunwei Dong, Yujiao Wu, and Qiao Ke. RRGcode: Deep hierarchical search-based code generation. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000256.

Gerling:2025:SPP

[Ger25] Lea Gerling. Syntaxpreserving program slicing for C-based software product lines. The Journal of Systems and Software, 219(??): ??, January 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224002991.

Golshanrad:2024:DAR

[GF24] Pouria Golshanrad and Fathiyeh Faghih. Deep-Cover: Advancing RNN test coverage and online error prediction using state machine extraction. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400030X.

Gutierrez-Fernandez:2024:VMS

[GFCLB^{+}24] Antonio M. Gutiérrez-Fernández, Ana Eva Chacón-Luna, David Benavides, Lidia Fuentes, and Rick Rabiser. Variability management and software product line knowledge in software companies. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001596.

Gkortzis:2021:SRC

[GFS21] Antonios Gkortzis, Daniel Feitosa, and Diomidis Spinellis. Software reuse cuts both ways: an empirical analysis of its relationship with security vulnerabilities. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301199.

Gesvindr:2020:ADE

[GGB20] David Gesvindr, Ondrej Gasior, and Barbora Buhnova. Architecture design evaluation of PaaS cloud applications using generated prototypes: PaaSArch Cloud Prototyper tool. The Journal of Systems and Software, 169(??):??, November 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301485.

Gao:2022:SRP

[GGB^{+}22] Hongcan Gao, Chenkai Guo, Guangdong Bai, Dengrong Huang, Zhen He, Yanfeng Wu, and Jing Xu. Sharing runtime permission issues for developers based on similar-app review mining. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002156.

Gortazar:2022:CEL

[GGMB^{+}22] Francisco Gortázar, Micael Gallego, Michel Maes-Bermejo, Iván Chicano-Capelo, and Carlos Santos. Costeffective load testing of WebRTC applications. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001388.

Gunatilake:2023:EMS

[GGMH23]  Hashini  Gunati-

lake, John Grundy, Ingo Mueller, and Rashina Hoda. Empathy models and software engineering — a preliminary analysis and taxonomy. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001425.

Gaitero:2021:SQS

[GGP21] Domingo Gaitero, Marcela Genero, and Mario Piattini. System quality and security certification in seven weeks: a multi-case study in Spanish SMEs. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000571.

Giamattei:2024:MTD

[GGP^{+}24] L. Giamattei, A. Guerriero, R. Pietrantuono, S. Russo, I. Malavolta, T. Islam, M. Dînga, A. Koziolek, S. Singh, M. Armbruster, J. M. Gutierrez-Martinez, S. Caro-Alvaro, D. Rodriguez, S. Weber, J. Henss, E. Fernandez Vogelin, and F. Simon Panojo. Monitoring tools for DevOps and microservices: a systematic grey literature review. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003011.

Giamattei:2024:AFR

[GGPR24] Luca Giamattei, Antonio Guerriero, Roberto Pietrantuono, and Stefano Russo. Automated functional and robustness testing of microservice architectures. The Journal of Systems and Software, 207(??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002522.

Guerra:2024:IIE

[GIP^{+}24] Eduardo Martins Guerra, André A. S. Ivo, Fernando O. Pereira, Romain Robbes, Andrea Janes, and Fábio Fagundes Silveira. Impermanent identifiers: Enhanced source code comprehension and refactoring. The Journal of Systems and Software, 216(??): ??, October 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224001821.

Giray:2021:SEP

[Gir21] Görkem Giray. A software engineering perspective on engineering machine learning systems: State of the art and challenges. The Journal of Systems and Software, 180(??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100128X.

Gao:2022:GSI

[GJW^{+}22] Xuejian Gao, Xue Jiang, Qiong Wu, Xiao Wang, Chen Lyu, and Lei Lyu. GT-SimNet: Improving code automatic summarization via multimodal similarity networks. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001716.

Garcia:2022:SJJ

[GKAHMO22] Boni García, Carlos Delgado Kloos, Carlos Alario-Hoyos, and Mario Munoz-Organero. Selenium-Jupiter: a JUnit 5 extension for Selenium WebDriver. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000516.

Garousi:2021:MBT

[GKB^{+}21] Vahid Garousi, Alper Buğra Keleş, Yunus Balaman, Zeynep Özdemir Güler, and Andrea Arcuri. Modelbased testing in practice: an experience report from the web applications domain. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001291.

Galindo:2024:FSP

[GKPS24] Carlos Galindo, Jens Krinke, Sergio Pérez, and Josep Silva. Field-sensitive program slicing. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003345.

Gemeinhardt:2025:GQP

[GKW25] Felix Gemeinhardt, Stefan Klikovits, and Manuel Wimmer. GeQuPI: Quantum Program Improvement with Multi-Objective Genetic Programming. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400267X.

Glasauer:2023:PMH

[Gla23] Christina Glasauer. The prevent-model:

Human and organizational factors fostering engineering of safe and secure robotic systems. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002242.

Gao:2025:LTL

[GLCH25] Donghui Gao, Changjian Liu, Ningjiang Chen, and Xiaochun Hu. LogGzip: Towards log parsing with lossless compression. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0164121225000172.

Goumopoulos:2020:FPC

[GM20] Christos Goumopoulos and Irene Mavrommati. A framework for pervasive computing applications based on smart objects and end user development. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302705.

Georgiou:2021:ESC

[GMCA21] Konstantinos Georgiou, Nikolaos Mittas, Alexandros Chatzigeorgiou, and Lefteris Angelis. An empirical study of COVID-19 related posts on Stack Overflow: Topics and technologies. The Journal of Systems and Software, 182(??):??, December 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221001862.

Garces:2021:TDSa

[GMFO^{+}21a] Lina Garcés, Silverio Martínez-Fernández, Lucas Oliveira, Pedro Valle, Claudia Ayala, Xavier Franch, and Elisa Yumi Nakagawa. Three decades of software reference architectures: a systematic mapping study. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001011.

Garces:2021:TDSb

[GMFO^{+}21b] Lina Garcés, Silverio Martínez-Fernández, Lucas Oliveira, Pedro Valle, Claudia Ayala, Xavier Franch, and Elisa Yumi Nakagawa. Three decades of software reference architectures: a systematic mapping study. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001011.

Garba:2022:SAM

[GMS22] Salisu Garba, Radziah Mohamad, and Nor Azizah Saadon. Selfadaptive mobile web service discovery framework for Dynamic Mobile Environment. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100217X.

Giordano:2022:UAI

[GPF22] Giammaria Giordano, Fabio Palomba, and Filomena Ferrucci. On the use of artificial intelligence to deal with privacy in IoT systems: a systematic literature review. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001613.

Gonzalez-Prieto:2023:RSE

[GPPDLF23] Ángel González-Prieto, Jorge Perez, Jessica Diaz, and Daniel López-Fernández. Reliability in software engineering qualitative research through Inter-Coder Agreement. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001024.

Garcia:2023:EWA

[GRdAL23] Boni García, Filippo Ricca, Jose M. del Alamo, and Maurizio Leotta. Enhancing Web applications observability through instrumented automated browsers. The Journal of Systems and Software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223001188.

Garousi:2020:STE

[GRLA20] Vahid Garousi, Austen Rainer, Per Lauvås, and Andrea Arcuri. Softwaretesting education: a systematic literature mapping. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300510.

Gualo:2021:DQC

[GRV^{+}21] Fernando Gualo, Moisés Rodriguez, Javier Verdugo, Ismael Caballero, and Mario Piattini. Data quality certification using ISO/ IEC 25012: Industrial experiences. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000352.

Georgiou:2020:EDI

[GS20] Stefanos Georgiou and Diomidis Spinellis. Energy-delay investigation of remote inter-process communication technologies. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302808.

Gokceoglu:2021:ADPa

[GS21a] Mustafa Gökçeoğlu and Hasan Sözer. Automated defect prioritization based on defects resolved at various project periods. The Journal of Systems and Software, 179(??): ??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412122100090X.

Gokceoglu:2021:ADPb

[GS21b] Mustafa Gökçeoğlu and Hasan Sözer. Automated defect prioritization based on defects resolved at various project periods. The Journal of Systems and Software, 179(??): ??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412122100090X.

Gharibi:2024:TEA

[GSF24] Reza Gharibi, Mohammad Hadi Sadreddini, and Seyed Mostafa Fakhrahmad. T5APR: Empowering automated program repair across languages through checkpoint ensemble. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001286.

Gnoyke:2024:EPS

[GSK24]  Philipp  Gnoyke,

Sandro Schulze, and Jacob Krüger. Evolution patterns of software-architecture smells: an empirical study of intra- and interversion smells. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002152.

Go:2023:SSI

[GSM^{+}23] Ken Russel Go, Sruthi Soundarapandian, Aparupa Mitra, Melina Vidoni, and Nicolás E. Díaz Ferreyra. Simple stupid insecure practices and GitHub’s code search: a looming threat? The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000936.

Gill:2020:TDL

[GTT^{+}20] Sukhpal Singh Gill, Shreshth Tuli, Adel Nadjaran Toosi, Felix Cuadrado, Peter Garraghan, Rami Bahsoon, Hanan Lutfiyya, Rizos Sakellariou, Omer Rana, Schahram Dustdar, and Rajkumar Buyya. ThermoSim: Deep learning based framework for modeling and simulation of thermal-aware resource management for cloud computing environments. The Journal of Systems and Software, 166(??):??, August 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300753.

Gu:2023:DHC

[Gu23] Zuguang Gu. On the dependency heaviness of CRAN/ Bioconductor ecosystem. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223000055.

Guevara-Vega:2024:RAH

[GVBC^{+}24]  Cathy  Guevara-

Vega, Beatriz Bernárdez, Margarita Cruz, Amador Durán, Antonio Ruiz-Cortés, and Martin Solari. Research artifacts for human-oriented experiments in software engineering: an ACM badgesdriven structure proposal. The Journal of Systems and Software, 218 (??):??, December 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002310.

Galster:2023:ERS

[GW23] Matthias Galster and Danny Weyns. Empirical research in software architecture — perceptions of the community. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121223000791.

Goodwin:2024:BDL

[GW24] Craig Goodwin and Sandra Woolley. Barriers to device longevity and reuse: a vintage device empirical study. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000347.

Herbold:2021:SMS

[HATG21] Steffen Herbold, Aynur Amirfallah, Fabian Trautsch, and Jens Grabowski. A systematic mapping study of developer social network research. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302077.

Horcas:2025:LBU

[HBK25] Jose-Miguel Horcas, Lola Burgueño, and Jörg Kienzle. Leveraging belief uncertainty for informed decision making in software product line evolution. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002796.

Ho:2025:EDE

[HBN^{+}25] Anh Ho, Anh M. T.

Bui, Phuong T. Nguyen, Amleto Di Salle, and Bach Le. EnseSmells: Deep ensemble and programming language models for automated code smells detection. The Journal of Systems and Software, 224(??):??, June 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000433.

Harrand:2022:ABE

[HBSV^{+}22] Nicolas Harrand, Amine Benelallam, César Soto-Valero, François Bettega, Olivier Barais, and Benoit Baudry. API beauty is in the eye of the clients: 2.2 million Maven dependencies reveal the spectrum of client–API usages. The Journal of Systems and Software, 184(??):??, February 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221002314.

Hachem:2020:MAP

[HCB^{+}20] Jamal EL Hachem, Vanea Chiprianov, Muhammad Ali Babar, Tarek AL Khalil, and Philippe Aniorte. Modeling, analyzing and predicting security cascading attacks in smart buildings systemsof-systems. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302584.

Huang:2022:SRG

[HCC22] Yeu-Shiang Huang, Kuei-Chen Chiu, and Wan-Ming Chen. A software reliability growth model for imperfect debugging. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000322.

Han:2023:LFL

[HDX^{+}23] Yongqi Han, Qingfeng Du, Jincheng Xu, Shengjie Zhao, Zhekang Chen, Li Cao, Kanglin Yin, and Dan Pei. LWS: a framework for logbased workload simulation in sessionbased SUT. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001309.

Heinrich:2020:ARM

[Hei20] Robert Heinrich. Architectural runtime models for integrating runtime observations and componentbased models. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122030159X.

Honel:2020:USC

[HELW20] Sebastian Hönel, Morgan Ericsson, Welf Löwe, and Anna Wingkvist. Using source code density to improve the accuracy of automatic commit classification into maintenance activities. The Journal of Systems and Software, 168(??): ??, October 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220301291.

Hejderup:2022:CWT

[HG22] Joseph Hejderup and Georgios Gousios. Can we trust tests to automate dependency updates? A case study of Java Projects. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001941.

Horcas:2023:MCT

[HGH^{+}23] Jose-Miguel Horcas, José A. Galindo, Ruben Heradio, David Fernandez-Amoros, and David Benavides. A Monte Carlo tree search conceptual framework for feature model analyses. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002278.

Hajiabadi:2024:DDA

[HGHK24] Hamideh Hajiabadi, Christopher Gerking, Lennart Hilbert, and Anne Koziolek. A data-driven active learning approach to reusing ML solutions in scientific applications. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000293.

Hirsch:2022:SLR

[HH22] Thomas Hirsch and Birgit Hofer. A systematic literature review on benchmarks for evaluating debugging approaches. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001303.

Henning:2024:BSS

[HH24] Sören Henning and Wilhelm Hasselbring. Benchmarking scalability of stream processing frameworks deployed as microservices in the cloud. The Journal of Systems and Software, 208(??):??, February 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223002741.

Hirsch:2025:BPE

[HH25] Thomas Hirsch and Birgit Hofer. Best practices for evaluating IRFL approaches. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S016412122500010X.

Huang:2020:CDL

[HHJ^{+}20] Yuan Huang, Xinyu Hu, Nan Jia, Xiangping Chen, Zibin Zheng, and Xiapu Luo. CommtPst: Deep learning source code for commenting positions prediction. The Journal of Systems and Software, 170(??): ??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301758.

Higo:2020:TJM

[HHK20] Yoshiki Higo, Shinpei Hayashi, and Shinji Kusumoto. On tracking Java methods with Git mechanisms. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300522.

Hussain:2023:BSC

[HHZW23] Yasir Hussain, Zhiqiu Huang, Yu Zhou, and Senzhang Wang. Boosting source code suggestion with selfsupervised Transformer Gated Highway. The Journal of Systems and Software, 196(??): ??, February 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222002291.

Hladik:2021:HFM

[HIDT21] Pierre-Emmanuel Hladik, Félix Ingrand, Silvano Dal Zilio, and Reyyan Tekin. Hippo: a formal-model execution engine to control and verify critical realtime systems. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001308.

Hofer:2021:PMS

[HJK^{+}21] Birgit Hofer, Dietmar Jannach, Patrick Koch, Konstantin Schekotihin, and Franz Wotawa. Product metrics for spreadsheets — a systematic review. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000078.

Heyn:2023:CAC

[HKP23] Hans-Martin Heyn, Eric Knauss, and Patrizio Pelliccione. A compositional approach to creating architecture frameworks with an application to distributed AI systems. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121222002801.

Hedenborg:2021:MEC

[HLL21]  Mathias  Heden-

borg, Jonas Lundberg, and Welf Löwe. Memory efficient contextsensitive program analysis. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000492.

Holtmann:2024:PMT

[HLS24] Jörg Holtmann, Grischa Liebel, and Jan-Philipp Steghöfer. Processes, methods, and tools in modelbased engineering

— a qualitative multiple-case study. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003382.

He:2023:NFA

[HLW^{+}23] Jiahao He, Shuangyin Li, Xinming Wang, Shing-Chi Cheung, Gansen Zhao, and Jinji Yang. Neural-FEBI: Accurate function identification in Ethereum Virtual Machine bytecode. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000225.

Hu:2023:MES

[HLZ^{+}23] Haize Hu, Jianxun Liu, Xiangping Zhang, Ben Cao, Siqiang Cheng, and Teng Long. A mutual embedded selfattention network model for code search. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121222002679.

Huang:2024:AMC

[HM24] Fuqun Huang and Henrique Madeira. Advancing modern code review effectiveness through human error mechanisms. The Journal of Systems and Software, 214(??): ??, August 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224001055.

Happe:2025:AIO

[HM25] Lucia Happe and Kai Marquardt. Authentic interdisciplinary online courses for alternative pathways into computer science. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122400284X.

Habchi:2021:ACS

[HMR21] Sarra Habchi, Naouel Moha, and Romain Rouvoy. Android code smells: From introduction to refactoring. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000613.

Heldal:2024:SCS

[HNM^{+}24]  Rogardt  Heldal,

Ngoc-Thanh Nguyen, Ana Moreira, Patricia Lago, Leticia Duboc, Stefanie Betz, Vlad C. Coroamă, Birgit Penzenstadler, Jari Porras, Rafael Capilla, Ian Brooks, Shola Oyedeji, and Colin C. Venters. Sustainability competencies and skills in software engineering: an industry perspective. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000219.

Hron:2022:WHS

[HO22] Michal Hron and Nikolaus Obwegeser. Why and how is Scrum being adapted in practice: a systematic review. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002077.

Hasan:2023:LSC

[HOAM23] Muhammad Hafiz Hasan, Mohd Hafeez Osman, Novia Indriaty Admodisastro, and Muhamad Sufri Muhammad. Legacy systems to cloud migration: a review from the architectural perspective. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000973.

Herrmann:2022:SES

[HOCK22] Marc Herrmann, Martin Obaidi, Larissa Chazette, and Jil Klünder. On the subjectivity of emotions in software projects: How reliable are pre-labeled data sets for sentiment analysis? The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001431.

Hora:2021:CTR

[Hor21] Andre Hora. Characterizing top ranked code examples in Google. The Journal of Systems and Software, 178(??): ??, August 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221000686.

Horcas:2023:MMR

[HPF23] Jose-Miguel Horcas, Mónica Pinto, and Lidia Fuentes. A modular metamodel and refactoring rules to achieve software product line interoperability. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002552.

Hevia:2024:QDI

[HPP24] José Luis Hevia, Guido Peterssen, and Mario Piattini. qSOA(R): Dynamic integration for hybrid quantum/classical software systems. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001067.

Hu:2020:NJA

[HPZ^{+}20] Gang Hu, Min Peng, Yihan Zhang, Qianqian Xie, and Mengting Yuan. Neural joint attention code search over structure embeddings for software Q&A sites. The Journal of Systems and Software, 170(??):??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301886.

Ho-Quang:2022:RSS

[HQNR^{+}22] Truong Ho-Quang, Arif Nurwidyantoro, Satrio Adi Rukmono, Michel R. V. Chaudron, Fabian Fröding, and Duy Nguyen Ngoc. Role stereotypes in software designs and their evolution. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000504.

Higo:2021:ISI

[HS21] Yoshiki Higo and Alexander Serebrenik. Introduction to special issue on source code analysis and manipulation. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000467.

Hyun:2023:TPB

[HSJB23] Sangwon Hyun, Jiyoung Song, Eunkyoung Jee, and Doo-Hwan Bae. Timed pattern-based analysis of collaboration failures in systemof-systems. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000080.

Harrand:2020:JDD

[HSVMB20] Nicolas Harrand, César Soto-Valero, Martin Monperrus, and Benoit Baudry. Java decompiler diversity and its application to metadecompilation. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301151.

Huang:2024:RAS

[HSW^{+}24] Ruiyao Huang, Qingni Shen, Yuchen Wang, Yiqi Wu, Zhonghai Wu, Xiapu Luo, and Anbang Ruan. ReenRepair: Automatic and semantic equivalent repair of reentrancy in smart contracts. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001523.

He:2021:SAM

[HTB21] TianZhang He, Adel N. Toosi, and Rajkumar Buyya. SLAaware multiple migration planning and scheduling in SDN-NFV-enabled clouds. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221000406.

Hsu:2023:EAC

[HTC^{+}23]  Tse-Chuan  Hsu,

Yao-Hong Tsai, William Cheng-Chung Chu, Shyh wei Chen, Hung-Lung Tsai, and Yu-Kang Chang. Exploration of advanced computer technology to address analytical and noise improvement issues in machine learning. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002157.

Hansen:2024:CSD

[HTG^{+}24] Simon Thrane Hansen, Casper Thule, Cláudio Gomes, Kenneth Guldbrandt Lausdahl, Frederik Palludan Madsen, Giuseppe Abbiati, and Peter Gorm Larsen. Co-simulation at different levels of expertise with Maestro2. The Journal of Systems and Software, 209(??): ??, March 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412122300300X.

Hassan:2025:CRD

[HTN^{+}25] Zara Hassan, Christoph Treude, Michael Norrish, Graham Williams, and Alex Potanin. Characterising reproducibility debt in scientific software: a systematic literature review. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003716.

Hu:2024:TMO

[Hu24]  Jian Hu.  Trace

matrix optimization for fault localization. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002959.

Hu:2025:CFD

[Hu25] Jian Hu. CG-FL:

a data augmentation approach using context-aware genetic algorithm for fault localization. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000275.

Hui:2025:UMT

[HWL^{+}25] Mingxuan Hui, Lu Wang, Hao Li, Ren Yang, Yuxin Song, Huiying Zhuang, Di Cui, and Qingshan Li.

Unveiling the microservices testing methods, challenges, solutions, and solutions gaps: a systematic mapping study. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002760.

Huang:2020:WTE

[HXJ^{+}20]  Yicheng  Huang,

Chang Xu, Yanyan Jiang, Huiyan Wang, and Da Li. WARDER: Towards effective spreadsheet defect detection by validitybased cell cluster refinements. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300935.

Huang:2024:EDF

[HYF^{+}24] Zijie Huang, Huiqun Yu, Guisheng Fan, Zhiqing Shao, Ziyi Zhou, and Mingchen Li. On the effectiveness of developer features in code smell prioritization: a replication study. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000116.

Hyrynsalmi:2025:COI

[Hyr25] Sonja M. Hyrynsalmi. Challenges and opportunities: Implementing diversity and inclusion in software engineering university level education in Finland. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002838.

He:2024:BTB

[HZ24] Xiao He and Tao Zan. BIT: a template-based approach to incremental and bidirectional model-to-text transformation. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001936.

Huang:2020:RTC

[HZT^{+}20] Rubing Huang, Quanjun Zhang, Dave Towey, Weifeng Sun, and Jinfu Chen. Regression test case prioritization by code combinations coverage. The Journal of Systems and Software, 169(??):??, November 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301540.

Ibias:2022:UMI

[Ibi22] Alfredo Ibias. Using mutual information to test from Finite State Machines: Test suite generation. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001108.

Irshad:2021:ABD

[IBP21] Mohsin Irshad, Ricardo Britto, and Kai Petersen. Adapting Behavior Driven Development (BDD) for large-scale software systems. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000418.

Imtiaz:2021:AMB

[IIK21] Javaria Imtiaz, Muhammad Zohaib Iqbal, and Muhammad Uzair Khan. An automated model-based approach to repair test suites of evolving web applications. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302314.

Illahi:2021:MLB

[ILUN21] Inam Illahi, Hui Liu, Qasim Umer, and Nan Niu. Machine learning based success prediction for crowdsourcing software projects. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000625.

Iqbal:2023:TCE

[IMTS23] Tahira Iqbal, James George Marshall, Kuldar Taveter, and Albrecht Schmidt. Theory of constructed emotion meets RE: an industrial case study. The Journal of Systems and Software, 197(??): ??, March 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222002205.

Islam:2023:RSP

[IPB23] Chadni Islam, Victor Prokhorenko, and M. Ali Babar. Runtime software patching: Taxonomy, survey and future directions. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300047X.

Isomottonen:2021:DGE

[IR21] Ville Isomöttönen and Emmi Ritvos. Digging into group establishment: Intervention design and evaluation. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000716.

Islam:2020:CED

[ISKB20] Muhammed Tawfiqul Islam, Satish Narayana Srirama, Shanika Karunasekera, and Rajkumar Buyya. Cost-efficient dynamic scheduling of big data applications in Apache Spark on cloud. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302894.

Indykov:2025:ATA

[ISW25] Vladislav Indykov, Daniel Strüber, and Rebekka Wohlrab. Architectural tactics to achieve quality attributes of machine-learning-enabled systems: a systematic literature review. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S016412122500041X.

Isomottonen:2023:SIS

[IT23] Ville Isomöttönen and Toni Taipalus. Status indicators in software engineering group projects. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000079.

Iammarino:2021:ESC

[IZAD21] Martina Iammarino, Fiorella Zampetti, Lerina Aversano, and Massimiliano Di Penta. An empirical study on the co-occurrence between refactoring actions and selfadmitted technical debt removal. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100073X.

Jahanshahi:2022:WMT

[JCNS^{+}22] Hadi Jahanshahi, Mucahit Cevik, José Navas-Sú, Ayşe Başar, and Antonio González-Torres. Wayback Machine: a tool to capture the evolutionary behavior of the bug reports and their triage process in open-source software systems. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000565.

Jorgensen:2020:SEE

[JH20] Magne Jørgensen and Torleif Halkjelsvik. Sequence effects in the estimation of software development effort. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302225.

Jevitha:2024:RVA

[JJS24] K. P. Jevitha, Bharat Jayaraman, and M. Sethumadhavan. Runtime verification on abstract finite state models. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001833.

Jolak:2025:EII

[JKD25] Rodi Jolak, Simon Karlsson, and Felix Dobslaw. An empirical investigation of the impact of architectural smells on software maintainability. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000500.

Janes:2023:OTT

[JLL23] Andrea Janes, Xiaozhou Li, and Valentina Lenarduzzi. Open tracing tools: Overview and critical comparison. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001887.

Jorgensen:2024:SLR

[Jør24] Magne Jørgensen. A systematic literature review on characteristics of the front-end phase of agile software development projects and their connections to project success. The Journal of Systems and Software, 216(??): ??, October 2024.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224002000.

Jolak:2022:CFS

[JRM^{+}22] Rodi Jolak, Thomas Rosenstatter, Mazen Mohamad, Kim Strandberg, Behrooz Sangchoolie, Nasser Nowdehi, and Riccardo Scandariato. CONSERVE: a framework for the selection of techniques for monitoring containers security. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002478.

Jin:2022:WBR

[JS22] Xianhao Jin and Francisco Servant. Which builds are really safe to skip? Maximizing failure observation for build selection in continuous integration. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121222000486.

Jiang:2025:JTS

[JSG25] Yuze Jiang, Beijun Shen, and Xiaodong Gu. Just-in-time software defect prediction via bimodal change representation learning. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002978.

Jiang:2023:DDL

[JST^{+}23] Yuan Jiang, Xiaohong Su, Christoph Treude, Chao Shang, and Tiantian Wang. Does deep learning improve the performance of duplicate bug report detection? An empirical study. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300002X.

Jiang:2022:HSA

[JSTW22] Yuan Jiang, Xiaohong Su, Christoph Treude, and Tiantian Wang. Hierarchical semantic-aware neural code representation. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000887.

Jin:2020:CLA

[JT20] Hao Jin and Tatsuhiro Tsuchiya. Constrained locating arrays for combinatorial interaction testing. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301874.

Jullien:2025:TFP

[JVZ25] Nicolas Jullien, Robert Viseur, and Jean-Benoît Zimmermann. A theory of FLOSS projects and open source business models dynamics. The Journal of Systems and Software, 224(??):??, June 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000512.

Jiang:2020:AMC

[JWZM20] Zijian Jiang, Ye Wang, Hao Zhong, and Na Meng. Automatic method change suggestion to complement multientity edits. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302158.

Jia:2024:PPM

[JYZ^{+}24] Kai Jia, Xiao Yu, Chen Zhang, Wenzhi Xie, Dongdong Zhao, and Jianwen Xiang. PMTT: Parallel multi-scale temporal convolution network and transformer for predicting the time to aging failure of software systems. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002127.

Jiang:2021:IRC

[JZM21] Zijian Jiang, Hao Zhong, and Na Meng. Investigating and recommending cochanged entities for JavaScript programs. The Journal of Systems and Software, 180(??): ??, October 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221001242.

Jia:2021:SCR

[JZW^{+}21] Li Jia, Hao Zhong, Xiaoyin Wang, Linpeng Huang, and Xuansheng Lu. The symptoms, causes, and repairs of bugs inside a deep learning library. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000327.

Kessel:2022:DDU

[KA22] Marcus Kessel and Colin Atkinson. Diversitydriven unit test generation. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001406.

Kessel:2024:CSE

[KA24] Marcus Kessel and Colin Atkinson. Code search engines for the next generation. The Journal of Systems and Software, 215(??):??, September 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224001109.

Kretsou:2021:CIA

[KAA^{+}21] Maria Kretsou, Elvira-Maria Arvanitou, Apostolos Ampatzoglou, Ignatios Deligiannis, and Vassilis C. Gerogiannis. Change impact analysis: a systematic mapping study. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030282X.

Khan:2023:SAQ

[KAW^{+}23] Arif Ali Khan, Aakash Ahmad, Muhammad Waseem, Peng Liang, Mahdi Fahmideh, Tommi Mikkonen, and Pekka Abrahamsson. Software architecture for quantum computing systems — a systematic review. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000778.

Kurian:2023:AGT

[KBBD23] Elson Kurian, Daniela Briola, Pietro Braione, and Giovanni Denaro. Automatically generating test cases for safety-critical software via symbolic execution. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000249.

Khorram:2024:LPT

[KBG^{+}24] Faezeh Khorram, Erwan Bousse, Antonio Garmendia, Jean-Marie Mottu, Gerson Sunyé, and Manuel Wimmer. A language-parametric test coverage framework for executable domain-specific languages. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224000207.

Koziolek:2024:FST

[KBP24] Heiko Koziolek, Andreas Burger, and Abdulla Puthan Peedikayil. Fast state transfer for updates and live migration of industrial controller runtimes in container orchestration systems. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121224000475.

Koziolek:2020:CFA

[KBPMJ20] Heiko Koziolek, Andreas Burger, Marie Platenius-Mohr, and Raoul Jetley. A classification framework for automated control code generation in industrial automation. The Journal of Systems and Software, 166(??):??, August 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122030056X.

Kramer:2024:CBD

[KBW^{+}24] Michel Krämer, Pascal Bormann, Hendrik M. Würz, Kevin Kocon, Tobias Frechen, and Jonas Schmid. A cloud-based data processing and visualization pipeline for the fibre rollout in Germany. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000517.

Khalil:2021:IRP

[KCMD21] Zeinab Abou Khalil, Eleni Constantinou, Tom Mens, and Laurence Duchien. On the impact of release policies on bug handling activity: a case study of Eclipse. The Journal of Systems and Software, 173(??):??, March 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302727.

Kazerouni:2021:FAI

[KDB^{+}21] Ayaan M. Kazerouni, James C. Davis, Arinjoy Basak, Clifford A. Shaffer, Francisco Servant, and Stephen H. Edwards. Fast and accurate incremental feedback for students’ software tests using selective mutation analysis. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000029.

Kadenic:2023:IRP

[KdJPK^{+}23] Maja Due Kadenic, Diego Augusto de Jesus Pacheco, Konstantinos Koumaditis, Gitte Tjørnehøj, and Torben Tambo. Investigating the role of Product Owner in Scrum teams: Differentiation between organisational and individual impacts and opportunities. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002364.

Koci:2023:WAE

[KFJA23] Rediana Koçi, Xavier Franch, Petar Jovanovic, and Alberto Abelló. Web API evolution patterns: a usagedriven approach. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000043.

Kim:2022:DPP

[KGL^{+}22] Kisub Kim, Sankalp Ghatpande, Kui Liu, Anil Koyuncu, Dongsun Kim, Tegawendé F. Bissyandé, Jacques Klein, and Yves Le Traon. DigBug — pre/post-processing operator selection for accurate bug localization. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000528.

Kanij:2024:EUA

[KGM24] Tanjila Kanij, John Grundy, and Jennifer McIntosh. Enhancing understanding and addressing gender bias in IT/SE job advertisements. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002140.

Kroher:2023:CIV

[KGS23] Christian Kröher, Lea Gerling, and Klaus Schmid. Comparing the intensity of variability changes in software product line evolution. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001322.

Kirschner:2025:RVB

[KGS^{+}25] Yves R. Kirschner, Moritz Gstür, Timur Sağlam, Sebastian Weber, and Anne Koziolek. Retriever: a viewbased approach to reverse engineering software architecture models. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003212.

Kim:2024:BBR

[KH24] Jeewoong Kim and Shin Hong. BugOss: a benchmark of real-world regression bugs for empirical investigation of regression fuzzing techniques. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122400164X.

Kopec-Harding:2023:IUC

[KHEC^{+}23] Kamilla Kopec-Harding, Sukru Eraslan, Bowen Cai, Suzanne M. Embury, and Caroline Jay. The impact of unequal contributions in student software engineering team projects. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002340.

Kokkoniemi:2023:SMS

[KI23] Mikko Kokkoniemi and Ville Isomöttönen. A systematic mapping study on group work research in computing education projects. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001905.

Kretschmer:2021:TAC

[KKE21] Roland Kretschmer, Djamel Eddine Khelladi, and Alexander Egyed. Transforming abstract to concrete repairs with a generative approach of repair values. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030279X.

Kasauli:2021:REC

[KKH^{+}21] Rashidah Kasauli, Eric Knauss, Jennifer Horkoff, Grischa Liebel, and Francisco Gomes de Oliveira Neto. Requirements engineering challenges and practices in large-scale agile system development. The Journal of Systems and Software, 172(??):??, February 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220302417.

Klare:2021:ECV

[KKL^{+}21] Heiko Klare, Max E.

Kramer, Michael Langhammer, Dominik Werle, Erik Burger, and Ralf Reussner. Enabling consistency in viewbased system development — the Vitruvius approach. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302144.

Karimi:2024:ACO

[KKRT24] Meysam Karimi, Shekoufeh Kolahdouz-Rahimi, and Javier Troya. Ant-colony optimization for automating test model generation in model transformation testing. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002777.

Kim:2020:URS

[KL20] Bong-Jae Kim and Seok-Won Lee. Understanding and recommending security requirements from problem domain ontology: a cognitive three-layered approach. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030145X.

Kropp:2020:SCA

[KMAB20] Martin Kropp, Andreas Meier, Craig Anslow, and Robert Biddle. Satisfaction and its correlates in agile software development. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300261.

Kotilainen:2025:ADA

[KMS^{+}25] Pyry Kotilainen, Niko Mäkitalo, Kari Systä, Ali Mehraj, Muhammad Waseem, Tommi Mikkonen, and Juan Manuel Murillo. Allocating distributed AI/ ML applications to cloud–edge continuum based on privacy, regulatory, and ethical constraints. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000019.

Kopczynska:2022:BPR

[KOPN22] Sylwia Kopczyńska, Mirosław Ochodek, Jakub Piechowiak, and Jerzy Nawrocki. On the benefits and problems related to using Definition of Done — a survey study. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121222001637.

Kokinda:2023:SSD

[KR23] Ella Kokinda and Paige Rodeghero. Streaming software development: Accountability, community, and learning. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223000250.

Kulyk:2023:PWR

[KRC23] Oksana Kulyk, Karen Renaud, and Stefan Costica. People want reassurance when making privacy-related decisions — not technicalities. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000158.

Khabou:2023:ISI

[KRD^{+}23] Nesrine Khabou, Ismael Bouassida Rodriguez, Khalil Drira, Paris Avgeriou, David C. Shepherd, Wing-Kwong Chan, and Raffaela Mirandola. Introduction to the Special Issue on Software-Intensive Autonomous Systems: Methods and applications. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200187X.

Kolahdouz-Rahimi:2020:CQF

[KRLS^{+}20] Shekoufeh Kolahdouz-Rahimi, Kevin Lano, Mohammadreza Sharbaf, Meysam Karimi, and Hessa Alfraihi. A comparison of quality flaws and technical debt in model transformation specifications. The Journal of Systems and Software, 169(??):??, November 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220301382.

Keshavarz:2024:JDG

[KRP24] Hossein Keshavarz and Gema Rodríguez-Pérez. JITGNN: a deep graph neural network framework for just-in-time bug prediction. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122400027X.

Kochanthara:2021:FSAa

[KRS^{+}21a] Sangeeth Kochanthara, Niels Rood, Arash Khabbaz Saberi, Loek Cleophas, Yanja Dajsuren, and Mark van den Brand. A functional safety assessment method for cooperative automotive architecture. The Journal of Systems and Software, 179(??): ??, September 2021. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221000881.

Kochanthara:2021:FSAb

[KRS^{+}21b] Sangeeth Kochanthara, Niels Rood, Arash Khabbaz Saberi, Loek Cleophas, Yanja Dajsuren, and Mark van den Brand. A functional safety assessment method for cooperative automotive architecture. The Journal of Systems and Software, 179(??): ??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221000881.

Kirchhof:2022:MMD

[KRSW22] Jörg Christian Kirchhof, Bernhard Rumpe, David Schmalzing, and Andreas Wortmann. MontiThings: Model-driven development and deployment of reliable IoT applications. The Journal of Systems and Software, 183(??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001849.

Karras:2020:RSP

[KSF20] Oliver Karras, Kurt Schneider, and Samuel A. Fricker. Representing software project vision by means of video: a quality model for vision videos. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302535.

Khandoker:2022:TLF

[KSG^{+}22] Azad Khandoker, Sabine Sint, Guido Gessl, Klaus Zeman, Franz Jungreitmayr, Helmut Wahl, Andreas Wenigwieser, and Roland Kretschmer. Towards a logical framework for ideal MBSE tool selection based on discipline specific requirements. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000553.

Keshani:2023:RMP

[KVP23] Mehdi Keshani, Simcha Vos, and Sebastian Proksch. On the relation of method popularity to breaking changes in the Maven ecosystem. The Journal of Systems and Software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223001334.

Ko:2022:FAC

[KZK22] Youngjoo Ko, Bin Zhu, and Jong Kim. Fuzzing with automatically controlled interleavings to detect concurrency bugs. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001042.

Kerdoudi:2022:NAS

[KZTS22] Mohamed Lamine Kerdoudi, Tewfik Ziadi, Chouki Tibermacine, and Salah Sadou. A novel approach for Software Architecture Product Line Engineering. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002685.

Lesoil:2023:ISP

[LABJ23] Luc Lesoil, Mathieu Acher, Arnaud Blouin, and Jean-Marc Jézéquel. Input sensitivity on the performance of configurable systems an empirical study. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000663.

Liu:2023:SFL

[LAL^{+}23] Jingyu Liu, Jun Ai, Minyan Lu, Jie Wang, and Haoxiang Shi. Semantic feature learning for software defect prediction from source code and external knowledge. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001486.

Lavazza:2023:EEC

[LALM23] Luigi Lavazza, Abedallah Zaid Abualkishik, Geng Liu, and Sandro Morasca. An empirical evaluation of the “Cognitive Complexity” measure as a predictor of code understandability. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002370.

Laurent:2022:MBA

[LATV22] Thomas Laurent, Paolo Arcaini, Catia Trubiani, and Anthony Ventresque. Mutation-based analysis of queueing network performance models. The Journal of Systems and Software, 191(??): ??, September 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222001078.

Leong:2024:TMR

[LB24] Iat Tou Leong and Raul Barbosa. Translating meaning representations to behavioural interface specifications. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000529.

Liu:2023:HCS

[LBCG23] Xiao Liu, Kelly Blincoe, Mohan Baruwal Chhetri, and John Grundy. Humancentric software engineering — approaches, technologies, and applications. The Journal of Systems and Software, 204(??): ??, October 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223001863.

Lee:2021:OBAa

[LBF^{+}21a] Seongmin Lee, David Binkley, Robert Feldt, Nicolas Gold, and Shin Yoo. Observationbased approximate dependency modeling and its use for program slicing. The Journal of Systems and Software, 179(??):??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221000856.

Lee:2021:OBAb

[LBF^{+}21b] Seongmin Lee, David Binkley, Robert Feldt, Nicolas Gold, and Shin Yoo. Observationbased approximate dependency modeling and its use for program slicing. The Journal of Systems and Software, 179(??):??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221000856.

Lee:2020:ELA

[LBG^{+}20] Seongmin Lee, David Binkley, Nicolas Gold, Syed Islam, Jens Krinke, and Shin Yoo. Evaluating lexical approximation of program dependence. The Journal of Systems and Software, 160 (??):??, February 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412121930233X.

Li:2024:RRS

[LBH^{+}24] Yufei Li, Liang Bao, Kaipeng Huang, Chase Wu, and Xinwei Li. RS-FIN: a Rule Searchbased Fuzzy Inference Network for performance prediction of configurable software systems. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003084.

Li:2025:CCS

[LBHW25] Yufei Li, Liang Bao, Kaipeng Huang, and Chase Wu. CSAT: Configuration structure-aware tuning for highly configurable software systems. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003601.

Lopez:2022:QMA

[LBMF^{+}22] Lidia López, Xavier Burgués, Silverio Martínez-Fernández, Anna Maria Vollmer, Woubshet Behutiye, Pertti Karhapää, Xavier Franch, Pilar Rodríguez, and Markku Oivo. Quality measurement in agile and rapid software development: a systematic mapping. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002661.

Lenarduzzi:2021:SLR

[LBT^{+}21] Valentina Lenarduzzi, Terese Besker, Davide Taibi, Antonio Martini, and Francesca Arcelli Fontana. A systematic literature review on technical debt prioritization: Strategies, processes, factors, and tools. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122030220X.

Li:2020:OBL

[LC20] Tong Li and Zhishuai Chen. An ontologybased learning approach for automatically classifying security requirements. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300479.

Lima:2021:IOE

[LCAC21] Stanley Lima, Jaime Correia, Filipe Araujo, and Jorge Cardoso. Improving observability in Event Sourcing systems. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001126.

Li:2023:GRR

[LCC^{+}23a] Zeyan Li, Junjie Chen, Yihao Chen, Chengyang Luo, Yiwei Zhao, Yongqian Sun, Kaixin Sui, Xiping Wang, Dapeng Liu, Xing Jin, Qi Wang, and Dan Pei. Generic and robust root cause localization for multidimensional data in online service systems. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001437.

Lin:2023:TFQ

[LCC^{+}23b] Hao Lin, Xiang Chen, Xuejiao Chen, Zhanqi Cui, Yun Miao, Shan Zhou, Jianmin Wang, and Zhan Su. TitleGen-FL: Quality predictionbased filter for automated issue title generation. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001893.

Lambiase:2024:EII

[LCP^{+}24] Stefano Lambiase, Gemma Catolino, Fabiano Pecorelli, Damian A. Tamburri, Fabio Palomba, Willem-Jan van den Heuvel, and Filomena Ferrucci. An empirical investigation into the influence of software communities’ cultural and geographical dispersion on productivity. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300273X.

Lim:2024:TCI

[LCS^{+}24] Jin Wei Lim, Thiam Kian Chiew, Moon Ting Su, Simying Ong, Hema Subramaniam, Mumtaz Begum Mustafa, and Yin Kia Chiam. Test case information extraction from requirements specifications using NLPbased unified boilerplate approach. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000487.

Liu:2022:SFM

[LCT22] Yin Liu, Breno Dantas Cruz, and Eli Tilevich. Secure and flexible messagebased communication for mobile apps within and across devices. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001492.

Luo:2024:GCI

[LCW^{+}24] Weilin Luo, Polong Chen, Hai Wan, Hongzhen Zhong, Shaowei Cai, and Zhanhao Xiao. Goal-conflict identification based on local search and fast boundary-condition verification based on incremental satisfiability filter. The Journal of Systems and Software, 213(??):??, July 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000797.

Li:2023:SEV

[LCY23] Chen Li, Dimitris Chrysostomou, and Hongji Yang. A speech-enabled virtual assistant for efficient human– robot interaction in industrial environments. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002133.

Li:2024:BPC

[LCY^{+}24] Zengyang Li, Guangzong Cai, Qinyi Yu, Peng Liang, Ran Mo, and Hui Liu. Bug priority change: an empirical study on Apache projects. The Journal of Systems and Software, 212(??):??, June 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000621.

Li:2022:BTN

[LDH22] Siqiao Li, Tadashi Dohi, and Okamura Hiroyuki. Burrtype NHPP-based software reliability models and their applications with two type of fault count data. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000942.

Liu:2022:OPW

[LDT22] Yin Liu, Siddharth Dhar, and Eli Tilevich. Only pay for what you need: Detecting and removing unnecessary TEE-based code. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000255.

Li:2025:ESA

[LDT^{+}25] Yinghua Li, Xueqi Dang, Haoye Tian, Tiezhu Sun, Zhijie Wang, Lei Ma, Jacques Klein, and Tegawendé F. Bissyandé. An empirical study of AI techniques in mobile applications. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002772.

Liu:2023:VSC

[LFFW23] Haiyang Liu, Yuqi Fan, Lin Feng, and Zhenchun Wei. Vulnerable smart contract function locating based on Multi-Relational Nested Graph Convolutional Network. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300170X.

Li:2022:CCM

[LFH^{+}22] Yuying Li, Yang Feng, Rui Hao, Di Liu, Chunrong Fang, Zhenyu Chen, and Baowen Xu.

Classifying crowdsourced mobile test reports with image features: an empirical study. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002181.

Lee:2024:SEC

[LFLS24] Sung Une Lee, Niroshinie Fernando, Kevin Lee, and Jean-Guy Schneider. A survey of energy concerns for software engineering. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003394.

Li:2025:CDP

[LGC^{+}25] Shuo Li, Haocheng Gao, Wei Chen, Yi Li, Haoxiang Tian, Chengwei Liu, and Dan Ye. Characterizing and detecting Python version incompatibilities caused by inconsistent version specifications. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000056.

Licorish:2022:USS

[LGKT22] Sherlock A. Licorish, Matthias Galster, Georgia M. Kapitsaki, and Amjed Tahir. Understanding students’ software development projects: Effort, performance, satisfaction, skills and their relation to the adequacy of outcomes developed. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/

                                        pii/S0164121221002466.

Lercher:2024:MAE

[LGMP24] Alexander Lercher, Johann Glock, Christian Macho, and Martin Pinzger. Microservice API evolution in practice: a study on strategies and challenges. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001559.

Levy:2023:SHH

[LGT^{+}23] Meira Levy, Eduard C. Groen, Kuldar Taveter, Daniel Amyot, Eric Yu, Lin Liu, Ita Richardson, Maria Spichkova, Alexandra Jussli, and Sébastien Mosser. Sustaining human health: a requirements engineering perspective. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001875.

Lathouwers:2024:SAG

[LH24a] Sophie Lathouwers and Marieke Huisman. Survey of annotation generators for deductive verifiers. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000153.

Levy:2024:LEU

[LH24b] Meira Levy and Irit Hadar. Learning to empathize with users through design thinking in hybrid mode: Insights from two educational case studies. The Journal of Systems and Software, 207(??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121223002261.

Lin:2022:AAS

[LHF22] Jhih-Sin Lin, Chin-Yu Huang, and Chih-Chiang Fang. Analysis and assessment of software reliability modeling with preemptive priority queueing policy. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000231.

Lelovic:2025:CIA

[LHG^{+}25] Luka Lelovic, Austin Huzinga, Gabriel Goulis, Anshpreet Kaur, Ricardo Boone, Umidjon Muzrapov, Amr S. Abdelfattah, and Tomas Cerny. Change impact analysis in microservice systems: a systematic literature review. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002851.

Liu:2024:RAS

[LHH^{+}24] Mugeng Liu, Xiaolong Huang, Wei He, Yibing Xie, Jie M. Zhang, Xiang Jing, Zhenpeng Chen, and Yun Ma. Research artifacts in software engineering publications: Status and trends. The Journal of Systems and Software, 213(??):??, July 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122400075X.

Lefticaru:2020:IRT

[LHN20] Raluca Lefticaru, Robert M. Hierons, and Manuel Núñez. Implementation relations and testing for cyclic systems with refusals and discrete time. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301679.

Liu:2024:DDS

[LHS^{+}24] Hezhen Liu, Chengqiang Huang, Ke Sun, Jiacheng Yin, Xiaoyu Wu, Jin Wang, Qunli Zhang, Yang Zheng, Vivek Nigam, Feng Liu, and Joseph Sifakis. Design for dependability — state of the art and trends. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000323.

Lan:2025:BFT

[LHY^{+}25] Wenwei Lan, Chen Huang, Tingting Yu, Li Li, and Zhanqi Cui. BaS-Fuzz: Fuzz testing based on difference analysis for seed bytes. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000081.

Lomio:2022:JTS

[LID^{+}22] Francesco Lomio, Emanuele Iannone, Andrea De Lucia, Fabio Palomba, and Valentina Lenarduzzi. Just-in-time software vulnerability detection: Are we there yet? The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000437.

Liu:2021:TSH

[Liu21] Shaoying Liu. A three-step hybrid specification approach to error prevention. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000728.

Liebel:2023:AMR

[LK23] Grischa Liebel and Eric Knauss. Aspects of modelling requirements in verylarge agile systems engineering. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000237.

Loch:2021:NBP

[LKP^{+}21] Wilton Jaciel Loch, Guilherme Piêgas Koslovski, Maurício Aronne Pillon, Charles Christian Miers, and Marcelo Pasin. A novel blockchain protocol for selecting microservices providers and auditing contracts. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001278.

Liu:2021:EQF

[LL21] Geng Liu and Luigi Lavazza. Early and quick function points analysis: Evaluations and proposals. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302788.

Liu:2023:UEP

[LL23] Pan Liu and Yihao Li. Using expression parsing and algebraic operations to generate test sequences. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121223001930.

Li:2025:HTB

[LL25] Yihao Li and Pan Liu. Hierarchical tree-based algorithms for efficient expression parsing and test sequence generation in software models. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000226.

Liu:2025:INM

[LLH^{+}25] Hengyuan Liu, Zheng Li, Baolong Han, Xiang Chen, Doyle Paul, and Yong Liu. Integrating neural mutation into mutation-based fault localization: a hybrid approach. The Journal of Systems and Software, 221(??):??, March 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/

science/article/
pii/S016412122400325X.

Liu:2021:CRE

[LLK^{+}21] Kui Liu, Li Li, Anil Koyuncu, Dongsun Kim, Zhe Liu, Jacques Klein, and Tegawendé F. Bissyandé. A critical review on the evaluation of automated program repair systems. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302156.

Lei:2022:DLA

[LLL^{+}22] Maggie Lei, Hao Li, Ji Li, Namrata Aundhkar, and Dae-Kyoo Kim. Deep learning application on code clone detection: a review of current knowledge. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002387.

Liu:2025:ADP

[LLL^{+}25] Yue Liu, Sin Kit Lo, Qinghua Lu, Liming Zhu, Dehai Zhao, Xiwei Xu, Stefan Harrer, and Jon Whittle. Agent design pattern catalogue: a collection of architectural patterns for foundation model based agents. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003224.

Luu:2021:TML

[LLNC21] Quang-Hung Luu, Man F. Lau, Sebastian P. H. Ng, and Tsong Yueh Chen. Testing multiple linear regression systems with metamorphic testing. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100159X.

Lenarduzzi:2020:DMM

[LLST20] Valentina Lenarduzzi, Francesco Lomio, Nyyti Saarimäki, and Davide Taibi. Does migrating a monolithic system to microservices decrease the technical debt? The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301539.

Liu:2023:GMC

[LLW^{+}23] Zhengli Liu, Bing Li, Jian Wang, Xiangfei Lu, and Yu Qiao. Goal model convergence and conflict detection for crossover services. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000201.

Li:2024:GGB

[LLW^{+}24] Zheng Li, Mingyu Li, Shumei Wu, Shunqing Xu, Xiang Chen, and Yong Liu. GBSR: Graphbased suspiciousness refinement for improving fault localization. The Journal of Systems and Software, 218 (??):??, December 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002334.

Li:2022:FTS

[LLWL22] Chunlin Li, Jun Liu, Min Wang, and Youlong Luo. Faulttolerant scheduling and data placement for scientific workflow processing in geodistributed clouds. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000073.

Lo:2022:APD

[LLZ^{+}22] Sin Kit Lo, Qinghua Lu, Liming Zhu, Hye-Young Paik, Xiwei Xu, and Chen Wang. Architectural patterns for the design of federated learning systems. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000899.

Liu:2023:SLR

[LLZ^{+}23] Yue Liu, Qinghua Lu, Liming Zhu, Hye-Young Paik, and Mark Staples. A systematic literature review on blockchain governance. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002527.

Liu:2021:FSA

[LM21] Shaoying Liu and Weikai Miao. A formal specification animation method for operation validation. The Journal of Systems and Software, 178(??): ??, August 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221000455.

Laigner:2022:CDI

[LMGK22] Rodrigo Laigner, Diogo Mendonça, Alessandro Garcia, and Marcos Kalinowski. Cataloging dependency injection anti-patterns in software systems. The Journal of Systems and Software, 184(??):??, February 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221002223.

Latif:2023:PEC

[LMR^{+}23] Saira Latif, Zaigham Mushtaq, Ghulam Rasool, Furqan Rustam, Naila Aslam, and Imran Ashraf. Pragmatic evidence of cross-language link detection: a systematic literature review. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002200.

Lopez-Martin:2020:TNN

[LMVRA^{+}20] Cuauhtémoc López-Martín, Yenny Villuendas-Rey, Mohammad Azzeh, Ali Bou Nassif, and Shadi Banitaan. Transformed k-nearest neighborhood output distance minimization for predicting the defect density of software projects. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300728.

Li:2023:AVD

[LMZ^{+}23] Xiaozhou Li, Sergio Moreschini, Zheying Zhang, Fabio Palomba, and Davide Taibi. The anatomy of a vulnerability database: a systematic mapping study. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000742.

Li:2022:EFM

[LMZT22] Xiaozhou Li, Sergio Moreschini, Zheying Zhang, and Davide Taibi. Exploring factors and metrics to select open source software components for integration: an empirical study. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000267.

Lenarduzzi:2021:DCQ

[LNST21] Valentina Lenarduzzi, Vili Nikkola, Nyyti Saarimäki, and Davide Taibi. Does code quality affect pull request acceptance? An empirical study. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302090.

Ly:2025:PWA

[LOBD25] Delina Ly, Michiel Overeem, Sjaak Brinkkemper, and Fabiano Dalpiaz. The power of words in agile vs. waterfall development: Written communication in hybrid software teams. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002875.

Lacerda:2020:CSR

[LPPG20] Guilherme Lacerda, Fabio Petrillo, Marcelo Pimenta, and Yann Gaël Guéhéneuc. Code smells and refactoring: a tertiary systematic review of challenges and observations. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300881.

Lenarduzzi:2023:CCS

[LPS^{+}23] Valentina Lenarduzzi, Fabiano Pecorelli, Nyyti Saarimaki, Savanna Lujan, and Fabio Palomba. A critical comparison on six static analysis tools: Detection, agreement, and precision. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002515.

Li:2022:EMP

[LQY^{+}22] Zengyang Li, Xiaoxiao Qi, Qinyi Yu, Peng Liang, Ran Mo, and Chen Yang. Exploring multi-programming-language commits and their impacts on software quality: an empirical study on Apache projects. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001844.

Le:2022:ARE

[LSB^{+}22] Ha Thanh Le, Lwin Khin Shar, Domenico Bianculli, Lionel Claude Briand, and Cu Duy Nguyen. Automated reverse engineering of rolebased access control policies of web applications. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002065.

Linsbauer:2021:CVC

[LSBG21] Lukas Linsbauer, Felix Schwägerl, Thorsten Berger, and Paul Grünbacher. Concepts of variation control systems. The Journal of Systems and Software, 171(??): ??, January 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220302041.

Lima:2025:HDL

[LSF^{+}25] Márcia Lima, Igor Steinmacher, Denae Ford, Grace Vorreuter, Ludimila Gonçalves, Tayana Conte, and Bruno Gadelha. How are discussions linked? A link analysis study on GitHub discussions. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002401.

Leite:2021:AMD

[LSFE21] Manoel Victor Rodrigues Leite, Lilian Passos Scatalon, André Pimenta Freire, and Marcelo Medeiros Eler. Accessibility in the mobile development industry in Brazil: Awareness, knowledge, adoption, motivations and barriers. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100039X.

Lesoil:2024:LIA

[LSG^{+}24] Luc Lesoil, Helge Spieker, Arnaud Gotlieb, Mathieu Acher, Paul Temple, Arnaud Blouin, and Jean-Marc Jézéquel. Learning input-aware performance models of configurable systems: an empirical evaluation. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002789.

Luo:2021:SIN

[LSSZ21] Xiapu Luo, Weiyi Shang, Xiaobing Sun, and Tao Zhang. Special issue on new generation of bug fixing. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121221000893.

Lenarduzzi:2020:SSI

[LST20] Valentina Lenarduzzi, Nyyti Saarimäki, and Davide Taibi. Some SonarQube issues have a significant but small effect on faults and changes. A large-scale empirical study. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301734.

Lienhardt:2024:PLD

[LtBD24] Michael Lienhardt, Maurice H. ter Beek, and Ferruccio Damiani. Product lines of dataflows. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121223003230.

Lonnfalt:2025:ITM

[LTG^{+}25] Albin Lönnfält, Viktor Tu, Gregory Gay, Animesh Singh, and Sahar Tahvili. An intelligent test management system for optimizing decision making during software testing. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002462.

Li:2020:EVC

[LTJ^{+}20] Nianyu Li, Christos Tsigkanos, Zhi Jin, Zhenjiang Hu, and Carlo Ghezzi. Early validation of cyber-physical space systems via multi-concerns integration. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

sciencedirect.com/
science/article/
pii/S0164121220301692.

Liu:2021:DCE

[LTZ^{+}21] Mingyi Liu, Zhiying Tu, Yeqi Zhu, Xiaofei Xu, Zhongjie Wang, and Quan Z. Sheng. Data correction and evolution analysis of the ProgrammableWeb service ecosystem. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001631.

Lee:2020:ETM

[LWC20] Jung-Chieh Lee, Yih-Tsyr Wang, and Chung-Yang Chen. The effect of transactive memory systems on process tailoring in software projects: the moderating role of task conflict and shared temporal cognitions. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300273.

Li:2020:HHM

[LWL20] Zheng Li, Haifeng Wang, and Yong Liu. HMER: a Hybrid Mutation Execution Reduction approach for mutation-based fault localization. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301230.

Li:2023:VVS

[LWL^{+}23] Zheng Li, Shumei Wu, Yong Liu, Jitao Shen, Yonghao Wu, Zhanwen Zhang, and Xiang Chen. VsusFL: Variablesuspiciousness-based fault localization for novice programs. The Journal of Systems and Software, 205(??):??, November 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223002170.

Li:2021:SSC

[LWP^{+}21] Zheng Li, Yonghao Wu, Bin Peng, Xiang Chen, Zeyu Sun, Yong Liu, and Deli Yu. SeCNN: a semantic CNN parser for code comment generation. The Journal of Systems and Software, 181(??):??, November 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001333.

Liu:2023:ARE

[LWSZ23] He Liu, Cheng Wei, Bo Sun, and Yinxue Zeng. Adaptive robustness evaluation for complex system prognostics and health management software platform. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001632.

Li:2020:ESC

[LWYW20] Yingling Li, Junjie Wang, Yun Yang, and Qing Wang. An extensive study of classlevel and methodlevel test case selection for continuous integration. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300923.

Leigh:2021:EER

[LWZ21] Andrew Leigh, Michel Wermelinger, and Andrea Zisman. Evaluating the effectiveness of risk containers to isolate change propagation. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000443.

Liu:2023:TAA

[LXL^{+}23] Pei Liu, Qingxin Xia, Kui Liu, Juncai Guo, Xin Wang, Jin Liu, John Grundy, and Li Li. Towards automated Android app internationalisation: an exploratory study. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002357.

Lin:2020:BBA

[LXLZ20] Junyu Lin, Lei Xu, Yingqi Liu, and Xiangyu Zhang. Black-box adversarial sample generation based on differential evolution. The Journal of Systems and Software, 170(??):??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301850.

Li:2020:TTR

[LXYL20] Can Li, Ling Xu, Meng Yan, and Yan Lei. TagDC: a tag recommendation method for software information sites with a combination of deep learning and collaborative filtering. The Journal of Systems and Software, 170(??):??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301941.

Li:2024:FSC

[LYG^{+}24] Xuan Li, Shuai Yuan, Xiaodong Gu, Yuting Chen, and Beijun Shen. Few-shot code translation via taskadapted prompt learning. The Journal of Systems and Software, 212(??):??, June 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000451.

Li:2023:FLB

[LYS^{+}23] Hongyan Li, Meng Yan, Weifeng Sun, Xiao Liu, and Yunsong Wu. A first look at bug report templates on GitHub. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001048.

Lu:2022:PDD

[LYZ^{+}22] Hai Lu, Ruyun Yu, Yan Zhu, Xiao He, Kaitai Liang, and William Cheng-Chung Chu. Policydriven data sharing over attributebased encryption supporting dual membership. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000346.

Long:2024:EIU

[LZ24] Peixun Long and Jianjun Zhao. Equivalence, identity, and unitarity checking in black-box testing of quantum programs. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000438.

Luo:2025:FQI

[LZ25] Junjie Luo and Jianjun Zhao. Formalization of Quantum Intermediate Representations for code safety. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121224002802.

Lesch:2023:LRI

[LZB^{+}23] Veronika Lesch, Marwin Züfle, André Bauer, Lukas Iffländer, Christian Krupitzer, and Samuel Kounev. A literature review of IoT and CPS — what they are, and what they are not. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000262.

Liu:2020:MPH

[LZJ20] Fang Liu, Lu Zhang, and Zhi Jin. Modeling programs hierarchically with stackaugmented LSTM. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300297.

Liu:2025:CFA

[LZL^{+}25] Jiaxin Liu, Yating Zhang, Yiwei Li, Tiecheng Ma, and Wei Dong. CIPAC: a framework of automated software construction based on collective intelligence. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000032.

Liu:2020:FAV

[LZP^{+}20] Ailun Liu, Huibiao Zhu, Miroslav Popovic, Shuangqing Xiang, and Lei Zhang. Formal analysis and verification of the PSTM architecture using CSP. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300418.

Mercan:2022:CDC

[MAC^{+}22] Hanefi Mercan, Atakan Aytar, Giray Coskun, Dilara Mustecep, Gülsüm Uzer, and Cemal Yilmaz. CITdaily: a combinatorial interaction testing-based daily build process. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000875.

Marsit:2021:REM

[MAK^{+}21] Imen Marsit, Amani Ayad, David Kim, Monsour Latif, Ji-Meng Loh, Mohamed Nazih Omri, and Ali Mili. The ratio of equivalent mutants: a key to analyzing mutation equivalence. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121221001369.

Marquez:2023:ATS

[MAK23] Gastón Márquez, Hernán Astudillo, and Rick Kazman. Architectural tactics in software architecture: a systematic mapping study. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002345.

Mamakou:2023:ICU

[Mam23] Xenia J. Mamakou. Intentions to continue using agile methods: the case of the Greek banking sector. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000808.

Magalhaes:2020:HHS

[MAP^{+}20] Claudio Magalhães, João Andrade, Lucas Perrusi, Alexandre Mota, Flávia Barros, and Eliot Maia. HSP: a hybrid selection and prioritisation of regression test cases based on information retrieval and code coverage applied on an industrial case study. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302043.

Masud:2022:ECM

[Mas22] Abu Naser Masud. Efficient computation of minimal weak and strong control closure. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121221002375.

Mujahid:2023:WCH

[MAS23] Suhaib Mujahid, Rabe Abdalkareem, and Emad Shihab. What are the characteristics of highlyselected packages? A case study on the npm ecosystem. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002643.

Mendonca:2024:FOT

[MAV24] Willian D. F. Mendonça, Wesley K. G. Assunção, and Silvia R. Vergilio. Feature-oriented test case selection and prioritization during the evolution of highly-configurable systems. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121224002024.

Molina:2025:ASC

[MBG^{+}25] Ana I. Molina, Crescencio Bravo, Jesús Gallardo, Carmen Lacave, and Miguel A. Redondo. Awareness support in collaborative programming tools: an evaluation based on programmer’s perception and eye tracking. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003200.

May:2024:PSS

[MBKL24] Richard May, Niklas Baron, Jacob Krüger, and Thomas Leich. Pandemic startup software engineering: an experience report on the development of a COVID-19 certificate verification system. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001511.

Munappy:2022:DMP

[MBO^{+}22] Aiswarya Raj Munappy, Jan Bosch, Helena Holmström Olsson, Anders Arpteg, and Björn Brinne. Data management for production quality deep learning models: Challenges and solutions. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000905.

Moonen:2020:ACR

[MBP20] Leon Moonen, David Binkley, and Sydney Pugh. On adaptive change recommendation. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300327.

Minhas:2023:CSD

[MBP23]  Nasir  Mehmood

Minhas, Jürgen Börstler, and Kai Petersen. Checklists to support decision-making in regression testing. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000924.

Masud:2020:MPC

[MC20] Abu Naser Masud and Federico Ciccozzi. More precise construction of static single assignment programs using reaching definitions. The Journal of Systems and Software, 166(??):??, August 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300704.

Mortara:2024:VOO

[MCDP24] Johann Mortara, Philippe Collet, and Anne-Marie Dery-Pinna. Visualization of objectoriented software in a city metaphor: Comprehending the implemented variability and its technical debt. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002716.

Menezes:2022:HFC

[MCH22] Gabriel Menezes, Bruno Cafeo, and Andre Hora. How are framework code samples maintained and used by developers? The case of Android and Spring Boot. The Journal of Systems and Software, 185(??):??, March 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002417.

Mashhadi:2024:ESB

[MCM^{+}24] Ehsan Mashhadi, Shaiful Chowdhury, Somayeh Modaberi, Hadi Hemmati, and Gias Uddin. An empirical study on bug severity estimation using source code metrics and static analysis. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002243.

Mascitti:2021:DPS

[MCMA21] Agostino Mascitti, Tommaso Cucinotta, Mauro Marinoni, and Luca Abeni. Dynamic partitioned scheduling of realtime tasks on ARM big.LITTLE architectures. The Journal of Systems and Software, 173(??):??, March 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302764.

Mostafa:2025:FTI

[MCRM25] Sakib Mostafa, Shamse Tasnim Cynthia, Banani Roy, and Debajyoti Mondal. Feature transformation for improved software bug detection and commit classification. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002498.

Mora-Cantallops:2020:CNA

[MCSAGB20] Marçal Mora-Cantallops, Salvador Sánchez-Alonso, and Elena García-Barriocanal. A complex network analysis of the Comprehensive R Archive Network (CRAN) package ecosystem. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301709.

Mahmud:2023:AIA

[MCY23]  Tarek  Mahmud,

Meiru Che, and Guowei Yang. Analyzing the impact of API changes on Android apps. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000596.

Ma:2020:CFD

[MCZX20] Ping Ma, Hangyuan Cheng, Jingxuan Zhang, and Jifeng Xuan. Can this fault be detected: a study on fault detection via automated test generation. The Journal of Systems and Software, 170(??): ??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301862.

Medeiros:2025:PEA

[MDAHS25] Raul Medeiros, Oscar Díaz, Mustafa Al-Hajjaji, and Michael Schulze. Portfolio evolution analysis for SPL scoping: Unveiling the dynamics with dedicated time-series dashboards. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003030.

Martou:2025:CTT

[MDML25] Pierre Martou, Benoît Duhoux, Kim Mens, and Axel Legay. Combinatorial transition testing in dynamically adaptive systems: Implementation and test oracle. The Journal of Systems and Software, 221(??):??, March 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003042.

Motta:2023:EBR

[MdOT23] Rebeca C. Motta, Káthia M. de Oliveira, and Guilherme H. Travassos. An evidence-based roadmap for IoT software systems engineering. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000754.

Mayr-Dorn:2024:ALW

[MDRdP^{+}24] Christoph Mayr-Dorn, Cosmina-Cristina Ratiu, Luciano Marchezan de Paula, Felix Keplinger, Alexander Egyed, and Gala Walden. Actionable light-weight process guidance.

The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001092.

Meyer:2022:IDI

[MdSKD22] Vinícius Meyer, Matheus L. da Silva, Dionatrã F. Kirchoff, and Cesar A. F. De Rose. IADA: a dynamic interferenceaware cloud scheduling architecture for latency-sensitive workloads. The Journal of Systems and Software, 194(??): ??, December 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121222001698.

Mayr-Dorn:2023:PAP

[MDVB^{+}23] Christoph Mayr-Dorn, Michael Vierhauser, Stefan Bichler, Felix Keplinger, Jane Cleland-Huang, Alexander Egyed, and Thomas Mehofer. ProCon: an automated processcentric quality constraints checking framework. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300122X.

Mayr-Dorn:2023:AIE

[MDWS^{+}23] Christoph Mayr-Dorn, Mario Winterer, Christian Salomon, Doris Hohensinger, and Harald Fürschuss. Assessing industrial end-user programming of robotic production cells: a controlled experiment. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002230.

Meckenstock:2024:SLD

[Mec24] Jan-Niklas Meckenstock. Shedding light on the dark side — a systematic literature review of the issues in agile software development methodology use. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000098.

Marcilio:2020:SAG

[MFBP20] Diego Marcilio, Carlo A. Furia, Rodrigo Bonifácio, and Gustavo Pinto. SpongeBugs: Automatically generating fix suggestions in response to static code analysis warnings. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122030128X.

Meixner:2024:VMP

[MFF^{+}24] Kristof Meixner, Kevin Feichtinger, Hafiyyan Sayyid Fadhlillah, Sandra Greiner, Hannes Marcher, Rick Rabiser, and Stefan Biffl. Variability modeling of products, processes, and resources in cyber– physical production systems engineering. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000505.

Melo:2022:IMR

[MFLS22] Ana Melo, Roberta Fagundes, Valentina Lenarduzzi, and Wylliams Barbosa Santos. Identification and measurement of Requirements Technical Debt in software development: a systematic literature review. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001650.

Mastropaolo:2024:LSG

[MFPB24] Antonio Mastropaolo, Valentina Ferrari, Luca Pascarella, and Gabriele Bavota. Log statements generation via deep learning: Widening the support provided to developers. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003424.

Matsubara:2022:SSM

[MGSC22] Patrícia Gomes Fernandes Matsubara, Bruno Freitas Gadelha, Igor Steinmacher, and Tayana Uchôa Conte. SEXTAMT: a systematic map to navigate the wide seas of factors affecting expert judgment software estimates. The Journal of Systems and Software, 185(??):??, March 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002429.

Mokadem:2020:DRS

[MH20] Riad Mokadem and Abdelkader Hameurlain. A data replication strategy with tenant performance and provider economic profit guarantees in Cloud data centers. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302213.

Maqbool:2024:PEE

[MH24] Bilal Maqbool and Sebastian Herold. Potential effectiveness and efficiency issues in usability evaluation within digital health: a systematic literature review. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002765.

Miryeganeh:2021:GUG

[MHH21] Nima Miryeganeh, Sepehr Hashtroudi, and Hadi Hemmati. GloBug: using global data in fault localization. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000583.

Mukhtar:2022:SDP

[MHJW22] Adil Mukhtar, Birgit Hofer, Dietmar Jannach, and Franz Wotawa. Spreadsheet debugging: the perils of tool over-reliance. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002168.

Mukhtar:2023:ESF

[MHJW23] Adil Mukhtar, Birgit Hofer, Dietmar Jannach, and Franz Wotawa. Explaining software fault predictions to spreadsheet users. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000717.

Mi:2022:TUV

[MHOM22] Qing Mi, Yiqun Hao, Liwei Ou, and Wei Ma. Towards using visual, semantic and structural features to improve code readability classification. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001467.

Mondal:2025:FCT

[MHR^{+}25] Amit Kumar Mondal, Mainul Hossain, Chanchal K. Roy, Banani Roy, and Kevin A. Schneider. FSECAM: a contextual thematic approach for linking feature to multi-level software architectural components. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002899.

Medeiros:2023:TMC

[MICV23] Nadia Medeiros, Naghmeh Ivaki, Pedro Costa, and Marco Vieira. Trustworthiness models to categorize and prioritize code for security improvement. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122300016X.

Muhammad:2023:HFD

[MKB23] Amna Pir Muhammad, Eric Knauss, and Jonas Bärgman. Human factors in developing automated vehicles: a requirements engineering perspective. The Journal of Systems and Software, 205(??):??, November 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223002054.

Mokos:2025:MBS

[MKB25] Konstantinos Mokos, Panagiotis Katsaros, and Preben Bohn. Model-based safety analysis of requirement specifications. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002759.

Moseler:2021:DFC

[MLBD21] Oliver Moseler, Felix Lemmer, Sebastian Baltes, and Stephan Diehl. On the diversity and frequency of code related to mathematical formulas in real-world Java projects. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302533.

Moreno-Lumbreras:2024:ICM

[MLGBRC24] David Moreno-Lumbreras, Jesus M. Gonzalez-Barahona, Gregorio Robles, and Valerio Cosentino. The influence of the city metaphor and its derivates in software visualization. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000281.

Molleri:2023:BIF

[MLJ23] Jefferson Seide Molléri, Casper Lassenius, and Magne Jørgensen. Backsourcing of IT with focus on software development

— a systematic literature review. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001668.

Marcen:2020:TLR

[MLPC20] Ana C. Marcén, Raúl Lapeña, Óscar Pastor, and Carlos Cetina. Traceability link recovery between requirements and models using an evolutionary algorithm guided by a learning to rank algorithm: Train control and management case.

The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300029.

Malavolta:2021:MGA

[MLS^{+}21] Ivano Malavolta, Grace A. Lewis, Bradley Schmerl, Patricia Lago, and David Garlan. Mining guidelines for architecting robotics software. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000662.

Ma:2025:STF

[MLYL25] Dongyu Ma, Zeyu Liang, Luming Yin, and Hongliang Liang. Symbolic testing of floating-point bugs and exceptions. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122400270X;
 https://papers.
ssrn.com/sol3/papers.
cfm?abstract_id=
4797965.

Mehta:2022:MIL

[MMB22] Anju Mehta, Nikhil Mehta, and Ishaan Bindal. Maximizing integrative learning in software development teams: a systematic review of key drivers and future research agenda. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000838.

Maniriho:2024:SLR

[MMC24] Pascal Maniriho, Abdun Naser Mahmood, and Mohammad Jabed Morshed Chowdhury. A systematic literature review on Windows malware detection: Techniques, research issues, and future directions. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003163.

Martou:2023:TSG

[MMDL23] Pierre Martou, Kim Mens, Benoît Duhoux, and Axel Legay. Test scenario generation for featurebased context-oriented software systems. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002461.

Malavolta:2022:SAA

[MMO22] Ivano Malavolta, Henry Muccini, and Ipek Ozkaya. Software architecture and artificial intelligence. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001376.

Michelon:2023:SBF

[MMSM^{+}23] Gabriela K. Michelon, Jabier Martinez, Bruno Sotto-Mayor, Aitor Arrieta, Wesley K. G. Assunção, Rui Abreu, and Alexander Egyed. Spectrum-based feature localization for families of systems. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002084.

Mertz:2021:TDF

[MN21a] Jhonny Mertz and Ingrid Nunes. Tigris: a DSL and framework for monitoring software systems at runtime.

The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000601.

Mondal:2021:HHC

[MN21b] Shouvick Mondal and Rupesh Nasre. Hansie: Hybrid and consensus regression test prioritization. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302405.

Mertz:2023:SRM

[MN23] Jhonny Mertz and Ingrid Nunes. Software runtime monitoring with adaptive sampling rate to collect representative samples of execution traces. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001036.

Macho:2024:DAV

[MOP24] Christian Macho, Fabian Oraze, and Martin Pinzger. DValidator : an approach for validating dependencies in build configurations. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003114.

Munoz:2023:TNF

[MPFB23] Daniel-Jesus Munoz, Mónica Pinto, Lidia Fuentes, and Don Batory. Transforming numerical feature models into propositional formulas and the universal variability language. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001656.

Mahmud:2022:IEI

[MPGB22] Redowan Mahmud, Samodha Pallewatta, Mohammad Goudarzi, and Rajkumar Buyya. iFogSim2: an extended iFogSim simulator for mobility, clustering, and microservice management in edge and fog computing environments. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000863.

Massacci:2024:ACE

[MPP24] Fabio Massacci, Aurora Papotti, and Ranindya Paramitha. Addressing combinatorial experiments and scarcity of subjects by provably orthogonal and crossover experimental designs. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000335.

Mariani:2020:PFM

[MPRX20] Leonardo Mariani, Mauro Pezzè, Oliviero Riganelli, and Rui Xin. Predicting failures in multi-tier distributed systems. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302389.

Mondal:2024:RIR

[MR24] Saikat Mondal and Banani Roy. Reproducibility of issues reported in Stack Overflow questions: Challenges, impact & estimation. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002036.

Marchezan:2022:SPL

[MRA^{+}22a] Luciano Marchezan, Elder Rodrigues, Wesley Klewerton Guez Assunção, Maicon Bernardino, Fábio Paulo Basso, and João Carbonell. Software product line scoping: a systematic literature review. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002673.

Moslehi:2022:USA

[MRA22b] Parisa Moslehi, Juergen Rilling, and Bram Adams. A user survey on the adoption of crowdbased software engineering instructional screencasts by the new generation of software developers. The Journal of Systems and Software, 185(??):??, March 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002405.

Mazuera-Rozo:2022:TSW

[MREVEA^{+}22] Alejandro Mazuera-Rozo, Camilo Escobar-Velásquez, Juan Espitia-Acero, David Vega-Guzmán, Catia Trubiani, Mario Linares-Vásquez, and Gabriele Bavota. Taxonomy of security weaknesses in Java and Kotlin Android apps. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000103.

Myllyaho:2021:SLR

[MRM^{+}21] Lalli Myllyaho, Mikko Raatikainen, Tomi Männistö, Tommi Mikkonen, and Jukka K. Nurminen. Systematic literature review of validation methods for AI systems. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001473.

Myllyaho:2022:MFT

[MRM^{+}22] Lalli Myllyaho, Mikko Raatikainen, Tomi Männistö, Jukka K. Nurminen, and Tommi Mikkonen. On misbehaviour and fault tolerance in machine learning systems. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100193X.

Mostaeen:2020:MLB

[MRR^{+}20] Golam Mostaeen, Banani Roy, Chanchal K. Roy, Kevin Schneider, and Jeffrey Svajlenko. A machine learning based framework for code clone validation. The Journal of Systems and Software, 169(??): ??, November 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220301394.

Mondal:2020:SCR

[MRS20] Manishankar Mondal, Chanchal K. Roy, and Kevin A. Schneider. A survey on clone refactoring and tracking. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302031.

Mandli:2025:CGC

[MRS25] Abhinav Reddy Mandli, Saurabhsingh Rajput, and Tushar Sharma. COMET:

Generating commit messages using delta graph context representation. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003510.

Macedo:2024:ZBE

[MRVS24] José Nuno Macedo, Emanuel Rodrigues, Marcos Viera, and João Saraiva. Zipperbased embedding of strategic attribute grammars. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000189.

Madeyski:2025:PTF

[MS25] Lech Madeyski and Szymon Stradowski. Predicting test failures induced by software defects: a lightweight alternative to software defect prediction and its industrial application. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000287.

Mumtaz:2021:SMS

[MSB21] Haris Mumtaz, Paramvir Singh, and Kelly Blincoe. A systematic mapping study on architectural smells detection. The Journal of Systems and Software, 173(??):??, March 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302752.

Mumtaz:2023:IRO

[MSB23] Haris Mumtaz, Paramvir Singh, and Kelly Blincoe. Identifying refactoring opportunities for large packages by analyzing maintainability characteristics in Java OSS. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001127.

Mass:2020:SOS

[MSC20] Jakob Mass, Satish Narayana Srirama, and Chii Chang. STEP-ONE: Simulated testbed for edgefog processes based on the opportunistic network environment simulator. The Journal of Systems and Software, 166(??):??, August 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300686.

Murphy:2025:STL

[MSDC25] Logan Murphy, Mahmood Saifi, Alessio Di Sandro, and Marsha Chechik. A structural taxonomy for lifted software product line analyses. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003248.

Mohamad:2024:MSE

[MSKS24] Mazen Mohamad, Jan-Philipp Steghöfer, Eric Knauss, and Riccardo Scandariato. Managing security evidence in safety-critical organizations. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001274.

Maartensson:2021:EEE

[MSMB21] Torvald Mårtensson, Daniel Ståhl, Antonio Martini, and Jan Bosch. Efficient and effective exploratory testing of large-scale software systems. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302806.

Mondal:2022:SSA

[MSRR22] Amit Kumar Mondal, Kevin A. Schneider, Banani Roy, and Chanchal K. Roy. A survey of software architectural change detection and categorization techniques. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001819.

Meding:2021:MMC

[MSS21] Wilhelm Meding, Miroslaw Staron, and Ola Söder. MeTeaM — a method for characterizing mature software metrics teams. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001035.

Meijer:2024:EEA

[MTA24] Willem Meijer, Catia Trubiani, and Aldeida Aleti. Experimental evaluation of architectural software performance design patterns in microservices. The Journal of Systems and Software, 218 (??):??, December 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002279.

Machuca-Villegas:2022:PHS

[MVGHPT22] Liliana Machuca-Villegas, Gloria Piedad Gasca-Hurtado, Solbey Morillo Puente, and Luz Marcela Restrepo Tamayo. Perceptions of the human and social factors that influence the productivity of software development teams in Colombia: a statistical analysis. The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001224.

Motogna:2023:EIE

[MVŞ23] Simona Motogna, Andreea Vescan, and Camelia Şerban. Empirical investigation in embedded systems: Quality attributes in general, maintainability in particular. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000730.

Maggi:2025:ECT

[MVSV25] Kevin Maggi, Roberto Verdecchia, Leonardo Scommegna, and Enrico Vicario. Evolution of code technical debt in microservices architectures. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003455.

Mauser:2025:CPA

[MW25] Lucas Mauser and Stefan Wagner. Centralization potential of automotive E/E architectures. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002644.

Mahmoud:2024:AUT

[MWD24]  May  Mahmoud,

Robert J. Walker, and Jörg Denzinger. API usage templates via structural generalization.

The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000177.

Mendes:2020:WUS

[MWFK20] Emilia Mendes, Claes Wohlin, Katia Felizardo, and Marcos Kalinowski. When to update systematic literature reviews in software engineering. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300856.

Mao:2022:UBP

[MWY^{+}22] Qun Mao, Weiwei Wang, Feng You, Ruilian Zhao, and Zheng Li. User behavior pattern mining and reuse across similar Android apps. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001825.

Moreschini:2024:ECT

[MYH^{+}24] Sergio Moreschini, Elham Younesian, David Hästbacka, Michele Albano, Jiří Hošek, and Davide Taibi. Edge to cloud tools: a multivocal literature review. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003370.

Newman:2020:GSS

[NAD^{+}20] Christian D. Newman, Reem S. Al-Suhaibani, Michael J. Decker, Anthony Peruma, Dishant Kaushik, Mohamed Wiem Mkaouer, and Emily Hill. On the generation, structure, and semantics of grammar patterns in source code identifiers. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301680.

Nashaat:2025:ETB

[NAEAK25] Mona Nashaat, Reem Amin, Ahmad Hosny Eid, and Rabab F. Abdel-Kader. An enhanced transformerbased framework for interpretable code clone detection. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000159.

Nikolaidis:2024:EOS

[NAV^{+}24] Nikolaos Nikolaidis, Elvira-Maria Arvanitou, Christina Volioti, Theodore Maikantis, Apostolos Ampatzoglou, Daniel Feitosa, Alexander Chatzigeorgiou, and Phillipe Krief. Eclipse Open Smart-CLIDE: an end-to-end framework for facilitating service reuse in cloud development. The Journal of Systems and Software, 207 (??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002728.

Nazar:2022:FBS

[NAZ22] Najam Nazar, Aldeida Aleti, and Yaokun Zheng. Featurebased software design pattern detection. The Journal of Systems and Software, 185(??): ??, March 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221002624.

Nunez:2020:CMD

[NBGC20a] Manuel Núñez, Daniel Bonhaure, Magalí González, and Luca Cernuzzi. Corrigendum to “A model-driven approach for the development of native mobile applications focusing on the data layer” [Journal of Systems and Software volume 161 (March 2020)]. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301928.

See [NBGC20b].

Nunez:2020:MDA

[NBGC20b] Manuel Núñez, Daniel Bonhaure, Magalí González, and Luca Cernuzzi. A modeldriven approach for the development of native mobile applications focusing on the data layer. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121219302638.

See corrigendum [NBGC20a].

Nazir:2024:AME

[NBP24] Roger Nazir, Alessio Bucaioni, and Patrizio Pelliccione. Architecting MLenabled systems: Challenges, best practices, and design decisions. The Journal of Systems and Software, 207 (??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002558.

Nguyen:2024:GCB

[NDD^{+}24] Phuong T. Nguyen, Juri Di Rocco, Claudio Di Sipio, Riccardo Rubei, Davide Di Ruscio, and Massimiliano Di Penta. GPTSniffer: a CodeBERTbased classifier to detect source code written by Chat-GPT. The Journal of Systems and Software, 214(??): ??, August 2024. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224001043.

Nguyen:2020:CSS

[NDDD20] Phuong T. Nguyen, Juri Di Rocco, Davide Di Ruscio, and Massimiliano Di Penta. CrossRec: Supporting software developers by recommending thirdparty libraries. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302341.

Nakamura:2022:WFA

[NdOdO^{+}22] Walter T. Nakamura, Edson Cesar de Oliveira, Elaine H. T. de Oliveira, David Redmiles, and Tayana Conte. What factors affect the UX in mobile apps? A systematic mapping study on the analysis of app store reviews. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001509.

Nguyen:2021:CNN

[NDP^{+}21] Phuong T. Nguyen, Davide Di Ruscio, Alfonso Pierantonio, Juri Di Rocco, and Ludovico Iovino. Convolutional neural networks for enhanced classification mechanisms of metamodels. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302508.

Napoleao:2021:SRK

[NdSR^{+}21] Bianca Minetto Napoleão, Érica Ferreira de Souza, Glauco Antonio Ruiz, Katia Romero Felizardo, Giovani Volnei Meinerz, and Nandamudi Lankalapalli Vijaykumar.

Synthesizing researches on knowledge management and agile software development using the meta-ethnography method. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000704.

Noferesti:2024:EES

[NEJ24] Morteza Noferesti and Naser Ezzati-Jivan. Enhancing empirical software performance engineering research with kernel-level events: a comprehensive system tracing approach. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001626.

Newton:2024:UPC

[NF24] Olivia B. Newton and Stephen M. Fiore. Understanding participation and corporatization in service of diversity in free/libre and open source software development projects. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002085.

Nosrati:2020:ULI

[NHA20] M. Nosrati, H. Haghighi, and M. Vahidi Asl. Using likely invariants for test data generation. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300315.

Nasir:2022:UIN

[NIJ22] Muhammad Nasir, Naveed Ikram, and Zakia Jalil. Usability inspection: Novice crowd inspectors versus expert. The Journal of Systems and Software, 183(??): ??, January 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221002193.

Nagaraju:2020:OTA

[NJF20] Vidhyashree Nagaraju, Chathuri Jayasinghe, and Lance Fiondella. Optimal test activity allocation for covariate software reliability and security models. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030114X.

Ni:2020:ABF

[NLS^{+}20] Zhen Ni, Bin Li, Xiaobing Sun, Tianhao Chen, Ben Tang, and Xinchen Shi. Analyzing bug fix for automatic bug cause classification. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300200.

Ndukwe:2023:HVS

[NLTM23] Ifeanyi G. Ndukwe, Sherlock A. Licorish, Amjed Tahir, and Stephen G. MacDonell. How have views on Software Quality differed over time? Research and practice viewpoints. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200200X.

Nashaat:2025:RSD

[NM25] Mona Nashaat and James Miller. Refining software defect prediction through attentive neural models for code understanding. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003108.

Nardone:2020:OBE

[NMG^{+}20] Roberto Nardone, Stefano Marrone, Ugo Gentile, Aniello Amato, Gregorio Barberio, Massimo Benerecetti, Renato De Guglielmo, Beniamino Di Martino, Nicola Mazzocca, Adriano Peron, Gaetano Pisani, Luigi Velardi, and Valeria Vittorini. An OSLC-based environment for systemlevel functional testing of ERTMS/ ETCS controllers. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302523.

Nadim:2022:EPC

[NMRS22] Md Nadim, Manishankar Mondal, Chanchal K. Roy, and Kevin A. Schneider. Evaluating the performance of clone detection tools in detecting cloned co-change candidates. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000085.

Nguyen:2023:AEA

[NMT^{+}23] Son Nguyen, Cuong Tran Manh, Kien T. Tran, Tan M. Nguyen, Thu-Trang Nguyen, Kien-Tuan Ngo, and Hieu Dinh Vo. ARist: an effective API argument recommendation approach. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001814.

Neto:2025:AME

[NMV25] Nuno Neto, Rolando Martins, and Luís Veiga. Atlas, a modular and efficient open-source BFT framework. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003613.

Nesic:2021:PLA

[NNG21] Damir Nesić, Mattias Nyberg, and Barbara Gallina. Product-line assurance cases from contract-based design. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121221000194.

Nguyen:2024:CCL

[NNV^{+}24] Son Nguyen, Thu-Trang Nguyen, Thanh Trong Vu, Thanh-Dat Do, Kien-Tuan Ngo, and Hieu Dinh Vo. Code-centric learning-based justin-time vulnerability detection. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000578.

Nijholt:2025:RDP

[NQT25] Max Nijholt, Giovanni Quattrocchi, and Damian Andrew Tamburri. Runtime defect prediction of industrial business processes: a focused look at real-life SAP systems. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003509.

Nocera:2025:SEE

[NRFS25] Sabato Nocera, Simone Romano, Rita Francese, and Giuseppe Scanniello. Software engineering education: Results from a training intervention based on Sonar-Cloud when developing web apps. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003522.

Nafi:2020:UCL

[NRRS20] Kawser Wazed Nafi, Banani Roy, Chanchal K. Roy, and Kevin A. Schneider. A universal cross language software similarity detector for open source software categorization. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302651.

Nasab:2021:AIS

[NSL^{+}21] Ali Rezaei Nasab, Mojtaba Shahin, Peng Liang, Mohammad Ehsan Basiri, Seyed Ali Hoseyni Raviz, Hourieh Khalajzadeh, Muhammad Waseem, and Amineh Naseri. Automated identification of security discussions in microservices systems: Industrial surveys and experiments. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001436.

Nasab:2023:ESS

[NSR^{+}23] Ali Rezaei Nasab, Mojtaba Shahin, Seyed Ali Hoseyni Raviz, Peng Liang, Amir Mashmool, and Valentina Lenarduzzi. An empirical study of security practices for microservices systems. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002394.

Ntentos:2025:UML

[NWZ25] Evangelos Ntentos, Stephen John Warnett, and Uwe Zdun. On the understandability of machine learning practices in deep learning and reinforcement learning based systems. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000111.

Nguyen:2025:APR

[NZA^{+}25] Thu-Trang Nguyen, Xiao-Yi Zhang, Paolo Arcaini, Fuyuki Ishikawa, and Hieu Dinh Vo. Automated program repair for variability bugs in software product line systems. The Journal of Systems and Software, 221(??):??, March 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224001973.

Ouni:2023:ISC

[OAH^{+}23] Ali Ouni, Eman Abdullah AlOmar, Oumayma Hamdi, Mel Ó Cinnéide, Mohamed Wiem Mkaouer, and Mohamed Aymen Saied. On the impact of single and cooccurrent refactorings on quality attributes in Android applications. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002121.

O:2023:ASE

[OBCR23] Martín G. Salido O., Gilberto Borrego, Ramón René Palacio Cinco, and Luis-Felipe Rodríguez. Agile software engineers’ affective states, their performance and software quality: a systematic mapping review. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001954.

Odu:2025:AIA

[OBW^{+}25] Oluwafemi Odu, Alvine B. Belle, Song Wang, Segla Kpodjedo, Timothy C. Lethbridge, and Hadi Hemmati. Automatic instantiation of assurance cases from patterns using large language models. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000214.

Olsson:2022:AMS

[OEW22] Tobias Olsson, Morgan Ericsson, and Anna Wingkvist. To automatically map source code entities to architectural modules with Naive Bayes. The Journal of Systems and Software, 183(??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001928.

Osei:2025:ABW

[OHM25] Samuel Banning Osei, Rubing Huang, and Zhongchen Ma. An attention-based wide and deep neural network for reentrancy vulnerability detection in smart contracts. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000299.

OBrien:2025:AGB

[OKG25] Lyndsey O’Brien, Tanjila Kanij, and John Grundy. Assessing gender bias in the software used in computer science and software engineering education. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002693.

Olianas:2025:STO

[OLR^{+}25] Dario Olianas, Maurizio Leotta, Filippo Ricca, Matteo Biagiola, and Paolo Tonella. STILE: a tool for optimizing E2E web test scripts parallelization. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003480.

Openja:2022:TDF

[OMA^{+}22] Moses Openja, Mohammad Mehdi Morovati, Le An, Foutse Khomh, and Mouna Abidi. Technical debts and faults in open-source quantum software systems: an empirical study. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001480.

Orts:2024:QCO

[OPF24] Francisco Orts, Remigijus Paulavičius, and Ernestas Filatovas. Quantum circuit optimization of an integer divider. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001365.

Ozener:2020:EFM

[ÖS20] O. Örsan Özener and Hasan Sözer. An effective formulation of the multi-criteria test suite minimization problem. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301059.

Overeem:2021:ECE

[OSJB21] Michiel Overeem, Marten Spoor, Slinger Jansen, and Sjaak Brinkkemper. An empirical characterization of event sourced systems and their schema evolution — lessons from industry. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121221000674.

Oliveira:2023:SLR

[OSM^{+}23] Delano Oliveira, Reydne Santos, Fernanda Madeiral, Hidehiko Masuhara, and Fernando Castor. A systematic literature review on the impact of formatting elements on code legibility. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001231.

Ouriques:2023:RKB

[OWGS23] Raquel Ouriques, Krzysztof Wnuk, Tony Gorschek, and Richard Berntsson Svensson. The role of knowledgebased resources in Agile Software Development contexts. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002485.

Opdebeeck:2021:PSV

[OZVRD21] Ruben Opdebeeck, Ahmed Zerouali, Camilo Velázquez-Rodríguez, and Coen De Roover. On the practice of semantic versioning for Ansible galaxy roles: an empirical study and a change classification model. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001564.

Pereira:2021:LSC

[PAM^{+}21] Juliana Alves Pereira, Mathieu Acher, Hugo Martin, Jean-Marc Jézéquel, Goetz Botterweck, and Anthony Ventresque. Learning software configuration spaces: a systematic literature review. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001412.

Poy:2025:IEC

[PÁngeles MoragaGC25] Olivia Poy, Ma.

Ángeles Moraga, Félix García, and Coral Calero. Impact on energy consumption of design patterns, code smells and refactoring techniques: a systematic mapping study. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003479.

Patrick:2020:ESR

[Pat20] Matthew T. Patrick.

Exploring software reusability metrics with Q&A forum data. The Journal of Systems and Software, 168(??): ??, October 2020.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301187.

Pasqua:2023:EES

[PBC^{+}23]  Michele  Pasqua,

Andrea Benini, Filippo Contro, Marco Crosara, Mila Dalla Preda, and Mariano Ceccato. Enhancing Ethereum smartcontracts static analysis by computing a precise Control-Flow Graph of Ethereum bytecode. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000481.

Pakonen:2024:EVP

[PBV24] Antti Pakonen, Igor Buzhinsky, and Valeriy Vyatkin. Evaluation of visual property specification languages based on practical model-checking experience. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001985.

Pauzi:2023:ANL

[PC23] Zaki Pauzi and Andrea Capiluppi. Applications of natural language processing in software traceability: a systematic mapping study. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000110.

Pereira:2020:SEL

[PCC^{+}20] Rui Pereira, Tiago Carção, Marco Couto, Jácome Cunha, João Paulo Fernandes, and João Saraiva. SPELLing out energy leaks: Aiding developers locate energy inefficient code. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302377.

Perez-Castillo:2023:DAQ

[PCJNP23] Ricardo Pérez-Castillo, Luis Jiménez-Navajas, and Mario Piattini. Dynamic analysis of quantum annealing programs. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122300078X.

Pecorelli:2020:LEA

[PDDD20] Fabiano Pecorelli, Dario Di Nucci, Coen De Roover, and Andrea De Lucia. A large empirical assessment of the role of data balancing in machinelearning-based code smell detection. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301448.

Perez:2020:SLR

[PDGMT20] Jorge Pérez, Jessica Díaz, Javier Garcia-Martin, and Bernardo Tabuenca. Systematic literature reviews in software engineeringenhancement of the study selection process using Cohen’s Kappa statistic. The Journal of Systems and Software, 168(??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301217.

Pinto:2023:CDD

[PdS23] Gustavo Pinto and Alberto de Souza. Cognitive driven development helps software teams to keep code units under the limit! The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122300225X.

Peixoto:2023:PBS

[PFC^{+}23] Mariana Peixoto, Dayse Ferreira, Mateus Cavalcanti, Carla Silva, Jéssyka Vilela, João Araújo, and Tony Gorschek. The perspective of Brazilian software developers on data privacy. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001996.

Paiva:2021:ASE

[PFdMF21] Débora Maria Barroso Paiva, André Pimenta Freire, and Renata Pontin de Mattos Fortes. Accessibility and software engineering processes: a systematic literature review. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302168.

Pigazzini:2021:SCB

[PFW21] Ilaria Pigazzini, Francesca Arcelli Fontana, and Bartosz Walter. A study on correlations between architectural smells and design patterns. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000819.

Pishchukhina:2024:DCM

[PGR24] Olga Pishchukhina, Daria Gordieieva, and Austen Rainer. Delivering computing module for the large part-time software development class from pre- to post-pandemic: an online learning experience. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000025.

Pekaric:2023:SRS

[PGW^{+}23] Irdin Pekaric, Raffaela Groner, Thomas Witte, Jubril Gbolahan Adigun, Alexander Raschke, Michael Felderer, and Matthias Tichy. A systematic review on security and safety of selfadaptive systems. The Journal of Systems and Software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223001115.

Panahandeh:2024:SAD

[PHLHM24] Mahsa Panahandeh, Abdelwahab Hamou-Lhadj, Mohammad Hamdaqa, and James Miller.

ServiceAnomaly: an anomaly detection approach in microservices using distributed traces and profiling metrics. The Journal of Systems and Software, 209(??): ??, March 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223003126.

Pietrantuono:2020:TRA

[Pie20] Roberto Pietrantuono. On the testing resource allocation problem: Research trends and perspectives. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302365.

Pallewatta:2024:MFS

[PKB24] Samodha Pallewatta, Vassilis Kostakos, and Rajkumar Buyya. MicroFog: a framework for scalable placement of microservicesbased IoT applications in federated fog environments. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003059.

Papoutsoglou:2022:AOS

[PKGA22] Maria Papoutsoglou, Georgia M. Kapitsaki, Daniel German, and Lefteris Angelis. An analysis of open source software licensing questions in Stack Exchange sites. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002107.

Popoola:2025:DSS

[PKS25] Saheed Popoola, Vineela Kunapareddi, and Hazem Said. Developing and sustaining a studentdriven software solutions center — an experience report. The Journal of Systems and Software, 220(??):??, February 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224003236.

Paradis:2024:ATB

[PKT24] Carlos Paradis, Rick Kazman, and Damian Tamburri. Analyzing the Tower of Babel with Kaiaulu. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000104.

Pan:2023:ESS

[PLL^{+}23] Jiaxin Pan, Zixuan Liu, Donglin Li, Lulu Wang, and Bixin Li. An empirical study of software architecture resilience evaluation methods. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001218.

Pan:2020:ETA

[PLP^{+}20] Minxue Pan, Yifei Lu, Yu Pei, Tian Zhang, Juan Zhai, and Xuandong Li. Effective testing of Android apps using extended IFML models. The Journal of Systems and Software, 159(??): ??, January 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219302079.

Peruma:2020:CRD

[PMDN20] Anthony Peruma, Mohamed Wiem Mkaouer, Michael J. Decker, and Christian D. Newman. Contextualizing rename decisions using refactorings, commit messages, and data types. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301503.

Pontillo:2025:EIC

[PMM^{+}25] Valeria Pontillo, Luana Martins, Ivan Machado, Fabio Palomba, and Filomena Ferrucci. An empirical investigation into the capabilities of anomaly detection approaches for test smell detection. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0164121224003649.

Pradhan:2021:LSQ

[PN21] Satya Pradhan and Venky Nanniyur. Large scale quality transformation in hybrid development organizations

— a case study.

The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302284.

Pham:2025:FML

[PNND25] Nga Pham, Hung Pham Ngoc, and Anh Nguyen-Duc. Fairness for machine learning software in education: a systematic mapping study. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002887.

Palma:2022:ALQ

[POWGH22] Francis Palma, Tobias Olsson, Anna Wingkvist, and Javier Gonzalez-Huerta. Assessing the linguistic quality of REST APIs for IoT applications. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000954.

Petry:2020:MBT

[POZ20] Kleber L. Petry, Edson OliveiraJr, and Avelino F. Zorzo. Model-based testing of software product lines: Mapping study and research roadmap. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300868.

Pascarella:2020:PML

[PPB20] Luca Pascarella, Fabio Palomba, and Alberto Bacchelli. On the performance of method-level bug prediction: a negative result. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302675.

Picchiami:2025:SSM

[PPL^{+}25] Leonardo Picchiami, Maxime Parmentier, Axel Legay, Toni Mancini, and Enrico Tronci. Scaling up statistical model checking of cyber-physical systems via algorithm ensemble and parallel simulations over HPC infrastructures. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002826.

Politowski:2021:GES

[PPM^{+}21] Cristiano Politowski, Fabio Petrillo, João Eduardo Montandon, Marco Tulio Valente, and Yann-Gaël Guéhéneuc. Are game engines software frameworks? A three-perspective study. The Journal of Systems and Software, 171(??): ??, January 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220302363.

Parizi:2022:HDT

[PPMC22] Rafael Parizi, Matheus Prestes, Sabrina Marczak, and Tayana Conte. How has design thinking being used and integrated into software development activities? A systematic mapping. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000024.

Ponce:2022:SRM

[PSAB22] Francisco Ponce, Jacopo Soldani, Hernán Astudillo, and Antonio Brogi. Smells and refactorings for microservices security: a multivocal literature review. The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122200111X.

Petrescu:2023:DNE

[PSGD23] Constantin Cezar Petrescu, Sam Smith, Rafail Giavrimis, and Santanu Kumar Dash. Do names echo semantics? A large-scale study of identifiers used in C++’s named casts. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000882.

Pan:2025:MCE

[PSL25] Yue Pan, Xiuting Shao, and Chen Lyu. Measuring code efficiency optimization capabilities with ACEOB. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002942.

Paulweber:2021:ULC

[PSZ21] Philipp Paulweber, Georg Simhandl, and Uwe Zdun. On the understandability of language constructs to structure the state and behavior in abstract state machine specifications: a controlled experiment. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000844.

Palomba:2021:PEC

[PT21] Fabio Palomba and Damian Andrew Tamburri. Predicting the emergence of community smells using socio-technical metrics: a machinelearning approach. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302375.

Passler:2025:ASA

[PtBD^{+}25] Juliane Päßler, Maurice H. ter Beek, Ferruccio Damiani, Einar Broch Johnsen, and S. Lizeth Tapia Tarifa. Analysing self-adaptive systems as software product lines. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003686.

Peischl:2022:TAS

[PTW22] Bernhard Peischl, Oliver A. Tazl, and Franz Wotawa. Testing anticipatory systems: a systematic mapping study on the state of the art. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200108X.

Paltoglou:2021:ARL

[PZDG21] Katerina Paltoglou, Vassilis E. Zafeiris, N. A. Diamantidis, and E. A. Giakoumakis. Automated refactoring of legacy JavaScript code to ES6 modules. The Journal of Systems and Software, 181(??): ??, November 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001461.

Pretel:2025:MNP

[PZVN^{+}25] Elena Pretel, Luis Zhinin-Vera, Elena Navarro, Víctor López-Jaquero, and Pascual González.

MAS4DT: a novel proposal for developing Digital Twins following a multi-agent system approach. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000123.

Qu:2024:DBA

[QHC^{+}24] Yubin Qu, Song Huang, Xiang Chen, Xingya Wang, and Yongming Yao. Detection of backdoor attacks using targeted universal adversarial perturbations for deep neural networks. The Journal of Systems and Software, 207 (??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002546.

Quin:2022:RLA

[QWG22] Federico Quin, Danny Weyns, and Omid Gheibi. Reducing large adaptation spaces in selfadaptive systems using classical machine learning. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000814.

Quin:2024:BTS

[QWGS24] Federico Quin, Danny Weyns, Matthias Galster, and Camila Costa Silva. A/B testing: a systematic literature review. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000542.

Qin:2023:MMC

[QWHH23] Rong Qin, Zeyu Wang, Sheng Huang, and Luwen Huangfu. MSTIL: Multi-cue shape-aware transferable imbalance learning for effective graphic API recommendation. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000456.

Qin:2024:CPC

[QZS^{+}24] Fangyun Qin, Zheng Zheng, Yulei Sui, Siqian Gong, Zhiping Shi, and Kishor S. Trivedi. Crossproject concurrency bug prediction using domain-adversarial neural network. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001225.

Ramli:2025:RAB

[RA25] Muhammad Rusyadi Ramli and Fredrik Asplund. Reference architectures as boundary objects in the coengineering of system properties: an empirical study. The Journal of Systems and Software, 220(??):??, February 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412122400308X.

Reyes-Anastacio:2020:KCC

[RAGCSS^{+}20] Hugo G. Reyes-Anastacio, J. L. Gonzalez-Compean, Victor J. Sosa-Sosa, Jesus Carretero, and Javier Garcia-Blas. Kulla, a container-centric construction model for building infrastructureagnostic distributed and parallel applications. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301254.

Restrepo:2021:SDA

[RATS21] Luisa Restrepo, Jose Aguilar, Mauricio Toro, and Elizabeth Suescún. A sustainable-development approach for selfadaptive cyber– physical system’s life cycle: a systematic mapping study. The Journal of Systems and Software, 180(??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001072.

Rio:2023:PCS

[RB23] Américo Rio and Fernando Brito e Abreu. PHP code smells in web apps: Evolution, survival and anomalies. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000390.

Rani:2023:DCC

[RBS^{+}23] Pooja Rani, Arianna Blasi, Nataliia Stulova, Sebastiano Panichella, Alessandra Gorla, and Oscar Nierstrasz. A decade of code comment quality assessment: a systematic literature review. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001911.

Russo:2024:GAS

[RBvB^{+}24] Daniel Russo, Sebastian Baltes, Niels van Berkel, Paris Avgeriou, Fabio Calefato, Beatriz Cabrero-Daniel, Gemma Catolino, Jürgen Cito, Neil Ernst, Thomas Fritz, Hideaki Hata, Reid Holmes, Maliheh Izadi, Foutse Khomh, Mikkel Baun Kjærgaard, Grischa Liebel, Alberto Lluch Lafuente, Stefano Lambiase, Walid Maalej, Gail Murphy, Nils Brede Moe, Gabrielle O’Brien, Elda Paja, Mauro Pezzè, John Stouby Persson, Rafael Prikladnicki, Paul Ralph, Martin Robillard, Thiago Rocha Silva, Klaas-Jan Stol, Margaret-Anne Storey, Viktoria Stray, Paolo Tell, Christoph Treude, and Bogdan Vasilescu. Generative AI in software engineering must be humancentered: the Copenhagen manifesto. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001602.

Rodrigues:2022:AMC

[RCA^{+}22] Gabriel Rodrigues, Ricardo Caldas, Gabriel Araujo, Vicente de Moraes, Genaína Rodrigues, and Patrizio Pelliccione. An architecture for mission coordination of heterogeneous robots. The Journal of Systems and Software, 191(??):??, September 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222000929.

Ren:2023:PUC

[RCA^{+}23] Ranci Ren, John W.

Castro, Santiago R. Acuña, Oscar Dieste, and Silvia T. Acuña. Perceived usability of collaborative modeling tools. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002029.

Ribeiro:2022:MFS

[RCT22] Victor Vidigal Ribeiro, Daniela Soares Cruzes, and Guilherme Horta Travassos. Moderator factors of software security and performance verification. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100234X.

Rezaalipour:2023:ABA

[RF23] Mohammad Rezaalipour and Carlo A. Furia. An annotationbased approach for finding bugs in neural network programs. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412122300064X.

Raselimo:2024:SBR

[RF24a] Moeketsi Raselimo and Bernd Fischer. Spectrumbased rule- and item-level localization of faults in context-free grammars. The Journal of Systems and Software, 215(??): ??, September 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121224001122.

Rossouw:2024:GBT

[RF24b] Christoff Rossouw and Bernd Fischer. Grammarbased test suite construction using coverage-directed algorithms over LRgraphs. The Journal of Systems and Software, 214(??): ??, August 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224001134.

Razian:2020:AAA

[RFB20] Mohammadreza Razian, Mohammad Fathian, and Rajkumar Buyya. ARC: Anomalyaware robust cloudintegrated IoT service composition based on uncertainty in advertised quality of service values. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122030039X.

Razian:2022:SCD

[RFB^{+}22] Mohammadreza Razian, Mohammad Fathian, Rami Bahsoon, Adel N. Toosi, and Rajkumar Buyya. Service composition in dynamic environments: a systematic review and future directions. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000474.

Romano:2021:RBS

[RFS^{+}21] Simone Romano, Davide Fucci, Giuseppe Scanniello, Maria Teresa Baldassarre, Burak Turhan, and Natalia Juristo. On researcher bias in Software Engineering experiments. The Journal of Systems and Software, 182(??):??, December 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001655.

Rodrigues-Filho:2025:EEM

[RFSP^{+}25] Roberto Rodrigues-Filho, Iwens Sene, Barry Porter, Luiz F. Bittencourt, Fabio Kon, and Fábio M. Costa. Exploring emergent microservice evolution in elastic deployment environments. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002966.

Rouland:2025:MDF

[RHJ25] Quentin Rouland, Brahim Hamid, and Jason Jaskolka. A model-driven formal methods approach to software architectural security vulnerabilities specification and verification. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002632.

Rosa:2023:DDA

[RJ23] Wilson Rosa and Sara Jardine. Datadriven agile software cost estimation models for DHS and DoD. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001346.

Rosa:2025:NME

[RJ25] Wilson Rosa and Sara Jardine. A novel method to early agile effort estimation through functional initiatives. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0164121224003467.

Ramakrishnan:2020:PEW

[RK20] Raghu Ramakrishnan and Arvinder Kaur. Performance evaluation of web service response time probability distribution models for business process cycle time simulation. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302547.

Rigas:2024:SIA

[RLK^{+}24] Emmanouil S. Rigas, Paris Lagakis, Makis Karadimas, Evangelos Logaras, Dimitra Latsou, Magda Hatzikou, Athanasios Poulakidas, Antonis Billis, and Panagiotis D. Bamidis. Semantic interoperability for an AI-based applications platform for smart hospitals using HL7 FHIR. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001389.

Riganelli:2024:ATR

[RMM24] Oliviero Riganelli, Daniela Micucci, and Leonardo Mariani. Automatic testing of runtime enforcers with Test4Enforcers. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003448.

Ramirez-Mora:2020:DIC

[RMOGA20] Sandra L. Ramírez-Mora, Hanna Oktaba, and Helena Gómez-Adorno. Descriptions of issues and comments for predicting issue success in software projects. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301242.

Ramac:2022:PCC

[RMT^{+}22] Robert Ramač, Vladimir Mandić, Nebojša Taušan, Nicolli Rios, Sávio Freire, Boris Pérez, Camilo Castellanos, Darío Correal, Alexia Pacheco, Gustavo Lopez, Clemente Izurieta, Carolyn Seaman, and Rodrigo Spinola. Prevalence, common causes and effects of technical debt: Results from a family of surveys with the IT industry. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002119.

Riesco:2022:ITS

[RO22] Adrián Riesco and Kazuhiro Ogata.

An integrated tool set for verifying CafeOBJ specifications. The Journal of Systems and Software, 189(??):??, July 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412122200053X.

Romero-Organvidez:2024:UFM

[ROGS^{+}24] David Romero-Organvidez, José A. Galindo, Chico Sundermann, Jose-Miguel Horcas, and David Benavides. UVLHub: a feature model data repository using UVL and open science principles. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400195X.

Runeson:2021:ODE

[ROL21] Per Runeson, Thomas Olsson, and Johan Linåker. Open Data Ecosystems — an empirical investigation into an emerging industry collaboration concept. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001850.

Recupito:2024:TDA

[RPC^{+}24] Gilberto Recupito, Fabiano Pecorelli, Gemma Catolino, Valentina Lenarduzzi, Davide Taibi, Dario Di Nucci, and Fabio Palomba. Technical debt in AI-enabled systems: On the prevalence, severity, impact, and management strategies for code and architecture. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001961.

Rani:2021:HIC

[RPL^{+}21] Pooja Rani, Sebastiano Panichella, Manuel Leuenberger, Andrea Di Sorbo, and Oscar Nierstrasz. How to identify class comment types? A multilanguage approach for class comment classification. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001448.

Restuccia:2022:APR

[RPM^{+}22] Francesco Restuccia, Marco Pagani, Agostino Mascitti, Michael Barrow, Mauro Marinoni, Alessandro Biondi, Giorgio Buttazzo, and Ryan Kastner. ARTe: Providing real-time multitasking to Arduino. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100265X.

Rajpathak:2022:ADR

[RPR22] Dnyanesh Rajpathak, Prakash M. Peranandam, and S. Ramesh. Automatic development of requirement linking matrix based on semantic similarity for robust software development. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100282X.

Rosa:2023:CES

[RPS^{+}23] Giovanni Rosa, Luca Pascarella, Simone Scalabrino, Rosalia Tufano, Gabriele Bavota, Michele Lanza, and Rocco Oliveto. A comprehensive evaluation of SZZ Variants through a developer-informed oracle. The Journal of Systems and Software, 202(??):

??, August 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223001243.

Retouniotis:2025:AMC

[RPS25] Athanasios Retouniotis, Yiannis Papadopoulos, and Ioannis Sorokos. Andromeda: a model-connected framework for safety assessment and assurance. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003005.

Razavian:2023:VDA

[RPT23] Maryam Razavian, Barbara Paech, and Antony Tang. The vision of on-demand architectural knowledge systems as a decision-making companion. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002369.

Ribeiro:2024:WMC

[RRFD24] Eduardo Ribeiro, André Restivo, Hugo Sereno Ferreira, and João Pedro Dias. WAS-MICO: Micro-containers in microcontrollers with WebAssembly. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001262.

Rosiak:2021:CTC

[RSL^{+}21] Kamil Rosiak, Alexander Schlie, Lukas Linsbauer, Birgit Vogel-Heuser, and Ina Schaefer. Customtailored clone detection for IEC 61131-3 programming languages. The Journal of Systems and Software, 182(??):??, December 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001679.

Rajan:2024:DAF

[RSLC24] Sai Sathiesh Rajan, Ezekiel Soremekun, Yves Le Traon, and Sudipta Chattopadhyay. Distributionaware fairness test generation. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001353.

Ricos:2023:DSM

[RSM^{+}23] Fernando Pastor Ricós, Arend Slomp, Beatriz Marín, Pekka Aho, and Tanja E. J. Vos. Distributed state model inference for scriptless GUI testing. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000407.

Raida:2024:SCS

[RSU^{+}24]  Maliha  Noushin

Raida, Zannatun Naim Sristy, Nawshin Ulfat, Sheikh Moonwara Anjum Monisha, Md. Jubair Ibna Mostafa, and Md. Nazmul Haque. A study on classifying Stack Overflow questions based on difficulty by utilizing contextual features. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002790.

Riegler:2023:MBM

[RSVW23] Michael Riegler, Johannes Sametinger, Michael Vierhauser, and Manuel Wimmer. A model-based mode-switching framework based on security vulnerability scores. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000286.

Rahimi:2022:VAI

[RV22] Mona Rahimi and Michael Vierhauser. Visualization of aggregated information to support class-level software evolution. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001297.

Ramos-Vidal:2025:SDS

[RVAC^{+}25] Delfina Ramos-Vidal, Wesley K. G. Assunção, Alejandro Cortiñas, Miguel R. Luaces, Oscar Pedreira, and Ángeles Saavedra Places. SPL-DB-Sync: Seamless database transformation during feature-driven changes.

The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003297.

Rong:2024:VIF

[RZLC24] Yuyang Rong, Chibin Zhang, Jianzhong Liu, and Hao Chen. Valkyrie: Improving fuzzing performance through deterministic techniques. The Journal of Systems and Software, 209(??): ??, March 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223002819.

Shahidi:2022:AEM

[SAZN22] Mahnoosh Shahidi, Mehrdad Ashtiani, and Morteza Zakeri-Nasrabadi. An automated extract method refactoring approach to correct the long method code smell. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000048.

Song:2023:CVA

[SB23] Jiyoung Song and Doo-Hwan Bae. Continuous verification with acknowledged MAPE-K pattern and time logic-based slicing: a platooning system of systems case study. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002352.

Shahin:2020:ADS

[SBC20] Mojtaba Shahin, M. Ali Babar, and Muhammad Aufeef Chauhan. Architectural design space for modelling and simulation as a service: a review. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301746.

Stievenart:2023:EEQ

[SBD23] Quentin Stiévenart, David Binkley, and Coen De Roover. An empirical evaluation of quasi-static executable slices. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000614.

Smiari:2020:ERP

[SBF20] Paraskevi Smiari, Stamatia Bibi, and Daniel Feitosa. Examining the reuse potentials of IoT application frameworks. The Journal of Systems and Software, 169(??): ??, November 2020. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301515.

Silva:2023:AFM

[SBM23] Públio Silva, Carla Bezerra, and Ivan Machado. Automating feature model maintainability evaluation using machine learning techniques. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002151.

Sas:2022:ASC

[SC22] Cezar Sas and Andrea Capiluppi. Antipatterns in software classification taxonomies. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121222000826.

Somasekaram:2022:HAC

[SCB22] Premathas Somasekaram, Radu Calinescu, and Rajkumar Buyya. High-availability clusters: a taxonomy, survey, and future directions. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002806.

SantaBarletta:2024:HQA

[SCD^{+}24] Vita Santa Barletta, Danilo Caivano, Mirko De Vincentiis, Anibrata Pal, and Michele Scalera. Hybrid quantum architecture for smart city security. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121224002061.

Sarmiento-Calisaya:2024:EAR

[SCdPL24] Edgar Sarmiento-Calisaya and Julio Cesar Sampaio do Prado Leite. Early analysis of requirements using NLP and Petri-nets. The Journal of Systems and Software, 208(??):??, February 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223002960.

Shin:2021:UAS

[SCN^{+}21] Seung Yeob Shin, Karim Chaouch, Shiva Nejati, Mehrdad Sabetzadeh, Lionel C. Briand, and Frank Zimmer. Uncertaintyaware specification and analysis for hardware-in-the-loop testing of cyberphysical systems. The Journal of Systems and Software, 171(??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

sciencedirect.com/
science/article/
pii/S0164121220302132.

Silva:2025:DAT

[SCPB25] Samira Silva, Ricardo Caldas, Patrizio Pelliccione, and Antonia Bertolino. Different approaches for testing body sensor network applications. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000044.

Sharma:2021:CSD

[SELS21]  Tushar  Sharma,

Vasiliki Efstathiou, Panos Louridas, and Diomidis Spinellis. Code smell detection by deep direct-learning and transfer-learning. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121221000339.

Steidl:2023:PCD

[SFR23] Monika Steidl, Michael Felderer, and Rudolf Ramler. The pipeline for the continuous development of artificial intelligence models — current state of research and practice. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000109.

Spieker:2020:AMT

[SG20] Helge Spieker and Arnaud Gotlieb. Adaptive metamorphic testing with contextual bandits. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300558.

Silva:2022:QAT

[SGG22] Camila Costa Silva, Matthias Galster, and Fabian Gilson. A qualitative analysis of themes in instant messaging communication of software developers. The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001133.

Silva:2024:AST

[SGG24] Camila Costa Silva, Matthias Galster, and Fabian Gilson. Applying short text topic models to instant messaging communication of software developers. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001560.

Sala:2025:OPA

[SGG^{+}25] Roberto Sala, Bruno Guindani, Enrico Galimberti, Federica Filippini, Hamta Sedghani, Danilo Ardagna, Sebastián Risco, Germán Moltó, and Miguel Caballer. OSCAR-P and aMLLibrary: Profiling and predicting the performance of FaaSbased applications in computing continua. The Journal of Systems and Software, 221(??): ??, March 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0164121224003261.

Salahirad:2023:MSE

[SGM23] Alireza Salahirad, Gregory Gay, and Ehsan Mohammadi. Mapping the structure and evolution of software testing research over the past three decades. The Journal of Systems and Software, 195(??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001947.

Shin:2022:ECF

[SGV22] Min Kyung Shin, Sudipto Ghosh, and Leo R. Vijayasarathy. An empirical comparison of four Java-based regression test selection techniques. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002582.

Shahin:2022:HSS

[SGW^{+}22] Mojtaba Shahin, Christabel Gonsalvez, Jon Whittle, Chunyang Chen, Li Li, and Xin Xia. How secondary school girls perceive Computational Thinking practices through collaborative programming with the micro:bit. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002041.

Sheghdara:2020:ARA

[SH20] Maged Sheghdara and Jameleddine Hassine. Automatic retrieval and analysis of high availability scenarios from system execution traces: a case study on hot standby router protocol. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412121930264X.

Shastri:2021:RPM

[SHA21] Yogeshwar Shastri, Rashina Hoda, and Robert Amor. The role of the project manager in agile software development projects. The Journal of Systems and Software, 173(??):??, March 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302612.

Sahar:2021:HIR

[SHB21] Hareem Sahar, Abram Hindle, and Cor-Paul Bezemer. How are issue reports discussed in Gitter chat rooms? The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302429.

Sembey:2024:ETH

[SHG24] Ruchi Sembey, Rashina Hoda, and John Grundy. Emerging technologies in higher education assessment and feedback practices: a systematic literature review. The Journal of Systems and Software, 211(??):??, May 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000311.

Sedki:2024:AAM

[SHLM24] Issam Sedki, Abdelwahab Hamou-Lhadj, and Otmane Ait Mohamed. AML: an accuracy metric model for effective evaluation of log parsing techniques. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001997.

Seifermann:2022:DVA

[SHWR22] Stephan Seifermann, Robert Heinrich, Dominik Werle, and Ralf Reussner. Detecting violations of access control and information flow policies in data flow diagrams. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002351.

Sun:2022:OMD

[SJC^{+}22] Yongqian Sun, Kunlin Jian, Liyue Cui, Guifei Jiang, Shenglin Zhang, Yuzhi Zhang, and Dan Pei. Online malicious domain name detection with partial labels for large-scale dependable systems. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200067X.

Sarker:2024:MVL

[SJH^{+}24] Orvila Sarker, Asangi Jayatilaka, Sherif Haggag, Chelsea Liu, and M. Ali Babar. A multivocal literature review on challenges and critical success factors of phishing education, training and awareness. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002947.

Sarker:2025:UPC

[SJH^{+}25] Orvila Sarker, Asangi Jayatilaka, Sherif Haggag, Chelsea Liu, and M. Ali Babar. Understanding practitioners’ challenges and requirements in the design, implementation, and evaluation of anti-phishing interventions. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122500024X.

Saieva:2022:UCT

[SK22a] Anthony Saieva and Gail Kaiser. Update with care: Testing candidate bug fixes and integrating selective updates through binary rewriting. The Journal of Systems and Software, 191 (??):??, September 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001054.

Salimi:2022:VVD

[SK22b] Solmaz Salimi and Mehdi Kharrazi. VulSlicer: Vulnerability detection through code slicing. The Journal of Systems and Software, 193(??): ??, November 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222001443.

Sharma:2024:SML

[SKG^{+}24]  Tushar  Sharma,

Maria Kechagia, Stefanos Georgiou, Rohit Tiwari, Indira Vats, Hadi Moazen, and Federica Sarro. A survey on machine learning techniques applied to source code. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003291.

Shehab:2024:CTD

[SKHLS24] Mohammed A. Shehab, Wael Khreich, Abdelwahab Hamou-Lhadj, and Issam Sedki. Committime defect prediction using one-class classification. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003096.

Sabbaghi:2020:FFE

[SKP20] Arash Sabbaghi, Mohammad Reza Keyvanpour, and Saeed Parsa. FCCI: a fuzzy expert system for identifying coincidental correct test cases. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301102.

Scheuner:2020:FSP

[SL20] Joel Scheuner and Philipp Leitner. Functionas-a-service performance evaluation: a multivocal literature review. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301527.

Sadiq:2020:SUA

[SLL20] Ayesha Sadiq, Yuan-Fang Li, and Sea Ling. A survey on the use of access permissionbased specifications for program verification. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302249.

Smite:2020:VSF

[SM20a] Darja Smite and Nils Brede Moe. Vendor switching: factors that matter when engineers onboard their own replacement. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301576.

Stray:2020:UCG

[SM20b] Viktoria Stray and Nils Brede Moe. Understanding coordination in global software engineering: a mixedmethods study on the use of meetings and Slack. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301564.

Staahl:2021:MPA

[SM21] Daniel Ståhl and Torvald Mårtensson. Mob programming: From avant-garde experimentation to established practice. The Journal of Systems and Software, 180(??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100114X.

Shen:2024:CBE

[SM24] Bowen Shen and Na Meng. ConflictBench: a benchmark to evaluate software merge tools. The Journal of Systems and Software, 214(??): ??, August 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224001298.

Sobhy:2020:RTE

[SMB^{+}20] Dalia Sobhy, Leandro Minku, Rami Bahsoon, Tao Chen, and Rick Kazman. Run-time evaluation of architectures: a case study of diversification in IoT. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412121930202X.

Steglich:2023:FAD

[SMdS^{+}23] Caio Steglich, Sabrina Marczak, Rodrigo Pereira dos Santos, Luiz Guerra, Luiz Mosmann, Marina Moreira, and Marcelo Perin. Factors that affect developers’ decision to participate in a Mobile Software Ecosystem. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002030.

Smite:2023:DDM

[ŠMF^{+}23] Darja Šmite, Nils Brede Moe, Marcin Floryan, Javier Gonzalez-Huerta, Michael Dorner, and Aivars Sablis. Decentralized decisionmaking and scaled autonomy at Spotify. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223000444.

Smite:2023:WHH

[SMH^{+}23] Darja Smite, Nils Brede Moe, Jarle Hildrum, Javier Gonzalez Huerta, and Daniel Mendez. Workfrom-home is here to stay: Call for flexibility in postpandemic work policies. The Journal of Systems and Software, 195(??):

??, January 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412122200228X.

Smite:2023:FWH

[ŠMKGH23] Darja Šmite, Nils Brede Moe, Eriks Klotins, and Javier Gonzalez-Huerta. From forced working-from-home to voluntary workingfrom-anywhere: Two revolutions in telework. The Journal of Systems and Software, 195(??): ??, January 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121222001856.

Sherin:2023:QES

[SMKI23] Salman Sherin, Asmar Muqeet, Muhammad Uzair Khan, and Muhammad Zohaib Iqbal. QExplore: an exploration strategy for dynamic web applications using guided search. The Journal of Systems and Software, 195(??):

??, January 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222001881.

Suciu:2023:TPB

[SMM23] Dan Mircea Suciu, Simona Motogna, and Arthur-Jozsef Molnar. Transitioning a projectbased course between onsite and online. An experience report. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002236.

Swarna:2024:IMS

[SMVC24] Karthik Chandra Swarna, Noble Saji Mathews, Dheeraj Vagavolu, and Sridhar Chimalakonda. On the impact of multiple source code representations on software engineering tasks — an empirical study. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003369.

Suarez-Otero:2023:CCF

[SOMSCT23] Pablo Suárez-Otero, Michael J. Mior, María José Suárez-Cabal, and Javier Tuya. CoDEvo: Column family database evolution using model transformations. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001383.

Spano:2024:FRC

[Spa24] Alvise Spanò. Flexible and reversible conversion between extensible records and overloading constraints for ML. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001869.

Shen:2025:DRD

[SPZW25] Yuanyuan Shen, Manman Peng, Fan Zhang, and Qiang Wu. Data race detection via few-shot parameter-efficient fine-tuning. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003339.

Silva:2021:IRP

[SRD^{+}21] Rodrigo F. Silva, Mohammad Masudur Rahman, Carlos Eduardo Dantas, Chanchal Roy, Foutse Khomh, and Marcelo A. Maia. Improved retrieval of programming solutions with code examples using a multi-featured score. The Journal of Systems and Software, 181(??):??, November 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221001606.

Souza:2025:TEI

[SRM^{+}25] Sabrina Souza, Eriky Rodrigues, Maria Meireles, Tanara Lauschner, Leandro Carvalho, José Carlos Maldonado, and Tayana Conte. Techniques for eliciting IoT requirements: Sensorina map and mind IoT. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003674.

Siitonen:2025:CTS

[SRS^{+}25] Valtteri Siitonen, Saima Ritonummi, Markus Salo, Henri Pirkkalainen, and Saija Mauno. Coping with technostress in the software industry: Coping strategies and factors underlying their selection. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000093.

Sanchez-Rivero:2025:AGE

[SRTGA^{+}25] Javier Sanchez-Rivero, Daniel Talaván, Jose Garcia-Alonso, Antonio Ruiz-Cortés, and Juan Manuel Murillo. Automatic generation of efficient oracles: the less-than case. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002474.

Sharma:2020:EOS

[SS20] Gaurav G. Sharma and Klaas-Jan Stol. Exploring onboarding success, organizational fit, and turnover intention of software professionals. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412121930216X.

Schneider:2023:AES

[SS23] Simon Schneider and Riccardo Scandariato. Automatic extraction of security-rich dataflow diagrams for microservice applications written in Java. The Journal of Systems and Software, 202(??): ??, August 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223001176.

Shams:2023:IEU

[SSO^{+}23] Rifat Ara Shams, Mojtaba Shahin, Gillian Oliver, Harsha Perera, Jon Whittle, Arif Nurwidyantoro, and Waqar Hussain. Investigating endusers’ values in agriculture mobile applications development: an empirical study on Bangladeshi female farmers. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000432.

Szalay:2021:PHI

[SSP21] Richárd Szalay, Ábel Sinkovics, and Zoltán Porkoláb. Practical heuristics to improve precision for erroneous function argument swapping detection in C and C++. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100145X.

Shatnawi:2020:RRS

[SSS^{+}20] Anas Shatnawi, Abdelhak Seriai, Houari Sahraoui, Tewfik Ziadi, and Abderrahmene Seriai. ReSIde: Reusable service identification from software families. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301722.

Sedighiani:2021:BFBa

[SSS21a] Kavan Sedighiani, Saeed Shokrollahi, and Fereidoon Shams. BASBA: a framework for Building Adaptable Service-Based Applications. The Journal of Systems and Software, 179(??):??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221000868.

Sedighiani:2021:BFBb

[SSS21b] Kavan Sedighiani, Saeed Shokrollahi, and Fereidoon Shams. BASBA: a framework for Building Adaptable Service-Based Applications. The Journal of Systems and Software, 179(??):??, September 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221000868.

Sunman:2022:AWA

[SSS22] Nezih Sunman, Yiğit Soydan, and Hasan Sözer. Automated Web application testing driven by pre-recorded test cases. The Journal of Systems and Software, 193(??): ??, November 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412122200139X.

Sodja:2020:CPM

[SSZ20] Anton Sodja, Igor Skrjanc, and Borut Zupancic. Cyberphysical modelling in Modelica with model-reduction techniques. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219302912.

Svensson:2024:ARP

[ST24]  Richard  Bernts-

son Svensson and Richard Torkar. Not all requirements prioritization criteria are equal at all times: a quantitative analysis. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003047.

Smite:2022:CPP

[STM^{+}22] Darja Smite, Anastasiia Tkalich, Nils Brede Moe, Efi Papatheocharous, Eriks Klotins, and Marte Pettersen Buvik. Changes in perceived productivity of software engineers during COVID-19 pandemic: the voice of evidence. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002715.

Sheikhaei:2023:SUR

[STW23] Mohammad Sadegh Sheikhaei, Yuan Tian, and Shaowei Wang. A study of update request comments in Stack Overflow answer posts. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222002667.

Sotiropolos:2020:DIF

[SV20] Panagiotis Sotiropolos and Costas Vassilakis. Detection of intermittent faults in software programs through identification of suspicious shared variable access patterns. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302298.

Sotiropoulos:2025:GRB

[SV25] Panagiotis Sotiropoulos and Costas Vassilakis. A generalized, rule-based method for the detection of intermittent faults in software programs. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002723.

Stamelos:2020:OSS

[SVAGB20] Ioannis Stamelos, Iraklis Varlamis, Dimosthenis Anagnostopoulos, and Jesus M. Gonzalez-Barahona. Open source systems: Enterprise software and solutions. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300236.

Serban:2024:SEP

[SvdBHV24] Alex Serban, Koen van der Blom, Holger Hoos, and Joost Visser. Software engineering practices for machine learning — adoption, effects, and team assessment. The Journal of Systems and Software, 209(??):??, March 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003023.

Scommegna:2024:UFU

[SVV24]  Leonardo  Scom-

megna, Roberto Verdecchia, and Enrico Vicario. Unveiling faulty user sequences: a modelbased approach to test three-tier software architectures. The Journal of Systems and Software, 212(??):??, June 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400058X.

Stievenart:2021:PWA

[SVVD21] Quentin Stiévenart, Noah Van Es, Jens Van der Plas, and Coen De Roover. A parallel worklist algorithm and its exploration heuristics for static modular analyses. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001394.

Silva:2020:GSC

[SWG^{+}20] Jefferson O. Silva, Igor Wiese, Daniel M. German, Christoph Treude, Marco A. Gerosa, and Igor Steinmacher. Google Summer of Code: Student motivations and contributions. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302614.

Shafiee:2020:SVR

[SWH^{+}20] Sara Shafiee, Yves Wautelet, Lars Hvam, Enrico Sandrin, and Cipriano Forza. Scrum versus Rational Unified Process in facing the main challenges of product configuration systems development. The Journal of Systems and Software, 170(??):??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220301643.

Shen:2020:ANN

[SWZ^{+}20] Qi Shen, Shijun Wu, Yanzhen Zou, Zixiao Zhu, and Bing Xie. From API to NLI: a new interface for library reuse. The Journal of Systems and Software, 169 (??):??, November 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122030162X.

Song:2022:CEI

[SXL^{+}22] Yi Song, Xiaoyuan Xie, Quanming Liu, Xihao Zhang, and Xi Wu. A comprehensive empirical investigation on failure clustering in parallel debugging. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001455.

Sun:2023:ASV

[SYB^{+}23] Xiaobing Sun, Zhenlei Ye, Lili Bo, Xiaoxue Wu, Ying Wei, Tao Zhang, and Bin Li. Automatic software vulnerability assessment by extracting vulnerability elements. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001851.

Singjai:2022:CAA

[SZ22] Apitchaka Singjai and Uwe Zdun. Conformance assessment of Architectural Design Decisions on API endpoint designs derived from domain models. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121222001352.

Swillus:2023:SOT

[SZ23] Mark Swillus and Andy Zaidman. Sentiment overflow in the testing stack: Analyzing software testing posts on Stack Overflow. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001991.

Sharbaf:2025:CCC

[SZS25] Mohammadreza Sharbaf, Bahman Zamani, and Gerson Sunyé. CoMPers: a configurable conflict management framework for personalized collaborative modeling. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121224002711.

Silveira:2022:EUD

[SZSV22] Sofia A. M. Silveira, Luciana A. M. Zaina, Leobino N. Sampaio, and Fábio L. Verdi. On the evaluation of usability design guidelines for improving network monitoring tools interfaces. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200005X.

Teixeira:2020:LAT

[TAF^{+}20] Sergio Teixeira, Bruno Alves Agrizzi, José Gonçalves Pereira Filho, Silvana Rossetto, Isaac Simões Araújo Pereira, Patrícia Dockhorn Costa, Adriano Francisco Branco, and Ruan Rocha Martinelli. LAURA architecture: Towards a simpler way of building situation-aware and business-aware IoT applications. The Journal of Systems and Software, 161(??):??, March 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219302687.

Taipalus:2020:EDC

[Tai20] Toni Taipalus. The effects of database complexity on SQL query formulation. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300571.

Taipalus:2024:DMS

[Tai24] Toni Taipalus. Database management system performance comparisons: a systematic literature review. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121223002674.

Tighilt:2023:MSM

[TAT^{+}23] Rafik Tighilt, Manel Abdellatif, Imen Trabelsi, Loïc Madern, Naouel Moha, and Yann-Gaël Guéhéneuc. On the maintenance support for microservice-based systems through the specification and the detection of microservice antipatterns. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001504.

Turin:2023:PRC

[TBD^{+}23]  Gianluca  Turin,

Andrea Borgarelli, Simone Donetti, Ferruccio Damiani, Einar Broch Johnsen, and S. Lizeth Tapia Tarifa. Predicting resource consumption of Kubernetes container systems using resource models. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001450.

terBeek:2023:SSP

[tBS23] Maurice H. ter Beek and Ina Schaefer. Systems and software product lines of the future. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000171.

Tondel:2022:CSS

[TC22] Inger Anne Tøndel and Daniela Soares Cruzes. Continuous software security through security prioritisation meetings. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001625.

Tan:2022:SES

[TCA22] Alvin Jian Jia Tan, Chun Yong Chong, and Aldeida Aleti. E-SC4R: Explaining software clustering for remodularisation. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002491.

Takerngsaksiri:2025:PDR

[TCTL25] Wannita Takerngsaksiri, Rujikorn Charakorn, Chakkrit Tantithamthavorn, and Yuan-Fang Li. Pytester: Deep reinforcement learning for text-to-testcase generation. The Journal of Systems and Software, 224(??):??, June 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000494.

Taipalus:2021:EMR

[TGG21] Toni Taipalus, Hilkka Grahn, and Hadi Ghanbari. Error messages in relational database management systems: a comparison of effectiveness, usefulness, and user confidence. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100131X.

Tuli:2022:HAB

[TGX^{+}22] Shreshth Tuli, Sukhpal Singh Gill, Minxian Xu, Peter Garraghan, Rami Bahsoon, Schahram Dustdar, Rizos Sakellariou, Omer Rana, Rajkumar Buyya, Giuliano Casale, and Nicholas R. Jennings. HUNTER: AI based holistic resource management for sustainable cloud computing. The Journal of Systems and Software, 184(??):??, February 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221002211.

Tanilkan:2024:MCU

[TH24] Sinan Sigurd Tanilkan and Jo Erskine Hannay. Managing the changing understanding of benefits in software initiatives. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002984.

Thabet:2022:SBO

[THB22] Marwa Thabet, Brahim Hnich, and Mouhebeddine Berrima. A sampling-based online Co-Location-Resistant Virtual Machine placement strategy. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000012.

Trautsch:2020:UIT

[THG20] Fabian Trautsch, Steffen Herbold, and Jens Grabowski. Are unit and integration test definitions still valid for modern Java projects? An empirical study on opensource projects. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121219301955.

Tahmooresi:2020:SRB

[THN20] Hamed Tahmooresi, Abbas Heydarnoori, and Reza Nadri. Studying the relationship between the usage of APIs discussed in the crowd and postrelease defects. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301606.

Tsoukalas:2020:TDF

[TKSC20] Dimitrios Tsoukalas, Dionysios Kehagias, Miltiadis Siavvas, and Alexander Chatzigeorgiou. Technical debt forecasting: an empirical study on opensource repositories. The Journal of Systems and Software, 170(??):??, December 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220301904.

Tong:2023:AAT

[TLXW23] Haonan Tong, Wei Lu, Weiwei Xing, and Shihai Wang. ARRAY: Adaptive triple featureweighted transfer Naive Bayes for cross-project defect prediction. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001164.

Torquato:2025:ETB

[TMV25] Matheus Torquato, Paulo Maciel, and Marco Vieira. Evaluation of timebased virtual machine migration as moving target defense against hostbased attacks. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002668.

Tokdemir:2022:SPD

[Tok22] Gul Tokdemir. Software professionals during the COVID-19 pandemic in Turkey: Factors affecting their mental well-being and work engagement in the home-based work setting. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000450.

Torres:2023:ICC

[TOO^{+}23] Adriano Torres, Caio Oliveira, Márcio Okimoto, Diego Marcílio, Pedro Queiroga, Fernando Castor, Rodrigo Bonifácio, Edna Dias Canedo, Márcio Ribeiro, and Eduardo Monteiro. An investigation of confusing code patterns in JavaScript. The Journal of Systems and Software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223001267.

Tiwari:2020:LSA

[TPGH20] Abhishek Tiwari, Jyoti Prakash, Sascha Groß, and Christian Hammer. A large scale analysis of Android–Web hybridization. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301898.

Torkashvan:2025:SFL

[TPV25] Reza Torkashvan, Saeed Parsa, and Babak Vaziri. SBFL fault localization considering faultproneness. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000317.

Tian:2025:TCE

[TQT^{+}25] Zhenzhou Tian, Rui Qiu, Yudong Teng, Jiaze Sun, Yanping Chen, and Lingwei Chen. Towards cost-efficient vulnerability detection with cross-modal adversarial reprogramming. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000330.

Tahir:2023:TFC

[TRD^{+}23] Amjed Tahir, Shawn Rasheed, Jens Dietrich, Negar Hashemi, and Lu Zhang. Test flakiness’ causes, detection, impact and responses: a multivocal review. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002327.

Tarimci:2022:MTP

[TS22] Arzu Behiye Tarımcı and Hasan Sözer. Mutation testing of PL/SQL programs.

The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001145.

Takahashi:2021:ESS

[TSLHS21] Aoi Takahashi, Natthawute Sae-Lim, Shinpei Hayashi, and Motoshi Saeki. An extensive study on smell-aware bug localization. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000832.

Taipalus:2020:UIS

[TSP20] Toni Taipalus, Ville Seppänen, and Maritta Pirhonen. Uncertainty in information system development: Causes, effects, and coping mechanisms. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301205.

Tuma:2021:FSTa

[TST^{+}21a] Katja Tuma, Christian Sandberg, Urban Thorsson, Mathias Widman, Thomas Herpel, and Riccardo Scandariato. Finding security threats that matter: Two industrial case studies. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100100X.

Tuma:2021:FSTb

[TST^{+}21b] Katja Tuma, Christian Sandberg, Urban Thorsson, Mathias Widman, Thomas Herpel, and Riccardo Scandariato. Finding security threats that matter: Two industrial case studies. The Journal of Systems and Software, 179 (??):??, September 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122100100X.

Tuna:2024:DCR

[TST24] Erdem Tuna, Carolyn Seaman, and Eray Tüzün. Do code reviews lead to fewer code smells? The Journal of Systems and Software, 215(??):??, September 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224001468.

Tang:2023:CDL

[TTB^{+}23] Wei Tang, Mingwei Tang, Minchao Ban, Ziguo Zhao, and Mingjun Feng. CS-GVD: a deep learning approach combining sequence and graph embedding for source code vulnerability detection. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000183.

Tian:2020:CRE

[TTL20] Yuli Tian, Jeff Tian, and Ning Li. Cloud reliability and efficiency improvement via failure risk based proactive actions. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300078.

Tan:2020:BSP

[TXW^{+}20] Youshuai Tan, Sijie Xu, Zhaowei Wang, Tao Zhang, Zhou Xu, and Xiapu Luo. Bug severity prediction using question-and-answer pairs from Stack Overflow. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300480.

Uludag:2022:RSA

[UPP^{+}22] Ömer Uludağ, Pascal Philipp, Abheeshta Putta, Maria Paasivaara, Casper Lassenius, and Florian Matthes. Revealing the state of the art of largescale agile development research: a systematic mapping study. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001601.

vanBladel:2021:CST

[vBD21] Brent van Bladel and Serge Demeyer. A comparative study of test code clones and production code clones. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221000376.

Vrzakova:2020:ARC

[VBMB20] Hana Vrzakova, Andrew Begel, Lauri Mehtätalo, and Roman Bednarik. Affect recognition in code review: an insitu biometric study of reviewer’s affect. The Journal of Systems and Software, 159(??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302080.

Venson:2024:ERS

[VCB24] Elaine Venson, Bradford Clark, and Barry Boehm. The effects of required security on software development effort. The Journal of Systems and Software, 207(??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002698.

Vidoni:2022:ITD

[VCF22] Melina Vidoni, Zadia Codabux, and Fatemeh H. Fard. Infinite technical debt. The Journal of Systems and Software, 190(??): ??, August 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222000772.

Viglianisi:2020:FSB

[VCT20] Emanuele Viglianisi, Mariano Ceccato, and Paolo Tonella. A federated society of bots for smart contract testing. The Journal of Systems and Software, 168 (??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301163.

vanDinten:2024:SFP

[vDDPZ24] Imara van Dinten, Pouria Derakhshanfar, Annibale Panichella, and Andy Zaidman. The slow and the furious? Performance antipattern detection in cyber– physical systems. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002996.

vandeGriend:2025:CQC

[vdG25] Arianne Meijer van de Griend. A comparison of quantum compilers using a DAG-based or phase polynomialbased intermediate representation. The Journal of Systems and Software, 221(??):??, March 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224002681.

Vacca:2021:SLR

[VDVC21] Anna Vacca, Andrea Di Sorbo, Corrado A. Visaggio, and Gerardo Canfora. A systematic literature review of blockchain and smart contract development: Techniques, tools, and open challenges. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302818.

Vierhauser:2023:GFM

[VGS^{+}23] Michael Vierhauser, Antonio Garmendia, Marco Stadler, Manuel Wimmer, and Jane Cleland-Huang. GRuM — a flexible modeldriven runtime monitoring framework and its application to automated aerial and ground vehicles. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001280.

Vogel-Heuser:2021:IET

[VHB21] Birgit Vogel-Heuser and Fandi Bi. Interdisciplinary effects of technical debt in companies with mechatronic products — a qualitative study. The Journal of Systems and Software, 171 (??):??, January 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302119.

Vancsics:2022:FLU

[VHSB22] Béla Vancsics, Ferenc Horváth, Attila Szatmári, and Árpád Beszédes. Fault localization using function call frequencies. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001339.

Vidoni:2022:URP

[Vid22] Melina Vidoni. Understanding Roxygen package documentation in R. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000310.

Vidal:2024:VDT

[VIL24] Fernando Richter Vidal, Naghmeh Ivaki, and Nuno Laranjeiro. Vulnerability detection techniques for smart contracts: a systematic literature review. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400205X.

Vincenzi:2025:MMT

[VKB^{+}25] Auri M. R. Vincenzi, Pedro H. Kuroishi, João Bispo, Ana R. C. da Veiga, David R. C. da Mata, Francisco B. Azevedo, and Ana C. R. Paiva. METFORD

— Mutation tEsTing Framework fOR anDroid. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003765.

Vianna:2023:GLR

[VKG^{+}23] Alexandre Vianna, Fernando Kenji Kamei, Kiev Gama, Carlos Zimmerle, and João Alexandre Neto. A grey literature review on data stream processing applications testing. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001395.

Vakkuri:2021:EMI

[VKJ^{+}21] Ville Vakkuri, Kai-Kristian Kemell, Marianna Jantunen, Erika Halme, and Pekka Abrahamsson. ECCOLA

— a method for implementing ethically aligned AI systems. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001643.

Verdecchia:2021:BET

[VKLM21] Roberto Verdecchia, Philippe Kruchten, Patricia Lago, and Ivano Malavolta. Building and evaluating a theory of architectural technical debt in softwareintensive systems. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000224.

Vishnubhotla:2024:EES

[VM24a] Sai Datta Vishnubhotla and Emilia Mendes. Examining the effect of software professionals’ personality & additional capabilities on agile teams’ climate. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000992.

Vishnubhotla:2024:ERB

[VM24b] Sai Datta Vishnubhotla and Emilia Mendes. Exploring the relation between personality traits and agile team climate: Aggregating results from a twice replicated study in a telecom company. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223003321.

Vishnubhotla:2021:UPR

[VML21] Sai Datta Vishnubhotla, Emilia Mendes, and Lars Lundberg. Understanding the perceived relevance of capability measures: a survey of Agile Software Development practitioners. The Journal of Systems and Software, 180(??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001102.

Vogelsang:2020:FDA

[Vog20] Andreas Vogelsang. Feature dependencies in automotive software systems: Extent, awareness, and refactoring. The Journal of Systems and Software, 160(??):??, February 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219302328.

vanRiet:2023:OAW

[vRMG23] Jasper van Riet, Ivano Malavolta, and Taher A. Ghaleb. Optimize along the way: an industrial case study on web performance. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002692.

Vayghan:2021:KCM

[VSTK21] Leila Abdollahi Vayghan, Mohamed Aymen Saied, Maria Toeroe, and Ferhat Khendek. A Kubernetes controller for managing the availability of elastic microservice based stateful applications. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121221000212.

Valderas:2022:MEI

[VTS22] Pedro Valderas, Victoria Torres, and Estefanía Serral. Modelling and executing IoT-enhanced business processes through BPMN and microservices. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002363.

Varela-Vaca:2023:FMB

[VVBGL^{+}23] Ángel Jesús Varela-Vaca, Diana Borrego, María Teresa Gómez-López, Rafael M. Gasca, and A. German Márquez. Feature models to boost the vulnerability management process. The Journal of Systems and Software, 195(??): ??, January 2023. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222002175.

Vierhauser:2023:ADS

[VWSCH23] Michael Vierhauser, Rebekka Wohlrab, Marco Stadler, and Jane Cleland-Huang. AMon: a domainspecific language and framework for adaptive monitoring of Cyber–Physical Systems. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001832.

Wiese:2023:IMP

[WB23] Marion Wiese and Klara Borowa. IT managers’ perspective on technical debt management. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122300095X.

Wu:2020:PBA

[WC20] Jianwei Wu and James Clause. A pattern-based approach to detect and improve nondescriptive test names. The Journal of Systems and Software, 168(??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301126.

Wu:2023:UBA

[WC23] Jianwei Wu and James Clause. A uniqueness-based approach to provide descriptive JUnit test names. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002169.

Wohlrab:2023:EQA

[WCGS23] Rebekka Wohlrab, Javier Cámara, David Garlan, and Bradley Schmerl. Explaining quality attribute tradeoffs in automated planning for self-adaptive systems. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200214X.

Wattanakriengkrai:2022:GRL

[WCH^{+}22] Supatsara Wattanakriengkrai, Bodin Chinthanet, Hideaki Hata, Raula Gaikovina Kula, Christoph Treude, Jin Guo, and Kenichi Matsumoto. GitHub repositories with links to academic papers: Public access, traceability, and evolution. The Journal of Systems and Software, 183 (??):??, January 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002144.

Wang:2023:ESR

[WCH^{+}23] Yilin Wang, Xiangping Chen, Yuan Huang, Hao-Nan Zhu, Jing Bian, and Zibin Zheng. An empirical study on real bug fixes from solidity smart contract projects. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001826.

Wu:2023:SLR

[WCL23] Zhiqiang Wu, Xin Chen, and Scott Uk-Jin Lee. A systematic literature review on Androidspecific smells. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121223000729.

Wang:2025:RRF

[WCZL25] Lei Wang, Shuhan Chen, Xikai Zhang, and Jiyuan Liu. Runtime reliability fractional distribution change analytics against cloud-based systems DDoS attacks. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003091.

Wang:2022:ESC

[WCZW22] Zehao Wang, Tse-Hsun (Peter) Chen, Haoxiang Zhang, and Shaowei Wang. An empirical study on the challenges that developers encounter when developing Apache Spark applications. The Journal of Systems and Software, 194 (??):??, December 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001674.

Wu:2022:WNV

[WDF^{+}22] Jin Wu, Jian Dong, Ruili Fang, Wen Zhang, Wenwen Wang, and Decheng Zuo. WDBT: Nonvolatile memory wear characterization and mitigation for DBT systems. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122200022X.

Wang:2025:MMT

[WDL^{+}25] Bangchao Wang, Yang Deng, Ruiqi Luo, Peng Liang, and Tingting Bi. MPLinker: Multitemplate prompttuning with adversarial training for issue-commit link recovery. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000196.

Wu:2024:LTS

[WDO24] Jingchi Wu, Tadashi Dohi, and Hiroyuki Okamura. Longterm software fault prediction with wavelet shrinkage estimation. The Journal of Systems and Software, 216(??): ??, October 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224001687.

Wang:2024:CFH

[WDT^{+}24] Hechen Wang, Peter Devine, James Tizard, Seyed Reza Shahamiri, and Kelly Blincoe. Conversation in forums: How software forum posts discuss potential development insights. The Journal of Systems and Software, 215 (??):??, September 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001535.

Wang:2023:CFY

[WDXX23] Wenjie Wang, Zihan Deng, Yinxing Xue, and Yun Xu. CC-Stokener: Fast yet accurate code clone detection with semantic token. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000134.

Wei:2024:AAC

[WFM^{+}24] Ran Wei, Simon Foster, Haitao Mei, Fang Yan, Ruizhe Yang, Ibrahim Habli, Colin O Halloran, Nick Tudor, Tim Kelly, and Yakoub Nemouchi. AC-CESS: Assurance Case Centric Engineering of Safetycritical Systems. The Journal of Systems and Software, 213(??):??, July 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000773.

Weber:2021:BAN

[WFR21]  Barbara  Weber,

Thomas Fischer, and René Riedl. Brain and autonomic nervous system activity measurement in software engineering: a systematic literature review. The Journal of Systems and Software, 178(??):??, August 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000431.

Wang:2022:MSF

[WGL^{+}22] Yihui Wang, Shanquan Gao, Xingtong Li, Lei Liu, and Huaxiao Liu. Missing standard features compared with similar apps? A feature recommendation method based on the knowledge from user interface. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001364.

Wang:2024:AMB

[WGMT24] Di Wang, Matthias Galster, and Miguel Morales-Trujillo. Application monitoring for bug reproduction in webbased applications. The Journal of Systems and Software, 207(??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002297.

Wang:2025:INB

[WGMT25] Di Wang, Matthias Galster, and Miguel Morales-Trujillo. Information needs in bug reports for web applications. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002747.

Wu:2020:SSS

[WGY^{+}20] Jintao Wu, Xing Guo, Guijun Yang, Shuhui Wu, and Jianguo Wu. Substructure similarity search for engineering servicebased systems. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300509.

Wang:2023:MAR

[WHK^{+}23] Lulu Wang, Peng Hu, Xianglong Kong, Wenjie Ouyang, Bixin Li, Haixin Xu, and Tao Shao. Microservice architecture recovery based on intraservice and interservice features. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001498.

Winters:2024:TA

[Win24] Titus Winters. Thoughts on applicability. The Journal of Systems and Software, 215(??):??, September 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224001316.

Wang:2024:EAD

[WJX^{+}24]  Rongcun  Wang,

Xingyu Ji, Senlei Xu, Yuan Tian, Shujuan Jiang, and Rubing Huang. An empirical assessment of different word embedding and deep learning models for bug assignment. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224000049.

Wohlrab:2020:WHB

[WKP20] Rebekka Wohlrab, Eric Knauss, and Patrizio Pelliccione. Why and how to balance alignment and diversity of requirements engineering practices in automotive. The Journal of Systems and Software, 162(??):??, April 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302900.

Wang:2024:EER

[WL24] Tong Wang and BiXin Li. EsAr-Cost: Estimating repair costs of software architecture erosion using slice technology. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002704.

Wang:2020:EED

[WLC^{+}20] Chuanqi Wang, Yanhui Li, Lin Chen, Wenchin Huang, Yuming Zhou, and Baowen Xu. Examining the effects of developer familiarity on bug fixing. The Journal of Systems and Software, 169(??):??, November 2020. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121220301266.

Wang:2025:RMU

[WLCC25] Zehao Wang, Wei Liu, Jinfu Chen, and Tse-Hsun (Peter) Chen. RPerf: Mining user reviews using topic modeling to assist performance testing: an industrial experience report. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003273.

Wimalasooriya:2022:SMS

[WLdCM22] Chathrie Wimalasooriya, Sherlock A. Licorish, Daniel Alencar da Costa, and Stephen G. Mac-Donell. A systematic mapping study addressing the reliability of mobile applications: the need to move beyond testing reliability. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100251X.

Wei:2025:EMM

[WLH^{+}25] Xueyin Wei, Jing Li, Xudong He, Weizhou Peng, Ying Zhu, Rongbin Gu, Yunlong Zhu, and Jun Huang. Extracting microservices from monolithic applications using consistent graph enhanced Graph Transformer. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000135.

Wan:2024:CGF

[WLL^{+}24] Xiaohui Wan, Tiancheng Li, Weibin Lin, Yi Cai, and Zheng Zheng. Coverageguided fuzzing for deep reinforcement learning systems. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224000062.

Wang:2024:DSV

[WLLJ24] Pingyan Wang, Shaoying Liu, Ai Liu, and Wen Jiang. Detecting security vulnerabilities with vulnerability nets. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002972.

Waseem:2020:SMS

[WLS20] Muhammad Waseem, Peng Liang, and Mojtaba Shahin. A systematic mapping study on microservices architecture in DevOps. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302053.

Waseem:2021:DMT

[WLS^{+}21] Muhammad Waseem, Peng Liang, Mojtaba Shahin, Amleto Di Salle, and Gastón Márquez. Design, monitoring, and testing of microservices systems: the practitioners’ perspective. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001588.

Wang:2024:VCS

[WM24] Quan Wang and Ignacio Lopez Moreno. Version control of speaker recognition systems. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001675.

Wolfart:2024:VDO

[WMA^{+}24] Daniele Wolfart, Jabier Martinez, Wesley K. G. Assunção, Thelma E. Colanzi, and Alexander Egyed. Variability debt in opportunistic reuse: a multi-project field study. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224000128.

Wong:2021:BAS

[WMAL21] W. Eric Wong, Nikolaos Mittas, Elvira Maria Arvanitou, and Yihao Li. A bibliometric assessment of software engineering themes, scholars and institutions (2013–2020). The Journal of Systems and Software, 180(??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001266.

Wang:2024:SPA

[WMH^{+}24] Xiajing Wang, Rui Ma, Wei Huo, Zheng Zhang, Jinyuan He, Chaonan Zhang, and Donghai Tian. SYNTONY: Potentialaware fuzzing with particle swarm optimization. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002753.

Wang:2022:TAM

[WMLM22] Yuqing Wang, Mika V. Mäntylä, Zihao Liu, and Jouni Markkula. Test automation maturity improves product quality — quantitative study of open source projects using continuous integration. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000280.

Wu:2023:MOE

[WMS23] Jie J. W. Wu, Thomas A. Mazzuchi, and Shahram Sarkani. A multiobjective evolutionary approach towards automated online controlled experiments. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000985.

Warnett:2025:MDM

[WNZ25] Stephen John Warnett, Evangelos Ntentos, and Uwe Zdun. A model-driven, metrics-based approach to assessing support for quality aspects in MLOps system architectures. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224003017.

Woods:2025:DRS

[Woo25] Eoin Woods. Dear researchers step 1: Find a team with a problem. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003625.

Wohlin:2022:ICS

[WR22] Claes Wohlin and Austen Rainer. Is it a case study?

— A critical analysis and guidance. The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001121.

Wen:2020:LSG

[WSL^{+}20] Melissa Wen, Rodrigo Siqueira, Nelson Lago, Diego Camarinha, Antonio Terceiro, Fabio Kon, and Paulo Meirelles. Leading successful governmentacademia collaborations using FLOSS and agile values. The Journal of Systems and Software, 164(??):??, June 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121220300303.

Wang:2023:LLO

[WTG23] Mingke Wang, Chuanqi Tao, and Hongjing Guo. LCVD: Looporiented code vulnerability detection via graph neural network. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001012.

Win:2023:EAP

[WTS23] Hsu Myat Win, Shin Hwei Tan, and Yulei Sui. Eventaware precise dynamic slicing for automatic debugging of Android applications. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

science/article/
pii/S0164121223000018.

Wang:2021:CWB

[WUK^{+}21] Dong Wang, Yuki Ueda, Raula Gaikovina Kula, Takashi Ishio, and Kenichi Matsumoto. Can we benchmark Code Review studies? A systematic mapping study of methodology, dataset, and metric. The Journal of Systems and Software, 180(??): ??, October 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121221001060.

Wang:2023:SCB

[WWC23] Tianlei Wang, Shaowei Wang, and Tse-Hsun (Peter) Chen. Study the correlation between the readme file of GitHub projects and their popularity. The Journal of Systems and Software, 205(??): ??, November 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223002017.

Wang:2020:IVS

[WXL^{+}20] Yingjie Wang, Guangquan Xu, Xing Liu, Weixuan Mao, Chengxiang Si, Witold Pedrycz, and Wei Wang. Identifying vulnerabilities of SSL/TLS certificate verification in Android apps with static and dynamic analysis. The Journal of Systems and Software, 167 (??):??, September 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030087X.

Wu:2023:AMT

[WXZL23] Lvyuan Wu, Zhiyu Xi, Zheng Zheng, and Xiaoli Li. Application of metamorphic testing on UAV path planning software. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001644.

Wang:2024:DWN

[WYFZ24] Ziyuan Wang, Min Yu, Yang Feng, and Weifeng Zhang. Do we need high-order mutation in faultbased Boolean-specification testing? The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122300328X.

Wei:2024:SEB

[WYG^{+}24] Changqing Wei, Xiangjuan Yao, Dunwei Gong, Huai Liu, and Xiangying Dang. Set evolution based test data generation for killing stubborn mutants. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.

sciencedirect.com/
science/article/
pii/S0164121224001663.

Wu:2020:CAL

[WZC^{+}20] Xiaoxue Wu, Wei Zheng, Xiang Chen, Fang Wang, and Dejun Mu. CVEassisted large-scale security bug report dataset construction method. The Journal of Systems and Software, 160 (??):??, February 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121219302304.

Wu:2025:MLR

[WZFD25] Di Wu, Hongyu Zhang, Yang Feng, and Zhenjiang Dong. MITU: Locating relevant tutorial fragments of APIs with multi-source API knowledge. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121224003406.

Wang:2025:IUO

[WZL^{+}25] Yong Wang, Huadong Zhou, Gui-Fu Lu, Cuiyun Gao, and Shuai Meng. Improving user-oriented fairness in recommendation via data augmentation: Don’t worry about inactive users. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122500055X.

Wang:2021:CDA

[WZZ21] Lei Wang, Yunqiu Zhang, and Xiaohu Zhu. Concept drift-aware temporal cloud service APIs recommendation for building composite cloud systems. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302922.

Xiaobo:2021:TRM

[XBS21] Yan Xiaobo, Liu Bin, and Wang Shihai. A test restoration method based on genetic algorithm for effective fault localization in multiplefault programs. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122030251X.

Xiao:2020:AAR

[XCH^{+}20] Jianmao Xiao, Shizhan Chen, Qiang He, Zhiyong Feng, and Xiao Xue. An Android application risk evaluation framework based on minimum permission set identification. The Journal of Systems and Software, 163(??):??, May 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300169.

Xin:2023:CCI

[XCZ23] Ruyue Xin, Peng Chen, and Zhiming Zhao. Causal-RCA: Causal inference based precise fine-grained root cause localization for microservice applications. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122300119X.

Xu:2022:EOC

[XDL^{+}22] Jia Xu, Ran Ding, Xiao Liu, Xuejun Li, John Grundy, and Yun Yang. EdgeWorkflow: One click to test and deploy your workflow applications to the edge. The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001479.

Xu:2021:CCS

[XLY^{+}21] Zhou Xu, Li Li, Meng Yan, Jin Liu, Xiapu Luo, John Grundy, Yifeng Zhang, and Xiaohong Zhang. A comprehensive comparative study of clustering-based unsupervised defect prediction models. The Journal of Systems and Software, 172(??):??, February 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220302521.

Xu:2021:UMR

[XTF^{+}21] Liming Xu, Dave Towey, Andrew P. French, Steve Benford, Zhi Quan Zhou, and Tsong Yueh Chen. Using metamorphic relations to verify and enhance Artcode classification. The Journal of Systems and Software, 182 (??):??, December 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001576.

Xiao:2024:CCN

[XXC^{+}24] Jianmao Xiao, Zhipeng Xu, Shiping Chen, Gang Lei, Guodong Fan, Yuanlong Cao, Shuiguang Deng, and Zhiyong Feng. Confix: Combining node-level fix templates and masked language model for automatic program repair. The Journal of Systems and Software, 216(??): ??, October 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121224001614.

Xu:2020:IML

[XZY^{+}20] Zhou Xu, Kunsong Zhao, Meng Yan, Peipei Yuan, Ling Xu, Yan Lei, and Xiaohong Zhang. Imbalanced metric learning for crashing fault residence prediction. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301837.

Xu:2025:BQE

[XZZ^{+}25] Yongrui Xu, Shang Zheng, Haitao Zou, Hualong Yu, and Shang Gao. Balancing quality and efficiency: an improved non-autoregressive model for pseudocodeto-code conversion. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002504.

Yarza:2021:LSM

[YAaO^{+}21] Irune Yarza, Mikel Azkarate-askatsua, Peio Onaindia, Kim Grüttner, Philipp Ittershagen, and Wolfgang Nebel.

Legacy software migration based on timing contract aware realtime execution environments. The Journal of Systems and Software, 172 (??):??, February 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302399.

Yilmaz:2024:FLB

[YB24] Rahime Yılmaz and Feza Buzluca. A fuzzy logic-based quality model for identifying microservices with low maintainability. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001882.

Yu:2023:SSL

[YCWD23] Siyu Yu, Ningjiang Chen, Yifan Wu, and Wensheng Dou. Self-supervised log parsing using semantic contribution difference. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000419.

Yuste:2022:EHA

[YDP22] Javier Yuste, Abraham Duarte, and Eduardo G. Pardo. An efficient heuristic algorithm for software module clustering optimization. The Journal of Systems and Software, 190(??): ??, August 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121222000851.

Yan:2023:EFL

[YJZZ23] Yue Yan, Shujuan Jiang, Yanmei Zhang, and Cheng Zhang. An effective fault localization approach based on PageRank and mutation analysis. The Journal of Systems and Software, 204(??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001942.

Yanakiev:2025:ASP

[YLC25] Ivan Yanakiev, Bogdan-Mihai Lazar, and Andrea Capiluppi. Applying SOLID principles for the refactoring of legacy code: an experience report. The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S016412122400298X.

Ye:2025:LSS

[YLG^{+}25] Jingwei Ye, Chunbo Liu, Zhaojun Gu, Zhikai Zhang, Xuying Meng, Weiyao Zhang, and Yujun Zhang. LogOW: a semi-supervised log anomaly detection model in open-world setting. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003492.

Yun:2024:PSC

[YLGS24] Shangbo Yun, Shuhuai Lin, Xiaodong Gu, and Beijun Shen. Project-specific code summarization with in-context learning. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001948.

Yang:2020:SSR

[YLHZ20] Yang Yang, Zheng Li, Liuliu He, and Ruilian Zhao. A systematic study of reward for reinforcement learning based continuous integration testing. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301965.

Yu:2025:AGN

[YLLY25] JunJie Yu, Yong Li, ZhanDong Liu, and QianRen Yang. Adaptive graph neural network protection algorithm based on differential privacy. The Journal of Systems and Software, 225(??):??, July 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000548.

Yang:2021:TEM

[YLW^{+}21]  Yingzhuo  Yang,

Zenan Li, Huiyan Wang, Chang Xu, and Xiaoxing Ma. Towards effective metamorphic testing by algorithm stability for linear classification programs. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001096.

Ye:2021:CSA

[YMDM21] He Ye, Matias Martinez, Thomas Durieux, and Martin Monperrus. A comprehensive study of automatic program repair on the QuixBugs benchmark. The Journal of Systems and Software, 171(??): ??, January 2021. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121220302193.

Ye:2022:VDV

[YML^{+}22] Jiaming Ye, Mingliang Ma, Yun Lin, Lei Ma, Yinxing Xue, and Jianjun Zhao. Vulpedia: Detecting vulnerable Ethereum smart contracts via abstracted vulnerability signatures. The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001236.

Yao:2023:FAB

[YOH^{+}23] Kundi Yao, Gustavo A. Oliva, Ahmed E. Hassan, Muhammad Asaduzzaman, Andrew J. Malton, and Andrew Walenstein. Finding associations between natural and computer languages: a case-study of bilingual LDA applied to the bleeping computer forum posts. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000468.

Yaman:2025:SVV

[YRC^{+}25] Sinem Getir Yaman, Pedro Ribeiro, Ana Cavalcanti, Radu Calinescu, Colin Paterson, and Beverley Townsend. Specification, validation and verification of social, legal, ethical, empathetic and cultural requirements for autonomous agents. The Journal of Systems and Software, 220(??):??, February 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224002735.

Yussupov:2021:FYD

[YSB^{+}21] Vladimir Yussupov, Jacopo Soldani, Uwe Breitenbücher, Antonio Brogi, and Frank Leymann. FaaSten your decisions: a classification framework and technology review of function-as-a-service platforms. The Journal of Systems and Software, 175(??):??, May 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.

                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000030.

Yilmaz:2022:USV

[YU22] Tolga Yilmaz and Özgür Ulusoy. Understanding security vulnerabilities in student code: a case study in a non-security course. The Journal of Systems and Software, 185(??):??, March 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221002430.

Yang:2024:CII

[YWCX24] Yilin Yang, Ziyuan Wang, Zhenyu Chen, and Baowen Xu. COPS: an improved information retrievalbased bug localization technique using context-aware program simplification. The Journal of Systems and Software, 207(??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/

science/article/
pii/S0164121223002637.

Yuan:2023:OSC

[YWLZ23] Dawei Yuan, Xiaohui Wang, Yao Li, and Tao Zhang. Optimizing smart contract vulnerability detection via multimodality code and entropy embedding. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000948.

Yang:2023:LHC

[YWS23] Kaiyuan Yang, Junfeng Wang, and Zihua Song. Learning a holistic and comprehensive code representation for code summarization. The Journal of Systems and Software, 203 (??):??, September 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001413.

Yasmin:2025:ESD

[YWTA25] Jerin Yasmin, Jiale Amber Wang, Yuan Tian, and Bram Adams. An empirical study of developers’ challenges in implementing Workflows as Code: a case study on Apache Airflow. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002929.

Yu:2022:DAP

[YWW22] Shiwen Yu, Ting Wang, and Ji Wang. Data augmentation by program transformation. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000541.

Yan:2021:SET

[YWY^{+}21] Rongjie Yan, Siqi Wang, Yixuan Yan, Hongyu Gao, and Jun Yan. Stability evaluation for text localization systems via metamorphic testing. The Journal of Systems and Software, 181 (??):??, November 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001370.

Yang:2024:QOT

[YXLH24] Huanhuan Yang, Ling Xu, Chao Liu, and Luwen Huangfu. Queryoriented two-stage attention-based model for code search. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003436.

Yoshimura:2025:MVL

[YYT25] Kentaro Yoshimura, Yuta Yamauchi, and Hideo Takahashi. Managing the variability of a logistics robotic system. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224003728.

Yu:2024:IFB

[YYW^{+}24] Ting Yu, Dongjin Yu, Dongjing Wang, Quanxin Yang, and Xueyou Hu. Iterative framework based on multi-task learning for service recommendation. The Journal of Systems and Software, 207(??): ??, January 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223002686.

Yang:2023:ETA

[YZC^{+}23] Guang Yang, Yu Zhou, Xiang Chen, Xiangyu Zhang, Tingting Han, and Taolue Chen. Exploit-Gen: Templateaugmented exploit code generation based on CodeBERT. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002539.

Yu:2022:EGG

[YZL^{+}22] Jiaojiao Yu, Kunsong Zhao, Jin Liu, Xiao Liu, Zhou Xu, and Xin Wang. Exploiting gated graph neural network for detecting and explaining self-admitted technical debts. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000036.

Yang:2025:DDL

[YZM^{+}25] Yanjing Yang, Xin Zhou, Runfeng Mao, Jinwei Xu, Lanxin Yang, Yu Zhang, Haifeng Shen, and He Zhang. DLAP: a Deep Learning Augmented Large Language Model Prompting framework for software vulnerability detection. The Journal of Systems and Software, 219(??): ??, January 2025. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224002784.

Yin:2022:SAS

[YZS22] Lu Yin, Junlong Zhou, and Jin Sun. A stochastic algorithm for scheduling bag-of-tasks applications on hybrid clouds under task duration variations. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122100220X.

Yang:2025:SEF

[YZS^{+}25] Xu Yang, Guofeng Zhang, Ziqi Shuai, Zhenbang Chen, and Ji Wang. Symbolic execution of floating-point programs: How far are we? The Journal of Systems and Software, 220 (??):??, February 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002863.

Yang:2024:CEJ

[YZZ^{+}24] Fengyu Yang, Guangdong Zeng, Fa Zhong, Peng Xiao, Wei Zheng, and Fuxing Qiu. CfExplainer: Explainable just-in-time defect prediction based on counterfactuals. The Journal of Systems and Software, 218 (??):??, December 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002267.

Zhang:2025:EAF

[ZAS25] Mingxing Zhang, Jun Ai, and Tao Shi. An empirical analysis of feature fusion task heads of ViT pre-trained models on OOD classification tasks. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000263.

Zhu:2022:DBV

[ZB22] Hong Zhu and Ian Bayley. Discovering boundary values of feature-based machine learning classifiers through exploratory datamorphic testing. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000097.

Zhao:2024:IPD

[ZCL24] Xiaofan Zhao, Tony Clear, and Ramesh Lal. Identifying the primary dimensions of DevSec-Ops: a multi-vocal literature review. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001080.

Zhang:2021:HCR

[ZCLP21] Fengyi Zhang, Bihuan Chen, Rongfan Li, and Xin Peng. A hybrid code representation learning approach for predicting method names. The Journal of Systems and Software, 180(??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221001084.

Zhao:2023:ESS

[ZCZL23] Yi Zhao, Hao Chen, Liang Zen, and Zhao Li. Effective software security enhancement using an improved PointNet++. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223001899.

Zaina:2022:PAB

[ZFC^{+}22] Luciana A. M. Zaina, Renata P. M. Fortes, Vitor Casadei, Leornardo Seiji Nozaki, and Débora Maria Barroso Paiva. Preventing accessibility barriers: Guidelines for using user interface design patterns in mobile applications. The Journal of Systems and Software, 186(??):??, April 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002831.

Zhang:2023:ETC

[ZFG^{+}23] Aiping Zhang, Liming Fang, Chunpeng Ge, Piji Li, and Zhe Liu. Efficient transformer with code token learner for code clone detection. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222002333.

Zabardast:2022:ASE

[ZFGH^{+}22] Ehsan Zabardast, Julian Frattini, Javier Gonzalez-Huerta, Daniel Mendez, Tony Gorschek, and Krzysztof Wnuk. Assets in software engineering: What are they after all? The Journal of Systems and Software, 193 (??):??, November 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001662.

Zhang:2022:TCP

[ZFS^{+}22] Quanjun Zhang, Chunrong Fang, Weisong Sun, Shengcheng Yu, Yutao Xu, and Yulei Liu. Test case prioritization using partial attention. The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001285.

Zabardast:2023:TAD

[ZGHG^{+}23] Ehsan Zabardast, Javier Gonzalez-Huerta, Tony Gorschek, Darja Šmite, Emil Alégroth, and Fabian Fagerholm. A taxonomy of assets for the development of softwareintensive products and services. The Journal of Systems and Software, 202(??):??, August 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000961.

Zhao:2024:TEH

[ZGLPA24] Tiange Zhao, Tiago Gasiba, Ulrike Lechner, and Maria Pinto-Albuquerque. Thriving in the era of hybrid work: Raising cybersecurity awareness using serious games in industry trainings. The Journal of Systems and Software, 210(??):??, April 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003412.

Zhao:2020:CRT

[ZGW^{+}20a] Shuai Zhao, Jorge Garrido, Ran Wei, Alan Burns, Andy Wellings, and Juan A. de la Puente. A complete run-time overhead-aware schedulability analysis for MrsP under nested resources. The Journal of Systems and Software, 159 (??):??, January 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121219302237.

Zheng:2020:IFP

[ZGW^{+}20b] Wei Zheng, Jialiang Gao, Xiaoxue Wu, Fengyu Liu, Yuxing Xun, Guoliang Liu, and Xiang Chen. The impact factors on the performance of machine learningbased vulnerability detection: a comparative study. The Journal of Systems and Software, 168(??):??, October 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220301229.

Zielske:2021:AAM

[ZH21] Malena Zielske and Tobias Held. Application of agile methods in traditional logistics companies and logistics startups: Results from a German Delphi study. The Journal of Systems and Software, 177(??):??, July 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221000479.

Zielske:2022:AMU

[ZH22] Malena Zielske and Tobias Held. Agile methods used by traditional logistics companies and logistics startups: a systematic literature review. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200070X.

Zhao:2023:HGW

[ZHLR23] Man Zhao, Rui Hou, Hui Li, and Min Ren. A hybrid grey wolf optimizer using oppositionbased learning, sine cosine algorithm and reinforcement learning for reliable scheduling and resource allocation. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001966.

Zhang:2023:ADV

[ZHM^{+}23] Lei Zhang, Sean Howard, Tom Montpool, Jessica Moore, Krittika Mahajan, and Andriy Miranskyy. Automated data validation: an industrial experience report. The Journal of Systems and Software, 197(??):??, March 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222002497.

Zhang:2024:SMM

[ZHS^{+}24] Weixing Zhang, Jörg Holtmann, Daniel Strüber, Regina Hebig, and Jan-Philipp Steghöfer. Supporting metamodel-based language evolution and rapid prototyping with automated grammar transformation. The Journal of Systems and Software, 214(??): ??, August 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224001146.

Zhang:2020:ADI

[ZJXG20] Yuwei Zhang, Dahai Jin, Ying Xing, and Yunzhan Gong. Automated defect identification via path analysis-based features with transfer learning. The Journal of Systems and Software, 166(??):??, August 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220300662.

Zhao:2024:TUI

[ZJY^{+}24]  Yangyang  Zhao,

Mingyue Jiang, Yibiao Yang, Yuming Zhou, Hanjie Ma, and Zuohua Ding. Towards an understanding of intradefect associations: Implications for defect prediction. The Journal of Systems and Software, 207 (??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223002534.

Zampetti:2022:ECS

[ZKDP22] Fiorella Zampetti, Ritu Kapur, Massimiliano Di Penta, and Sebastiano Panichella. An empirical characterization of software bugs in opensource Cyber–Physical Systems. The Journal of Systems and Software, 192 (??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222001315.

Zhou:2023:RPP

[ZLC^{+}23] Xin Zhou, Shanshan Li, Lingli Cao, He Zhang, Zijia Jia, Chenxing Zhong, Zhihao Shan, and Muhammad Ali Babar. Revisiting the practices and pains of microservice architecture in reality: an industrial inquiry. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001972.

Zheng:2025:ECV

[ZLD^{+}25] Wei Zheng, Chang Liu, Peiran Deng, Xiang Chen, and Xiaoxue Wu. Enhancing concurrency vulnerability detection through AST-based static fuzz mutation. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121225000202.

Zolduoarrati:2025:CCA

[ZLG25] Elijah Zolduoarrati, Sherlock A.

Licorish, and John Grundy. A crosscontinental analysis of how regional cues shape top stack overflow contributors. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000068.

Zhou:2020:ISB

[ZLS20] Cheng Zhou, Bin Li, and Xiaobing Sun. Improving software bug-specific named entity recognition with deep neural network. The Journal of Systems and Software, 165(??):??, July 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220300534.

Zolduoarrati:2022:IIC

[ZLS22] Elijah Zolduoarrati, Sherlock A. Licorish, and Nigel Stanger. Impact of individualism and collectivism cultural profiles on the behaviour of software developers: a study of Stack Overflow. The Journal of Systems and Software, 192(??):??, October 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222001327.

Zolduoarrati:2023:SSH

[ZLS23] Elijah Zolduoarrati, Sherlock A. Licorish, and Nigel Stanger. Secondary studies on human aspects in software engineering: a tertiary study. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000493.

Zolduoarrati:2025:SOH

[ZLS25] Elijah Zolduoarrati, Sherlock A˘2024 Licorish, and Nigel Stanger. Stack Overflow’s hidden nuances: How does zip code define user contribution? The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000421.

Zhou:2023:LML

[ZLSY23] Cheng Zhou, Bin Li, Xiaobing Sun, and Sheng Yu. Leveraging multilevel embeddings for knowledge-aware bug report reformulation. The Journal of Systems and Software, 198(??):??, April 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223000122.

Zhang:2023:DTG

[ZLW^{+}23] Fengji Zhang, Jin Liu, Yao Wan, Xiao Yu, Xiao Liu, and Jacky Keung. Diverse title generation for Stack Overflow posts with multiple-sampling-enhanced transformer. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121223000675.

Zhang:2024:EEF

[ZLW^{+}24] Yilei Zhang, Haoyu Liao, Zekun Wang, Bo Huang, and Jianmei Guo. EFACT: an External Function Auto-Completion Tool to strengthen static binary lifting. The Journal of Systems and Software, 215(??):??, September 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121224001377.

Zhao:2022:GRA

[ZLX^{+}22] Kunsong Zhao, Jin Liu, Zhou Xu, Xiao Liu, Lei Xue, Zhiwen Xie, Yuxuan Zhou, and Xin

Wang. Graph4Web:

a relation-aware graph attention network for web service classification. The Journal of Systems and Software, 190(??):??, August 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000681.

Zhou:2025:EPT

[ZLZ^{+}25] Xiyu Zhou, Peng Liang, Beiqi Zhang, Zengyang Li, Aakash Ahmad, Mojtaba Shahin, and Muhammad Waseem. Exploring the problems, their causes and solutions of AI pair programming: a study on GitHub and Stack Overflow. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002486.

Zhao:2023:BBE

[ZMLZ23]  Pengzhan  Zhao,

Zhongtao Miao, Shuhan Lan, and Jianjun Zhao. Bugs4Q: a benchmark of existing bugs to enable controlled testing and debugging studies for quantum programs. The Journal of Systems and Software, 205 (??):??, November 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002005.

Zakeri-Nasrabadi:2023:SLR

[ZNPR^{+}23]  Morteza  Zakeri-

Nasrabadi, Saeed Parsa, Mohammad Ramezani, Chanchal Roy, and Masoud Ekhtiarzadeh. A systematic literature review on source code similarity measurement and clone detection: Techniques, applications, and challenges. The Journal of Systems and Software, 204 (??):??, October 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223001917.

Zheng:2021:MOO

[ZPL21] Bingbing Zheng, Li Pan, and Shijun Liu. Marketoriented online biobjective service scheduling for pleasingly parallel jobs with variable resources in cloud environments. The Journal of Systems and Software, 176(??):??, June 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221000315.

Zhang:2024:HFE

[ZPSW24] Fan Zhang, Manman Peng, Yuanyuan Shen, and Qiang Wu. Hierarchical features extraction and data reorganization for code search. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002911.

Zhou:2021:ESO

[ZRGJ21] Zhide Zhou, Zhilei Ren, Guojun Gao, and He Jiang. An empirical study of optimization bugs in GCC and LLVM. The Journal of Systems and Software, 174(??):??, April 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220302740.

Zheng:2022:IAJ

[ZSCD22] Wei Zheng, Tianren Shen, Xiang Chen, and Peiran Deng. Interpretability application of the Just-in-Time software defect prediction model. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000218.

Zou:2023:DLF

[ZSF^{+}23] Yinglong Zou, Haofeng Sun, Chunrong Fang, Jiawei Liu, and Zhenping Zhang. Deep learning framework testing via hierarchical and heuristic model generation. The Journal of Systems and Software, 201(??): ??, July 2023. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223000766.

Zhou:2022:ASC

[ZSZ^{+}22] Yu Zhou, Juanjuan Shen, Xiaoqing Zhang, Wenhua Yang, Tingting Han, and Taolue Chen. Automatic source code summarization with graph attention networks. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121222000279.

Zhang:2023:AMV

[ZTK^{+}23] Miaomiao Zhang, Yu Teng, Hui Kong, John Baugh, Yu Su, Junri Mi, and Bowen Du. Automatic modelling and verification of Autosar architectures. The Journal of Systems and Software, 201(??):??, July 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223000705.

Zhang:2024:ICF

[ZWC^{+}24] Lingyu Zhang, Huiyan Wang, Chuyang Chen, Chang Xu, and Ping Yu. Incrementalconcurrent fusion checking for efficient context consistency. The Journal of Systems and Software, 207(??):??, January 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/

                                        science/article/
                                        pii/S0164121223002479.

Zhang:2025:DKC

[ZWF^{+}25] Shiyu Zhang, Xingya Wang, Lichao Feng, Song Huang, Zhenyu Chen, and Zhihong Zhao. DeepKernel: 2D-kernels clustering based mutant reduction for costeffective deep learning model testing. The Journal of Systems and Software, 219(??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002917.

Zhu:2024:MIL

[ZWL^{+}24] Yuhan Zhu, Jian Wang, Bing Li, Yuqi Zhao, Zekun Zhang, Yiming Xiong, and Shiping Chen. MicroIRC: Instancelevel root cause localization for microservice systems. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001900.

Zhou:2024:RCC

[ZWP^{+}24] Daihong Zhou, Yijian Wu, Xin Peng, Jiyue Zhang, and Ziliang Li. Revealing code change propagation channels by evolution history mining. The Journal of Systems and Software, 208 (??):??, February 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223003072.

Zhang:2022:RHE

[ZWY^{+}22] Zhiyi Zhang, Ziyuan Wang, Fan Yang, Jiahao Wei, Yuqian Zhou, and Zhiqiu Huang. Random or heuristic? An empirical study on path search strategies for test generation in KLEE. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121222000334.

Zhang:2023:SVD

[ZX23] Chunyong Zhang and Yang Xin. Static vulnerability detection based on class separation. The Journal of Systems and Software, 206(??):??, December 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121223002273.

Zhang:2024:SAR

[ZXH^{+}24] Chen Zhang, Jianwen Xiang, Rui Hao, Wenhua Hu, Domenico Cotroneo, Roberto Natella, and Roberto Pietrantuono. SGT: Agingrelated bug prediction via semantic feature learning based on graphtransformer. The Journal of Systems and Software, 217 (??):??, November 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224002012.

Zhou:2020:RAD

[ZXW20] Siwei Zhou, Jianwen Xiang, and W. Eric Wong. Reliability analysis of dynamic fault trees with spare gates using conditional binary decision diagrams. The Journal of Systems and Software, 170 (??):??, December 2020. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121220301849.

Zhu:2024:DSS

[ZXZS24] Jianfei Zhu, Guanping Xiao, Zheng Zheng, and Yulei Sui. Deep semisupervised learning for recovering traceability links between issues and commits. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224001547.

Zhu:2025:DBM

[ZYK^{+}25] Wenqing Zhu, Norihiro Yoshida, Toshihiro Kamiya, Eunjong Choi, and Hiroaki Takada. Development and benchmarking of multilingual code clone detector. The Journal of Systems and Software, 219 (??):??, January 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121224002590.

Zhao:2022:PLS

[ZYL^{+}22] Zhehao Zhao, Bo Yang, Ge Li, Huai Liu, and Zhi Jin. Precise learning of source code contextual semantics via hierarchical dependence structure and graph attention networks. The Journal of Systems and Software, 184 (??):??, February 2022. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121221002053.

Zhu:2023:SGA

[ZYZ^{+}23] Kuiyu Zhu, Ming Yin, Dan Zhu, Xiaogang Zhang, Cunzhi Gao, and Jijiao Jiang. SCGRU: a general approach for identifying multiple classes of selfadmitted technical debt with text generation oversampling. The Journal of Systems and Software, 195 (??):??, January 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S016412122200190X.

Zhu:2021:SDP

[ZYZZ21] Kun Zhu, Shi Ying, Nana Zhang, and Dandan Zhu. Software defect prediction based on enhanced metaheuristic feature selection optimization and a hybrid deep neural network. The Journal of Systems and Software, 180 (??):??, October 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121221001230.

Zhang:2025:MUP

[ZZH^{+}25] Jianzhang Zhang, Jialong Zhou, Jinping Hua, Nan Niu, and Chuang Liu. Mining user privacy concern topics from app reviews. The Journal of Systems and Software, 222(??):??, April 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL https://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121225000238.

Zhang:2025:NDL

[ZZJ25] Xinrui Zhang, Pincan Zhao, and Jason Jaskolka. Navigating the DevOps landscape. The Journal of Systems and Software, 223(??):??, May 2025. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL https://www.
sciencedirect.com/
science/article/
pii/S0164121224003753.

Zhang:2024:DPD

[ZZK^{+}24] Fengji Zhang, Zexian Zhang, Jacky Wai Keung, Xiangru Tang, Zhen Yang, Xiao Yu, and Wenhua Hu. Data preparation for deep learning based code smell detection: a systematic literature review. The Journal of Systems and Software, 216 (??):??, October 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001766.

Zhong:2023:MCB

[ZZL^{+}23] Chenxing Zhong, He Zhang, Chao Li, Huang Huang, and Daniel Feitosa. On measuring coupling between microservices. The Journal of Systems and Software, 200(??):??, June 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121223000651.

Zhu:2021:HKT

[ZZP21] Qianqian Zhu, Andy Zaidman, and Annibale Panichella. How to kill them all: an exploratory study on the impact of code observability on mutation testing. The Journal of Systems and Software, 173(??):??, March 2021. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

                                        URL http://www.
                                        sciencedirect.com/
                                        science/article/
                                        pii/S0164121220302545.

Zhang:2023:MMC

[ZZTC23]  Xiaopan  Zhang,

Furong Zhang, Zheng Tang, and Xingjun Chen. A MILP model on coordinated coverage path planning system for UAV-ship hybrid team scheduling software. The Journal of Systems and Software, 206 (??):??, December 2023. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121223002492.

Zhang:2024:CAC

[ZZY^{+}24]  Xiangyu  Zhang,

Yu Zhou, Guang Yang, Tingting Han, and Taolue Chen. Context-aware code generation with synchronous bidirectional decoder. The Journal of Systems and Software, 214(??):??, August 2024. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.
sciencedirect.com/
science/article/
pii/S0164121224001110.